A Bibliography of Publications in International Journal of Foundations of Computer Science

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 June 2024 Version 1.99

Title word cross-reference

#P [Zan91]. #P-Completeness [Zan91].

(−β) [Dom12]. (1, 0) [ZK19]. (1, 2) [BZ13]. (2 + p) [ZG13]. (2, 2) [ST16]. (2 · t) [PT19]. (3 · t) [PT19]. (3k + 1) [DZ00]. (A, B) [JL01]. (δ, α) [CCF09]. (δ, γ, α) [FG08]. (δ, κ_{δ}, α) [FG08]. (n, k) [Fen22, WC13, CDX21, CC98, CHYT14, HLHH06, YCL11]. (n, n(n + 1)) [NS98]. (n − 2) [XZW^{+}21]. (r, t) [WNF20]. 1 [CHWX09, Dic93, LR04, TCT14]. 11 [LJ17]. 2 [AV96, BYP95, CSN21, FFMW19, HKT00, HJP^{+}13, JZ16, JW08, Leo03, Pri06, TSFZRP17, XZS16, XCX17, ŽM11]. 2^{n} [CKZ17]. 2m [ZWCL14]. 3 [BYP95, DH96, JSPD03, KPS18, LJ17, SJ04, ST93, Tsi06].

3 log_{2} n [Far20]. 4 [XC15, ZZC15]. 7/3 [DSS15]. ^{2} [HK23]. ^{⋆} [MTVM15, ZLL20]. _{2} [Joh00]. A [XBE02]. ab ∗ c [KL03]. AQ_{n} [XZZY19]. ASPACE(log log n) [GP13]. β [Shu11]. c [CDFK19]. C^{1} [XBE02]. C_{k} [Yan21]. CT L [MTVM09]. CT L∗ [MTVM09]. I_{2} [BW14]. J [BL14]. R [BL14]. D [HLY^{+}04, AE99, DG98, RS01, YW20]. ℓ [DDHL11]. f [DGL93]. F_{p} + νF_{p} [WGF16]. \frac{7}{3} [Ram05]. g [LHD^{+}24, ZLL20]. G(2^{m}, 2) [YCTW10]. G^{xy+} [AT15]. G^{xy−} [BTO17]. GF(2^{n}) [WXF16]. H [GMU15]. K [BT07, CHWX09, PV98, ZBS05, Aku06, AAI^{+}20, AE99, BJD20, CSN21, DDHL11, DG98, DGL93, ESS20, EHS15, GWF^{+}24, IZN99, INY07, KPS13, LMZC20, LZ12, MXY^{+}04, MO23, Nak04, RS04, TCLS10, YTN01, YW20, Yan21, ZZZ16, ZK19]. K_{1,r} [RS22]. K_{m,m} [Kan15]. L [ADD^{+}18, PSS12]. L(2, 1) [LLW18]. L(j, k) [Cal15]. L_{p} [CMR07]. M [Jun14, PT18, PT19, Teh16a, Teh16b]. F_{2^{2m}}^{∗} [ZWCL14]. F_{d} [YW20]. Fq + uFq [YZP21]. Z_{p^{2}} [HSS19]. GF(2) [BB99]. UG_{b}(n, n(n + 1)) [Noc98]. μ [DL12]. N [AM09, Bed18, JM03, MGCVdlP20, PV98, Far20, INY07, LMZC20, LLW21, Yan21]. O [Mal07]. O(1) [ST99]. O(n) [MM97]. O(n^{2}) [Bad09]. ω [COT12, DI02, DÉK22, EJ23, Fin12, Fin21, Hon02, Hon07, KSV03, KMM06, Sel08, Sta05]. P [AFO06, ARV07, BGMV08, BCC^{+}11, BFM06, CD06, CCFS07, CVPV08, DI05, FOP05, Fre05, FO07, FIO08, FH11, GH07, IYD05, IW07, Iba11, ILT11, LZGN06, Luc09, Mad03, MDAPHPJ^{+}11, NSVA12, PDPPJ11, Pău00, PPJR06, PPJR07, PPJS07, PPRPS11, PBMZ06, PLMZ11, RCTC^{+}09, Sbu06, SRPC11, YDI08]. P_{4} [MR99, RRT99]. P_{6} [GV03]. P_{n,k} [YTN01]. π/3 [TH22]. q [BM16, BCMS20, FBK05]. R [FZCFB08, Wan21]. s [Dic93]. Σ_{2}^{p} [KL00]. σ_{f} = 2^{2n} + 2^{n+3}(n ≥ 3) [ZWW^{+}14]. Z [SMS92].

-Abelian [KPS13]. -Adic [XZS16]. -Ambiguous [BCMS20]. -Ary [AE99, DG98, LMZC20, Yan21, DZ00, RS01, YW20, PV98]. -Automata [KSV03]. -Calculus [DL12]. -Center [BJD20]. -Chains [DI02]. -Channel [Nak04]. -Clique [CSN21]. -Closed [AAI^{+}20]. -Collapsing [Pri06]. -Contamination [CSN21]. -Covering [ZBS05]. -Cube [LMZC20, LLW21, Yan21]. -Cubes [DG98]. -D [CHWX09, JW08, SJ04]. -Decomposition [Dic93, Joh00]. -Dependent [DGL93]. -DFA [AV96]. -Differences [Aku06]. -Dimensional [AE99, JZ16, LR04]. -Disjoint [BT07]. -Drawings [ADD^{+}18]. -Edge-Connectivity [Tsi06].

-Edge-Labeling [Cal15]. -Edge-Labelings [LLW18]. -Equivalence [Hon07, Hon02, PT18, PT19, Teh16a]. -Equivalent [Teh16b]. -error [YW20]. -Extra [Wan21, ZLL20]. -Fault-Tolerant [XZW^{+}21]. -Fibonacci [ESS20]. -Free [Bed18, GV03, RS22]. -Good-Neighbor [LHD^{+}24]. -Gram [FBK05]. -Hamiltonian [BZ13]. -Heap [Jun14]. -Independent [CSN21, TCLS10]. -integers [Dom12]. -Intersection [EHS15]. -Language [Fin12]. -Languages [COT12, Sel08]. -Like [HK11]. -LRC [WNF20]. -Matching [CCF09, FG08]. -Matchings [DGL93]. -Matrices [BM16, BCMS20]. -Means [CHWX09]. -Mesh [FZCFB08]. -Modified-Bubble-Sort [CDX21]. -out-of- [DDHL11]. -Packing [TSFZRP17]. -Partners [RRT99]. -Patches [XBE02]. -Periodic [CKZ17]. -Planarity [CDFK19]. -Plateaued [XCX17]. -Power [Sta05]. -Power-Free [DSS15, RS04]. -Powers [Shu11, Ram05]. -Pushdown [DÉK22]. -Qubit [JM03]. -Queens [MGCVdlP20]. -Rational [Fin21]. -Regular [EJ23, KMM06]. -Resilient [TCT14]. -Rooted [GWF^{+}24]. -Round [LJ17]. -SAT [ZG13, ZK19]. -Search [ZZZ16]. -Sided [ST93]. -Space [JZ16]. -Star [CC98, CHYT14, Fen22, HLHH06, WC13, YCL11]. -State [KPS18]. -Subgraph [GMU15]. -Substitution [Mal07]. -Super [ZK19]. -Systems [PSS12]. -Temporal [SMS92]. -th [YTN01]. -Tree [LZ12]. -Trees [IZN99, YTN01, JL01, PV98]. -Trivial [BL14]. -Truck [MXY^{+}04]. -Uniform [XC15, ZZC15]. -Union [EHS15]. -variable [ZWCL14]. -Vertex [Far20]. -Way [AM09]. -Words [ST16].

160 [WLC12]. 19 [Ibr22].

2-Adic [KK19]. 2-Designs [WDFN21]. 2-Dimensional [Che22b]. 2012 [SSS13].

2018 [Câm20]. 2021 [MR23]. 25th [MR23]. 2CCC [BE95]. 2ETIME [ABH17]. 2ETIME-Complete [ABH17]. 2NFAs [KM17].

3-Ary [LLW21]. 3-Disjoint [Par23b]. 3-Edge-Connected [ST11]. 3-Repetitions [GS12b].

60th [CVM20].

7 [DE08]. 7-Colourings [JP08].

’98 [GJV00a, HO00]. ’99 [MS99b, Pal01a].

ABE [HLC^{+}19, YMC^{+}17]. Abelian [AILR16, BSCH22, CRSZ11, CK16, CCI12, DR12, DMSS16, GRRS14, IMS03, KPS13, PP11, SS01]. Abstract [DG09, TZ91, WPX^{+}21]. Abstraction [ADHR09, ACV13, BPZ07, CFH^{+}03, MH06, NTSH06, WM13]. Accelerating [BIIN04, SNB24]. Acceleration [IN05, IN08]. Acceptance [GQZ15, Mer08]. Accepting [Das19, Dom04, DM08, HH20, IIT91]. Acceptors [BvdB18, IR14, Iba15]. Access [DCS13, Rud15, SK04, Sun00]. Accountable [YMC^{+}17]. ACD [Mar92]. ACD-Ground [Mar92]. ACE [YM19]. Achieving [JW08]. Across [CM12]. Action [HFLD09]. Active [DV11, JK14a, JK14b, PDPPJ11, PLMZ11, Qua07]. Activity [BGMV08]. Acyclic [AMR08, BPR09, FZFDCHB05, GVL07, KLB13, KS19, ZWS96]. Ad [AWF03, CIS03, LBJ03, SB12, WLF03, WD03]. Ad-Hoc [CIS03]. Adapting [CFG12]. Adaptive [BKS12, CLT14, CHYT14, KG11, LX94, LBJ03, SW09, TL99, Tse16, VJDT05]. Add [ANDZM09]. Addition [Wan04]. Additional [Dan22]. Additive [BLS20, SS07a]. Adic [KK19, XZS16]. Adjacent [AKS14]. Adjustable [HZZT12, WY05]. Adjusting [KSJ08]. Advanced [Qua07]. Advances [CDFK19, DRS23, HO00]. Adversary [BHK^{+}18b]. Advertisements [NH02]. Advice [BBB^{+}18, FH05, KSY14]. Aerial [Ami05]. Affine [BKP18, IKPY21, NKP^{+}22, Rov00]. Affirmative [PHPJRN^{+}11]. AFL [BJ07a]. Against [BCFR07, BHK^{+}18b, HMZ05, HCETPL^{+}12, HLC^{+}19, KMZS19, LWS^{+}20, TCT14, Uen13]. Agent [BF07, BDDN01, EH12, MM07, NH02]. Agents [DSS08, FHL07, LK11, LCVLV09, LRT92, MCS08]. Agglomeration [BYİT21, KB20, NZZ24]. Agglomeration-Based [BYİT21]. Aggregation [RGR11]. Agreement [BVM00, Gua21, KMZS19, MNS11]. Agreements [Tru08]. Aid [CMWZ19]. Alberto [SCIS15]. Algebra [GC15, GB03, Hea11, Lar99]. Algebraic [BM16, BMW91, BÉ11, CD21, FH05, GW24, HLC^{+}19, Kri97, TCT14, TJZ13, ZWCL14]. Algebras [ALR04, Ali16, BE92, BE93, JPŠ19, KLB13, MRT95, Ole92, SN13, TST01a]. Algorithm [ATK12, ANDZM09, ARS11, BV08, BB04, BKS12, CPY02, CF06, CFRD08, CDJ09, CTZ01, CL03, CLT14, CHYT14, DGN07, DN16, DG98, FL09, FZAM08, FJ12, Fri10, Fuj17, GKSZ19, GLV14, Gro03, GD12, GWL^{+}17, HKV17, Hei97, HO99, HM04, HW17, Hut02, IST05, IZN99, JHK08, KK10, Kar99, Kör03, KTT20, LW93, Li01, LJH^{+}17, LCL06, MDAPHPJ^{+}11, MOSZ18, MTNN99, MS19, MC13, NGHK15, NWHL22, Nis07, Okh06, PRN13, PYTH10, PR00, PS22, Pym92, QFL^{+}15, RN22, SRN^{+}20, SH22, SW09, SS07b, ST99, SKW08, Tor13, TSFZRP17, Tsi06, WG17, Won96, Won01, XS11, XWL^{+}22, YW22, ZSG^{+}22, ACM11, CCM11]. Algorithmic [BS12, CFMR05, DGMM15, GGR14, HPV99, Riv04, SR21].

Algorithms [AFB96, Aku06, AILR16, AC05, AMR05, AMR11, ADD^{+}18, AE02, AE05, Ars15, AMOZ07, BT07, BRM07, BH02, BCFL12, Bur12b, CD15, CCM97, CCF09, CFG12, CGKN08, CHWX09, CD20, CHA^{+}92, CPC99, CHZ06, CCG^{+}11, DP90, DPS99, DD13, DGL93, DWS15, DMSS16, ERW04, ECY02, FK19, FZ15, FZEBB05, FPPS03, FA06, GO09, GHJS05, Gol90, GM19, GKS^{+}19, HL06, HP09b, HLW09, IMP12, INY07, IMS03, JMSO05, JZ16, KSMMT18, KKH90, LTW02, Leu04, Li12a, LMM^{+}12, MPS99, Mas04, MPS24, Moh02, Moh03, Nak04, NB18, OSZ92, RLWW96, SRR15, Sah01, SK01, SK20, SK03, SL21, SJ04, SG04, Ste93, TV07, Tor15, TL99, Tse16, WRNK03, WM05, WH03, ZBS05, Zom03, FG08]. Alignment [AES18, AE02, BBM^{+}12, CK08b, FM96, GD12, PYTH10, TFF18]. Alignment-to-Alignment [FM96]. Alignments [CCP18]. Alive [BC12]. Allocation [BRSRC11, NWK06, PS22, WG17]. Allowing [Asa23]. Almost [BN20, BKST18, Far20, HJ13, Kur20, PS12a, PP11]. Almost-Equivalence [HJ13]. Almost-Group [BN20]. Almost-Universal [BKST18]. Alphabet [Dom12, GNP^{+}06, JMR91, JJS08, Jir11, KMRY20, Mas19, Pig15]. Alphabet-Independent [GNP^{+}06]. Alphabet-Invariant [KMRY20]. Alphabets [CTS18, Leu16, Mas13, NR21]. Alternate [ESS21]. Alternating [AK14, BCPR07, CLLL08, GZY24, HIIW01, HIR^{+}92, IIT91, JK19, MO10, Slo95, ZZZ23]. Alternation [HK23]. Alternative [dSMOC18, Set08]. Ambiguity [AMR11, Iba15, KMK11, Leu05, MS04, MPJ07, Şer09, SL17]. Ambiguous [BCMS20, Mig90, Pau24]. Amenable [Ble21]. American [SGZ02]. Amiable [Ata07]. Among [DDPS19]. Amount [BGRY16]. Amplitudes [Nis03]. Analog [LWJ^{+}10]. Analog/Mixed [LWJ^{+}10]. Analog/Mixed-Signal [LWJ^{+}10]. Analyses [KPM15, Tse16, ZPXX17]. Analysis [AHL^{+}13, AT23, AH07, BYP95, BV98a, Bee95, BAK12, BCB12, BYİT21, BET03, DN16, DES09, EH12, FK19, FSWF11, FZAM08, FBK05, Gol90, HP09b, HM04, IDR97, Ibr22, KR97, KM23, Leo03, LCY12, Li12a, LC22, LN08, LPP92, Lug11, MH06, MGGP08, NAK^{+}15, OM96, PV98, RWZ01, ROK08, Set08, TY03, TY23, TV94, Wan04, WR16, Yam03, YLZ14, YB06, Yen08, ZZZ16, ZL22, ZWC^{+}22, ZZC22]. Analytic [BMMR11]. Analyzing [CCP18, DW04]. Anarchy [FFMW19]. And/Or [FIO08, DW04]. Angle [MB17]. Annealed [SA22]. Annotated [KSJ08]. Annotation [BDL08]. Announcement [CIS16, IS21]. Anonymity [TFS19, ZYZ^{+}19]. Anonymous [AOSY10, FDFZB12, Špr09, XS06]. Answer [PHPJRN^{+}11]. Ant [KAPF05, dMLBPP20]. Antennae [AC05]. Antennas [TH22]. Anti [BJ07a, KMG11]. Anti-AFL [BJ07a]. Anti-Spikes [KMG11]. Antidictionary [Shu14]. Antimirov [AMR09]. Antiport [AFO06, ARV07]. Any [PS12b, TSFZRP17]. Anytime [CD15]. Aperiodic [BS92, BS15, DJR18, Sel08]. Apices [MAN06]. APN [XC15, ZH13]. Apostolico [SCIS15]. Appearances [DDD18]. Application [Cas05, MNS11, PB20, RN22, SB01, URS07, ZH06]. Applications [BKST18, CK08a, Câm20, CCF09, CHWX09, CW11, CB09, CK18, DI02, Fin12, GC15, GGR14, HBN08, KL03, KKS05b, KMS11, KM90, Li07, LL23, MM97, PRS98, PYTH10, Suc90, Zom01c]. Applied [dMLBPP20]. Approach [BET03, BMMR11, CLMP16, CMMR04, CMWZ19, EAB^{+}16, GSD03, HMZ05, IMP^{+}05, Kri97, KSM22, LW06b, MG14, MGGP08, Qua07, SGZ02].

Approximability [DJL^{+}07]. Approximate [BH02, JLL23, MRRV06, NRS18, ORS08, WKS^{+}08, ZBS05]. Approximated [BB04]. Approximating [BR08, BVM00, BDG^{+}11, Fre02, Gol14, HL01, LZ12, Rya15, YJ05]. Approximation [AE02, AP90, ABDP05, BLS20, CS93, CCG^{+}11, GY12, GM19, HJP^{+}13, HW17, JMSO05, JSO10, KK10, LTW02, MPS24, NB18, SK20, SL21, SS07b, Ste93, Tei17, WG17, XS11]. Approximations [RV22, Shu07]. Aquatic [YLX22]. Arbitrage [DLW02]. Arbitrarily [BSOR10]. Arbitrary [EZ01, GS12a, HKV17, Hei97, JWB03, LOPR18, NGHK15, XHLF02]. Arc [GP17, KHLC12]. Architecture [MDL97, YLZ14]. Architectures [AP92b, CPJ06, PR23]. Arcs [MM97, RR18]. Area [CR14]. Arithmetic [BB03a, FMC04, FT11, GK11, ŠM05]. Arithmetical [Okh05]. Arity [CL07b, DZ00]. Arrangement [FWZ15, LX17]. Arrangements [KL05]. Array [CE98, FS06, GPC09, Jun14, LC18, ZYYH14]. Arrays [AE99, Fre05, MMP10, PA98, SMAN13, WH03]. Arthur [CCPS04, Vin05]. Articles [FRS24]. Articulation [Kar99]. Artin [AR16]. Ary [AE99, DG98, LMZC20, LLW21, PV98, Yan21, DZ00, RS01, YW20]. Asian [HO00, GJV00a]. Aspects [BM16, BRST07, HK09a, Riv04, SR21]. Assembly [BHR09, BKLS20, IPR07, IP08, JK14a, JK14b, Rog09, RCTC^{+}09, SW17]. Asset [XWL^{+}22]. Assignment [Bar90, DGN07, GSD03, Hir91, NSVA12, WD90]. Associated [MSMR22, Sal11]. Association [TBGP20]. Associations [YZY^{+}18]. Assume [LSWW13]. Assume-Guarantee [LSWW13]. Assumptions [GKS17]. Asymmetric [Gol14, KMŠ21, WR16]. Asymmetry [FPS02]. Asymptotic [FY08, PR12, Szw95]. Asymptotically [CDPR11]. Asynchronous [Ott15, Yue13].

Asynchrony [SR00a]. ATM [GKKP99]. Atomic [Anc02]. Atomicity [WPX^{+}21]. Atoms [BT13, EKKS18]. Attack [DS02, DEKZ11, HCETPL^{+}12, LJ17, WLC12]. Attacks [BNBN20, DEKZ11, HLC^{+}19, LWS^{+}20, TCT14]. Attention [ZZC22, ZSG^{+}22]. Attraction [HKRS19]. Attribute [BV08, TYM^{+}17, WHLH17]. Attribute-Based [TYM^{+}17, WHLH17]. Auditing [LWS^{+}20]. Augmentation [NS13, YH11]. Augmented [GRI24, XZZY19, ZLL23]. Augmenting [GKS^{+}19]. Authenticated [LHT09, LH11, MMS17]. Authentication [BKST18, Gua21, HCETPL^{+}12, LB04, YTP11]. Author [Ano97, Ano98, Ano99, Ano00, Ano01a, Ano02, Ano03a, Ano04a, Ano05a, Ano06, Ano07, Ano08, Ano09, Ano11, Ano12, Ano13, Ano14, Ano15, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21, Ano22]. Authorized [WZCH19]. Auto [CGKN08]. Auto-Intersection [CGKN08]. Autocorrelation [KYZS17]. Automata [AHK07, ABH^{+}09, AK14, AMR11, ACMP20, AMR08, AR16, ACFE09, ABH17, AHK17, BBP11, BK24, Bed18, BHK19, BH20, Ber13, BN20, BMP03, BCD14, BMP15, BHK18a, BCPR07, BCHK09, BCV23, BHK07, BRST07, BKM11, BKM12, BKM15, BW14, BMMR11, BMMR12, BKW02, CFM12, CFM13, CPY02, CLW09, CL15, Cha02, CLOZ04, CC05, CCR^{+}90, CFY16, CG06, CR15, CMR07, CMRR08, CVMVMV00, CKK02, CTS18, DJ12, Dom04, Dro92, DK98, DM11, DP14, DÉK22, DFK23, D’s03, Dub95, ÉM11, Ési12, FGS^{+}90, FTT10, FMR20b, Fin19, FHKK23, Fre08, FK13, Fuj17, GLV14, GI22, GHWZ05, GVL07, Glö07, Glö10, GSZ99, GH13, GH15, GQZ15, GC18, GPP20, Gus13, GP15, HHNP23, HMZ05, HW05, HK09b, HJ13, HJ17, HJK18, HKMW22, HKKŠ13, IJT^{+}93, IM21, IJMP21, JM13]. Automata [JJS08, JJŠ18, JK19, JO07, JK07, KZ10, Kop21, Kör03, KR16, KBH99a, KSV03, KMS06, KSY14, Kud07, KL11, KMM06, KR08, KMO10, KO13, KMW14b, KMW14a, KMW16, KO18b, LL20, LP19, Löd15, LT24, Loh10, Mac96, MS20, MMR20, Mal05, MR11, MvZ22, Mar08b, MVMM02, Mar97, Mar09, Mas13, MHT09, MZ12, MO07, MO09, MS18, Mod21, Moh03, Moh13, MP91, MPJ07, Nak18, NKP^{+}22, NTSH06, NWK05, NWK06, NCC^{+}07, OS19, Oli13, Ott15, Pau24, PI95, Pig09, PP14, Pig15, POM22, PM13, SYS19, SS07a, Sao92, SY12, ŠM07, Sir15, Slo95, SVF09, Sut03, Tam08, Tor13, Tor15, TY15, Vor16, Vor18, WM13, WKS^{+}08, YDI08, YW06, YBI11, ZHZ11, ZZ18, ZQL12, dBDZ19, CV13, Câm20]. Automata-Based [Tor13]. Automated [CGR13, KM02, Pen93, TW09]. Automatic [ADR11, BCDP08, BK16, CRS12, DMSS16, GHS13, GRRS14, LD01, Loh05, LBL06, MH06, RS15, SS12a, SF07]. Automaticity [MRSS19]. Automaton [AČ11, AMZ20, BGK^{+}20, CZOdlH17, CL14, CC05, CGL12, IT13, JHK08, KPS18, MOSZ18, Okh03, Pol05, Prů17]. Automaton-Based [Okh03]. Autonomous [BFMBS11]. Auxiliary [DZ00, KR16, LMG20]. Auxiliary-Input [LMG20]. Average [BLP18, BGN10, BMMR11, BMMR12, BMMR19, CS93, DN16, FZAM08, KMIS09]. Average-Case [BLP18]. Averaging [CM12, Ste11]. Avoidance [Sha04, SH22]. Avoiding [AGM19, CRSZ11, GS12b, KMŠ21, ORS08, Ram05, WAG^{+}06]. Aware [LBJ03]. Axiomatic [Bur12b]. Axioms [HST01]. aying [FMV13].

B [Lag17, LF96, OM96]. B-Trees [Lag17, LF96, OM96]. Babai [GGJ^{+}19]. Back [Asa23, GH15]. Backbone [FPPS03]. Backtracking [MT95b]. Backward [FL09]. Backward-Oracle-Matching [FL09]. Bad [KMZS19]. Badger [NWHL22]. Balance [JL01, LF96, MMR10]. Balanced [CZTH13, CS00a, EJ23, Fle96, GW24, Lag14, LW93, LX19, LL16, LW21, MX11, RAB15, YTP11, ZWW^{+}14]. Balanced-by-Construction [EJ23]. Balancedness [LLS21]. Balancing [Hei97, MD00, ST01]. Banded [BL01]. Bandwidth [GR03]. Banishing [HJV93]. Banyan [KR97]. Barrier [GM19, Uen13]. Base [DRDN08, FZ03, Hon06, MP91]. Base-Station [DRDN08]. Based [ADR11, ARS11, AEMY21, ABL^{+}11, AH07, BCB12, BYİT21, BK95, BBE24, BNR99, BDDN01, BKS12, CCM11, CP06, CDPT16, CCD07, CST^{+}17, CK18, CVDV10, DPS93, DEZ01, FDFZB12, FZT14, GWL02, GR03, HK02, HO99, HW10, JC03, JK07, KMZS19, KB20, LYX^{+}19, LHT09, LTZ12, LC22, LH11, LYHW19, LYY^{+}21, Luc09, MLO17, MM07, MMS17, MMS05, ND02, NKW08, NZH22, NSVA12, NZZ24, Okh03, PRN13, PR23, Qua07, RHN^{+}22, RK09, RN22, RR04, SB12, ST01, SH22, SL17, SZ22, TWZ11, TYM^{+}17, TBGP20, Tor13, Tor15, Tse16, TFS19, VG01, Ver09, WNF19, WWT20, WNF20, WHLH17, WD03, WZCH19, XHLF02, XCX16, YTLC02, YW22, YW06, ŽM11, ZYZ^{+}19, ZZC22, ZZN23, ZSG^{+}22, ZPXX17, ZGCZ18, vLW15, FBK05, WLZT21, ZWCL14]. Basic [BV08, Vor18]. Basis [Sub90a, Sub90b]. Basketball [SH22]. Batch [DFLL02, LLQ06, PY04, ZPXX17]. Bayesian [ZLW^{+}17]. BDD [FBK05]. BDD-based [FBK05]. Be [AAV00]. Becomes [KM07b]. Beeps [EP17]. Before [BSS12]. Begins [BSS12]. Behavior [AC05, EH12, LWW22, SB01, TCT14]. Behavioral [BCB12]. Behaviors [PQ06]. Behaviour [BMMR19, PR12]. Belated [Tse16]. Belief [RHN^{+}22]. Benford [Rav08]. Bent [XCX17, ZLL11]. Bernays [RS95]. Beta [CS18, Kop21]. Beta-Shifts [CS18, Kop21]. Better [KH21]. Between [BCV23, CLT09, CNT22, Faz08, Fia08, GGJ^{+}19, HKS13, HN10, KA18, Láz13, Nag21, Sal13, YZY^{+}18, ZWS96, LL20]. Beyond [FGH^{+}07, HJ13, RKRR02]. Bi [GV03, NS13]. Bi-Cographs [GV03]. Biautomata [HJ14, HJ16]. Bichromatic [MB17]. Bideterministic [Tam08]. Bidirectional [BGM^{+}18, GMNS15]. Bifurcation [APMP17]. Big [MLO17, MMS17, ZLW^{+}17]. Bilateral [YLX22]. Bimonoids [DP14]. Bimorphisms [MT10]. Bin [BDI^{+}11, FFMW19, HJP^{+}13, JZ16, MV11]. Binary [Ata07, BMS18, CdBD23, CRSZ11, CDJ09, CKZ17, CS00a, DSS15, FLFR19, GWF^{+}24, HH12, HH11, HH24, Har24, HFLD09, Hol11, IN08, JS03, KYZS17, KK90, LZGN06, Mas19, NWHL22, OW92, PS12b, RAB15, Sal07, Sha04, Smy12, Vor16, WD20, XZS16, YB06]. Binding [AES18, AB17b]. Binoid [GN11]. Binomial [ZZC15]. Bio [DH05, MB06]. Bio-Computation [MB06]. Bio-Operation [DH05]. Bioinformatics [KKS05b]. Biological [LJH^{+}17]. Biology [RCTC^{+}09]. Bipartite [FGV99, GV03, LMZC20, LV08, Par23b, Toš06, WQY16, Won96, Won01]. Bipartitioning [HT95]. Bipartization [LL23]. Bird [Ami05]. Birthday [CVM20]. Bisemigroup [GN11]. Bisimulation [AHK07, ABH^{+}09, MC13]. Bisplit [GV03]. Biswapped [NAS22]. Bit [BT17, CF06, CCF09, DD13, DES09, HN06]. Bit-Parallel [CF06, CCF09, DD13, HN06]. Bit-Split [DES09]. BiTCN [ZZC22]. BiTCN-Attention [ZZC22]. Bitonic [INY07]. Bitwise [FNI16]. Bivariate [TWZ11]. Black [CS96, DSS08, HHP17, MC02]. Black-Box [HHP17]. Blackbox [WCD^{+}14]. Blackwell [GZ12]. Block [BLLS03, FLM^{+}21, LJ17, MRRV06]. Blocking [Dai97]. Bloom [Sal18]. Blow [JJS08]. Blow-Ups [JJS08]. Blum [Câm14].

BNF [dMLBPP20]. Bond [KKS05a]. Bond-Free [KKS05a]. Bonsai [PPR18]. Boolean [BB99, BJY90, BLY12, CM92, CH15, Car11, CLMP16, DQFL12, ÉK07, FY11, GW24, Hea11, HSS07, IP08, JK19, KY90, KSM22, LO10, LHG11, LT24, Okh06, dSMOC18, PP11, Sch10, SS01, SFL17, SH17, TCT14, TJZ13, YKCW23, ZWCL14, ZWW^{+}14]. Boosting [AKMW20]. Bootstrap [DVG03]. Bordered [GRRS14, KM07a, KM08]. Borders [BSCH22, ŠM07]. Bottlenecks [JYF91]. Bottom [AMZ20, FSM11, Gaz06, Mal15]. Bottom-Up [AMZ20, FSM11, Gaz06, Mal15]. Bound [BBP11, CE98, FY08, HPP99, SSF20, Uen13, ZSW14, ZG13]. Boundaries [DFK23]. Boundary [DRDN08, EH15, Fre02]. Bounded [AMT20, AEMY21, BLM04, CFM12, CRSZ11, DDD18, De 06, DFLL02, DGMM15, FLM^{+}21, FV24, FCS05, IJT^{+}93, IS12, JLL23, JZ16, LNP16, LZ93, MMP10, Mee12, Pet11, PZX07, Vik96, WLF03]. Boundedness [vdM00]. Bounds [AF20, ADD^{+}18, BKM15, BE19, DH18, Dom04, DSS15, FKM^{+}21, Gus13, HHH07, JWB03, KM22, LHG11, MV11, PL23, SNJ11, Uen13, WNF19, XCMT20, YS13, ZK19, dBDZ19]. Box [HHP17]. Boyer [CFG12]. BP [SZ22]. Brain [RHN^{+}22]. Branch [HPP99]. Branch-and-Bound [HPP99]. Branching [Bed18, KS19, PSA17]. Brane [CP06]. Breadth [CCR^{+}90]. Breaking [Uen13]. Bridge [Láz13]. Bridges [GD98]. Bringing [Asa23]. Broadcast [Anc02, CFMS15, LAHN14, Nak04, PZX07, RMZW19]. Broadcasting [CYS^{+}12, HT09, PP06, WD03, XLC^{+}04]. Broken [AAV00]. Brownian [Nis07]. Browsing [DE08]. Bruijn [BGM^{+}18, KX12, Noc98, NS98, WWT20, WRNK03]. Brute [CCP05]. Brzozowski [DN16, GLV14, SKW08]. Bubble [CDX21, ZH19]. Bubble-Sort [ZH19]. Büchi [FKV06, KL11, LL20]. Buffer [DLC^{+}14]. Bulk [CCG^{+}11, FNI16]. Bundles [LWW00]. Burnside [KPS18]. Bursty [SK04, SB17]. Buses [BT00, Mat04, PA98, WH03]. Buy [CCG^{+}11]. Buy-At-Bulk [CCG^{+}11]. Byzantine [CGK^{+}21, PP06].

Cache [Leo03]. Caching [BLR09]. Cactus [TSFZRP17]. Calculi [AH07]. Calculus [BDSV06, CP06, DL12, Kri92, Oga00, PT90, Pym92, RS95, Yue13]. Can [AAV00]. Cannot [KPS18]. Canonical [BJ05, BJ06, BJ07b, CC05, FGV99, GSZ09, MAN05, WM13]. Cantor [Ési12, Sta05]. Capability [Gua21]. Capacities [Li12b]. Capacity [AF20, BKM12, DST10, FL97, Li12b, Zet11]. Capital [NZH22]. Captures [DW03]. Capturing [FW90, ISAZ08]. Cardinality [NWHL22]. Cardinals [Fin19]. Care [Ros03]. Careful [Vor16]. Carpi [Ber11]. Carriers [GH07]. Cartesian [MRT95, Ole92]. Cartesian-Closed [MRT95]. Cascade [WGD18]. Cascading [Sal18, Sub05, Wan14]. Case [BLP18, BMS12, BDC90, Das19, DN16, FK06, Fle96, KP10b, Lag17, Mal24, PSA17, YH11, ZSW14]. Cases [BCR11]. Catalytic [HFLD09]. Categorical [Sak01]. Categories [MOM91, Oli13, RGR11]. Category [ÉM11, MRT95, Ole92]. Catenation [CLMP16, CGKY11, CGKY12]. Catenation-Intersection [CGKY11]. Catenation-Reversal [CGKY12]. Catenation-Star [CGKY12]. Catenation-Union [CGKY11]. Catenations [Mel93]. Caterpillars [AB17a]. CATT [ZWC^{+}22]. Cauchy [Ruo96]. Cauchy-Peano [Ruo96]. Causal [BCB12]. Cayley [BK16, CP99, CL07a, LLL22, XZY19]. CCA [SZFX20]. CCA-Secure [SZFX20]. CCZ [BH11]. CCZ-Equivalence [BH11]. Cd [FO08, BCVVH07, CVDV10, MMK22, MO09, Sun05]. CD-Systems [MO09]. CDF [RN22]. CDF-DWT [RN22]. CDS [Fuj16]. Cell [AFO06, RCTC^{+}09]. Cell/Symbol [AFO06]. Cellular [BK24, DJ12, Dub95, FZ03, GSD03, JHK08, Kop21, Mar08b, Mod21, Sir15]. Center [BJD20]. Centralized [Ott13]. Cerný [GGJ^{+}19, Ste11]. Certain [GRI24, KRK16, Sal11, Won01]. Certificate [ZGCZ18]. Certificate-Based [ZGCZ18]. Certificateless [DZH16, SZFX20]. Certify [GHWZ05]. Chain [CCQ24, GV23, GSZ99, JSO10, YLX22]. Chain-Free [GV23]. Chains [DI02, DHR08]. Chandra [KMW12]. Change [CTS18]. Changes [LZ93, Vik96]. Channel [BBL^{+}12, BNS03, GSD03, HLH19, NN93, Nak04]. Channels [MG14, YBM11]. Chaos [EMRB12]. Characterisation [D’s03]. Characteristic [Çev20, IB12, LCXS19, WNF19]. Characteristics [OS01]. Characterization [ÉI14, KM17, MM05, MCS08, Mar08b, Okh05, OS93, Ric19, RW11, YTN01]. Characterizations [IS12, IM21, JM03, KSV00, OY11, PPJY08]. Characterize [MMK22]. Characterizing [IW07, JC03]. Charts [EGPS10]. Checking [CGR13, CFH^{+}03, EHK06, HW10, IM21, LD01, Sch10, YW06]. Checkpoint [PNN^{+}10]. Checkpointing [GCK08, MM07, YSM^{+}00a]. Chemical [HFLD09, KPM15]. China [NZH22, SZ22, ZL22]. Chinese [ZZC22]. Choffrut [BMY17]. Chomsky [DV14, PPJY08]. Chongqing [SZ22]. Choose [INY07]. Chord [CCF08]. Chordal [FHL07, NS13]. Christian [BMY17]. chunk [AP92a]. Church [AD12, KM07b]. Ciliate [DH05]. Ciliates [BHR09]. Cipher [LJ17]. Ciphertext [SZFX20, YM19]. Circle [Klo96a]. Circuit [Bir11, LWJ^{+}10, RVT06, Vin05]. Circuit-Size [Bir11]. Circuits [FGH^{+}07, GB03, GRB03, IP08, PRS98, SUZ13, YB06]. Circulant [HSS19, SZQS18, YCTW10]. Circular [Asv07, DS96, GP17, MM97, MMR10]. Circular-Arc [GP17]. Circulating [SK01]. Circulation [GS12a]. Circumscription [Lis93]. City [SA22]. CKY [BIIN04]. Class [AGM14, BS92, CPJ06, ERW04, Has00, Jai95, KK19, MR11, MN00, Oka99, Sch13, TCT14, WDFN21]. Classes [Arv97, AP90, ABDP05, CCPS04, CM92, Cap96, DY19, GO09, Géc07, GR00, HT12, HK95, KSV00, LV08, NCC^{+}07, SH17, UU07, XZS16, XCX17, vLW15]. Classic [IN13]. Classical [BMP15, Fia08, Oga00, ZQL12, CV13]. Classification [ATK12, RHN^{+}22, SKL03, ZSG^{+}22, ZLW^{+}17]. Classifying [SWZ97]. Claus [HHH07]. Clauses [FGL^{+}90, SN13]. Clique [BLM04, CSN21, DJL^{+}07, GR00, LV08, MR99, MM97, Ste93, SK20]. Clique-Width [BLM04, GR00, LV08, MR99]. Clock [D’s03]. Close [Fre02]. Closed [AAI^{+}20, MRT95, Ole92, TBGP20, TW09]. Closed-Set-Based [TBGP20]. Closeness [AO11, Dan11, Dan22, YB19, ZLG21]. Closure [CK08a, DMMM14, HIIW01, LT24]. Closures [BGS11]. Cloud [MLO17, MGJ19, WHLH17, YMC^{+}17, ZLW^{+}17]. Cluster [ABL^{+}11, BBP11, Ber13, BNR99, IN08, URS07]. Cluster-Based [ABL^{+}11, BNR99]. Cluster-Dot [IN08]. Clustered [Che22a, CDFK19, FPP03]. Clustering [BKS12, CL03, CHWX09, ECY02, FPPS03, LC22, MMS05, ZC05]. Clustering-Based [LC22]. Clusters [BLMR05, CFMR05, CVOV11, LCVLV09, SK03]. CMP [For10]. Co [BLM04]. Co-Gem-Free [BLM04]. Coalgebras [Oli13]. Coarse [MS99a]. Cobham [Kre21, MRSS19]. COCOON’02 [IZ04]. Code [DK12, KHS21, LYHW19, ND02, PR11, Rud15, TY23]. Code-Based [LYHW19]. Codebooks [SSF20]. coded [GP13]. Codes [AGM14, BKST18, Bur12a, CCQ24, CFPR03, DT20, FLFR19, GMNS15, GRB03, HS11, HSS19, Kun16, Leo03, LZ15, SZQS18, WGF16, WF17, WNF19, WDFN21, WF21, YTP11, YZP21]. Codewords [Arn17]. Coding [CIY01, CK08a, KKS05b, ŠM05]. Cographs [GV03]. Collaborative [SP04]. Collage [IST05]. Collapsing [APV06, BZ10, Pri06]. Collection [CVM20]. Collision [Nak04]. Colloquium [DRS23]. Colonies [MCS08]. Colony [KAPF05, dMLBPP20]. Colored [AFB96]. Coloring [Bod91, BHK^{+}18b, CKK02, SG04]. Colorings [GHJS05, IZN99]. Colouring [SS99]. Colourings [JP08]. Combination [HW17]. Combinations [CB09]. Combinatorial [ACDL18, CCF08, DD06, MM05, TV07]. Combinatorics [BS12, BMMR11, EMR10, GHS13, IZ04]. Combinatory [RS95]. Combined [CLMP16, CGKY11, CGKY12, SY07, ACM11]. Combining [Bar90]. Committed [Çev20]. Common [AMT20, AILR16, AE05, DD13, IMP^{+}05, KS10, LW05, LW06a]. Communicating [BKM11, BKM12, BKM15, CCFS07, CVMVMV00, DPS97, Kri92, LRT92, MS07, MVMM02, Ott13, Ott15, Tru08]. Communication [Ada10, BV98a, BF97, BKM15, DHIÖ97, DDPS19, FL97, HYLF20, LC18, Nak04, PPR02, Špr09, YBM11, ZC13, ZYYH14]. Communications [CCM97, RVT06]. Community [ROK08]. Commutative [BH11, CD21, MR91]. Commutativity [IDR97, MS12]. Commuting [Cai94, MSMR22]. Compact [BMS12, LYX^{+}19, PPR18, YM19]. Comparative [OM96, ZL22]. Comparing [Sal07]. Comparison [FA06, FV24, HT12, KA18, LHD^{+}24]. Compatible [MIN11]. Compensation [Sem20]. Competence [BCVVH07, CVDV10]. Competence-Based [CVDV10]. Competitive [Leu04, ZZZ16]. Competitiveness [Pal03]. Compiler [DVG03]. Complement [Jir14, O’N15]. Complementary [CSN21, CCQ24]. Complementation [Bed18, FKV06, JJS05, JPŠ19, RC05]. Complements [HP09b]. Complete [ABH17, BGI^{+}18, BA24, DK11, GWF^{+}24, HW10, LD01, MW05, RWZ01, RS01, ZYLW12, GP13, GI19]. Completely [BCV23, DVG03]. Completeness [ABDP05, FOP05, HJV93, LBL06, Nag20, Zan91]. Completing [BCHK09]. Completion [BZ13, DFLL02, DK11, LLQ06, MMY10, PY04]. Completions [ST16]. Complex [Brz13, BD19]. Complexities [HH20, Jir14, KK19, Sch02, TY15]. Complexity [Ada10, AFO06, AKK19, AOSY10, AP92b, Arv97, AP90, BGN10, BHK19, BAK12, BPT16, BFL02, Bod91, BT17, BHNR04, BMMR11, BLY12, BL12, BT13, BL14, BKLS20, BCC13, CSR12, CK08a, Câm14, CLMP16, CRSZ11, CK16, CDM13, CS93, CGKY11, CGKY12, Dai97, Das04, Das19, Das21, DLW02, DG98, DM08, DK12, EH15, EHS15, FH05, FZ13, FL97, GY12, GI22, GPS14, GH15, HS08, HKNS16, HT12, Hol11, HK03, HK09b, HK11, HJ14, HJM19, IDR97, IR14, IJMP21, IYZ04, JS02, JMR91, JJS05, JM11, Jür08, KN21, KEH16, KLH16, KSV00, KLS05, KO13, Leu05, Lis93, Loh05, LMW08, Lüc18, MNS18, MvZ22, Mas19, MTVM09, MTVM15, MT95b, MB06, NRS18, NRS19, NB18, O’N15, OS19, PS02, PR11, Prů17, Rao08, RR18, Rya15, SS07a, SY07, SMS90, Sch10]. Complexity [SW17, SD16, Sun05, Toš06, TL99, VW93, WD20, WAG^{+}06, Wid12, WP08, XZS16, YS13, YTLC02, YW20, YWY94, Yen08, ZZT91].

Complexity-Theoretic [FH05].

Component

[GV23, GCH20, GZZX21, IN10, LZZN22, NB18, PR23, ZZZ23, ZYZX18, ZGL^{+}22]. Component-Based [PR23]. Components [BGMV08, CVOV11, DL12, JHK08, LCY12, MMK22, Mas09, Ott13, ST11]. Composed [ABH^{+}09]. Composite [AO10, YB23]. Composition [AM09, ARS11, BCDP08, LZZN22, Wan04]. Compositional [TW09, WM13]. Compositionality [FT09]. Compositions [BM23, Mal18, Mal24, Teh18]. Compressed [HI18, IST05, IB12, KS06, KSS08, Loh10, MHT09, WF17]. Compressible [PL23]. Compression [CDLW05, CK08b, DM05, De 06, KM90, KK05, Sal18]. Computability [Bur12b, Gra90, LS98]. Computable [BS92, CZ11, FW24, SS12a, Sch02]. Computation [AHR02, BDL08, CMRR08, DW03, EL13, FNI16, GO09, GRV10, GS12a, GR03, HL04, HN06, Lüc18, LLW21, MB06, Nis03, PDPPJ11, RZ12, RS17, SA22, ST11, SP04, SZQ^{+}17, VP99]. Computational [BKM12, BZ10, DLW02, FOP05, GKS17, HK09b, Ibr22, IPR07, JWB03, JS02, LMM^{+}12, MT95b, NB18, SD16, Sir15, WAG^{+}06]. Computations [Bee95, CD15, CE98, CK18, DK98, HK09a, HFLD09, HK19, KLP20, LD01, Mee12, YSM^{+}00a]. Computer [DRS23, TH01]. Computers [Rya15, Sah01]. Computing [AETZ05, AO10, BMSMT11, BFL02, Cai94, CZOdlH17, CLW09, CMMR04, CMWZ19, EAB^{+}16, FS21, FJ12, FKT07, FT11, GPPJR13, GCK08, Hea11, HO00, IZ04, LTZ12, Li00b, MLO17, MDL97, Obt01, Obt06, Pal01b, Pău00, PPR02, PPJR07, RS00, RR04, RC11, SRN^{+}20, SVSN01, SGZ02, Sto92, SUZ13, TZ11, UU07, WP08, XFJ03, Yue13, ZZT91, Zom03]. Concatenation [JJS05, Okh07]. Concentration [Dai97, WY22]. Concept [BOV08, DE08, Jai98, ROK08]. Concerning [CCF08, Hon02, IR14]. Concise [LP19]. Concurrency [Luc09]. Concurrent [BPT16, BET03, Dro92, DK98, MM07, PQ06, SKW08]. Condition [MP07, Mel93, Pal08, WWT20, ZWW^{+}14]. Conditional [GTCV19, LW05, LW06a, LYH^{+}15, LYG17, LX19, LHD^{+}24, MLO17, ZLL20, ZCX12]. Conditions [FT09, FO08, LBL06, Oka00, WFG15]. Conference [IZ04, MR23, SNJ11]. Confidentiality [SZQ^{+}17]. Confidentiality-Preserving [SZQ^{+}17]. Configuration [Par23a, WC04]. Configurations [ZL22]. Conflicts [MSR06]. Conformations [FKM^{+}21]. CONGESTED [SK20]. Congestion [GKKP99, KKP97, ZYYH14]. Congruential [YZZ22]. Conjecture [AV96, BMY17, Ber11, SFL17, PHPJRN^{+}11, Ste11]. Conjectures [GGJ^{+}19, RS04]. Conjugates [BMR^{+}14]. Conjunctive [AK14, DR94, Jeż08, Okh03]. Connected [AWF03, DWS15, ET14, Iba02, IN10, JHK08, KK10, KPS18, Li01, MTNN99, MNN06, ST11, Tor15, WAF03, ZH22]. Connection [WGD18]. Connections [DM08]. Connectivity [CV14, Che22b, FP04, GCH20, GZZX21, HLHH06, JLL23, LLY13, LX17, LLL22, LZZN22, MGL23, NPSY00, TH22, Tsi06, WFG15, WLZT21, XZY19, Yan21, YZZ22, ZZZ23, ZLL23, ZYZX18, ZH19, NS13, WC13]. CoNP [RWZ01]. CoNP-Complete [RWZ01]. Consensus [BvdB18, RS13, SK01]. Consequence [BK95]. Conservativity [Sel98]. Considering [YLX22]. Consistency [ADR11]. Consistent [YSM^{+}00a]. Constant [ANDZM09, BM23, CL98, FZCFB08, FT11, GPP20, JYF91, Lag17, LZ15, NS18, OW92, Smy12, Sun00, WF21, WQ97].

Constant-Degree [CL98]. Constant-Free [NS18]. Constant-Height [GPP20]. Constant-Memory [Smy12]. Constant-Weight [WF21]. Constant-Width [JYF91]. Constrained [AE05, BJD20, CFM13, CHWX09, GD12, NWHL22, NCC^{+}07, RAB15, Tor13]. Constraint [MZ01]. Constraints [ADR11, AE02, BB03a, Com90, FTT10, FM01, FS98, GR03, JSO10, JLL23, LL20, LTW02, LOPR18, MN00, NN93, PYTH10]. Construct [GKSZ19]. Constructing [AAA^{+}09, CPY02, CC05, DH96, MC02, PS12b, TJZ13, WWT20, XC15, YCTW10, ZH13, ZWCL14]. Construction [BF07, CGL12, DD08, DÉK22, EJ23, FZT14, FLFR19, HYT15, HHP17, JLL23, KKS05a, LW06b, MOSZ18, MDL97, Sak01, SNB24, Set08, SKW08, WF17, WZ15, WF21, XCMT20, Zho02]. Constructions [DQFL12, LL16, SNJ11, Sal13, SSF20, WPZ16, WNF20, WKS^{+}08]. Constructive [BRSRC11, Fre08, Oga00]. Constructivizing [Arv97]. Constructors [Huy91]. Constructs [HST01]. Contact [ZWC^{+}22]. Containment [NRT00]. Contamination [CSN21]. Contended [SB01]. Content [Cig04, GSZ09]. Context [Asv07, BMS92, BCR11, BCD14, BESW07, BHK05, BIIN04, DV14, EIM18, ÉO13, FLST12, GKRS10, HKS13, HW10, KK07, Kog18, KRK16, KM07b, LO13, MMK22, Mig90, Ott13, Pal08, Rav08, Rei07, Sao92, Tei17, Tra02, Tru08, YLX22]. Context-Free [Asv07, BCR11, BCD14, BESW07, BHK05, BIIN04, DV14, EIM18, ÉO13, FLST12, GKRS10, HKS13, HW10, KK07, KRK16, LO13, Mig90, Pal08, Rav08, Rei07, Sao92, Tei17, Tra02, Tru08]. Context-Freeness [Kog18]. Context-Sensitive [Ott13]. Contexts [CFRD08, Has00]. Continuous [CZ11, GFK98, RHS10]. Continuous-Space [CZ11]. Contour [BLL06]. Contribution [Rov00]. Control [BCP22, BV20, DCS13, ES01, FK06, HST01]. Controllability [MH06]. Controllable [SSS09]. Controlled [DST10, DEKZ11, MVM07]. Convergecast [AHL^{+}13, MG20]. Convergence [MV11]. Convergent [ECY02]. Convertible [LHT09, LH11]. Convex [BJD20, CLW09, DRDN08, MAN06, MNN06, SRN^{+}20]. Convolutional [LWW22, WY22]. Convolutions [Zha17]. Conway [FNI16]. Cooking [GW18]. Cooperating [FFH15, GV23, Kar09, Mas09, MO07]. Cooperation [ARV07, SB12]. Cooperative [FZ02]. Cooperativeness [MH06]. coordinate [ACM11]. Coordinated [GCK08]. Coordination [YLX22]. Core [MMSV23, Teh15]. Cores [MX11]. Correcting [GRB03]. Correctness [Bee95]. Correlation [EAB^{+}16, GK11]. Correspond [BLS^{+}05]. Correspondence [DRS14, Fin12, HH11]. Corrigendum [MS16a]. Cost [ACMP20, DGN07, FH05, For10, HI18, OW92, TV94, WHLH17, WO03]. Cost-Effective [WHLH17]. Cost-Optimal [WO03]. Countable [Bed18, RC05]. Counter [EIM18, IJT^{+}93, IDY08, NKP^{+}22, Pet11, SY12]. Counterexample [CFH^{+}03]. Counterexample-Guided [CFH^{+}03]. Counters [CR15, INY07, Rao08]. Counting [AČ11, CP03, CCP18, GO09, MR11, SJ04, Toš06, ZSW14]. Cover [CPY02, CGH05, CYS^{+}12, HW17, Kör03, LW21]. Coverability [GRV10]. Coverage [CDM13, FK13, GM19]. Covering [DS96, GGR14, YB06, ZBS05]. Coverings [TSS13]. Covers [CCP05, ER06, LMZC20, MPS24, Par23b]. Covid [Ibr22]. Covid-19 [Ibr22]. CP [YMC^{+}17]. CP-ABE [YMC^{+}17]. CPS [Oga00]. CPS-Calculus [Oga00]. CPU [CYZ14]. CQn [XZW^{+}21]. Cramer [LYY^{+}21]. Crawlers [LKM02]. Credit [Tse16]. Credit-Based [Tse16]. Crick [KM08, MMR20]. Criteria [HL04]. Critical [AA19, DW04, HB06, SS12a, Sun11]. Crochemore [FJ12]. Cross [WM05]. Cross-Pollinating [WM05]. Crossed [LC18, NZT^{+}24, Tru08, XZW^{+}21, ZFL^{+}17]. Crossing [BPT06, ST16]. Crosstalk [KAPF05]. Crowd [Sir15]. CRS [KLP20]. Cryptanalysis [LYY^{+}21]. Cryptographic [DQFL12, FY11]. Cryptography [CST^{+}17, YYW19]. Cryptosystem [LHT09]. Cryptosystems [LYY^{+}21]. CTL [MTVM15]. Ctl* [CZ11]. Cube [CX98, GRI24, HYLF20, LMZC20, LC18, LLW21, NZT^{+}24, PS12b, Yan21, ZYYH14, ZFL^{+}17]. Cube-Free [PS12b]. Cube-Of-Rings [CX98]. Cubes [CLT14, DG98, ESS20, ESS21, XZZY19, XZW^{+}21, ZCX12]. Cubic [GWF^{+}24]. Cuboids [JSPD03]. Curve [Fre02]. Customizing [LX94]. Cyber [SA22]. Cyber-Physical [SA22]. Cycle [Dan22, GP15, KB20, LLL21, LCXS19, NS98, RS22, Ros00, Won96, Noc98]. Cycle-Related [KB20]. Cycle-Stealing [Ros00]. Cycles [APMP17, DH18, GKSZ19, LX19, Won01, ZFL^{+}17]. Cyclic [DESW05, YZP21]. Cyclotomic [KK19, XZS16]. Cyclotomy [XCX16]. Cylindrical [ZWC^{+}22].

D [CHWX09, FFMW19, HJP^{+}13, JSPD03, JW08, Leo03, LJ17, SJ04, ŽM11]. D0L [Hon02, Hon06, Hon07, Sal07]. DAGs [CR14, PRS98]. D’Alessandro [Ber11]. Dassow [BRST07]. Data [ATK12, BSG03, KY96, LOD07a, LOD07b, LC22, Lin08a, MLO17, MMS17, MGJ19, Oka99, Oka00, RGR11, RN22, RR06, Ros00, SKL03, Sal18, SH22, TV14, TZ91, WHLH17, YZY^{+}18, YMC^{+}17, ZPXX17, ZLW^{+}17]. Data-Parallel [Ros00]. Database [HMZ05, Lin08b, SEE99]. Databases [Lar98, MT95b, VS93]. Datacube [Poo04]. Datalog [vdM00]. Datawords [MR11]. Date [KS10]. Davidson [HO99]. DCC [YZZ22]. DDOS [DEKZ11]. De-Quantisation [CCM11]. Deadlines [PZX07]. Deadlock [BDC90]. Dealer [Sun00]. Death [EMR10]. Debates [YSD16]. Decaying [FIO08]. Decentralized [MMS05, YM19]. Decidabilities [BKM15]. Decidability [AT12, BHK19, BH20, BAK12, BCD14, Bur12b, DS05, DFK23, DK12, Dur13, FM13, Gaz06, Loh05, RHS10, Yen08]. Decidable [AGM14, CRS12, HH24, Man15]. Decide [BK24, DK11]. Deciding [Dai97]. Deciphering [GMNS15]. Decision [CMWZ19, DH05, DMSS16, IR14, MVM07, Mod21, ZB00, ZB02]. Decisions [Cig04]. Decoder [BBFZM06]. Decoding [GMNS15, OSZ92]. Decomposable [FGV99]. Decomposition [CFPR03, Dic93, FGV99, Joh00, MAN05, SVF09, Yen09, ZWCL14]. Decompositions [AACR18, CSV02, DS05, PR00]. Decontaminating [FHL07]. Decontamination [LPS07]. Decryption [CCD07]. Dedicated [BRST07, FRS24]. Deduplication [MGJ19]. Definability [BV98b, ES01, MSV23]. Definable [DK98]. Defined [DH05, EMR11, Hut02, JP06, KMRY20]. Definitions [Kam95, Moh03]. Degenerate [BRM07, IMP12, LJH^{+}17]. Degradation [ZWC^{+}22]. Degree [AMT20, ABT16, Asl16, AHK17, AO10, AA13, BTO17, BB04, CL98, DH96, GW24, HL01, HLY^{+}04, KSM22, KA18, LDLW17, Tor13, WLF03, WQ97]. Degree- [DH96, HLY^{+}04]. Degrees [EKKS18, Won96]. Del [KRK16]. Delaunay [Dev02]. Delay [GMNS15, JS97]. Delays [LLZ07]. Delegators [RS07]. Deleting [KO13]. Deletion [AB91, De 06, GMU15, KS11, MGJ19, PPJY08]. Deletions [WAG^{+}06]. Delivery [AF20]. Delta [BLS^{+}05, KSS08, dSMOC18]. Demand [HT09, IZN05, PZX07]. Dembowski [WDFN21]. Demonic [Tha91]. Dendric [BDD^{+}18]. Dense [MX11]. Density [DSS15, EIM18]. Dependability [ABL^{+}11]. Dependencies [BV08, Lin08b]. Dependency [YJ05]. Dependent [DGL93, WG17]. Deployment [FZ03]. Depth [BS92, BLS^{+}05, CCR^{+}90, KL12, Pro96]. Deque [CCR^{+}90]. Derandomized [SS01]. Derandomizing [Vin05]. Derivation [DFP99, dMLBPP20, Mas09]. Derivations [CVDV10]. Derivative [BMMR11, BMMR12, SL17]. Derivative-Based [SL17]. Derived [GLV14, WD20]. descent [ACM11]. Description [GM90, KRK16]. Descriptional [Câm14, Das04, DM08, GH15, HT12, HK03, HK09b, HJ14, HJM19, KN21, KO13, Leu05, LMW08, MvZ22, Mer08, Sun05]. Descriptions [DK12, Pin12]. Descriptive [CS93, GI22, WP08]. Design [AMR05, CCG^{+}11, KR97, LC22, LL23, XWL^{+}22]. Designated [Ver09]. Designed [GD12]. Designing [FZEBB05, SK03]. Designs [PRS98, WDFN21]. Detailed [ZPXX17]. Detection [EP17, Nak04, San13, YW22]. Detectors [Huy91]. Determination [AHR02]. Determine [FSWF11]. Determined [Far20, Géc07]. Determinism [CL15, HKKŠ13]. Deterministic [Ada10, AMR08, AHK17, BBK17, BV20, CDPR11, CCFS07, EP17, FKM^{+}21, GLV14, Gia11, Glö10, HPP99, HJK18, IS12, JM13, JJS08, JK19, KZ10, LO13, LMG20, Löd15, LT24, Mas13, MO09, MC13, Nag21, OS01, Pig09, Slo95, WF17]. Determinization [CCP05]. Deutsch [CCM11]. Developable [Fre02]. Development [McN90, YLX22]. Developments [GVL07, MR23, Sek20]. Deviation [DPS99, MPS99]. DFA [AV96, NKW08, SKW08]. DFA-Based [NKW08]. DFCA [CP03]. DFT [SEE99]. Diagnosabilities [LHD^{+}24]. Diagnosability [DXZ20, Wan21, ZLL20, ZCX12, ZGL^{+}22]. Diagnosis [BCB12, SL17]. Diagnostic [CLT14]. Diagram [WGD18]. Diameter [GKS^{+}19, NZT^{+}24, NS98, Noc98]. Diameter-Optimally [GKS^{+}19]. Dichotomy [RS22]. Dickson [Kog21]. Dictionary [AE04, De 06]. Difference [BMP03, CZTH13, LL16, MvZ22, Van05, YTP11]. Differences [Aku06]. Different [GJKS18, Leu05]. Differential [ABDP05]. Differentially [XC15, ZZC15]. Diffusion [BCC^{+}11, Hei97]. Digit [HP09b]. Digital [KPS93]. Digitized [SMAN13]. Digraphs [QFL^{+}15]. Dimension [DG98, LZ15]. Dimensional [AGM14, AE99, BT00, BKP18, CdL04, Che22b, DJ12, Dub95, JZ16, JP06, KPSC08, LR04, MS20, MP22, Mod21, NR18, Prů17, SKL03, SMAN13]. Dimensionality [BHL^{+}97]. Dimensions [KKH90, Poo04]. Diophantine [CE98, IDY08]. Direct [Kop21, Nag20, SB12]. Directed [ADD^{+}18, BPR09, FZFDCHB05, KLB13, RR18]. Direction [BF07, FS98]. Directional [TH22]. Directory [ADR11]. Directory-Based [ADR11]. Disambiguation [Moh13]. Discontinuous [ÜS02]. Discord [EGPS10]. Discounting [CM12]. Discovery [TBGP20]. Discrepancies [EGPS10]. Discrete [BDG^{+}11, BLL06, CZ11, DPR07, JRPIP08, Yen08]. Discrete-Time [CZ11]. Discretized [AEMY21]. Disequilibrium [VJDT05]. Disjoint [BT07, DH18, GSZ09, HKV17, LPC11, LMZC20, LW21, Par23b, RLWW96]. Disjoint-Paths [LPC11]. Disjunctive [DR94]. Disk [CYS^{+}12, Fuj16]. Dispatching [KD99]. Distance [AE04, CZOdlH17, CB09, CMR07, HKS13, HL01, HLY^{+}04, HI18, Li07, Moh03, NRS19, PRN13, YHK14, ZWS96]. Distances [ST99]. Distinct [LZGF16, PL23]. Distributed [AETZ05, AHR02, ABL^{+}11, BCB12, BB04, BKS12, CLT14, Cig04, DCS13, DEMT05, FFH15, FBHH01, GV23, HPP99, KK10, KG11, KBH99a, KSV03, LTZ12, Mas09, MO07, MV11, Pal01b, SK01, SK20, San13, SF07, SP04, Tsi06, WLF03, WC04, WRNK03, XS11, YSM^{+}00a, ZC05]. Distribution [AS18, BBM^{+}12, Cas95, DG98, MMR10, PNN^{+}10, RR06, Rav08, SNWW06, SNJ11]. Distributions [Gol90]. Diverse [BGI^{+}18]. Diversity [Qua07]. Diversity-Based [Qua07]. DLOG [Gre96]. DLT [MR23]. DNA [ANDZM09, CK08a, DW03, FMC04, FK05, FKT07, IMP12, Sal18]. Does [MCM^{+}11]. Domain [CGH05]. Domains [Dro92]. Dominance [SJ04]. Dominating [AWF03, DWS15, KK10, NGHK15, Tor15, WAF03]. Domination [AA19, HKT00, LLW^{+}22, SR21, TK19]. Dominoes [RR99]. Dot [BS92, BLS^{+}05, IN08, JP06, KL12]. Dot-Depth [BS92, BLS^{+}05, KL12]. Double [AMR11, CHA^{+}92, CS99, HKMW22, HSS19, JSKM20, LOZ98, MB03]. Double-Ended [CS99]. Double-Head [HKMW22]. Double-Tape [AMR11]. Doubling [APMP17, Wid12]. Doubly [Lin08a]. Doubly-Linked [Lin08a]. Down [BCC^{+}96, KM23, LW93, LT24, MSV23]. Download [Li12b]. DP [CV13]. Dragon [SSS13]. Drawing [DEKW06, Pat06, ZH06]. Drawings [ADD^{+}18, MAN06, MNN06]. Drip [CP06]. Driven [BESW07, DS02, HKMW22, NKW08, OS19]. DSMS [ST01]. Dual [CLT14, DRS14, HL04, HSS19, LPC11, Okh07, SZQS18, ZCX12, ACM11]. Dual-Cubes [CLT14, ZCX12]. Dual-Net [LPC11]. Due [KS10]. Duplication [DGMM15, FMR20a]. Duplications [Sem20]. Duval [HN04]. DWT [RN22]. Dyck [BSCH22]. Dynamic [BV98a, BV20, BDC90, CFMS15, Cas95, CZ11, DEZ01, DGR24, GWL02, GR03, Hei97, HI18, JP07, KG11, KK90, Lag14, LOD07a, LOD07b, Li00a, Lug11, MO94, MD00, NWK05, NWK06, PPR18, PFG^{+}01, Rud15, SK04, SH22, TZ11, Wan14, XFJ03]. Dynamical [PBMZ06, Toš06]. Dynamically [CVPV08, LCVLV09]. Dynamics [Kop21, MB06].

e-Normalization [Moh02]. e-Removal [Moh02]. E-Unification [GJV00b]. Earliest [FSM11]. Earliness [KS10]. Earliness-Tardiness [KS10]. Early [PPJS07]. Easier [Lug11]. Eco [LK11, LCVLV09]. Eco-Grammar [LK11, LCVLV09]. Economic [NZH22]. Eden [Toš06]. Edge [AB91, AJM^{+}21, BAK12, BS16, Cal15, CV14, DJL^{+}07, ET14, GMU15, GZZX21, KA18, LDLW17, LX19, LLW18, LZZN22, NPSY00, Par23b, ST11, Tsi06, WFG15, XZW^{+}21, YZZ22, ZLL23, ZYZX18, LLL22]. Edge-Connectivity [ZLL23]. Edge-Deletion [AB91]. Edge-Pancyclicity [XZW^{+}21]. Edge-Path-Replacement [LLW18]. Edges [DEKW06]. Edit [AE04, CZOdlH17, CB09, HKS13, HI18, Moh03, PRN13, YHK14]. Edit-Distance [HKS13, Moh03]. Edited [SS24]. Editing [FM96, ZWS96]. Editor [Zom01c]. Editorial [AETZ05]. Editors [Hsu98, NO99]. EDZL [WR16]. Effect [CL07b, FPS02, NZH22]. Effective [BK24, Fin21, Ruo96, SS12b, WHLH17]. Effectively [YMC^{+}17]. Efficiency [EH12, ZL22, ZSG^{+}22]. Efficient [ADHR09, AAI^{+}20, ARS11, Anc02, BBFZM06, BRM07, BS01, BB03a, CPY02, CF06, CCF09, CCD07, CDJ09, CL10, DHIÖ97, DCS13, DZH16, ERW04, FL09, FZFDCHB05, FLP13, FG08, GLV14, GRV10, GSD03, GS12a, GRB03, HH22, HYT15, Huy91, INY07, IMS03, Kör03, KB20, LF96, LOD07a, LOD07b, Li01, LYHW19, MD00, MIN11, MHT09, MOSZ18, MS19, MC13, NGHK15, Okh03, PT14, Ros03, SRN^{+}20, SK04, SNB24, SUZ13, TWZ11, TFF18, Tsi06, WKS^{+}08, WRNK03, WY05, ZZ18, ZC05]. Eigenvalues [QD03]. ELAN [BKKR01]. Election [AOSY10, FDFZB12, FZAM08, XS06]. Electronic [FK06]. Elegant [PRN13]. Elementary [Rog09]. Elements [AES18, KNR18, LLY13, LMZC20, VW93]. ElGamal [HLH19, LHT09, RN22]. ElGamal-like [HLH19]. Embeddability [CLT09]. Embeddable [BPT06]. Embedded [CDFK19, ZLL23]. Embedding [DLT06, GRI24, GPP20, GWF^{+}24, Mar97, RAB15, RN22, WXF16, ZFL^{+}17]. Embeddings [Li00a, LLL21]. Emerging [CVPV08]. Emptiness [ABH17]. Ems1 [PRN13]. Emulated [YBM11]. Enable [AF20]. Encoded [Câm14, CFG12]. Encoding [CK18, KSS08, OSZ92]. Encodings [CG09]. Encrypted [ZLW^{+}17]. Encryption [BB03b, GKS17, HLH19, LHT09, LMG20, LH11, MLO17, MMS17, SZFX20, TFS19, WLC12, WZ15, WHLH17, WZCH19, ZYZ^{+}19]. Ended [CS99, Tsu01, TST01b]. Ending [CD15]. Endomorphisms [Ric19]. Energy [Jür08, Nak04, QFL^{+}15, SUZ13, WY05]. Energy-Efficient [SUZ13, WY05]. Enforcing [PQ06]. Enhanced [LW06b, NWHL22]. Enhancement [NWK05]. Enhancing [Qua07]. Ensure [Bee95]. Entangled [LB04]. Entropy [CMRR08]. Enumerable [vLW15]. Enumerating [CC05]. Enumeration [CKZ17, CRS12, DMSS16]. Environment [MLO17]. Epidemic [Ibr22]. Epigenetic [BDL08]. Episturmian [JP04]. Equality [BMW91, HH12, Hon12, Mel93, Sel98, Szw95, WZCH19]. Equals [RS13]. Equation [HSS07, MOSZ18]. Equational [BE95, Pin12]. Equations [CHKL07, CK07, ELS15, IDY08, LP11, LS98, LO11, MNS18, NS18, Okh05, PT90].

Equivalence [BDSV06, BH11, CD21, CMR07, DHR08, HJ13, Hon02, Hon07, IJT^{+}93, KL03, Man15, NTSH06, Pau24, PT18, PT19, Teh16a, WGD18]. Equivalences [BJ05, BJ07b, HJ97, BJ06]. Equivalent [GVL07, Teh16b, ZB00]. Erasing [Zet11]. Erasure [LZGF16]. Erbil [Ibr22]. Errata [BJ06, Tsu01]. Erratum [HT04a, LW06a, MTVM15, Ata11]. Error [GRB03, HL04, IKPY21, YW20]. Error-Correcting [GRB03]. Error-Free [IKPY21]. Errors [AACR18, HJ13, HJ17, KHS21]. Ésik [Fül17]. Essential [CL07b]. Estimation [CTZ01, SY07, SEE99]. Estuarine [LR04]. Eulerian [Ber13, Gus13]. Evacuation [Sir15]. Evaluating [KY90, Li00a]. Evaluation [ABL^{+}11, BLY12, Cha02, DZ00, HYLF20, Li12a, SK01, TH01, YH11, ZLL23]. Even [Faa19, GI19, GW24]. Even-Odd [Faa19]. Even-Variable [GW24]. Event [D’s03, Yen08]. Every [Far20]. Evidence [BK95, SZ22]. Evolution [EH12, Riv04, SZFX20]. Evolutionary [DT20, DM08, FRV19, HL01, MGCVdlP20]. Exact [AMR08, BBM^{+}12, EL13, GQZ15, KL00, LLZ07, NKP^{+}22, ZSW14]. Exactly [Cai94]. Example [CHKL07, GRRS14]. Examples [CM92]. Exchange [CST^{+}17, TYM^{+}17]. Exchangeable [LYY^{+}21]. Exchanged [LC18, LLW21, NZT^{+}24, WLZT21, ZFL^{+}17]. Exclusion [KG11, DDHL11]. Execution [FZAM08, Wan04, ZC05]. Execution-Time [Wan04]. Exhaustive [IN05, IN08]. Existence [DI02, RS07, Ruo96, Shu11]. Existential [Szw95]. Existing [FZ03]. Expected [CZOdlH17, Li00a]. Experience [CFMR05]. Experiments [BCP22, DES09]. Explicit [DDPS19, KN93, Kam98, vdHM92]. Exploiting [BDSV06]. Exploration [CP16, ER14, HZZT12, PT14]. Explore [CFRD08]. Exploring [Gia11]. Explosion [DS02]. Exponent [BNBN20, SS12a].

Exponential

[AA19, BCFR07, ÇA18, Fri10, GO09, Gol14]. Exponential-Time [GO09]. Exponentiation [HP09b]. Exponents [KMIS09]. Expressibility [MT95b]. Expressible [AB91]. Expression [CKW09, HW05, Han13, Kog18, NR21, SL17]. Expressions [BR20, BMMR19, CSY03, Cha02, CLOZ04, DM11, GH13, GH15, GC18, HWW06, HK11, KNR21, KMRY20, Loh10, TV14, YZ07]. Expressive [Hen02, RHS10]. Expressiveness [Yue13]. Expspace [ZYLW12]. Expspace-Complete [ZYLW12]. Extended [BM23, BHK07, DG98, FIO08, Mal24]. Extending [Pat06]. Extension [BMS18, EL13, Hen02, KM02]. Extensions [BLY12, DM12, HN04, Ver09, XLC^{+}04]. Externalities [SL21]. Extra [Che22b, GZZX21, Wan21, ZLL20]. Extractable [Kun16]. Extremal [CdBD23, ZLG21]. Eye [Ami05].

FA [CKW09]. Face [RLWW96]. Facility [XS11]. Factor [AES18, CISH07, JPŠ19, MM05]. Factor- [JPŠ19]. Factorial [Shu07]. Factorization [BGI^{+}18, BOV08, DD08, Hit20]. Factorizations [CL14, Kop21]. Factors [AILR16, HN10, PAS08, XWL^{+}22]. Failure [FWZ15, NTSH06, PNN^{+}10]. Fair [MSR06]. Faithful [APP91]. Families [DH05, DD08, HJK12, IM20, KY96, MRS97, MAG09, OY11, SRPC11, WDFN21]. Family [BKST18, ESS20, KSMMT18, LYY^{+}21, MMK22]. FAS [JRPIP08]. Fast [Ars15, BLP18, BOV08, ECY02, FPPS03, FNI16, FA06, GO09, GKS^{+}19, IML04, Kan15, LCL06, NWK06, PP06, SJ04, TCT14, Zha17]. Fastest [CFMS15, Hut02, XFJ03]. Fat [DEKW06]. Father [AMZ20]. Fault [CL07a, CHYT14, FZEBB05, GZY24, GWL^{+}17, HY97, KR97, LPC11, LYH^{+}15, LYG17, LX19, NZT^{+}24, XS11, XZY19, XZZY19, XZW^{+}21, YZZ22, ZZN23, ZCX12]. Fault-Free [GWL^{+}17, LX19]. Fault-Tolerance [ZZN23]. Fault-Tolerant [CHYT14, LPC11, XS11, XZY19, XZZY19, YZZ22]. Faults [KNR18, LX19, NPSY00, PP06, Par23b, WCD^{+}14, YBM11, YCL11]. Faulty [CP16, GKKP99, GWL^{+}17, LLY13, LMZC20, LLL21, MGL23]. Feature [MN00, SRR15]. Feedback [GB03, HG11, KHLC12, XLZ19, YB06]. Feedback-Free [GB03, YB06]. Feferman [HK95]. Few [GJKS18, MR99]. FHE [CK18, KLP20]. FHE-Based [CK18]. FHSs [XCMT20]. Fibonacci [DMSS16, ESS20]. Fibonacci-Automatic [DMSS16]. Field [RW11, YW20]. Fields [LCXS19, WNF19, WNF20]. Fighter [KLS^{+}19]. Fighting [FLP13]. File [Li12b, NN93]. Files [KSS08, WRNK03]. Filter [ARS11, MCM^{+}11]. Filter-Based [ARS11]. Filtered [DM08]. Filtering [DEKZ11]. Filters [DT20, FBK05, Sal18]. Financial [LC22]. Find [GI19, Gia11, MTNN99]. Finding [DGL93, ET14, Fuj16, GKRS10, GHWZ05, HKV17, HCG96, IMP^{+}05, IB12, IZN99, Kar99, MM97, NRT00, PR00, VW93, Won96, Won01, ZB00]. Fine [Sel08, BSOR10, KPS13]. Finite [AM09, ARS11, AMR11, AMR08, AMR15, AHK17, BGN10, BHK19, BH20, BBL^{+}12, BMW91, BHK07, BKM11, BKM12, BKM15, CSR12, CZOdlH17, CPY02, CLOZ04, CGH05, CGKN08, CFY16, CL07b, CGL12, CTS18, DL12, Das19, DGK08, Dom04, DFK23, FFH15, Fin19, Fle20, FHKK23, Fre08, GLV14, GHWZ05, GMNS15, GH13, GH15, GQZ15, HS08, HK23, HN10, HK09b, HJ17, HJK18, Iba15, JJS08, JJŠ18, JK19, KZ10, KL03, Kör03, KLS05, KSY14, KMW14b, KMW14a, LCXS19, Mac96, MS20, MMR20, MM17, Mar08a, MVMM02, MZ12, Mel93, Moh13, NWK05, NWK06, RW11, RV22, SS07a, SMS92, SD16, Shu14, ŠM07, SS01, SN13, Vor16, Vor18, WNF19, WNF20, ZQL12]. Finite-Image [DFK23]. Finite-Memory [KZ10]. Finite-State [AM09, ARS11, AMR11, CSR12, CZOdlH17, CGKN08, Mac96, SN13]. Finite-Valuedness [Iba15]. Finitely [AK10, AM03, GI22, Pau24]. Finiteness [AK06]. Fire [FLP13, KLS^{+}19]. Firing [GLP07]. First [AB91, BB04, DGK08, DZ00, Has00, IMP^{+}05, KKH90, Lin08a, MN00, Rov00, Ueh99]. First-Class [Has00, MN00]. First-Fit [KKH90]. First-Order [AB91, DGK08, DZ00, Lin08a]. Fit [KKH90]. Five [CH15]. Five-Valued [CH15]. Fixed [DS96, FL97, HL06, JJS08, LOZ98, MB17, Poo04, QLWL06, SW17, Toš06]. Fixed-Height [SW17]. Fixed-Length [QLWL06]. Fixed-Parameter [HL06]. Fixpoint [ELS15]. Flat [CDFK19, MT95b, Oka99]. Flexible [FMN06, JMSO05]. Flipping [LRR08, ZG13]. Flips [AAH02]. Flooding [CIS03, LBJ03]. Floundering [BM90]. Flow [LLZ07, Mas04, SS07b]. Flows [DW04]. FM [GNP^{+}06, IN05, IN08]. FM-Index [GNP^{+}06]. FO [HK23]. Fold [RKRR02]. Folded [DHIÖ97, MGL23]. Football [CKL15]. Forbidden [HK23, LLL21, WAG^{+}06, Yah12, Yen08]. Forbidding [Mas09]. Force [CCP05]. Forecasts [CL10]. Foremost [CFMS15, XFJ03]. Forest [Ali16, GO09, LZ12]. Forests [ERW04, Yah12]. Forever [HJM19]. Foreword [BNR05a, BNR05b, Hol05, Hol06, Hol08, Hol09, Hsu98]. Forgetting [Glö07, Glö10]. Form [BMMR19, Ési12, FSM11, GJV00b, LZGN06, Lin08a, VS93, Asv07]. Formal [BGS11, CSY03, CD21, CFRD08, DM05, DK12, ILT11, MDAPHPJ^{+}11, McN90, MT95b, ROK08]. Formalism [dSMOC18]. Formalisms [HJW11]. Formalization [HK95]. Formalizations [KKS05a]. Formed [LCVLV09]. Forms [Cai94]. Formula [DS02, Uen13]. Formula-Driven [DS02]. Formulæ [HKKŠ13]. Formulas [CE98, Sch10]. Forums [XCC16]. Forums-Oriented [XCC16]. Forward [CD95, Lug11, WHLH17]. Foundations [HBN08]. Four [KK19, MTNN99, MNN06, SZQS18, SH17]. Four-Connected [MNN06]. Four-Valued [SH17]. Fourth [VS93]. FPGA [DEZ01, IN08, IN10]. FPGA-Based [DEZ01]. FPSOLVE [ELS15]. FPTAS [KS10]. Fractional [Sha04]. Fragment [HCG96, MW05]. Fragments [DGK08, MTVM09, MTVM15]. Framework [GGR14, LTZ12, Lin07, NS13, NWK05, TST01b, Tsu01]. Free [Asv07, Bed18, BMS92, BCR11, BCD14, BESW07, BHK05, BIIN04, BLM04, BL12, CD06, CR15, DV14, DSS15, EH15, EHS15, EIM18, ÉO13, FLST12, GS18, GKRS10, GV23, GB03, GV03, GWL^{+}17, HWW06, HS11, HKS13, Han13, HW10, HLH19, IKPY21, JM11, JPŠ19, Kam95, KKS05a, KK07, KEH16, KRK16, KM07b, LO13, LX19, MR91, Mig90, Nag21, NS18, Pal08, PS12b, Rav08, Rei07, RS22, RS04, Sao92, Sta07, Tei17, TSZ16, Tra02, Tru08, WZCH19, YB06, YJ05]. Freeness [Kog18, Nag21]. Frege [HK95]. Frequencies [CK16]. Frequency [CZTH13, WPZ16, XCX16]. Frequency-Hopping [WPZ16, XCX16]. Frequent [BLM15]. Frictional [DLW02]. Frontier [AT12, CHZ06]. Frontiers [GPPJR13]. FTSM [LLL22]. FTSM- [LLL22]. Fujisaki [TFS19]. Full [Bur12a, WLC12, ZHZ11]. Full-Text [ZHZ11]. Fullness [CdL04]. Fully [IST05, MC13]. Function [BR20, BKST18, CJ20, MMS17, dSMOC18, PS02, Sta05]. Functional [Ano01c, BV08, BKKR01, HST01, Hin01, Moh13, Pre01, Sal13, Wil91]. Functions [BB99, BMS92, BLY12, BH11, CM92, CH15, Car11, CGH05, CL07b, DQFL12, EMR11, FY11, FW24, Fin21, FK05, GW24, HK95, HG11, HI18, Jai95, KM02, KY90, KSV00, KSM22, Kur20, LHG11, LL16, NAK^{+}15, Obt01, PP11, Ros03, Rya15, SS01, SFL17, SH17, SUZ13, TST01a, TCT14, Teh18, TJZ13, WDFN21, XC15, XCX17, Yam03, YTP11, YKCW23, ZH13, ZLL11, ZWW^{+}14, ZWCL14]. Functorial [DD12]. Further [CD06, Das21, Sbu06, ZYLW12]. Fusing [TV07]. Fusion [SH22]. Fuzzy [BOV08, ÉK07, LC22, Par23a, SH22, ZL22]. Fuzzy-Set [ZL22].

GA [SZ22, VJDT05, Sun11]. GA-PSO-BP [SZ22]. Gain [MM11]. Galerkin [ÜS02]. Game [BvdB18, Fia08, FL12, GC15, GW18, FNI16]. Games [AT12, BFL02, Bod91, CM12, COT12, FZ02, FZ12, FFMW19, Fri10, GZ12, GJMP06, KL10, Vin05]. Gandy [Obt06]. Gang [BS01]. Gap [CNT22, FM96, SNB24]. Gap-Greedy [SNB24]. Gapped [FBK05, HMZ05, PAS08]. Gapped-Factors [PAS08]. Gaps [AACR18, IMP^{+}05]. Garbage [Nak18]. Gardens [Toš06]. Gas [WY22]. Gathering [BCDM23]. Gear [AT11, ZWC^{+}22]. Gem [BLM04]. Gem- [BLM04]. Gemmating [FOP05]. Gene [ATK12, BHR09, DM05, IPR07, IP08, MGGP08, Rog09]. General [AMR11, BK95, BB04, Dic93, FPP03, HI18, Leu16, MS20, MD00, Moh03, TL99, ZGL^{+}22]. Generalization [GMNS15, HW05]. Generalizations [CLLL08, IM21, KHS21, LD04]. Generalized [Arn17, CDX21, Dai97, Dan11, GWL^{+}17, HH11, HH24, HW05, KKH90, KK19, KM19, Kur20, LL16, MGCVdlP20, Nak03, NS98, Okh06, Rao08, Sch02, Tho06, WD20, WM13, WC13, XZS16, YKCW23, ZYYH14, ZH19, ZGCZ18, Noc98]. Generalized-Concentration [Dai97]. Generate [IN08, Jeż08, KPS18]. Generated [AK10, CL07a, JPŠ19, KMG11, LLL22, LWJ^{+}10, XZY19]. Generating [Asv07, BBC00, BMS92, BS92, CCP18, Dom12, LT21, RS04, Tru08]. Generation [AMR08, BBE24, KMS06, LBL06, Smy12, TV07, ÜS02, Wan14]. Generative [DST10, Zet11]. Generator [Rya21]. Generators [HBN08, NAK^{+}15]. Generic [BET03, ELS15, LW06b, MZ01, Moh02]. Genetic [ATK12, AC05, LMM^{+}12, Nis07, WM05]. Genome [IMP12, SSK96]. Genomic [BBM^{+}12, Sem20]. Geo [SS12b]. Geometric [CHWX09, CCG^{+}11, GGR14, GS09, MRS97, PSS12]. Geometrical [CDJ09]. Geometry [RS17]. Girod [GMNS15]. Given [CC05]. Global [FTT10, JHK08]. Globally [Slo95]. Glushkov [BMMR12, ZZ18]. Goals [BM90]. Godan [ZZZ23]. Goedel [Szw95]. Golomb [BMP03]. Good [DQFL12, FY11, LHD^{+}24, TCT14]. Goodby [SSS13]. Gowers [YKCW23]. GPU [CYZ14, FNI16]. GPUs [GD12]. Graded [BV08]. Grained [MS99a]. Gram [FBK05]. Grammar [AMR05, BCVVH07, CVDV10, CVOV11, DPS97, FFH15, FO08, LK11, LCVLV09, Láz13, MS07, MMK22, Mas09, Ott13, Sun05, Tru08]. Grammars [AK14, Asv07, BCFR07, BESW07, BIIN04, BCC^{+}96, CCR^{+}90, DPS93, DFP99, DST10, Fer07, GV23, GSZ99, GPP20, Jeż08, KK07, KM15, LO10, LX94, dMLBPP20, MVM07, MS16a, MS16b, MO10, Nag20, Okh06, Pal08, Wil91, YJ05, Zet11]. Granularity [Kri97]. Graph [ADR11, AAV00, AB91, AMOZ07, AJMO11, AJM^{+}21, AT15, BBC00, BBE24, BDI^{+}11, BHK^{+}18b, CDX21, CC98, CHYT14, DLT06, Far20, FW90, FL97, GO09, GR00, HO99, HZZT12, KLB13, LLW18, LL23, LWW22, LOPR18, MSMR22, Oka98, RK09, RZ12, TSFZRP17, UU07, WY22, ZH06, ZZN23]. Graph-Bin [BDI^{+}11]. Graphs [AES18, AFB96, AP92a, AMT20, ABT16, ADD^{+}18, Asl16, AA19, AO10, AT11, AB17b, BTK13, BTO17, BPR09, BO97, BGM^{+}18, BHL^{+}97, BB04, BCV23, BS16, BPT06, BLM04, BHR09, CP16, CV14, Çev20, CL07a, CLLL08, CDFK19, CPC99, ÇA18, DL12, Dan22, DP90, DH18, DW04, ERW04, EL13, EZ01, FK19, FWZ15, FP04, FGV99, Fuj16, GV03, GRI24, GP09, GS09, GP17, HKT00, HBIT08, HLHH06, HY97, JWB03, KKB24, Klo96a, KPM15, KHLC12, KA18, LWYL14, LDLW17, LX17, LLL22, LWW00, LOZ98, LLW^{+}22, LV08, MR99, MG20, MTNN99, MAN05, MAN06, MNN06, NGHK15, NPSY00, NS98, OS93, Par23b, RS22, RLWW96, RRT99, RR99, SS99, SR21, SG04, ST99, TV14, Toš06, WAF03, WFG15, Wan21, WQY16, Won96, Won01, XZY19, YCTW10, YB19, YZZ22]. Graphs [ZWS96, ZZZ23, ZH19, Noc98, WC13, YCL11]. Greedy [BR18, FKM^{+}21, Fuj16, GKSZ19, SNB24]. Green [YLX22]. Greibach [Asv07]. Grey [CDLW05]. Grid [BFMBS11, BCDM23, BE19, JP08, KKB24, LMM^{+}12, MNN06, ST93, Cas05, PT14, YLZ14]. Grids [Cal15, MM17, NR18]. Ground [Mar92]. Group [BK24, BN20, CLLL08, DM12, FZ15, GZY24, HYT15, KPS18, ZZZ23]. Grouping [Lar99]. Groups [Ble21, CD20, MSMR22, PP11, SS01]. Grover [KNR18]. Growth [GKRS10, NZH22, Shu14]. Grzegorczyk [Cap96]. GSM [LO10]. Guarantee [LSWW13]. Guaranteed [DPR07, Ros00, YSM^{+}00a]. Guaranteeing [MPV04]. Guarantees [Pal03]. Guarded [FGL^{+}90]. Guess [FSWF11]. Guest [AETZ05, NO99, Zom01c]. Guided [CFH^{+}03, DDM07, HZZT12, WY22]. Guidelines [Ros00]. GVW [HLC^{+}19].

Hairpin [DK11, MMY10, PRY01, ST16]. Half [Kam95]. Half-Monotone [Kam95]. Halting [FO07]. Hamilton [DH18]. Hamiltonian [BZ13, CP16, LX19, LLL21, Noc98, NS98, RS22]. Hamiltonicity [LYG17]. Handling [BCHK09]. Harary [ABT16]. Hard [BLLS03, BVM00, Dic93, ZB00]. Harder [CKL15]. Hardest [MO23]. Hardness [AMT20, LWW00, SL21]. Hardware [For10, IN05, INY07]. Harmonic [CCF08]. Harmony [LTZ12]. HAS-160 [WLC12]. Hash [BKST18, LYX^{+}19, NAK^{+}15]. Hashes [Wan14]. Hashing [CKW09, LPP92, MB03]. Hausdorff [Sta05]. Having [GI22]. Head [HKMW22, KMW14b, KMW14a, ZSG^{+}22]. Heads [IT13]. Heap [BSG03, Jun14, Pro96]. Hedges [BOV08]. Height [GPP20, Rei07, SW17]. Helmet [YW22]. Helping [AKS95]. Hessenberg [MS19]. Heterogeneity [RC11]. Heterogeneous [BLMR05, CFMR05, CYS^{+}12, EZ01, OS01]. Heuristic [CHYT14, CDLW05, De 06, LY94, PS22, WAF03]. Hexagonal [GSD03]. Hidden [FZ13, IMS03, KTT20]. Hiding [RN22]. Hierarchical [FK19, GM90, GCH20, GWF^{+}24, JS02, Loh10, SVSN01, SK03, SP04, WC04, WHLH17, ZYZ^{+}19]. Hierarchies [BLS^{+}05, BKM15, DH05, IM21, KP10a, Sch02]. Hierarchy [BGK^{+}20, BKM11, BZ10, BJY90, CSR12, Dev02, DZ00, GV23, HW00, Okh05, PPJY08, Rei07, Sel08, YZY^{+}18]. High [CH15, Fin12, KR97, KKP97, Li12b, LKM02, YLX22]. High-Capacity [Li12b]. High-Performance [LKM02]. High-Quality [YLX22]. High-Speed [KKP97]. Higher [BYP95, CCPS04, GW24]. Higher-Order [BYP95]. Highly [BCFR07, PL23]. Highly-Polynomial [BCFR07]. Highways [AAA^{+}09]. Hirschberg [JHK08]. Historical [MP93]. Histories [Faz08]. Hit [WPZ16]. Hits [HM04]. Hoare [HV02]. Hoc [AWF03, CIS03, CL03, LBJ03, SB12, WLF03, WD03]. Hole [DSS08]. Holes [RR99]. Holonomic [BMS92]. Home [ST01]. Home-Based [ST01]. Homing [SYS19]. Homogeneous [JSPD03]. Homomorphic [CK18, MLO17, RMZW19]. Homomorphism [Suc90]. Homomorphisms [LO13]. Honey [NWHL22]. Honeycombs [Sib97]. Honour [CVM20]. Hop [AF20, KKP97]. Hop-Congestion [KKP97]. Hopping [CZTH13, WPZ16, XCX16]. Horn [FGL^{+}90, SN13]. Housing [SZ22]. Huffman [CFG12]. Huge [NR21]. Hulls [CLW09]. Hurry [FZ12]. HV [MP22]. HV-Palindromes [MP22]. Hybrid [BHK07, CFH^{+}03, DPR07, FK06, FFH15, FK13, KSMMT18, LMM^{+}12, NWHL22, Smi95, SW09, XBE02]. Hybridization [ATK12]. Hyper [Bad09, CFMR05, HJ16, JM13, MQ11, MQ12, Wan21]. Hyper-Clusters [CFMR05]. Hyper-Minimal [HJ16]. Hyper-Minimization [JM13, MQ11, MQ12, Bad09]. Hyperbolic [Mar08b, Mar08a]. Hypercube [BV98a, GWL^{+}17, WC04, WLZT21, WRNK03, ZLL23]. Hypercubes [DXZ20, Li00a, LX19, LLL21, MGL23, Nak03, Zaj09, ZLL20, ZYZX18]. Hypercubest [LW21]. Hypermesh [LYH^{+}15, LHD^{+}24].

I/Os [MMSV23]. ID [CCD07]. ID-Based [CCD07]. Ideal [APP91, Mas19]. Ideals [DT20]. Idempotency [Leu16]. Idempotent [KP10b]. Identical [LLQ06]. Identifiable [Jai95]. Identification [AAI^{+}20, CL07b, Jai98, KB20, NZZ24]. Identifying [AES18, Arn17, CIRS08, CCI12, ZL22]. Identity [Faa19, FZT14, LYX^{+}19, LH11, MLO17, TFS19, WZCH19, ZYZ^{+}19, ZPXX17]. Identity-Based [FZT14, LYX^{+}19, LH11, MLO17, TFS19, WZCH19, ZYZ^{+}19, ZPXX17]. IDPM [LJH^{+}17]. II [BJ07b, Ros00]. III [DMSS16]. Image [DFK23, MPV04]. Images [CDLW05, DE08, FRS06, IN08, KS06, PS12a, SY10]. Immediate [AHR02]. Immunity [EAB^{+}16, GW24, LPS07, TCT14, TJZ13, ZWCL14]. Implement [Cha02]. Implementation [BCPR07, DK12, HST01, LPP92, MHT09, NWK05, NKW08]. Implementations [BBFZM06, Câm20, DEMT05]. Implementing [JHK08]. Implication [Lin08b]. Implications [BV08]. Implicit [Cha02, vdHM92]. Importance [BYİT21, FCS05, KB20]. Imprecise [HL04]. Impreciseness [CTZ01]. Imprecision [Cha97]. Improved [AMT20, DGN07, Dom04, Gro03, Han13, HW17, JZ16, Leu04, LJH^{+}17, PR00, Sal18, SS07b, WLC12, YW22]. Improvement [BC12, EG02]. Improving [PS22]. IMRT [CHWX09]. In-Network [BRSRC11]. In-Place [GPC09]. Inclusion [AEMY21, BCR11, CTZ01]. inclusive [DXZ20]. Incompatible [Jan93]. Incomplete [KHLC12]. Incompleteness [Fin19]. Incremental [DZ00, PNN^{+}10]. Indegrees [ZH22]. Independence [ÇA18, HKT00]. Independent [AWF03, CSN21, CK07, GNP^{+}06, MTNN99, NGHK15, SR21, TCLS10, Ueh99, YCTW10]. Indeterminate [SW09]. Index [Ano97, Ano98, Ano99, Ano00, Ano01a, Ano02, Ano03a, Ano04a, Ano05a, Ano06, Ano07, Ano08, Ano09, Ano11, Ano12, Ano13, Ano14, Ano15, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21, Ano22, BO97, FFH15, GNP^{+}06, LYY^{+}21]. Index-Shuffle [BO97]. Indexed [BC06]. Indexing [PAS08, ŽM11, ZHZ11]. Indicators [MS04]. Indices [KM22]. Induced [AWF03].

Induction [TY03]. Inductive

[BCC13, Kam95, Vik96, Wan04]. Industrial [FGH^{+}07]. Industry [ZL22]. Inequalities [Faz08, FM13, LW05, LW06a]. Inequality [FP04, GRS21]. Inexactitude [CMMR04]. Infection [FLP13]. Inference [BRSV13, MN00, Vik96]. Infinite [BCDM23, BHNR04, CK16, COT12, CNT22, CTS18, DM12, Dom12, DK98, DSS15, EKKS18, FW24, Fin04, Fin12, HHNP23, HK23, IBS01, Jai95, KPS18, Löd15, Mel93, PI95, Ric19, Sao92, Sha04, Sta05, WDFN21]. Infinite-State [IBS01]. Infix [HWW06]. Infix-Free [HWW06]. Influential [NZZ24]. Info [GI19]. Information [BB03b, CCF09, CST^{+}17, DG90, Jür08, Li07, SB01, TWZ11]. Informational [GSZ09]. Informed [MD00]. Infrastructures [DW04]. Inhibition [XCC16]. Inhibitors [Sbu06]. Initial [Mee12]. Initiality [BE95]. Initiations [MM07]. Inner [DMMM14]. Input [DZ00, FK05, HKMW22, LZGN06, LMG20, Lin07, MS19, Moh02, OS19, YLX22]. Input-Driven [HKMW22, OS19]. Input/Output [MS19]. Ins [KRK16]. Ins-Del [KRK16]. Insertable [Kun16]. Insertion [CW11, EG02, HKNS16, KS11, KMK11, OY11, PPJY08]. Insertion-Deletion [PPJY08]. Insertion-Query [CW11]. Insights [RC11]. Inspired [AETZ05]. Instance [BIIN04]. Instance-Specific [BIIN04]. Instances [HHH07, IMS03]. Instruction [TV94]. Insulated [LH11]. Insurance [ZL22]. Integer [FZ02, HHH07, HHNP23, HH22, Hit20, PA98]. Integers [SMS92, Dom12]. Integrity [BTK13, TK19]. Intelligence [dMLBPP20, Zho02]. Intelligent [DE08, LKM02, NH02]. Intention [ZSG^{+}22]. Interacting [BCB12]. Interaction [JWB03, Yue13]. Interactions [JWB03]. Interactive [AKS95]. Interchange [HL01]. Interconnection [CP99, CX98, CD09, Hsu98, LYH^{+}15, LHD^{+}24, LLW21, QD03, WQ97, ZLL23]. Interconnections [BF97]. Interesting [HPV99]. Interface [DE08]. Internal [Che22a]. International [MR23, DRS23]. Internetworking [GD98]. Interplay [BCV23, GGJ^{+}19, Kop21]. Interpolate [Fre02]. Interpolating [Ibr22]. Interprocedural [TY03]. Intersection [BCD14, CGKN08, CGKY11, EHS15, Fle20, HS08, YJ05]. Interstage [SS07b]. Interval [CPC99, EL13, GP17, NTSH06, SS99, ST99]. Intra [DDM07]. Intra-Molecular [DDM07]. Intractable [YHK14]. Intramolecular [IPR07]. Intransitive [WO03]. Intricacies [CHKL07]. Introduction [BBM90, NO99]. Intruder [ISAZ08]. Intuitionistically [TW09]. Invariant [KMRY20]. Inverse [ACFE09, FK13]. Inverses [Bir11]. Invertible [Dub95, Sut14]. Investigations [Pig15]. Involution [BCN12]. Involutively [KM07a]. Involving [Pan91]. Iota [dSMOC18]. Iota-Delta [dSMOC18]. IP [LOD07a, LOD07b]. Irreducible [WXF16]. Irregular [MS99b]. Irreversibility [AHK17]. Irreversible [GLPP22]. Iso [KTT20]. Iso-Oriented [KTT20]. Isolated [YB22, YB23]. Isomorphic [BVM00]. Isomorphism [AV96, Gre96, RK09]. Isoperimetric [AE99, BS16, RZ12, WFG15]. Isotopisms [BH11]. Isotropic [WNF20]. ISPAN [Pal01a]. Issue [Ano01c, BRST07, CD02, CVM20, DRS23, FRS24, Hin01, HO00, Hsu98, LC02, MR23, Pal01b, Pre01, RS00, Sek20, TY02, Yu02, YYW19, Zom01a]. Issues [Ami05, BF97, Cas05, RHS10, vdHM92]. Items [BLM15]. Iterated [BvdB18, Sta05]. Iteration [BE92, BE93, CLW09, FL12, Sut14]. Iterative [KPSC08, MMP10, SL21, ST16, Smy12].

Jacobsthal [PS02]. Job [BS01, JMSO05, Li01]. Jobs [CYZ14, FCS05, Jan93, JSO10, LY94, Zaj09]. Join [CGKN08, SEE99, YB19]. Joint [Coo17]. Jordan [Cai94]. Journeys [XFJ03]. JPEG [KS06]. Jumbled [BCFL12]. Jumping [BHK19, BH20, CFY16, KM15, MS20, MMR20, MZ12, Vor18]. Jürgen [BRST07]. Justification [VS93].

k-Isoperimetric [WFG15]. kernels

[ACM11]. Key

[GKS17, Gua21, HLH19, KMZS19, LYX^{+}19, LMG20, LH11, MNS11, SNWW06, SNJ11, TYM^{+}17, WLC12, WZ15]. Key-Insulated [LH11]. Keyed [MMS17]. Keyed-Function [MMS17]. Keyword [HLH19]. Kindhearted [SSS13]. Kinds [HG11]. Kinetics [HFLD09]. Kintala [KMW12]. Kit [HPV99]. Kite [XHLF02]. Kleene [BC06, GN11, HSS07]. Knapsack [KS10]. Knödel [BHL^{+}97]. Knot [San13]. Knowledge [BLR09, Pan91, ROK08, WCD^{+}14, vdHM92]. Known [XC15, ZH13]. Kolmogorov [Jai95, Sch02]. Kronecker [CV14]. Kuratowski [BGS11, JPŠ19]. Kurdistan [Ibr22].

l [FMV13]. Label [HZZT12].

Label-Guided [HZZT12]. Labeled [AMT20, DHR08, Fuj17]. Labeling [Cal15, IN10, MAN05]. Labelings [LLW18]. Labelling [NCC^{+}07]. Labels [HZZT12, KMRY20]. Laceability [LLY13]. Lagrange [Ibr22]. Lambda [Hir91, TST01a, PT90]. lambda-Calculus [PT90]. Lambda-Representable [TST01a]. lambdaPi [Pym92]. lambdaPi-Calculus [Pym92]. LAN [GD98]. Language [AEMY21, BRST07, BV98b, CC05, CDJ09, Cos90, DH05, DGMM15, ES01, Fin12, GKRS10, HKS13, HJK12, IR14, MM05, MRS97, McN90, Mer08, Mod21, MR23, MO23, Okh05, OY11, PS02, Pri06, Rov00, Sek20, YS13]. Languages [Ada10, AK06, AK10, AA20, AT16, BGN10, BLS20, BMS92, BCR11, BCD14, BC06, BBE24, BJ07a, BHK05, BCC^{+}96, BKW02, BGS11, BL12, BT13, Brz13, BL14, BD19, CPY02, CSV02, CL14, CD21, COT12, DK11, DES09, DJ12, Dom04, DK98, DV14, DPS97, EJ23, EH15, EHS15, EIM18, ÉO13, Faz11, FLST12, Fin04, FS21, FV24, GN11, GTCV19, Géc07, GI22, Gia11, Glö07, Gol90, GLPP22, HWW06, HS08, HS11, HK03, HH20, Huy91, IJT^{+}93, IW07, IS12, IM20, Jeż08, JM11, Jir14, JPŠ19, JP06, KKS05a, KP10a, KP10b, KEH16, KLH16, KHS21, KY96, Kog18, Kör03, KMG11, KMS06, KM19, KRK16, LNP16, LZ93, LO13, Leu16, MP07, MMK22, Mas19, Mig90, Nag21, ND02, Ogi94, Oka99, Okh03, OY11, PRY01, PPJY08, Pig09]. Languages [PP14, Pig15, Pin12, Rav08, RS12, Rei07, RV22, Sch13, Sel08, Shu07, Shu14, SR00a, SWZ97, Sta05, Sta07, Tei17, TSZ16, Tra02, YJ05, YZ07, ZQL12, vLW15, GP13, Ata11]. Laplacian [QFL^{+}15]. Large [BIIN04, BS15, DCS13, DEMT05, FPPS03, Fin19, FGH^{+}07, HH12, MDL97, Sha04, WRNK03, Won96]. Large-Scale [DCS13]. Late [LY94]. Latency [IN10]. Lattice [BSCH22, MLO17]. Lattice-Based [MLO17]. Latticed [KL10]. Lattices [BNBN20, BOV08, DE08, FV24, LYX^{+}19, LMG20]. Laws [BE95]. Layout [CP99, Nak03]. Layouts [GKKP99]. LCD [FLFR19, HSS19]. LDPC [BBFZM06]. Leader [AOSY10, FDFZB12, FZAM08, XS06]. Leaf [BV98b]. Leakage [HHP17, ZYZ^{+}19]. Leakage-Resilient [ZYZ^{+}19]. Learnability [KY96, Oka00]. Learnable [Oka99]. Learner [ZSG^{+}22]. Learning [CM92, CJS92, Cha97, KL00, LZ93, PFG^{+}01, SS01, Tor13, Tor15]. Left [BCHK09, CNT22, KH21]. Left-Infinite [CNT22]. Left-Linear [BCHK09]. Leftmost [DFP99, MS16a, MS16b]. Leibniz [Sel98]. Lemma [GTCV19, Kog21]. Length [AE02, DS96, FLFR19, Gus13, KMŠ21, Mar09, Pro96, QLWL06, SRN^{+}20]. Lengths [BR18, FT09, GP15, dBDZ19]. Lessness [FH05]. Letter [KP10b, Wid12]. Letters [CK16, LRR08]. Level [PS12b]. Levels [BLS^{+}05, BHK05]. Lexicographic [ZH22]. Lexicographically [FS21, Ueh99]. LFSR [WGD18]. Library [AMR05, RR06]. Life [EMR10, Rya15, ZL22, FNI16]. Light [Hea11, Rov00]. Lightweight [Gua21, HCETPL^{+}12]. Like [CFG12, CVPV08, HV02, HK11, LYY^{+}21, Par23b, HLH19]. Limit [APMP17, Gol90, Oka99, Oka00, Sch02]. Limitations [HJ91, LO11]. Limited [GPP20, HT12, KAPF05, Mas13, PP14, RRT99]. Limiting [AP90, CJS92, RS17, Vik96]. Limits [Ueh99]. Lindenmayer [BV20, Das04, DV11, HT12]. Lindström [BV98b]. Line [CGL12, CGK^{+}21, FPS02, KL05, Mas04, Pat06, Prů17]. Linear [AK14, AMR15, BC06, BÉ11, BCHK09, CCQ24, CFPR03, DPR07, DI02, DGN07, DGR24, FZ02, FKM^{+}21, GV03, Gra90, KK19, LC18, MOM91, MTNN99, MGCVdlP20, Nak03, Okh03, RLWW96, RC05, SFL17, Tei17, WGF16, WD20, WDFN21, YW20, YZZ22, ZYYH14, vdM00]. Linear-Time [CFPR03, MTNN99, RLWW96]. Linearizability [WPX^{+}21]. Linearly [CM92, YCL11]. Link [BY18, BA24, FWZ15, YB19, ZLG21]. Linkable [LW06b]. Linkage [OW92, VJDT05]. Linked [ACV13, KK07, Lin08a]. Links [Dre07, GKKP99, WP08]. List [Nak04]. Literally [KP10b]. Liveness [BHK18a, JC03]. LKH [SNWW06]. LL [MO23]. Load [Hei97, Li00a, MD00, ST01].

Local [AE02, Ars15, CYS^{+}12, CTS18, FL12, HN06, IN05, IN08, JP06, LSWW13, LPS07, RS13, XZY19]. Local-Connectivity [XZY19]. Localities [Cas95, LZGF16]. Locality [RR04]. Locally [Fri10, HYLF20, HJ91, RS12]. Locate [DSS08]. Locating [NAS22]. Location [MG14, Pre90, TZ11, XS11]. Locations [NR18]. Löf [Tsu01, TST01b]. Log [GWL02, MM11, TV94]. Log-Gain [MM11]. Logic [Ano01c, AH11, BM90, DGK08, DGR24, FMR20b, FMC04, FT11, GN04, GSZ99, HV02, HS95, Hin01, Lin08a, Luc09, Lüc18, MOM91, Oga00, Pre01, Rov00, RKRR02, Sal13, SMS92, Sub90a, Sub90b]. Logic-Based [Luc09]. Logical [D’s03, HKKŠ13, KM17, MCS08, RW11]. Logically [DK98]. Logics [DP14, LRT92, Par23a, Pen93]. LogP} [BNR99]. LogPQ [TH01]. Logspace [HJ97]. Longest [AILR16, AE05, DD13, KKB24, UU07, Won01]. Look [AE04]. Look-Up [AE04]. Lookahead [Fuj16, RS07]. Lookup [SK04]. Loop [BAK12, CHA^{+}92, JS97, Leo03]. Loopless [TV07]. Loops [Sem20]. Lossless [CDLW05, KK05, XHLF02]. Lossy [PRS98]. Low [IN10, KPSC08, WPZ16]. Low-Dimensional [KPSC08]. Low-Hit-Zone [WPZ16]. Low-Latency [IN10]. Lower [CE98, FY08, Gus13, LHG11, Uen13, ZK19, dBDZ19]. LR [FZCFB08, Okh06]. LR-Mesh [FZCFB08]. LRC [WNF20]. LRU [De 06]. LSC [HK02]. LSP [Ric19]. LTL [DPR07, MW05]. Lucas [ESS21]. Łukasiewicz [Sta07]. Lyndon [SY10, Suc90].

M [BSG03, PPR18]. m-Bonsai [PPR18]. M-Heap [BSG03]. Machine [HFLD09, HW17, KS10, LLZ07, Mal18, PY04, PFG^{+}01, Rud15, SSS09, SS07b, vLW15]. Machines [AKMW20, BCP22, Cap96, CGKN08, Dub95, FPP03, FBHH01, GJKS18, HIIW01, HHW99, HPP99, HJ17, HIR^{+}92, IJT^{+}93, Iba02, IDY08, IS12, IIT91, IIK^{+}04, Jan93, Kap05, LLQ06, Mer08, Pet11, Slo95, YS13]. Made [FKV06]. Magic [HJK12, Jir11, Van05]. Magnus [FRS24, SS24, Ste24]. Makespan [DLC^{+}14]. Making [vdHM92]. Malleable [LTW02]. Management [LWW22, SVSN01, TZ11]. Manufacturing [PFG^{+}01]. Many [BSOR10, GS18, GI22, LMZC20, LW21, MRT95, Ole92, YCL11, Zan91]. Many-One [Zan91]. Many-Sorted [MRT95, Ole92]. Many-to-Many [LMZC20, LW21]. Map [Wid12]. Mapping [AP92b, Ata11, EZ01, Hei97, IMP12, Teh15]. Mappings [LO10]. MapReduce [AS18]. Maps [BFM06, BKP18, HCG96, KPSC08]. Marginal [KHS21]. Marked [KNR18, NR18]. Market [DLW02]. Markov [DHR08]. Markovian [HJW11, MGGP08]. Martin [Tsu01, TST01b]. Martin-Löf [Tsu01, TST01b]. Mass [HFLD09]. Mass-Action [HFLD09]. Massively [AP92b]. Master [DPR^{+}08, GS12a, LYX^{+}19]. Master-Slave [GS12a]. Master-Worker [DPR^{+}08]. Match [HMZ05]. Matching [Aku06, BLP18, BH02, BZ13, BCFL12, CCFG12, CF06, CCF09, CLLL08, CB09, CPC99, CHZ06, DES09, FL09, FPPS03, Fia08, GW18, Han13, IST05, KS06, KLH16, LJH^{+}17, LZZN22, LCL06, MHT09, ND02, NRS18, Prů17, SKL03, SW09, WH03, XZL^{+}19, Zha17, FG08]. Matchings [DGL93, HCG96]. Mate [CP06]. Mate/Drip [CP06]. Material [ZWC^{+}22]. Mathematical [BCC13, Ibr22, NAK^{+}15]. Matrices [BM16, BMS18, BCMS20, BL01, Cai94, CJ20, CSAT20, HHH07, HN06, MS12, Oli13, PT18, SY10, Şer09, SHN09, SMAN13, Teh16a, WF17]. Matrix [Ata11, DFP99, DPR^{+}08, HT04a, HT04b, KRK16, MS04, MS16a, MS16b, Teh15, WXF16, Zet11].

Matrogenic [AP92a]. Matter [MCM^{+}11]. Max [Mas04, Pau24, Poo04, HW00]. Max-Plus [Pau24]. Maximal [AWF03, Bur12a, DD08, DGL93, FY08, Luc09, PL23, PR12, TSFZRP17, Ueh99, XZY19]. Maximality [KKS05a]. Maximally [WFG15]. Maximization [CS93]. Maximize [AJMO11, CR14]. Maximizing [Ros00, SRN^{+}20]. Maximum [AMT20, AMOZ07, BT07, BL01, BVM00, CPC99, DJL^{+}07, FKT07, GW24, MM97, Wan04, Won96]. MCFLs [ÉI14]. Mealy [CG06, KPS18]. Mean [BR08, GZ12]. Mean-Payoff [GZ12]. Meaning [HKKŠ13]. Means [CCP05, CHWX09, PPJY08]. Measure [AKK19, CS93, Sta05, Ueh99]. Measures [AT15, BLM15, BCC13, KS19, PSA17, RR04, Sch02]. Measuring [MKB^{+}11]. Mechanism [ZSG^{+}22]. Mechanisms [Obt06]. Meet [LJ17]. Meet-in-the-Middle [LJ17]. Meeting [BCDM23, SSF20]. Meets [BSS12, FFH15]. Megabase [BBM^{+}12]. Mem [CP06]. Membership [AK06, Arv97, Fuj17, Loh10, MS20, Nag20]. Membrane [BMSMT11, CMMR04, DI05, FT11, GPPJR13, MB06, Nis07, Obt01, Obt06]. Membranes [PDPPJ11, Pău00, PPR02, PPRPS11, PLMZ11]. Memoriam [Fül17, KMW12, Ste24]. Memory [BLR09, FBHH01, FRS24, HPP99, KZ10, Mor10, Smy12]. Menger [MGL23, WLZT21, YZZ22]. Merge [WO03]. Mergeable [CS99]. Merged [DD13]. Merger [INY07]. Merging [CP03]. Merlin [CCPS04, Vin05]. Mesh [EG02, FZCFB08, ISAZ08, Li01, RM98, ÜS02, WC04]. Meshes [BT00, FZEBB05, JW08, Mat04, XHLF02]. Message [EGPS10, FBHH01]. Messages [MN00]. Meta [PS22, SVSN01]. Meta-Computing [SVSN01]. Meta-Heuristic [PS22]. Metaheuristic [HCETPL^{+}12, LTZ12, SS12b]. Metalinear [MS07, Sun05]. Metalogic [Cos90]. Method [ACFE09, BNBN20, EH12, FK13, GMNS15, IN08, KM02, Li00a, RN22, SNB24, TFF18, ÜS02]. Methods [CCM97, Fre08, KKS05a, MZ01]. Metric [CLT09, MSMR22, XS11]. Meyniel [RR99]. Microarray [ATK12]. Middle [LJ17, VW93]. Millionaire [GKS17]. Min [KR97, Tor13, HW00]. Min-Degree [Tor13]. Mind [LZ93, Vik96]. Mine [WY22]. Minima [MS99a]. Minimal [ARV07, AMR08, BBC00, CIY01, CPY02, CP03, DWS15, GRV10, HBN08, HN04, HT04a, HT04b, HJ16, HJK18, Jai95, Jai98, JS97, JMR91, JJS08, MB17, Shu11, Suc90, Szw95, TA17, Teh18]. Minimality [Tam08]. Minimalizations [Pol05]. Minimax [HL04]. Minimization [AHK07, FSM11, GLV14, JM13, KLB13, LL23, MQ11, MQ12, ND02, Vin05, Bad09]. Minimize [AMOZ07, LRR08, Mas04]. Minimizing [DFLL02, GKKP99, HJ13, HJ17, KS10, Kör03, LY94, LLQ06, MMSV23, PY04]. Minimum [AJMO11, BGRY16, BBB^{+}18, BB04, BRSV13, CYS^{+}12, DGN07, DJL^{+}07, DLC^{+}14, FPPS03, Fuj16, GMU15, GWF^{+}24, GCK08, KK10, KHLC12, MPV04, MAN06, QFL^{+}15, Tor13, WAF03, Wan04, ZH22]. Minimum-Process [GCK08]. Mining [GWL02, LC22]. Minor [NRT00]. Miss [Leo03]. Mitrana [CVM20]. Mixed [CYZ14, DI02]. Mixed-Signal [LWJ^{+}10]. ML [Has00]. MM [ZLL20]. MM* [DXZ20]. Mobile [BFMBS11, BF07, BT17, BDDN01, CIS03, DSS08, FPPS03, FHL07, GCK08, HT09, IML04, LBJ03, MM07, SB12, TZ11, WP08, Zom03]. Mod [HKT00, SUZ13]. Mod- [HKT00]. Modal [DL12]. Mode [DI05, Fre05, Mas09, WLC12]. Model [ACDL18, BCB12, BNR99, BMS12, CFMR05, CGR13, Çev20, CFH^{+}03, DW03, DXZ20, EHK06, FZFDCHB05, HW10, LYX^{+}19, LAHN14, LYG17, LLL21, LR04, Nak04, Sak01, Sch10, SP04, Špr09, SZ22, Tha91, TH01, YW06, ZLL20, ZZC22]. Model-Based [BCB12]. Model-Checking [CGR13]. Modeled [CLT14]. Modeling [BCC^{+}11, Cas05, JRPIP08, KSS08, LCY12, PSS12, Sun11, XBE02]. Modelled [HFLD09]. Modelling [AH07, BDL08, DM05, Ibr22, PR23, SK01]. Models [APP91, BBFZM06, BZ10, CTS18, DRS23, DEMT05, For10, HJ97, HJW11, IJMP21, IP08, KPM15, LHD^{+}24, LWJ^{+}10, LW06b, dMLBPP20, Lüc18, Mal18, RCTC^{+}09, RS17, Sah01, Suc90, WY05]. Modes [FFH15]. Modest [Ros90]. Modification [Rud15]. Modifications [AJM^{+}21]. Modified [BSG03, BHL^{+}97, CDX21, Ibr22, IIT91, KYZS17]. Modified-Bubble-Sort [CDX21]. Modifiers [AG01]. Modifying [TY23]. Modular [BPZ07, DS02, Hit20, RCTC^{+}09]. Modules [BJ07b]. Modulo [CGR13]. Modulus [WD20]. Molecular [DDM07, EHK06]. Molecules [FMC04, FK05]. Monadic [SMS92, vdM00]. Monitoring [LWW22]. Monogenic [LV08]. Monoid [KM08, KLS05]. Monoids [BR08, BS92, Bur12a, DM11, Géc07, Loh05, MR91]. Monomial [Kur20]. Monotone [DDD18, Kam95]. Monotonic [ADHR09, ACV13, TY15]. Monotonicity [JC03]. Moore [CFG12]. Moore-Like [CFG12]. Morphic [Dur13, FRS06, Hon12, NP09, OY11, PS12a]. Morphism [Ram05]. Morphisms [HH24, Hol11, JP04, Kar09, PPJR07, RS04, Teh16b]. Morse [DSS15, Ram05]. Mosaic [BRSV13]. Mosses [AMR09]. Most [Brz13, BD19, SKL03]. Most-Specific-Rule [SKL03]. Motif [PRN13]. Motifs [IMP^{+}05]. Move [FM96]. MP [MM11]. MPEG [DE08]. MPEG-7 [DE08]. Muller [Arn17, FZ12]. Multi [AKS14, ABH17, APMP17, BCC^{+}96, CCD07, CGKN08, HP09b, JF18, KMW14b, KMW14a, LMG20, LWW22, Mal15, MX11, NCC^{+}07, RR06, SK01, SH22, TYM^{+}17, TFS19, Ver09, WM05, YBI11, ZC13, ZSG^{+}22]. Multi-Behavior [LWW22]. Multi-Cores [MX11]. Multi-Exponentiation [HP09b]. Multi-Head [KMW14b, KMW14a, ZSG^{+}22]. Multi-Objective [WM05]. Multi-Party [TYM^{+}17]. Multi-Processor [RR06]. Multi-Push-Down [BCC^{+}96]. Multi-Pushdown [AKS14, ABH17]. Multi-Receiver [CCD07, TFS19]. Multi-Secret [ZC13]. Multi-Sensor [SH22]. Multi-Sequential [JF18]. Multi-Stability [APMP17]. Multi-Tape [CGKN08, NCC^{+}07]. Multi-Tokens [SK01]. Multi-Track [YBI11]. Multi-use [LMG20]. Multicast [FPS02, SNWW06]. Multicasting [Gon01, XLC^{+}04]. Multicomputers [MS99a]. Multicounter [Iba02]. Multicriteria [CMWZ19]. Multidigraphs [Fuj17]. Multidimensional [KPS93, Tho06]. Multienvironment [MDAPHPJ^{+}11]. Multihead [Mac96, Slo95]. Multihop [CYS^{+}12]. Multikey [KLP20]. Multilayer [RHN^{+}22]. Multilingual [CK08b]. Multimessage [Gon01]. Multioperator [SVF09]. Multipartite [BA24]. Multiparty [KLP20]. Multiple [CF06, FK05, GD12, Lin07, LZGF16, MB03, Mat04, NR18, RVT06, XBE02, XWL^{+}22, YCTW10]. Multiple-Sided [XBE02]. Multiplication [MX11]. Multiply [ACV13, WF21]. Multiply-Linked [ACV13]. Multiprocessor [BLR09, CD09, SS12b, YH11]. Multiprocessors [WR16]. Multipseudoperiodic [MDGH13]. Multiresolution [XHLF02]. Multisequencer [SK01]. Multiset [BPT16, BMR^{+}14, CG06, Faa19]. Multisets [AFIV22, Bas97, CG09]. Multistage [KAPF05]. Multitape [IT13].

Multitriangle [WQ97]. Multivalued [Lin08b]. Music [CCF09, FMN06]. Musical [CCF08, CIRS08]. Mutants [MCS08]. Mutex [LCY12]. Mutual [KG11]. Mutually [YSM^{+}00a].

NAAP [LBJ03]. Naïve [ZLW^{+}17]. Name [CB09]. Nameless [Kam98]. Natural [Cha97]. Nature [AETZ05]. Nature-Inspired [AETZ05]. Near [BW14, HT09, XCX17]. Near-Bent [XCX17]. Near-Optimal [HT09]. Nearest [HL01]. Nearly [BJ07a, KS19, SSF20]. Necessary [WWT20, ZWW^{+}14]. Negative [CS18]. Neighbor [ABT16, BTK13, BTO17, HL01, KA18, LHD^{+}24, WQY16, LBJ03]. Neighbourhoods [DP90, NRS18]. Nerode [SMS90]. Nested [CZTH13, DP14, FGL^{+}90, Gre96, HLW09, RT16]. Net [LPC11]. Nets [AH11, BCB12, GRV10, JC03, MOM91, Muk92, RHS10, YWY94, Yen09]. Network [BRSRC11, Cas05, CL98, CX98, CCG^{+}11, DR05, FZ03, JLL23, KR97, Klo96b, LYG17, LOZ98, LPS07, Lug11, LLW21, MKB^{+}11, NZZ24, Oka98, RHN^{+}22, RR18, SZ22, WQ97, WY22, ZYYH14]. Network-Based [RHN^{+}22]. Network-Guided [WY22]. Networks [AWF03, AOSY10, AHL^{+}13, AO11, AT23, BV98a, BY18, BYİT21, BA24, BNS03, BLR20, BDDN01, CP99, CDPT16, CIS03, CFMS15, CL03, CYS^{+}12, CHA^{+}92, Che22b, Cig04, CD95, CD09, DHIÖ97, DGN07, DCS13, DT20, DM08, Fen22, FPPS03, FRV19, GKKP99, GZY24, GSD03, GCH20, GNC^{+}03, GZZX21, GWF^{+}24, HKV17, Hei97, Hsu98, ISAZ08, JS97, KAPF05, KKP97, KB20, Láz13, Li12a, LYH^{+}15, LMZC20, LHD^{+}24, LBJ03, LC18, LZZN22, LWW22, MMS05, MCM^{+}11, MGCVdlP20, NAS22, PPR02, QD03, Ros00, SB12, SL21, SP04, TH22, TL99, WLF03, WD03, WY05, WZCH19, XLC^{+}04, XFJ03, Yan21, YB22, YB23, ZC13, ZZZ23, ZLL23, ZGL^{+}22, DDHL11]. Neural [FIO08, IW07, KMG11, LWW22, PPJR06, PPJR07, PPJS07, SRPC11, SZ22]. Newcomb [Rav08]. NFA [JMR91, Leu05, Pol05, RS07]. NFAs [CCP05, DESW05, KS19, KH21, Van05]. NFSR [WGD18]. NL [DK11]. NL-Complete [DK11]. NLC [Joh00]. No [Nak04]. Node [BYİT21, BA24, HKV17, KB20, NZZ24, WQ97, WY05]. Node-Disjoint [HKV17]. Nodes [BCDM23, IML04]. Noisy [MG14]. Non [AG01, Ada10, AS18, BM90, BCHK09, BD19, CD15, CCQ24, CK07, Dai97, DPR07, DXZ20, DESW05, ES01, FLST12, FHKK23, Fre08, GJV00b, GRB03, GPP20, HL01, IMS03, Jeż08, KZ10, Kap05, Kut05, MvZ22, MSMR22, MC13, PP11, TY15]. Non-Abelian [IMS03, PP11]. Non-Blocking [Dai97]. Non-Boolean [PP11]. Non-Chain [CCQ24]. Non-Commuting [MSMR22]. Non-Constructive [Fre08]. Non-Definability [ES01]. Non-Deterministic [Ada10, KZ10, MC13]. Non-Ending [CD15]. Non-Floundering [BM90]. Non-inclusive [DXZ20]. Non-Linear [DPR07]. Non-Periodic [CK07]. Non-Primitive [FLST12]. Non-Qubit [GRB03]. Non-Recursive [Kap05, Kut05]. Non-Regular [Jeż08]. Non-Returning [BD19]. Non-Self-Embedding [GPP20]. Non-Standard [AG01]. Non-Symmetric [GJV00b]. Non-Synchronizing [TY15]. Non-Unary [MvZ22]. Non-Uniform [AS18, FHKK23]. Non-Uniform-Degree [HL01]. Non-Uniqueness [DESW05]. Nonblocking [WM13]. Nonce [KMZS19]. Nonce-Based [KMZS19]. Noncounting [KY96]. Nondeterminism [HKKŠ13, KH21, PSA17]. Nondeterministic [BKW02, Cha02, CC05, GPS14, HK03, HK09b, HJ14, HJ17, HK19, JRPIP08, JJS08, KO18b, Mar09, Nag21, Sao92, Tha91, Vin05]. Nondeterministically [HHN^{+}95]. Nonenumerable [Sch02]. Nonexistence [ZLL11]. Nonlinear [HG11, Kur20, PP11]. Nonlinearity [CH15, Car11, GW24, LHG11]. Nonregular [Mer08, YS13]. Nonsingular [XLZ19]. Nonstandard [Bee95, BSBZ08]. Nonterminal [Das21]. Nonterminals [KK07]. Norm [YKCW23]. Normal [Asv07, BMMR19, Cai94, Ési12, FSM11, Lin08a, RKRR02, Rya21, VS93]. Normalish [Ble21]. Normalization [Moh02]. Note [AHR02, BB99, BHL^{+}97, BS16, CKK02, FM13, GMU15, Har24, IIK^{+}04, LZ15, Mac96, Mas13, Szw95, YB19, Zaj09]. Notes [Okh07]. Notion [Gra90]. Notions [IYD05, SNJ11]. Novel [DCS13, KSM22, LYX^{+}19, LH11, SRR15, SGZ02]. NP [BGI^{+}18, Dic93, GP13, GI19, GSZ09, MW05, Nag20, SL21]. NP-Complete [GI19, BGI^{+}18, MW05, GP13]. NP-Completeness [Nag20]. NP-Hard [Dic93]. NP-Hardness [SL21]. NP-Pairs [GSZ09]. Number [AMR15, AB17b, AE99, BLS20, CdBD23, CP03, ÇA18, CFIJ10, DV11, Dom04, FY08, FRV19, FT11, GRS21, GRRS14, HB06, HJK12, JWB03, KA18, LZ93, LY94, NAS22, Pan91, PR12, RS01, RRT99, Vik96, WQY16]. Numbering [MNS11]. Numberings [Jai95]. Numbers [BS16, BPT06, CK18, HFLD09, Jir11, LO11, PDPPJ11, RS15, Van05, Wan04]. Numeration [JP04]. Numerical [CCM97, SGZ02].

O [Fle96, KKB24, OM96]. O-Shaped [KKB24]. O-Trees [OM96]. OBDDs [IKPY21]. Object [HK02, LX94, MT95a, YZ07]. Object-Oriented [LX94, YZ07]. Objective [WM05, YTLC02]. Observable [AT12]. Observer [CCM11]. Observer-Based [CCM11]. Observing [Cas95]. Obstacle [SH22]. Obtained [BMS18, CP03]. Occurrence [JSKM20]. Occurrences [CFIJ10, MS04, Sal07, SY10]. OCR [CB09]. Octal [GJMP06]. Odd [Faa19, KKB24, TJZ13]. Odd-Sized [KKB24]. Off [KL05, Mas04, KM18]. Off-Line [KL05, Mas04]. Offline [CW11]. Offs [Kap05, KKP97, Kut05]. Okamoto [TFS19]. omega [SMS90, CL14]. omega-Tree [SMS90]. On-Demand [PZX07]. On-Line [CGL12, FPS02, KL05, Mas04, Prů17]. One [AK14, BBP11, BH20, Ber13, BMP15, BKP18, CFY16, DI05, Dub95, HJP^{+}13, HK19, HIR^{+}92, IS12, KL12, KMW14b, KMW14a, LP11, Mod21, NS18, Obt01, PB20, POM22, SKL03, Slo95, TYM^{+}17, Zan91, ZWW^{+}14]. One-Cluster [BBP11]. One-Dimensional [BKP18, Dub95, Mod21, SKL03]. One-Membrane [DI05]. One-Round [TYM^{+}17]. One-Time [HK19]. One-Turn [AK14]. One-Variable [NS18]. One-Way [BH20, BMP15, CFY16, HIR^{+}92, IS12, KMW14b, KMW14a, Obt01, POM22, Slo95]. Online [BBB^{+}18, BLM15, BHK^{+}18b, CYZ14, DLC^{+}14, FCS05, JP07, JLL23, JZ16, Pal03, WZCH19, ZZZ16, ZSG^{+}22]. Onto [EZ01]. Ontologies [Zho02]. Open [GPPJR13, Tsu01, TST01b]. Open-Ended [Tsu01, TST01b]. Operating [DI05, ZL22]. Operation [BHK05, CK08a, CLMP16, DH05, MR91, NZZ24, OS19, YB19]. Operational [BMSMT11, BHK19, Das19, Das21, ÉI14, KEH16]. Operations [AP92a, BGN10, CP06, CS96, CGKY11, CGKY12, FM96, FMC04, FT11, GNC^{+}03, HH20, JJŠ18, KKS05b, PS02, SY07, SEE99, SD16]. Operator [AT16, BMS18, HJM19]. Operators [HW00, PR11]. Opportunities [Zom03]. Optical [BF97, KAPF05, LYH^{+}15, LHD^{+}24, LC18, PA98, Sah01, WH03]. Optically [BT00]. Optimal [AAA^{+}09, AC05, BF07, BCDM23, CZTH13, CP99, Cal15, CDPR11, CS96, DH18, DSS15, FZ03, FM01, FOP05, FLFR19, GD98, GZ12, GW24, HT09, KLP20, KK90, KTT20, KR08, Lag17, LZ15, Lüc18, MQ11, Nak04, OS01, OSZ92, Poo04, TCT14, TJZ13, WPZ16, WO03, WH03, XCX16, XCMT20, ZZT91, ZWCL14]. Optimally [AAV00, GKS^{+}19]. Optimization [JS02, KM90, KAPF05, MZ01, NWHL22, PS22, SSS09, WM05, YTLC02]. Optimizations [GV03]. Optimize [GSZ99]. Optimizing [ZSG^{+}22]. Optimum [CD95]. Option [SGZ02]. Optoelectronic [Sah01]. Oracle [FL09]. Oracles [CISH07, FZT14, IN13, KL00, MM05]. Order [AES18, AB91, BYP95, BGM^{+}18, DG98, DGK08, DZ00, EGPS10, KK19, Lar98, LHG11, Lin08a, Lug11, Set08, Szw95, ZH22]. Ordered [AKS14, ABH17, Bas97, KL11, KO18b, Pro96, Yah12, ZB02]. Ordering [Com90]. Orderings [BC06, BÉ11, GHJS05, RC05]. Orderly [MAN05, ZH06]. Ordinal [Ési12]. Organizing [Láz13]. Orientation [AMOZ07, AJMO11, AJM^{+}21, ZH22]. Oriented [DSS08, KTT20, LX94, XCC16, YZ07]. Oritatami [FKM^{+}21, HKRS19]. Orthoconvex [ST93]. Orthogonal [DKSS11, WNF19]. Oscillating [HFLD09]. Ostrom [WDFN21]. Other [DH96, PSA17, RS13]. Out-of-Core [MMSV23]. Outdegree [AMOZ07, AJMO11]. Outer [LLW^{+}22, MAN06, SR21]. Outer-Independent [SR21]. Outer-Paired [LLW^{+}22]. Output [MS19, Ros00, Rya21]. Outputs [FMR20b, RT16]. Outsourced [YMC^{+}17]. Overcoming [DEKZ11]. Overhead [OM96]. Overlap [BHR09, BKLS20, CCM97, CNT22, DSS15, HS11, LOPR18].

Overlap-Free [DSS15, HS11]. Overlapping [HT95]. Overlaps [AGM19]. Overlay [CDPT16]. Overview [BMSMT11]. Own [GW18].

P [FMV13, AFIV22, CV13, KMG11, PB20]. P2P [Li12b]. Packaging [FBHH01]. Packed [Zha17]. Packet [DES09, GFK98, MMS05, SKL03]. Packing [BDI^{+}11, FFMW19, HJP^{+}13, JZ16, LOPR18, MV11, Nag06, TSFZRP17]. Packings [CZTH13]. Pair [DÉK22]. Paired [LLW^{+}22, Par23b]. Pairing [CST^{+}17, Ros03, Ver09, WZCH19]. Pairing-Based [CST^{+}17, Ver09]. Pairing-Free [WZCH19]. Pairs [CCQ24, GSZ09, ST99]. Palindromes [DD06, MP22]. Palindromic [AACR18, BGI^{+}18, BHNR04, BR18, Çev20, DMMM14, FLST12]. PAMA [LCL06]. Panconnectivity [XZZY19]. Pancyclicity [XZW^{+}21]. Pansiot [GS12b]. paper [Tsu01]. Papers [CS02, CS00b, CVM20, Elb01, KMS02, KBH99b, Pal01a, SR00b, YSM^{+}00b]. Paradigm [Sir15]. Parallel [AC05, AP92b, BS01, BCVVH07, BF97, BKM11, BKM12, BKM15, BBM^{+}12, BZ10, CCM97, CF06, CCF09, CPJ06, CPC99, CR14, CVMVMV00, DP90, DD13, DGL93, DPS97, EAB^{+}16, FBHH01, FNI16, GD12, HB06, HH22, Hea11, HS95, HW17, HN06, IMP12, Kan15, KS11, KSMMT18, LTZ12, LLQ06, LMM^{+}12, LPP92, LLW21, MS07, MIN11, MVMM02, MS99a, MDL97, OS01, OSZ92, Ott13, Ott15, Pal01b, Ros00, Sah01, SS99, SK03, ŠM05, TH01, Tru08, VG01, VJDT05, WM05, WH03, Zaj09, Zom03, ZC05]. Parallelism [BV20, IYD05]. Parallelizing [LR04]. Parameter [AT11, HL06, RZ12]. Parameterization [DD12]. Parameterized [ADHR09, CFRD08, RR18]. Parameters [KPS93]. Parametric [ACFE09, CE98, FK13, NTSH06, PR23].

Parametrized [FK19]. Parent [Lag14]. Parenthesis [Lag14]. Parikh [Ata11, AT16, BM16, BMS18, BCMS20, CFM12, CSAT20, Hon06, Kog21, MS12, PT18, SY10, Şer09, SHN09, SMAN13, Teh15, Teh16a]. Parity [Fri10, FL12, GW18]. Parsing [Bas97, BIIN04, Kog18, Okh06]. Part [Ano01c, CS00b, Elb01, GJV00a, Hin01, JK14a, JK14b, KBH99b, Li00b, MS99b, Pal01a, Pre01, SR00b, YSM^{+}00b, Zom01a, BJ07b, HT12]. Partial [AES18, BSOR10, BS12, BMMR11, BMMR12, FO07, GS18, IZN99, Lin08b, MPS24, MRT95, PRS98, Pat06, PHPJRN^{+}11, Smi95, dBDZ19]. Partial-Total [Smi95]. Partially [AT12, Bas97, KL11, Lag17, MR91]. Parties [XZL^{+}19]. Partition [CZTH13, DJL^{+}07, HPV99, KMŠ21]. Partition-Type [CZTH13]. Partitionable [Li01]. Partitioned [Mat04]. Partitioning [HO99, IZN05, JSPD03]. Partitions [BMS12]. Partners [RRT99]. Party [TYM^{+}17]. Passbits [MB03]. Passenger [GH07]. Past [Gur16]. Patches [XBE02]. Path [AH11, AHL^{+}13, BLL06, FT09, GVL07, HB06, JW08, KM18, KKB24, LMZC20, LLW18, LW21, MVM07, Par23b, Pro96, Yen09]. Path-Controlled [MVM07]. Path-Equivalent [GVL07]. Paths [BSCH22, DPS99, GR03, GKS^{+}19, HKV17, LPC11, MPS99, RLWW96, UU07, YTN01]. Pathway [BCC^{+}11, JRPIP08]. Pattern [BLP18, BCFL12, CCFG12, CHZ06, DPS97, FMR20b, FS05, IST05, KS06, LJH^{+}17, MHT09, ND02, NRS18, SW09, ZYYH14, Zha17, ZZN23]. Pattern-Matching [SW09]. Patterned [SW17]. Patterns [BCN12, DPS93, HK23, LC18, Prů17, SK04, XWY^{+}22]. Paun’s [PHPJRN^{+}11]. Payoff [GZ12]. PC [CVOV11]. PCP [HH24]. Peano [Ruo96]. Pebbles [KMW14b]. Peer [AF20]. Peer-to-Peer [AF20]. Peers [Li12b]. PEI [VP99]. Penalties [WG17]. Perfect [AFB96, GR00, Kur20, PP11, Sun00]. Performance [BLM15, For10, HYLF20, KR97, Li12a, LKM02, NWK05, NKW08, PV98, Qua07, SK01, TZ11, TH01, WR16, YLZ14, YH11]. Period [APMP17]. Period-Doubling [APMP17]. Periodic [CKZ17, CK07, HH24]. Periodicity [BSBZ08, HN10]. Periods [BSOR10, CCI12, GRS21, HG11, KPS13]. Permitting [GTCV19]. Permutation [Nag20, RM98, Wid12, ZZC15]. Permutational [Oka98]. Permutations [CS18, Faa19, GKSZ19, LCXS19, QLWL06, Teh18, XC15]. Persistent [HK09a, Lag17]. Personnel [WD90]. Perspective [TV94]. Petersen [DHIÖ97, Wan21]. Petri [JC03, AH11, BCB12, GRV10, MOM91, Muk92, RHS10, YWY94, Yen09]. PFAs [CdBD23]. Phantoms [JSPD03]. Phase [ZYLW12]. Phenomenon [Kut05]. Photographs [Ami05]. Phrase [MO10]. Phrase-Structure [MO10]. Phylogenies [HLY^{+}04]. Phylogeny [AFB96]. Physical [AD12, JWB03, RS17, SA22]. Pi [Yue13]. Pi-Calculus [Yue13]. Picture [AGM14, BESW07, Gia11, SMAN13]. Pictures [AGM19, Fin04]. Piecewise [BKP18, KP10a, XC15]. PIM [NZH22]. Pipelined [BT00]. Pipelining [FM01]. Pitching [ÜS02]. PKI [AH07]. PKI-Based [AH07]. Place [GPC09]. Placement [AC05, DRDN08, URS07]. Planar [BPT06, KLB13, MTNN99, Pre90, RLWW96, Toš06]. Planarity [CDFK19, DOR06, HL06]. Plane [AAV00, Mar08b, Mar08a, MAN05, MAN06, MNN06]. Plateaued [XCX17]. Platforms [DPR^{+}08, DEMT05, KSMMT18]. Platoon [PS22]. Playing [FZ12]. Plays [GW18]. PLC [XWY^{+}22]. Plus [Pau24]. PMC [LHD^{+}24]. PN [ZH13]. Point [Aku06, DD12, MB17, Pre90, RAB15, ZC13]. Point-To-Point [ZC13]. Points [DLT06, Kar99, SSK96, Toš06]. polar [ZWCL14]. Polarizations [FRV19].

Polarized [MGCVdlP20]. Pollinating [WM05]. Polling [TL99, Tse16]. Polygon [BJD20, KM18, SRN^{+}20]. Polyhedral [AAH02]. Polymorphic [APP91]. Polynomial [AAV00, AP90, BCFR07, BB99, BLS^{+}05, Cai94, Dic93, GKRS10, GO09, HH24, HW00, HT04a, HT04b, Ibr22, IZN99, Joh00, MX11, PLMZ11, Shu07, Tra02, WD20]. Polynomial-Time [IZN99]. Polynomials [EKKS18, RW11, TWZ11, XLZ19, ZZC15]. Polytime [Cap96]. POPS [DR05]. Popular [Dar13]. Population [HJW11, Sun11]. Port [NN93]. Portfolio [NWHL22, YTLC02]. Posets [Bed18, Yah12]. Position [AMZ20, MCM^{+}11]. Positioned [LK11]. Positive [CM92, HJ91, KY96, MAG09, Oka99, Oka00]. Possession [ZPXX17]. Post [DRS14, Fin12, HH11]. Potential [AES18, GQZ15]. Power [BMP15, BCP22, CCFS07, DSS15, Fuj16, HIR^{+}92, IPR07, JWB03, Kar09, Mal15, MRS97, Mer08, RHS10, RS04, Sal11, Slo95, SRPC11, Sta05, Sto92, Sut03, WD20, LBJ03]. Powerful [ACMP20]. Powers [CRSZ11, CFIJ10, Faz11, Sha04, Shu11, YTN01, Ram05]. Practical [CSY03, Fuj17, PPR18, TH01, ZLW^{+}17]. Practice [BCFR07, CCFG12]. PRAM [FPP03, For10, JHK08, TV94]. PRAM-Algorithm [JHK08]. Precedence [JSO10, KD99, LTW02]. Preclusion [CLLL08, Fen22]. Precoloring [EL13]. Predecessors [AHR02]. Predicate [vdM00]. Predicates [SWZ97]. Predicational [ES01]. Predict [SB01]. Prediction [BDC90, SZ22]. Predictor [WY22]. Preemptive [HL04, HLW09]. Preface [ASTZ12, AY99, Ano01b, Ano03b, Ano03c, Ano03d, Ano03e, Ano04b, Ano05b, BC14, BRST07, BN07, BN08, BFN10, BNF11, BFN12, BP11, Cal05, CP19, Câm20, Cha03, CLR19, CVV08, CVÉ10, CVM20, CVDV22, DR06, DP13, Den02, DN11, DW11, DS08, DS11, DÉ12, DLMS12, Ési15, FSTY16, FGM^{+}11, FKN11, FRS24, GP08, GJ07, GH09, HP08, HP09a, HS17, HRS17, HK08, Hol12, HK15, HK21, HY06, IY07, IR09, IV18, Ito10, JS21, JR14, KO18a, LJF22, MH12, MBR18, ML12, MP12, MNP12, MR13, MR23, NW03, NW04, NB06, NY10, PPJ06, PT07, PV13, PS18, SY05, Sek20, Shu16, Sos09, Wan06, YN08, YI13, Yu11, YYW19, Zom01c, DRS23]. Preference [FDFZB12]. Preference-Based [FDFZB12]. Prefix [AGM14, CDPT16, CFPR03, DGMM15, EH15, EHS15, FMR20a, Han13, JPŠ19, OM96, Sta07]. Prefix- [JPŠ19]. Prefix-Free [EH15, EHS15, Han13, Sta07]. Prefix-Suffix [DGMM15]. Prefixing [GM90]. Premaximal [PS12b]. Preorders [GW18]. Preprocessing [AKMW20]. Presemifields [BH11]. Presence [Cig04, LPS07]. Present [Gur16]. Preserving [Gaz06, LO13, Mal15, NTSH06, Ric19, SZQ^{+}17, XZL^{+}19]. Prices [FFMW19, SZ22]. Pricing [SL21]. Prime [CFPR03, KYZS17, WD20, YW20]. Prime-Power [WD20]. Primer [BRM07]. Primitive [CJ20, DR12, FLST12, KMS11]. Principal [Hir91]. Principles [AMR05, AD12, Obt06]. Prints [Şer09]. Priority [CS99, Elm06, GZ12, GNC^{+}03]. Prisms [CSN21]. Privacy [XZL^{+}19]. Privacy-Preserving [XZL^{+}19]. Private [BNBN20]. Privileged [FJPS16]. Probabilistic [CZOdlH17, CHYT14, CMR07, CMRR08, DTY15, DY19, Fre08, HV02, HIIW01, IKPY21, Mad03, MDAPHPJ^{+}11, PBMZ06, RHN^{+}22]. Probabilities [Szw95]. Probability [CJ20]. Probably [MMS17]. Probing [Li12b]. Problem [AA20, AP92b, Asa23, BJD20, BLR09, BCR11, BCD14, BHK18a, BB04, BL01, BDG^{+}11, BLM15, BDI^{+}11, CF06, CCF09, Che22a, CKK02, DDD18, DGN07, DRDN08, DRS14, DD13, FPS02, FZ13, FP04, Fin12, Fle20, Fuj17, GKS17, GLP07, GD12, HH11, HL04, HJK12, HO99, Hon02, Hon06, Hon07, IMS03, KL03, KKB24, KLS^{+}19, KMŠ21, LAHN14, LW05, LW06a, LZ12, Lin07, LL23, MXY^{+}04, MS20, Mar92, Mar08a, MGCVdlP20, Nag20, NSVA12, NB18, NAK^{+}15, Pan91, RC11, SB17, SS07b, Ste93, Tor13, Tor15, Vin05, WD90, YTLC02, ZZZ16, Ueh99]. Problems [AK06, AMT20, AE05, AB91, BPR09, BHK05, BKP18, BCC13, CCF08, CHWX09, CCI12, CD95, CR15, CS93, DH05, DJL^{+}07, FZ15, GC15, GGR14, GPPJR13, Gol90, Gon01, Hut02, IDY08, Iba11, Iba15, JMSO05, Kar09, KPSC08, Lar98, Löd15, Loh10, LOPR18, Man15, MVM07, RWZ01, RLWW96, TY15, WG17, Yen08, ZYLW12]. Procedure [GN04]. Procedures [BET03, FMC04, FK05, FKT07, Sal11]. Process [AH07, DD12, GCK08, Kri97, SA22, SN13]. Processes [Cas95, FGH^{+}07, HW10, SMS92]. Processing [BRSRC11, CW11, HS95, HLW09, KBH99a, SSS09]. Processor [CE98, Leu04, RR06]. Processors [DT20, DM08, FRV19, HB06, LY94, MCM^{+}11, MGCVdlP20, NKW08]. Product [DPR^{+}08, MS12]. Production [Wil91]. Products [BK16, CV14, CR15, TSS13, YLX22]. Prof [SSS13]. Profile [Car11]. Program [RR04, Rud15, Wan04, XWY^{+}22]. Program-Based [RR04]. Programmed [Fer07]. Programming [Ano01c, Cos90, FZ02, GN04, Hin01, ND02, NWHL22, Pre01, RR06, Rov00, Sub90a, Sub90b]. Programs [ACV13, BM90, BAK12, BET03, CIY01, CJS92, HB06, HV02, Jai95, RKRR02, Sao92, Sto92, Tha91, Vik96]. Progress [APV06, Pal03]. Projection [LL20]. Projections [BK24, TZ91]. Prolog [HST01, MT95b]. Prolongeable [CDJ09]. Promoters [Sbu06]. Promoters/Inhibitors [Sbu06]. Proof [AKS95, GN04, GM90, Kog21, Kre21, LYX^{+}19, Nag20]. Proofs [Arv97]. Propagating [MMK22]. Proper [MM97]. Properties [AB91, BK24, BMS18, BLL06, Çev20, CRS12, CC98, Dai97, DPR07, DH96, DD08, DD06, DQFL12, DMSS16, DK12, FH05, FY11, Fin21, GK11, JC03, KMS11, Kun16, LOZ98, MT10, MMR10, MSMR22, NPSY00, Pri06, RS13, Sak01, TW09, Vor18]. Property [DFK23, Elm06, Gaz06, HIIW01, Ric19, WM13]. Proportional [GPS14]. Proposal [Špr09]. Propositional [Par23a, Pla96, Sal13]. Protect [YMC^{+}17]. Protein [HMZ05]. Proteins [PPRPS11]. Protocol [BV98a, Gua21, GCK08, HCETPL^{+}12, HT09, KMZS19, XZL^{+}19]. Protocols [ADR11, CIS03, LWS^{+}20]. Provable [YYW19, ZPXX17]. Provably [GH13, RMZW19]. Proving [GHS13, GRRS14, Sak01]. Proxy [DZH16, LMG20, MLO17]. Pruning [WD03]. Pseudo [KMS11, Rya21, ST93]. Pseudo-Primitive [KMS11]. Pseudo-Random [Rya21]. Pseudorandom [NAK^{+}15]. Pseudovarieties [Ali16]. PSO [SZ22]. PSPACE [JYF91, vdM00, DW03]. PTAS [DFLL02, GJKS18]. Public [GKS17, HLH19, LYX^{+}19, LMG20, NZH22, WZ15, YMC^{+}17]. Public-Key [GKS17]. Publicly [SZQ^{+}17]. Pumping [GTCV19, MP07]. Pure [JM03, Mal07]. Pursuit [IML04]. Push [BCC^{+}96]. Pushdown [AK14, AKS14, ABH17, AKMW20, CVMVMV00, DÉK22, GPP20, HKMW22, IJT^{+}93, KMO10, LNP16, Löd15, Lug11, Mas13, Nak18, OS19, Ott15, PI95, Pig09, RT16, Sao92, Set08]. Pushout [ALR04]. PVsub [AP92a].

Q3Ap [LMM^{+}12]. QoS [PS22, XLC^{+}04]. Qsort [MIN11]. Quadratic [BBP11, CCI12, KS10, NWHL22, NSVA12, XCX17]. Qualitative [CMWZ19, ZL22]. Quality [MKB^{+}11, PS22, YLX22]. Quantifiers [BV98b, Lüc18]. Quantifying [AS18, EGPS10]. Quantisation [CCM11]. Quantitative [DV14, DRS23]. Quantum [ATK12, Arn17, AD12, BMP03, BCD14, BMP15, BB03b, FZ15, Fia08, GRB03, GJMP06, Gro03, GQZ15, IMS03, IN13, KR03, Kud07, LB04, NR18, Nak18, Nis03, SY12, YSD16, Yam03, ZQL12]. Quasi [Ber13, MT10]. Quasi-Eulerian [Ber13]. Quasi-One-Cluster [Ber13]. Quasi-Relabeling [MT10]. Qubit [GRB03, JM03]. Queens [MGCVdlP20]. Queries [Arn17, Ars15, Cig04, GSZ99, Lag14]. Query [CW11, Lag17, Mee12, ST99, VG01]. Query-Based [VG01]. Query-Optimal [Lag17]. Querying [TV14]. Questions [IR14, Shu14]. Queue [AKMW20, Elm06, Iba02]. Queue-Connected [Iba02]. Queueing [YLZ14]. Queues [CS99, Fer07]. Quickest [GR03]. Quickheaps [NPPS11]. Quine [RS95]. Quine-Bernays [RS95]. Quirky [Lüc18]. Quotient [BL12, OS19, WD20].

R. [Ble21]. Rabbit [FSWF11]. Radical [BW14]. Radio [DGN07]. Radius [Coo17, DESW05]. Ramsey [PDPPJ11]. Random [BT17, BKS12, FZT14, KPM15, Li12a, MD00, NR21, NPSY00, Rud15, Rya21, Sub05, ZK19, ZG13]. Random-Access [Rud15]. Randomized [BDDN01, BHK^{+}18b, DR05, FDFZB12, Li00b, MD00, RS00, SRR15]. Randomness [KMZS19, Sun00]. Range [DGN07, MS99a, Poo04, RGR11]. Range-Aggregation [RGR11]. Ranges [HH20, Jir14, WY05]. Rank [GI19, KM19, Sun00, TA17]. Ranking [BPZ07, DPS99, ERW04, MPS99, Nak04]. RAQM [DRS23]. Rate [GKRS10, Pal03]. Ratio [FCS05, HZZT12]. Rational [AK06, BGN10, CK18, FW24, Fin12, Fin21, GC18, KMŠ21, RC05, RS15, Shu07, TWZ11, ZC13, ACM11]. Rationale [CFMR05]. Re [LMG20, MLO17, RR06]. Re-Distribution [RR06]. Re-Encryption [LMG20, MLO17]. Reachability [BKP18, FT09, GJV00b, HBIT08, IBS01, IDY08, Kar09, KPSC08, LN08, Mar09, Set08, SN13, TY23]. Reachable [BCV23]. Reaction [APMP17, BFM06, BLR20, BEMR11, EMR10, EMR11, EMRB12, ER14, Sal13, TA17, Teh18]. Reactions [HFLD09]. Reactive [SR00a]. Read [AS18]. Real [KD99, Leu04, LCY12, Pal03, Rya15, SK01, XWY^{+}22, YS13]. Real-Life [Rya15]. Real-Time [KD99, Leu04, LCY12, Pal03, XWY^{+}22, YS13]. Realistic [DVG03]. Realizability [LBL06]. Realizer [MAN05]. Realizing [LC18]. Reals [Mee12]. Realtime [DY19]. Rearrangements [SSK96]. Reasonable [BHK18a, Kre21]. Reasoning [DN07, EN03, LSWW13, MT95a, TW09]. Reassignment [KZ10]. Rebalancing [LF96, MO94]. Rebound [IIK^{+}04]. Receiver [CCD07, TFS19]. Recipient [ZYZ^{+}19]. Reciprocity [SB12]. Reciprocity-Based [SB12]. Recoding [GPC09]. Recognition [DP90, GV03, JP07, Mod21, NWK06, Okh03, YS13]. Recognizable [DK98, Fin04, FV24, Gia11, RW11]. Recognize [CR15]. Recognized [MM05, ZQL12]. Recognizer [SRPC11]. Recognizing [BM90, FS21, GI19, LT21, LWW00]. Recombinants [BRSV13]. Recombination [DDM07]. Recommendation [GWL02, LWW22]. Recomputation [NZH22]. Reconfigurable [BBFZM06, BT00, FZEBB05, FZFDCHB05, MDL97, PA98, RM98, WH03]. Reconstructing [FLM^{+}21, FS06]. Recovering [IN13]. Recovery [WZ15]. Rectangle [Uen13, WLC12]. Rectangles [KTT20, Nag06]. Recurrence [Dur13, LS98]. Recurrent [MO94, NP09].

Recursion [JK14b]. Recursive [APP91, AT12, KM02, Kap05, Kut05, LZ93, LPC11, Sal11, YCTW10]. Recursively [vLW15]. Red [CS96, MC02]. Red-Black [CS96, MC02]. Redex [FW90]. Reduce [CKW09, Li12b]. Reduced [GI22, Sut03]. Reducibilities [DR94]. Reducibility [HJ97]. Reducing [BCFR07, KH21]. Reduction [BHR09, DG09, FMR20a, HH11, Hit20, MS19]. Reductions [AV96, HJ91, Zan91]. Reducts [Wan14]. Redundancy [VS93]. Redundant [WXF16]. Reed [Arn17]. Reference [IMP12]. Refinement [CFH^{+}03, HPV99, MH06]. Regex [Sch13]. Region [DRDN08, YW06]. Regional [NZH22]. Register [ACMP20, HFLD09]. Registers [HG11, XLZ19]. Regression [MM11]. Regular [Ada10, AK06, AK10, AB17a, BR20, BLS20, BS16, BMMR19, BT13, Brz13, BL14, BD19, Cal15, CSV02, CSY03, Cha02, CLOZ04, CDJ09, COT12, CS02, CS00b, CKW09, Coo17, CFPR03, DK11, DM11, EJ23, Elb01, EH15, EHS15, Faz11, FS21, FO08, GKRS10, GV23, GH13, GH15, GLPP22, GZZX21, HWW06, HKS13, Han13, HK03, HK11, IW07, Jeż08, JM11, Jir14, KMS02, KEH16, KLH16, KMRY20, KBH99b, KMM06, Loh10, NR21, NPSY00, PP14, PT90, RS12, Sel08, SR00b, SL17, TV14, Tei17, TW09, YSM^{+}00b, YJ05, Fin12]. Regular-Expression [Han13]. Regularity [BKW02, Mal15, Pal08, RS13, ST16]. Regularity-Preserving [Mal15]. Regulation [BDL08]. Regulatory [AES18]. Relabeling [MT10]. Relabelings [Kan15]. Related [AO11, AB17b, BPR09, CHZ06, Iba11, KB20, TY15, WDFN21, WLC12]. Related-Key [WLC12]. Relating [BT00, Mal05]. Relation [GZZX21, HK95, HN10]. Relational [Lar98, Lar99, Tha91, VS93, YBI11]. Relations [BK95, DI02, DZ00, Fin12, JF18, KMŠ21, KL10, Lin08b, TZ91, WGD18]. Relative [CMRR08]. Relaxed [JL01, LF96]. Relaxing [De 06]. Relay [CIS03]. Relevant [CCI12]. Reliability [Jai98, ZWC^{+}22, ZLL23]. Reliable [MGJ19, YBM11]. Remarks [BSBZ08, CSN21, Das21, FJPS16, Hon02, Kud07, MMY10, Tru08, VG01]. Removal [HKRS19, KTT20, Moh02]. Removals [GPS14]. Rendezvous [CDPR11, EP17]. Repair [LZGF16]. Repeated [Cig04]. Repeats [Riv04]. Repetition [VG01]. Repetitions [CdL04, FJ12, GS12b, IYZ04, PL23]. Replacement [LLW18]. Replication [Qua07]. Report [APV06]. Reporting [SJ04]. Representable [TST01a]. Representation [BB99, BJ05, BJ06, BJ07b, O’N15, ROK08, WXF16, XHLF02, Zho02]. Representations [BB03a, BK16, HP09b, LP19, PPJY08, ZZ18]. Representative [TBGP20]. Representing [HKKŠ13, Smy12]. Requests [CVPV08]. Required [Sun00]. Requires [Fri10]. Research [FH11, GPPJR13, SZ22, XCC16, Zom03]. Resemble [KMS06]. Reservations [KL05]. Reset [Gus13, GP15, Mas19]. Residual [AO11, Dan11, YB19, ZLG21]. Resiliency [CL07a]. Resilient [SNWW06, TCT14, YBM11, ZYZ^{+}19]. Resolution [Pla96]. Resource [BRSRC11, BDG^{+}11, CTZ01, FM01, PS22, SVSN01, WG17, YH11]. Resources [RS17, SB01]. Respect [RR18]. Restarting [JO07, KR08, KMO10, KO13, KO18b, MO07, MO09, MPJ07, POM22, PM13]. Restricted [BMS18, BFL02, BBE24, BE19, CSAT20, DP90, DS05, GWL^{+}17, MNS18, Nis03]. Restriction [FFH15, HCG96, HLW09]. Restriction-Fragment [HCG96]. Restrictive [PB20]. Result [CP06, ES01, LD01]. Resulting [HH20]. Results [AA13, BGRY16, BKM11, CD06, CKZ17, DGMM15, FOP05, HK09b, LS98, MSV23, RS04, SYS19, Sbu06, YKCW23, YWY94, ZLG21]. Retrieval [CCF09, FMN06]. Returning [BKM15, BD19]. Reusability [KR03]. Reusing [FZ03]. Reveal [LKM02]. Reversal [CGKY12, Jir14, Rao08]. Reversals [QLWL06]. Reversibility [Iba11]. Reversible [AKMW20, GI22, HJK18, KPS18, KM23, LP19, RN22]. Reviews [ZZC22]. Revisited [AMR09, DR94, FJ12, GV23, KS11, KX12, LT21, Nag21, Pre90, TA17]. Revisiting [DPR^{+}08]. Revocable [SZFX20]. Revocation [HYT15]. Rewrite [AMR09]. Rewriting [Bar90, BCVVH07, BPT16, BKKR01, FW90, GHWZ05, KMS06, Luc09, Mad03, ND02]. Rewriting-Based [ND02]. RFID [HCETPL^{+}12]. Rhythms [CIRS08]. Rich [PS12a]. Right [BH20, CNT22, FLM^{+}21, KH21]. Right-Bounded-Block [FLM^{+}21]. Right-Infinite [CNT22]. Rigid [GJV00b]. Rigidity [BDD^{+}18]. Ring [CL98, CCQ24, DSS08, GS12a, LW06b, Mar97, Sub90a, Sub90b, ZGCZ18]. Ring-Theoretic [Sub90a, Sub90b]. Rings [BW14, CX98, EN03, FHL07, GLP07, YWY94]. RLE [HI18]. RLE-Compressed [HI18]. RNG [CIS03]. Road [CKK02]. Robot [SH22]. Robots [BFMBS11, BT17, CGK^{+}21, DDPS19]. Robust [DPR07, DW03, ECY02, HJ91, HJV93]. Robustness [AB17a, MCS08]. Roman [SR21]. Root [CHZ06]. Root-To-Frontier [CHZ06]. Rooted [GWF^{+}24, Yah12]. Rosser [KM07b]. Rostering [MZ01]. Rotation [GW24, SFL17]. Rotations [MO94]. Rotator [KHLC12]. Rough [TSS13]. Round [CLT14, KLP20, LJ17, LZZN22, TYM^{+}17]. Round-Optimal [KLP20]. Route [GR03].

Routed [PV98]. Router [LOD07a, LOD07b, MMS05]. Router-Based [MMS05]. Routing [BDC90, BDDN01, CHA^{+}92, CHYT14, Cig04, FPS02, GD98, GFK98, GP17, JW08, KAPF05, LPC11, OS01, PA98, RM98, RS01, RVT06, Sib97]. Row [MS20, WAG^{+}06]. RP [BJY90]. RSA [BNBN20]. Rule [Fer07, dMLBPP20, PB20, SKL03]. Rulers [BMP03]. Rules [AFO06, BCHK09, TBGP20, Zet11]. Rumors [XCC16]. Run [LD01, MHT09]. Run-Time [LD01, MHT09]. Runs [FY08, FJ12, KMIS09]. Runtime [Rud15]. Rupture [ABT16, Asl16, AO10, AA13, BTO17, KA18, LDLW17, YB22, YB23].

Safe [Cap96]. Safety [CHYT14, IBS01]. Salesman [BL01]. Salesmen [Klo96b]. Sampling [CCP18, MM17]. Sanitizers [YM19]. SAT [HW10, YW06, ZG13, ZK19]. SAT-Based [HW10, YW06]. Satisfiability [DDD18, MTVM09, MTVM15, ZSW14]. Sato [RKRR02]. SBN [KR97]. SC-Expressions [YZ07]. SC320 [MDL97]. Scalable [BBFZM06, Hei97, WHLH17, WH03]. Scale [CDLW05, DCS13, DEMT05, MDL97]. Scales [CM12]. Scan [JP08, PRS98]. Scanning [DES09]. Scattered [Bed18, DSS08, ÉO13, ÉI14, MMK22, RC05]. Scattering [BFMBS11, BT17, KA18, WQY16]. Scenario [YTLC02]. Scenario-Based [YTLC02]. Schedulability [WR16]. Schedule [CD95, RWZ01]. Scheduler [TSFZRP17]. Scheduling [BV98a, BS01, BLMR05, BNR99, BDG^{+}11, BE19, Cas05, CTZ01, CYZ14, CR14, DFLL02, DEZ01, DLC^{+}14, DEMT05, FL97, FBHH01, FCS05, GJKS18, Gro03, HB06, HL04, HW17, HLW09, Jan93, JSO10, KSMMT18, Klo96b, KD99, LAHN14, LTZ12, LTW02, LLZ07, Li01, MXY^{+}04, MMSV23, Mas04, NN93, Pal03, PY04, PZX07, PFG^{+}01, RC11, SSS09, SS07b, Sun11, SS12b, WY05, WR16, YH11, Zaj09, Zom01b, Zom01c]. Schema [KS11]. Scheme [DCS13, DZH16, FPP03, Fuj16, HHP17, HLH19, LD04, LHT09, LH11, LYHW19, MD00, TWZ11, ZC13, ZGCZ18]. Schemes [FL12, GP17, JSO10, MMS17, PNN^{+}10, SNWW06, Sun00, WGF16]. Schnyder [MAN05]. Schützenberger [DV14]. Science [DRS23, HO00]. Scientific [RR04]. Scope [LNP16]. Scope-Bounded [LNP16]. Score [HN06]. Screening [IN08, IN05]. Search [ACDL18, BRM07, Brz13, CS00a, CGK^{+}21, Fle96, HM04, HLH19, IN05, IN08, JS03, KK90, KNR18, LTZ12, PRN13, WM05, ZZZ16]. Searching [Ami05, CFG12, DE08, KPS93, MP93, ST93]. Seat [KL05]. Seating [KL05]. Second [LHG11, Set08, Szw95]. Second-Order [Szw95]. Secrecy [BKST18]. Secret [LD04, MNS11, Sun00, TWZ11, WGF16, ZC13]. Secure [HLH19, KLP20, LYHW19, MLO17, MG14, MMS17, MGJ19, RMZW19, SNWW06, SNJ11, SZFX20, TWZ11, ZLW^{+}17]. Securing [CST^{+}17, SA22]. Security [DLW02, LW06b, LWS^{+}20, NAK^{+}15, SNJ11, WHLH17, YYW19]. Seeking [MD00]. Segmentation [RHN^{+}22]. Selected [Che22a, Pal01a]. Selected-Internal [Che22a]. Selection [ATK12, CD20, NB18, SA22, SRR15, WRNK03]. Selective [HHN^{+}95]. Self [CDPT16, DDHL11, DTY15, DWS15, FDFZB12, FZAM08, GHJS05, GS12a, GPP20, HKRS19, HHW99, HSS19, JK14a, JK14b, KK10, Kar99, Láz13, MvZ22, NGHK15, ST11, San13, SW17, SZQS18, TY23, TSFZRP17, WD03, XS06]. Self-Assembly [JK14a, JK14b, SW17]. Self-Attraction [HKRS19]. Self-Dual [HSS19, SZQS18]. Self-Organizing [Láz13]. Self-Pruning [WD03]. Self-Similar [JK14b]. Self-Specifying [HHW99].

Self-Stabilizing

[CDPT16, DWS15, FDFZB12, FZAM08, GHJS05, GS12a, KK10, Kar99, NGHK15, ST11, TSFZRP17, XS06, DDHL11]. Self-Verifying [MvZ22]. Selfish [FFMW19, MV11]. Semantics [AG01, BMSMT11, BKKR01, CZ11, Cos90, Kri97, Luc09, MT95b]. Semi [GTCV19, KN21, KK05, SF07]. Semi-Automatic [SF07]. Semi-Conditional [GTCV19]. Semi-Lossless [KK05]. Semi-Simple [KN21]. Semiautomata [BJ05, BJ06, BJ07b]. Semicomputable [TZ91]. Semifeasible [FH05]. Semiformal [Špr09]. Semigroups [AK10, BGK^{+}20, BS15, Fle20, TSS13]. Semilinear [IS12]. Semilinearity [IM20, Yen09]. Semirings [ELS15]. Semisimple [AR16]. Sender [WZ15]. Sending [Asa23]. Sense [BF07, FS98]. Sensing [AKK19, WF17]. Sensitive [MMK22, Ott13]. Sensitivities [POM22]. Sensitivity [ZWC^{+}22]. Sensor [AHL^{+}13, BNS03, DCS13, MKB^{+}11, SH22, SP04, TH22, WY05]. Sentences [Szw95]. Sentiment [ZZC22]. Separability [JM03, Teh16b]. Separable [CM92, KMŠ21, Mat04]. Separating [AAV00, DZ00, MB17, vLW15]. Separation [AA20, Fia08, JSKM20]. Separations [BJY90]. Separators [BBC00]. Sequence [CZTH13, CW11, EGPS10, GD12, HMZ05, KYZS17, Lin07, PYTH10, Rya21, WPZ16, XCX16]. Sequences [Ars15, BLP18, BBM^{+}12, CCF08, CKZ17, CRS12, Coo17, DN07, Dur13, GK11, Hon12, IMP12, KX12, KK19, LJH^{+}17, NP09, Sal07, SS12a, Tho06, WWT20, WD20, WO03, XZS16, YW20]. Sequencing [Sal18]. Sequential [CCFS07, DI05, Fre05, JF18, Kan15, LRT92, Toš06]. Sequentiality [Pau24]. Serializable [Ogi94]. Series [CD21, CR14, Mal05]. Servers [OS01, URS07]. Service [BS01, BCDP08, Li12b, PS22]. Services [SA22]. Set [Aku06, AWF03, BRSV13, CSN21, CGL12, Elm06, FZ15, GRV10, HLW09, KK10, KLS05, KMW16, LLL21, MM97, RAB15, TBGP20, Tor15, Ueh99, WAF03, XCMT20, ZL22]. Sets [AK06, AGM19, BMW91, BMP03, BLL06, CZTH13, CJ20, CYS^{+}12, CL07b, DLT06, DGL93, DWS15, DS05, DR94, ÉK07, FH05, FV24, HT95, HHN^{+}95, Hon06, Hon12, KHLC12, LO11, Mel93, MB17, NGHK15, Prů17, RW11, RC05, Ros90, RS15, SMS90, Sto92, TCLS10, TV94, WPZ16, XCX16]. Setting [BV08, HST01, HHP17, LMG20, TYM^{+}17]. Several [HLC^{+}19, LD04, SH17, XCX17]. Shamir’s [LD04]. Shape [Gaz06]. Shaped [KKB24]. Shapes [MC02]. Shared [BLR09, Mor10, RR18]. Shared-Memory [Mor10]. Sharing [BDG^{+}11, LD04, Li12b, Sun00, TWZ11, WGF16, WHLH17, ZC13]. Sharpened [FP04]. Sheng [CISS12, SSS13]. Shift [HG11, XLZ19]. Shifts [Asv07, BK24, CS18, JP04, Kop21]. Shop [JMSO05, SS07b]. Shops [LLZ07]. Short [FLFR19, IMP12]. Shorter [GH13]. Shortest [AHL^{+}13, CFMS15, DPS99, Hut02, JW08, KM18, LW05, LW06a, MPS99, ST99, XFJ03]. Shortest-Path [JW08]. Should [Ros03]. Shoup [LYY^{+}21]. Shrinking [JO07]. Shuffle [BO97, BMS18, CSV02, CL98, DKSS11, DS05]. Shuffle-Ring [CL98]. Shuffling [EH12]. Siblings [LL20]. Side [SRN^{+}20]. Sided [ACDL18, ST93, XBE02]. Sidel’nikov [YW20, KYZS17]. Signal [BCC^{+}11, LWJ^{+}10]. Signature [DZH16, HHP17, LW06b, LYHW19]. Signatures [HYT15, Ver09]. Signcryption [FZT14, RMZW19, ZGCZ18]. Signed [HP09b, QLWL06]. Similar [FA06, JK14b]. Similarity [Ars15, BOV08, DSS15, HN06, RV22]. Simple [AFB96, BCFR07, CDLW05, CHKL07, Fle96, GNP^{+}06, HH12, HYT15, Huy91, IST05, Jun14, KN21, KM18, MS16a, MS16b, Oka99, SNB24, WAF03]. Simple-Algorithms [AFB96]. Simple-Yet-Efficient [HYT15]. Simplification [Löd15]. Simplifications [KNR21]. Simplified [XCMT20]. Simulate [Dub95]. Simulating [CPJ06, FZCFB08, JWB03]. Simulation [BCDP08, FGS^{+}90, FPP03, FZFDCHB05, FNI16, GB03, KL10, LWJ^{+}10, MDAPHPJ^{+}11, Mat04, Qua07, SVSN01, YB06]. Simulations [ÉM11, KR08, KMW14a, Pet11]. Simultaneous [Sha04]. Since [McN90]. Sinecure [FK06]. Single [ALR04, BNS03, GH07, KS10, SSS09]. Single-Channel [BNS03]. Single-Pushout [ALR04]. Sink [EG02]. SINR [LAHN14]. Siphon [JC03]. Siphon-Based [JC03]. Site [AES18]. Six [EAB^{+}16]. Size [BBP11, BHK18a, Bir11, BMMR12, CSR12, CKW09, De 06, FKM^{+}21, GS12a, KO13, SEE99, Sun11, Uen13, YM19, vLW15]. Size-Computation [GS12a]. Sized [KKB24]. Sizes [ZB02]. Slave [GS12a]. SLDNF [Pla96]. SLDNF-Resolution [Pla96]. SLMAP [HCETPL^{+}12]. Small [AKM^{+}11, ARV12, AE04, CdBD23, CD20, CGL12, CD09, DL12, DGK08, FRV19, HIR^{+}92, KM17, KS10, Leu16, Mer08, PR00, UU07, YSD16, ZB00]. Smallest [FS21, NRT00, SRN^{+}20]. Smart [SA22]. SMP [SK03]. SNQ [PB20]. Soccer [CKL15]. Social [SL21, WZCH19]. Sofic [Sut03]. Soft [Nag06]. Software [BJ07b, FM01, KR03, LX94, Qua07, ST01]. Solid [HS11, ST93]. Soliton [BJ07a, JK07]. Solution [Anc02, MGCVdlP20, NSVA12, Pan91]. Solutions [BIIN04, CK07, Ruo96, ZZT91, ZK19]. Solver [ELS15]. Solving [Com90, Fri10, FL12, GGR14, Gon01, HSS07, Lin07, LMM^{+}12, MNS18, MZ01]. Some [AA19, AA13, AT23, BM16, BCR11, BE95, Bod91, CCF08, Çev20, CKZ17, ÇA18, For10, FH11, GC15, Gol90, GR00, HH20, IR14, IMS03, KM22, KPS93, KNR18, Kud07, Kun16, LL16, MMY10, Mee12, Oka00, Pri06, Shu14, TL99, TY15, YWY94, ZQL12, ZZC15, vdHM92]. Sort [CDX21, Lar98, ZH19]. Sorted [MRT95, Ole92, WO03]. Sorting [BLLS03, BMR^{+}14, BNS03, DR05, FS05, HH22, MRRV06, MIN11, PA98, QLWL06, RM98, WRNK03]. Soundness [Kam98]. Source [GR03]. Source-Based [GR03]. Space [AOSY10, BGRY16, CF06, CZ11, Fre02, HIR^{+}92, IJMP21, JZ16, KM18, Kör03, KTT20, MMP10, PLMZ11, SSK96, Sta05, ÜS02, WNF19, WNF20, YS13, ZZ18]. Space-Efficient [ZZ18]. Space-Optimal [KTT20]. Space-Time [ÜS02]. Spaces [Câm14, CLT09, CMWZ19, HIIW01]. Spanner [SNB24]. Spanners [AWF03, DH96, GS09, WLF03]. Spanning [BBB^{+}18, BB04, Dar13, ERW04, ET14, Fuj17, HLHH06, LLY13, LX17, LZ12, MTNN99, MAN05, Tor13, YCTW10]. Sparse [DR94, ET14, VP99]. Sparseness [DH96]. Special [Ano01c, BRST07, CD02, CVM20, DRS23, FRS24, Hin01, HO00, Hsu98, LC02, MR23, Pal01b, Pre01, RS00, Sek20, Smy12, TY02, Yu02, YYW19, YLX22, Zom01a]. Species [MCS08]. Specific [BIIN04, FV24, LKM02, SKL03]. Specification [BJ07b, SKW08, XWY^{+}22]. Specifications [BMW91, FW24, HK02, LSWW13, SR00a]. Specified [KNR21, Teh18]. Specifying [HHW99, HJW11]. Spectra [CH15, SH17]. Spectral [Coo17]. Spectrum [RK09]. Speed [KKP97, RS17, WH03]. Speed-Up [WH03]. Speedup [BR08]. Spi [BDSV06]. Spike [PPJR06]. Spikes [FIO08, KMG11, PB20]. Spiking [FIO08, IW07, KMG11, PPJR06, PPJR07, PPJS07, SRPC11]. Spin [ILT11]. SpliceTAPyR [TFF18]. Splicing [ARV12, KN21, LMW08]. Split [DES09, GLV14, RS22]. Split-Minimization [GLV14]. Splits [CB09]. Splitting [PRS98]. Spreading [XCC16]. Squad [GLP07]. Square [GS18, JK19]. Square-Free [GS18]. Squarefree [JP07]. Squares [GLP07, Har24, MMR10, ORS08, PR12, Sha04]. ST [MNS11]. ST-Numbering [MNS11]. Stability [AA19, APMP17, EMRB12, KD99]. Stabilization [DTY15, San13]. Stabilizing [CDPT16, DWS15, FDFZB12, FZAM08, GHJS05, GS12a, HKV17, KK10, Kar99, NGHK15, ST11, TSFZRP17, XS06, DDHL11]. Stable [Hol11]. Stack [BBK17, IM21, IJMP21]. Stacked [RHN^{+}22]. Stage [ZZZ16]. Standard [AG01, BPR09, LYX^{+}19, MIN11, PR12, ZC13]. Star [BMMR19, BL12, CC98, CHYT14, CGKY12, DH18, Fen22, GCH20, HLHH06, HY97, Jir14, JPŠ19, MR91, OY11, YJ05, ZZN23, ZH19, WC13, YCL11]. Star-Free [BL12, YJ05]. Start [FO08]. State [AM09, ARS11, AMR11, BGN10, BLMR05, BHK19, BMMR11, BKLS20, CSR12, CZOdlH17, CK08a, CLMP16, CCP05, CGKN08, CGKY11, CGKY12, Das19, DS02, EH15, EHS15, GY12, GPS14, HS08, HKNS16, HK02, HH20, IBS01, JJS05, Jir14, KPS18, KEH16, KLH16, KLS05, Mac96, NRS18, NRS19, OS19, PS02, PR11, SS07a, SY07, SMS92, SN13, WGD18, Yen08]. State-Based [HK02]. State-Size [CSR12]. Stateless [KMO10, KMW14b, Mas13, YDI08]. States [BLR09, BMP15, CdBD23, CP03, HKKŠ13, JM03, LB04, MVMM02, NWK06, ZQL12]. Static [BET03, Câm14, Cas95, TZ11]. Station [DRDN08]. Stationary [PT14]. Stations [FZ03]. Statistical [GK11, Mal18, MG14]. Stay [BC12]. Steady [BLMR05]. Steady-State [BLMR05]. Stealing [Ros00]. Steganography [RN22]. Steinby [FRS24, SS24, Ste24]. Steiner [Che22a, RR18, SK20, SSK96, SB17, Tor15]. Stencil [Leo03]. Step [LOZ98, Muk92, ZYLW12]. Steps [FT11, JWB03]. Stepwise [KN93, MM11]. Stevens [Fri10]. Stevens-Stirling-Algorithm [Fri10]. Stigmergic [DDPS19]. Still [ACMP20]. Stirling [Fri10]. Stochastic [Li12b, SB01, Tor13]. Stoichiometric [MM11]. Storage [OM96, WHLH17]. Store [CD95]. Stored [Rud15]. Stored-Program [Rud15]. Straight [Pat06]. Straight-Line [Pat06]. Strategies [BRSRC11, BKKR01, Fia08, GZ12, Rog09, TZ11]. Strategy [BC12, FL12]. Stream [BRSRC11]. Streaming [AF20, BLM15]. Streams [Lin07]. Strength [MS18, ZWC^{+}22]. Strict [RS13, WPX^{+}21]. Strictly [Dai97, MAG09, RS12, XCMT20]. String [BH02, CZOdlH17, CF06, CCI12, DJR18, DS96, FY08, GHWZ05, KMG11, KMIS09, LRR08, LCL06, NWK06, NKW08, YBI11]. Strings [AAI^{+}20, BCFL12, CFIJ10, DD08, FS05, Fre05, FRS06, IN13, JP07, Lag14, PL23, Smy12, SW09, TCLS10, ZBS05, Zha17]. Strong [BMMR19, BJY90, DP14, GM90, Iba11, MGL23, NGHK15, PT18, PT19, Teh16a, WLZT21, YZZ22]. Stronger [NPPS11]. Strongly [GLPP22, HHP17, ZH22]. Structural [BCB12, JK14b, XWY^{+}22]. Structure [AK10, BSG03, CCF08, Che22b, CISH07, GZY24, HK95, IIT91, JMR91, LKM02, MGGP08, MO10]. Structures [ACV13, Cha02, ER14, JK14b, LOD07a, LOD07b, Lin08a, RGR11, SKL03, Sun00, SFL17, WRNK03]. Study [CSY03, CSAT20, FK06, VJDT05]. Sturmian [BPR09, DD06, Mig90, PR12, Tho06]. Style [RKRR02]. Sub [PL23, Yan21]. Sub- [PL23, Yan21]. Subalgorithm [Nis07]. Subarrays [BT07]. Subclasses [BHK05, Gia11, TSZ16]. Subcube [MGL23]. Subcubic [SG04]. Subdivision [XHLF02]. Subdivision-Based [XHLF02]. Subgraph [AMT20, AB91, GMU15, WLZT21, XZL^{+}19, ZZN23]. Subgraph-based [WLZT21]. Subgraphs [ESS20, ET14, Far20, LT21]. Subgroup [FZ13, IMS03]. Subgroups [Ble21]. Sublinear [FMN06, Mod21]. Sublinear-Time [Mod21]. Sublinearly [MMP10]. Sublogarithmic [HIIW01]. Submatrices [WAG^{+}06]. Submodular [SSS09]. Subnetwork [Fen22]. Suboptimal [GD98]. Suboptimal-Optimal [GD98]. Subregular [HJK12]. Subregularly [DST10]. Subsequence [AE05, DD13]. Subsequences [YW20]. Subsequential [AM03]. Subset [CIS03, Mar09, Vor16]. Subshifts [MM17]. Subspace [WNF19]. Subspaces [WNF20]. Substitution [KN93, Kam98, Mal07, MCM^{+}11]. Substitutions [Dom12, KL03, Tho06]. Substitutive [BDD^{+}18]. Substrings [DS96, IB12]. Subtree [BVM00, Gre96, HLY^{+}04, KEH16]. Subtree-Free [KEH16]. Subversion [LWS^{+}20]. Subword [BPR09, CK08a, Čer08, Faz08, FM13, JPŠ19, MS04, Sal07, SY10, TSZ16]. Subword-Free [JPŠ19, TSZ16]. Subwords [AČ11]. Successful [Rog09]. Successors [FS21]. Succinct [BMP03, HBN08, KRK16, ROK08]. Sufficient [KL00, Oka00, WFG15, ZWW^{+}14]. Suffix [DGMM15, FMR20a, FS06, GPC09, HBIT08, Hol11, JPŠ19, LJA09, MM05, NRS19, PL06]. Suffix- [JPŠ19]. Suffixes [BMR^{+}14, FS05]. Suggestions [FH11]. Suites [BMS12]. Sum [KMIS09]. Summary [GH15]. Sums [Sal11]. Super [CV14, LLY13, LX17, Yan21, ZK19].

Supercompilation [LN08]. Supernode [JS03]. Superstring [LW05, LW06a]. Supertrees [NRT00]. Supply [IZN05, YLX22]. Support [LRR08]. Surface [BPT06, KTT20]. Surfaces [AAH02, Fre02]. Surveillance [MKB^{+}11]. Survey [DGK08, Man15, MOM91, PPJS07, PPRPS11, Riv04]. Survives [JYF91]. SVMs [ACM11]. Swaps [CCFG12]. Swarm [dMLBPP20]. Sweep [GM19]. Switched [RVT06]. Switches [GFK98]. Switching [GP09, KG11]. Symbol [AFO06, JSKM20, NCC^{+}07]. Symbolic [BB03a, Bee95, BCPR07, Com90, MC13, MB06, Set08]. Symbols [DV11]. Symmetric [GJV00b, GW24, MvZ22, O’N15, SFL17, TWZ11, TH22, Van05, KR97]. Symmetries [BDSV06]. Symmetry [Čer08, MRS97]. Symplectic [WNF20]. Symport [AFO06, ARV07]. Symport/Antiport [AFO06, ARV07]. Synchronization [FMV13, GLP07, Vor16]. Synchronize [BGMV08, IT13]. Synchronized [AK14, CKK02, HIR^{+}92, Slo95]. Synchronizing [AR16, BBP11, Ber13, BN20, CJ20, TY15, dBDZ19]. Synchrony [SR00a]. Syntactic [BL14, KM08, Sak01]. Syntax [KM23]. Synthesis [BBL^{+}12, SF07, XWY^{+}22]. Synthesizing [FW24, HK02]. System [AMR09, BGMV08, CLT14, EZ01, FK06, GWL02, GM90, HK95, LYX^{+}19, LC22, NSVA12, SK01, TA17]. Systematic [JP08]. Systems [ADHR09, AFO06, ARV07, AKM^{+}11, ARV12, AFIV22, ABL^{+}11, AF20, AKS95, AKS14, APMP17, AH07, Bar90, BCVVH07, BLR09, BF97, BCC^{+}11, BFM06, BV20, BLR20, BEMR11, CE98, CD06, CCFS07, CFH^{+}03, CZ11, CVMVMV00, CVPV08, CVDV10, CVOV11, CK07, DPR07, DI05, Das04, DV11, DG09, DEZ01, DZ00, DG90, DPS97, EMR10, EMR11, EMRB12, ER14, FFH15, FOP05, Fre05, FO07, FO08, FIO08, FH11, GH07, GHWZ05, GM90, GCK08, HKRS19, HT12, HK02, IBS01, IYD05, IW07, Iba11, IST05, ILT11, JP06, JP04, KN21, KNR21, Kri92, KMG11, KMS06, KMK11, KRK16, LK11, LCVLV09, Láz13, LTZ12, LZGN06, Leu04, Li01, LCY12, Li12b, LWW22, LMW08, dMLBPP20, Luc09, Lug11, Mad03, MS07, MM11, MVMM02, MDAPHPJ^{+}11, MMK22, MT95a, Mas09, MO07, MO09]. Systems [MDL97, Mor10, Muk92, MC13, MB06, OY11, Ott13, Ott15, PDPPJ11, Pău00, PPJR06, PPJR07, PPJS07, PPJY08, PPRPS11, PB20, Pen93, PBMZ06, PT90, PR23, PLMZ11, PFG^{+}01, PSS12, Qua07, RCTC^{+}09, SA22, Sal13, SVSN01, Sbu06, Set08, SRPC11, Sta05, Sun05, Sut03, TA17, Teh18, Toš06, Tru08, WC04, Wil91, YDI08, Yen08, ZC05]. Systolic [FGS^{+}90, MP91].

Table [BESW07, LWW00, NKW08].

Table-Driven [BESW07, NKW08]. Tables [HI18, LOD07a, LOD07b]. Tags [HMZ05]. Tally [DR94]. Tamaki [RKRR02]. Tandem [Riv04]. Tape [AMR11, CGKN08, Nak18, NCC^{+}07]. Tapes [KSY14]. Tardiness [KS10]. Target [DEKZ11]. Target-Controlled [DEKZ11]. Task [BNR99, DEZ01, EZ01, FL97, FBHH01, MMSV23, RR06, Sun11, YH11]. Tasks [HL04, LTW02, MZ01, ZC05]. Taxonomies [KSJ08, ROK08]. Taxonomy [CFRD08, Glö10]. Technique [EL13, RN22]. Techniques [FZ02, HPV99, RK09, SEE99]. Technology [SH22]. Telecommunications [AC05]. Temperature [JK14a]. Template [DDM07, WH03]. Template-Guided [DDM07]. Templates [ER06]. Temporal [GN04, LRT92, MG20, PQ06, Pen93, SMS92]. Tenacity [LWYL14]. Tents [ÜS02]. Term [Bar90, FW90, TST01a]. Terminating [Mas09]. Termination [CGR13, DPR07, DG09, GHWZ05, KM02]. Terms [CSAT20, Hir91, JC03, OY11, YTN01].

Ternary [Jir11, XCX17]. Tessellation [Prů17]. Test [AKM^{+}11, BMS12, CDJ09, FK13, WZCH19]. Testability [RS13]. Testable [KP10a, RS12]. Testing [AMR11, BDSV06, CLT09, CL10, CDFK19, HL06, MSR06, Mer08, WCD^{+}14, Yah12]. Tests [KY90]. Tetration [Hit20]. Text [CK08b, KK05, ZHZ11]. Texts [CFG12, CIRS08, IB12]. th [YTN01]. Their [CLLL08, CK18, HJ14, KM08, KMS11, KP10b, KY96, LO11, MS16a, MS16b, POM22, QD03, SY07]. Theorem [BC06, BSOR10, BGS11, DV14, GN11, GHS13, GRRS14, Kog21, Kre21, MRSS19, Ruo96, SMS90, VG01, KPS13]. Theorem-Proving [GHS13, GRRS14]. Theorems [Fin19, Suc90]. Theoretic [Çev20, DGMM15, FH05, FZ15, GC15, Pan91, Sub90a, Sub90b]. Theoretical [Ami05, HBN08]. Theoretically [TWZ11]. Theories [CGR13, Mar92]. Theory [AR16, AD12, BLS20, BK95, BRST07, Bur12b, Kam95, Láz13, McN90, MR23, SMS92, Sek20, Smi95, Suc90, Tor15, Tsu01, TST01b, Wan04, YLZ14, Zom01c]. Thesis [AD12]. Thompson [Ble21]. Thorny [YB22]. Thoughts [Mee12]. Three [Cha02, CLT14, CK07, ET14, Fin12, KKH90, NS18, Tse16]. Three-Edge-Connected [ET14]. Three-Round [CLT14]. Three-Variable [NS18]. Three-Vertex-Connected [ET14]. Threshold [CCD07, SUZ13, WD20]. Thresholds [GP15]. Throttle [FK06]. Thue [DSS15, Ram05]. Tight [AF20, AHL^{+}13, BE19, HJP^{+}13, PL23, PZX07, YS13]. Tighter [FKV06]. Tightness [CD09]. Tile [JK14a, JK14b, SW17]. Tiled [Leo03]. Tiling [Gia11, Mar08a, PM13]. Tilings [Mar08b]. Time [AAV00, ANDZM09, BCFR07, Cai94, CD06, CM12, CCI12, CZ11, CFPR03, DPR07, DFLL02, EH12, FZAM08, FZCFB08, Fle96, FMN06, Fri10, GKRS10, GO09, GV03, Gol14, Gra90, HH24, HK19, HG11, IR14, IZN99, JWB03, Joh00, KM18, Kör03, KD99, Kri97, Lag17, LD01, Leo03, Leu04, LLQ06, LCY12, LWW00, MM97, Mas04, MHT09, MTNN99, MV11, Mod21, MGCVdlP20, Nak04, NTSH06, Pal03, Pet11, PY04, RLWW96, SK01, ST99, Sun11, ÜS02, WG17, Wan04, XWY^{+}22, YS13]. Time-Bounded [Pet11]. Time-Critical [Sun11]. Time-Free [CD06]. Time-Interval [NTSH06]. Time-Shuffling [EH12]. Time-Space [KM18]. Timed [AEMY21, ACFE09, Kri92, NTSH06]. Times [Li12b, SSS09]. Tissue [AFO06, ARV07, AFIV22, CVPV08, FOP05, NSVA12]. Tissue-Like [CVPV08]. TLC [Hen02]. Token [DG98, GS12a, PT14]. Tokens [DSS08, SK01]. Tolerance [FWZ15, GZY24, HY97, KR97, LYH^{+}15, LZGF16, ZZN23]. Tolerant [CHYT14, FZEBB05, LPC11, XS11, XZY19, XZZY19, XZW^{+}21, YZZ22]. Tool [HPV99]. Top [KM23, LW93, LT24, MSV23]. Top-Down [KM23, LW93, LT24, MSV23]. Topic [LKM02]. Topic-Specific [LKM02]. Topics [GPPJR13]. Topological [CC98, FS98, KM22, Kop21]. Topologically [HCG96]. Topology [FH11, Hei97, KG11, Oka98]. Tori [FHL07, LLY13, Sib97]. Torus [BF07, Che22b, ISAZ08, LYG17, Mar97, Par23b]. Torus-Like [Par23b]. Toruses [GLP07]. Total [ALR04, DFLL02, FIO08, IZN99, KS10, LLQ06, LWYL14, PY04, SR21, Smi95]. Totally [FGV99, WNF20]. Tour [BEMR11]. TPR [IML04]. Trace [BR08, Gol90, KM19, Pen93]. Traceability [HCETPL^{+}12]. Traces [LWJ^{+}10]. Track [YBI11]. Tractable [BCR11, HL06, YHK14]. Trade [Kap05, KM18, KKP97, Kut05]. Trade-off [KM18]. Trade-Offs [Kap05, KKP97, Kut05]. Trading [XWL^{+}22]. Traffic [DEKZ11]. Trains [PPJR06]. Trajectories [DKSS11, DS05, KKS05b]. Transactional [SK01]. Transcription [AES18]. Transcriptome [TFF18]. Transcriptomics [AS18]. Transducers [AM03, AM09, ARS11, AMR11, AMR15, BR20, BBL^{+}12, BBK17, BM23, CGH05, DJR18, FSM11, Gaz06, Iba15, KMRY20, KMŠ21, LLS21, Mal05, Mal15, Mal24, Man15, MSV23, Moh02, Moh13, RT16]. Transduction [BCC^{+}11]. Transductions [BvdB18, Sut14]. Transfer [HLY^{+}04]. Transfers [NN93]. Transfinite [DN07]. Transform [KSM22]. Transformation [ALR04, AT15, AT23, BTK13, BTO17, TSS13, TFS19]. Transformations [KLS05, MRS97, PT19, RKRR02]. Transient [BLY12, YBM11, YB06]. Transients [GB03]. Transition [Muk92, Tam08]. Transitions [CTS18, ZYLW12]. Transitive [DI02]. Transitivity [JP06]. Translation [Mal18]. Transmission [JS97]. Transparent [GD98, YSD16]. Transportation [Asa23]. Transporter [SS07b]. Transposition [LLL22]. Transpositions [CL07a, XZY19]. Traveling [BL01]. Trawling [DEKZ11]. Tree [AHK07, ABH^{+}09, AA20, AMZ20, BBB^{+}18, BBE24, BM23, BB04, BCHK09, BKW02, CDPT16, CCP18, Che22a, CS00a, CHZ06, DL12, DST10, ÉM11, FGS^{+}90, FTT10, Fle96, FSM11, Fuj17, FV24, Gaz06, GV23, Géc07, GC18, HH11, HBIT08, JM13, KM90, KM18, KEH16, KLH16, KK90, LL20, Li00a, LZ12, LJA09, LLS21, LT24, Lüc18, MO94, Mal05, MT10, Mal15, Mal18, Mal24, MG20, Man15, MSV23, MC02, MMSV23, MS18, MOSZ18, MP91, Pau24, PR00, PAS08, RAB15, Rei07, RVT06, SK20, SMS90, SB17, SVF09, Tei17, Tor13, XS06, YHK14, ŽM11, DDHL11]. Tree-Based [BBE24, ŽM11]. Tree-Height [Rei07]. Tree-to-Tree [Mal18].

Tree-to-Word [LLS21]. Tree-Width [Fuj17]. Trees [BYP95, BCV23, CS96, Dar13, DOR06, ERW04, FDFZB12, FA06, GI19, GRI24, Gre96, GKS^{+}19, GWF^{+}24, HL01, IML04, IZN99, IZN05, JL01, JS03, JK07, Lag17, LW93, LF96, LLL22, MM17, MTNN99, MAN05, MPS24, OSZ92, OM96, OW92, PI95, PV98, PL06, Pro96, RS01, Sao92, Smy12, XHLF02, YTN01, YZY^{+}18, YCTW10, ZB00, ZB02, ZH06]. Treewidth [AMT20, Klo96a]. Trellis [FGS^{+}90]. Trémaux [DOR06]. tri [NS13]. Triangle [FP04, SRN^{+}20, XHLF02]. Triangles [AAV00, MB17, Sib97]. Triangulating [AFB96]. Triangulation [DPT02]. Triangulations [Fre02]. Trick [Ste11]. Trie [AČ11, PPR18]. Tries [KPS93]. Trinomial [ZZC15]. Trinomials [WXF16]. Triple [DÉK22, JS97, LOZ98, LCXS19, YZP21]. Triple-Cycle [LCXS19]. Triple-Pair [DÉK22]. Trivalent [CP99]. Trivial [BL14]. Truck [MXY^{+}04]. TSP [Gol14]. Tube [AKM^{+}11]. Tumor [RHN^{+}22]. Tunable [BBM^{+}12]. Turing [AD12, Cap96, Dub95, HIIW01, HJV93, IIT91, IIK^{+}04, Mer08, Slo95]. Turn [AK14]. Tutte [GO09]. TVDH [AKM^{+}11]. Twisted [HYLF20, ZLL20]. Two [AF20, AGM14, Ars15, ACDL18, BR20, BHK18a, BSBZ08, BT00, BKW02, CH15, CL15, CdL04, CHZ06, CGKY11, CGKY12, CTS18, DLT06, DJ12, Fin21, FS05, FHKK23, FL12, GP15, HKV17, HH24, HJP^{+}13, HL06, HKKŠ13, HG11, IJT^{+}93, IS12, JP06, JM03, Kap05, KYZS17, KKH90, KP10b, Klo96b, KL11, KMO10, LY94, Leu04, LLZ07, LCXS19, MS20, MP22, MMK22, Mel93, MSMR22, NR18, OS01, Prů17, RWZ01, RLWW96, SS07b, Ste93, SMAN13, WO03, XZS16, XZL^{+}19, YZY^{+}18, ZZZ16, ZQL12, ZG13]. Two-Dimensional [AGM14, BT00, CdL04, DJ12, JP06, MS20, MP22, NR18, Prů17, SMAN13]. Two-Face [RLWW96]. Two-Hop [AF20].

Two-Machine [LLZ07, SS07b]. Two-Pattern [FS05]. Two-Prime [KYZS17]. Two-Processor [Leu04]. Two-Pushdown [KMO10]. Two-Sided [ACDL18]. Two-Way [BR20, BHK18a, BKW02, CL15, FHKK23, HKKŠ13, IJT^{+}93, IS12, Kap05, KL11, ZQL12]. Type [Bar90, BYİT21, CZTH13, Hir91, Kam95, MM17, MN00, PB20, PI95, Smi95, Tsu01, TST01b]. Type-Free [Kam95]. Typeness [KMM06]. Types [APP91, GJKS18, TZ91].

U2 [YKCW23]. UGB [NS98]. Ultralinear [MP07]. Un-Kleene [HSS07]. Un-Oriented [DSS08]. Unambiguity [CL15, Pau24]. Unambiguous [CFM13, FRS06, JJŠ18, Mal24, MS18, Mor10, Rav08]. Unary [AK10, BCN12, Das19, DESW05, GP13, Glö07, Jeż08, JMR91, KMW14a, MvZ22, PS02, Pig09, Pig15]. Unbordered [HN10]. Unbounded [Car11, HB06]. Unbreakable [OS93]. Uncertainty [PR23]. Unconditionally [SNJ11]. unconventional [CV13]. Uncountable [DY19]. Undecidabilities [BKM15]. Undecidability [BKM11, Fin12, HHH07]. Undecidable [Kog18, Mar92, Mar08a]. Understanding [Zet11]. Undirected [FZFDCHB05, XLC^{+}04, ZWS96]. Unfold [RKRR02]. Unfold/Fold [RKRR02]. Unforgeable [HHP17]. Unicast [GWL^{+}17]. Unidirectional [GS12a]. Unification [GJV00b, Pym92]. Unified [BNBN20, CLMP16, NS13]. Uniform [AS18, Anc02, BFMBS11, BMW91, Dur13, FHKK23, Fuj17, HL01, KNR21, KSV00, LPP92, MM17, XC15, ZZC15]. Uniformity [CdL04]. Uniformly [NP09]. Union [CGKY11, EHS15, GI19, GNC^{+}03, HS08, JM11, Nag21]. Union-Free [JM11, Nag21]. Union-Freeness [Nag21]. Unique [DD08, Ruo96]. Uniqueness [DESW05]. Unit [Fuj16, FCS05, Zaj09]. Unitary [HN06, IKPY21]. Universal [AKM^{+}11, ARV12, BKST18, CL14, DG90, GKSZ19, Lis93, Pol05, Sch02, Ver09]. Universality [Bur12b, CP06, JK14a, PB20]. Universally [Tra02]. Universe [MAG09]. Unknown [LP11]. Unknowns [CK07]. Unordered [FA06, YHK14]. Unpaired [LMZC20, LW21]. Unranking [ERW04]. Unrelated [GJKS18, Jan93]. Unreliable [KY90]. Unsolvability [BHK05]. Unweighted [MQ12]. Update [Fle96, GPC09, Lag17, LOD07a, LOD07b]. Update-Efficient [LOD07a, LOD07b]. Updating [LW93, OW92]. Upload [AF20]. Upper [BBP11, PL23, ZSW14, ZG13]. Ups [JJS08]. Upward [HL06]. Use [BCC^{+}11, SS12b, LMG20]. Used [LKM02]. Useful [BGRY16]. Usefulness [BPR09]. User [DE08, KMRY20]. User-Defined [KMRY20]. Using [AES18, AČ11, AS18, AH07, BBFZM06, BNBN20, BS01, Bee95, BC12, CTZ01, CST^{+}17, CK08b, DW04, DSS08, DZ00, DE08, EP17, FGH^{+}07, FHL07, FK13, FNI16, GD98, HHH07, HV02, HP09b, HFLD09, Ibr22, IML04, ILT11, INY07, IN08, IN10, Jai95, KAPF05, KS10, LX94, LB04, LWJ^{+}10, LWW22, MO94, MGCVdlP20, PB20, PAS08, Pol05, PS22, RN22, RCTC^{+}09, SA22, SKL03, SB01, SN13, Wan14, WXF16, WM13, WHLH17, XHLF02, YBI11, ZL22]. Usual [ES01].

V4 [YW22]. Valid [HCG96]. Valuation [DM11]. Value [KMIS09]. Valued [CH15, SH17]. Valuedness [Iba15]. Values [BFL02]. VANETs [CST^{+}17]. Var [YTLC02]. Variable [BGM^{+}18, CL07b, DDD18, GW24, NS18, TY03, TJZ13, ZWCL14]. Variable-Order [BGM^{+}18]. Variables [EAB^{+}16, Kam98, ZG13]. Variant [Pău00, VG01]. Variants [CVDV10, FL09, FRV19, HLC^{+}19, JL01, MS16a, MS16b, NB18]. Variations [DRDN08, YHK14]. Varieties [KP10b]. Various [BLM15, IYD05]. Varying [HG11]. Vector [BH02, CHYT14, SYS19]. Vectorial [Car11, DQFL12, FY11]. Vectors [PL06]. Vehicular [PS22]. Verbal [ZSG^{+}22]. Verifiability [YMC^{+}17]. Verifiable [SZQ^{+}17]. Verification [ADHR09, ADR11, AEMY21, BB03a, BDSV06, DPR07, FK06, FK13, Iba02, ILT11, LD01, LN08, LWJ^{+}10, MG14, MDAPHPJ^{+}11, Pen93, WM13, YBI11]. Verified [DVG03]. Verifier [Ver09]. Verifiers [YSD16]. Verifying [FGH^{+}07, HV02, MvZ22]. Version [CMWZ19, Faa19, Jun14]. Versions [BSBZ08]. Versus [COT12, DPS97, CV13]. Vertex [AT11, ET14, Far20, FP04, HW17, Kan15, MPS24, PRS98, RZ12, SS99, WQY16, ZLG21]. Vertex-Connectivity [FP04]. Vertex-Neighbor-Scattering [WQY16]. Vertices [DW04, GWL^{+}17, RR18]. Very [FPPS03, FGH^{+}07]. Vesicles [AFIV22]. VH [ZWC^{+}22]. VH-CATT [ZWC^{+}22]. Via [BCDP08, Kar09, KL05, LL23, LN08, YLZ14, BLS20, Kog21, Zan91]. Victor [CVM20]. Video [AF20, HT09]. Video-On-Demand [HT09]. View [Ami05, DD12]. Viewed [Wil91]. Viral [DM05]. Virtual [BCC^{+}11, GNC^{+}03, LJA09]. Visibly [RT16]. Visitors [ECY02]. VLSI [LL23]. Volume [Ano97, Ano98, Ano01a, Ano02, Ano03a, Ano04a, Ano05a, Ano06, Ano07, Ano08, Ano09, Ano11, Ano12, Ano13, Ano14, Ano15, Ano16, Ano17, Ano18, Ano19, Ano20, Ano21, Ano22, SS24]. Volumes [BCC^{+}11]. Vs [SR00a, HKKŠ13]. vs. [DTY15]. VTLoE [MT95a]. Vulnerabilities [DW04]. Vulnerability [AT11, AT15, AT23, BY18, BA24].

Walk [BKS12, Li12a]. Walking [DPT02]. Walks [NR18, Sub05]. Walsh [CH15, KSM22, SH17]. Walsh-Transform [KSM22]. Watson [KM08, MMR20]. Way [AM09, BR20, BH20, BMP15, BHK18a, BKW02, CL15, CFY16, FHKK23, HIR^{+}92, HKKŠ13, IJT^{+}93, IS12, Kap05, KL11, KMW14b, KMW14a, Obt01, POM22, Slo95, ZQL12]. WDM [LC18, XLC^{+}04]. Weak [AA20, ACMP20, Asl16, BSOR10, DTY15, GV03, HLC^{+}19, KR08]. Weak-Bisplit [GV03]. Weak-Rupture [Asl16]. Weakly [AWF03, DWS15, GLPP22]. Weakly-Connected [AWF03]. Web [ECY02, HM04, NH02, Zho02]. Weber [BCDM23]. Wedderburn [AR16]. Weibull [PNN^{+}10]. Weight [CS00a, FPPS03, LW93, WF21]. Weight-Balanced [LW93]. Weighted [AMR05, AM09, AJMO11, BLP18, BM23, CL15, CLOZ04, CGKN08, DM11, DP14, DÉK22, DFK23, DGR24, ÉM11, FV24, GVL07, HHNP23, HI18, IMP12, JC03, KS10, LLQ06, Mal05, MQ12, Mal15, Mal24, Moh02, Moh03, Oli13, PYTH10, SS07a, SVF09, Tei17, ZHZ11]. Weights [HN06, KR16]. Welch [SSF20]. Well [Hut02, RT16, ZH06]. Well-Defined [Hut02]. Well-Nested [RT16]. Well-Orderly [ZH06]. Wheel [AB17b, BYİT21]. Wheel-Type [BYİT21]. Wheels [AO11]. Where [WCD^{+}14]. Whether [CDJ09, DK11]. Which [ERW04]. While [GPC09]. Whiteman [KK19]. Whose [Rya21]. Wide [NZT^{+}24]. Width [BLM04, DL12, Fuj17, GR00, JYF91, LV08, MR99, PR00, RVT06]. Wildcard [DES09]. Wildcards [GS18, Zha17]. Wilf [BSOR10, KPS13]. Windmill [GRI24]. Window [KO13]. Winning [Fia08]. Wirelength [GWF^{+}24]. Wireless [AWF03, AHL^{+}13, BNS03, BDDN01, CYS^{+}12, DCS13, FPPS03, Li12a, MG14, MKB^{+}11, SP04, TH22, WLF03, WP08]. Within [AE04]. Without [CC05, FZT14, GI19, GKS17, KLP20, Lug11]. Word [BBP11, CR15, CHKL07, CK07, DP14, DSS15, GRS21, Hon12, IYZ04, JF18, LP11, LLS21, MNS18, Mar92, NS18, dBDZ19]. Words [APV06, Ata07, BPR09, BMS18, BCMS20, BDD^{+}18, BC06, BSBZ08, BSOR10, BS12, BSCH22, BMR^{+}14, BHNR04, BJ05, BJ06, BJ07b, CdL04, CRSZ11, CK16, Čer08, CNT22, CGL12, DGK08, DM12, DR12, DSS15, DMSS16, EKKS18, ÉO13, FW24, Fin19, FS21, FLM^{+}21, FJPS16, GPC09, GS18, GHS13, GRRS14, GS12b, Gus13, HH12, HHNP23, Har24, HK23, HN10, IYZ04, JSKM20, JP04, KM07a, KM08, KMS11, Löd15, MP22, MMR10, MPV04, MDGH13, Mel93, Mig90, PS12a, PS12b, PR12, Pri06, Ram05, RS04, Ric19, Sal11, Sha04, ST16, Teh15, Teh16b]. Work [BBM^{+}12]. Worker [DPR^{+}08]. Workflows [LBL06]. Working [Elm06, Fre05, PLMZ11]. Working-Set [Elm06]. Worksharing [RC11]. Workstations [Ros00]. Wormhole [PV98]. Worst [Fle96, Lag17, PSA17, YH11, ZSW14]. Worst-Case [Fle96, Lag17, YH11, ZSW14]. Wreath [BK16]. WWW [LKM02].

Yao [GKS17]. Yen [AH11]. Yield [ER14]. YOLO [YW22]. Yu [CISS12, SSS13].

Zero [KMŠ21, LL16, WCD^{+}14]. Zero-Avoiding [KMŠ21]. Zero-Difference [LL16]. Zero-Knowledge [WCD^{+}14]. Zoltán [Fül17]. Zonal [CL03]. Zone [WPZ16]. Zoom [ER14].

References

Aytac:2013:SRR

[AA13] Aysun Aytaç and Hanife Aksu. Some results for the rupture degree. International Journal of Foundations of Computer Science (IJFCS), 24(8):1329–??, December 2013. CODEN IFC-SEN. ISSN 0129-0541.

Ayta:2019:EDC

[AA19] Aysun Aytaç and Betül Atay Atakul. Exponential domination critical and stability in some graphs. International Journal of Foundations of Computer Science (IJFCS), 30(5):781– 791, August 2019. ISSN 0129-0541. URL https:

                                                       //www.worldscientific.
                                                       com/doi/10.1142/S0129054119500217.

Alirezazadeh:2020:WSP

[AA20] Saeid Alirezazadeh and Khadijeh Alibabaei. Weak separation problem for tree languages. International Journal of Foundations of Computer Science (IJFCS), 31(05):583– 593, August 2020. ISSN 0129-0541. URL https:

                                                       //www.worldscientific.
                                                       com/doi/10.1142/S0129054120500276.

Ahn:2009:COH

[AAA^{+}09] Hee-Kap Ahn, Helmut Alt, Tetsuo Asano, Sang Won Bae, Peter Brass, Otfried Cheong, Christian Knauer, Hyeon-Suk Na, Chan-Su Shin, and Alexander Wolff. Constructing optimal highways. International Journal of Foundations of Computer Science (IJFCS), 20(1):3– 23, February 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Adamczyk:2018:PDG

[AACR18] Michałl Adamczyk, Mai Alzamel, Panagiotis Charalampopoulos, and Jakub Radoszewski. Palindromic decompositions with gaps and errors. International Journal of Foundations of Computer Science (IJFCS), 29(8):1311–1329, December 2018. ISSN 0129-0541.

Aichholzer:2002:FPS

[AAH02] O. Aichholzer, L. S. Alboul, and F. Hurtado. On flips in polyhedral surfaces. International Journal of Foundations of Computer Science (IJFCS), 13(2):303–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Alamro:2020:EIC

[AAI^{+}20] Hayam Alamro, Mai Alzamel, Costas S. Iliopoulos, Solon P. Pissis, Wing-Kin Sung, and Steven Watts. Efficient identification of k-closed strings. International Journal of Foundations of Computer Science (IJFCS), 31(05):595– 610, August 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500288.

Accornero:2000:AST

[AAV00] A. Accornero, M. Ancona, and S. Varini. All separating triangles in a plane graph can be optimally “broken” in polynomial time. International Journal of Foundations of Computer Science (IJFCS), 11(3):405– 422, 2000. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Arvind:1991:EDG

[AB91] V. Arvind and S. Biswas. Edge-deletion graph problems with first-order expressible subgraph properties. International Journal of Foundations of Computer Science (IJFCS), 2(2):83– 100, June 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Aytac:2017:RRC

[AB17a] Aysun Aytaç and Zeynep Nihan Odabaş Berberler. Robustness of regular caterpillars. International Journal of Foundations of Computer Science (IJFCS), 28 (7):835–??, November 2017. CODEN IFCSEN. ISSN 0129-0541.

Aytac:2017:BNW

[AB17b] Vecdi Aytaç and Zeynep Nihan Berberler. Binding number and wheel related graphs. International Journal of Foundations of Computer Science (IJFCS), 28 (1):29–38, January 2017.

CODEN IFCSEN. ISSN 0129-0541.

Ausiello:2005:CDA

[ABDP05] Giorgio Ausiello, Cristina Bazgan, Marc Demange, and Vangelis Th. Paschos. Completeness in differential approximation classes. International Journal of Foundations of Computer Science (IJFCS), 16(6):1267– ??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Abdulla:2009:CBT

[ABH^{+}09] Parosh Aziz Abdulla, Ahmed Bouajjani, Lukáš Holík, Lisa Kaati, and Tomáš Vojnar. Composed bisimulation for tree automata. International Journal of Foundations of Computer Science (IJFCS), 20(4):685– 700, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Atig:2017:EOM

[ABH17] Mohamed Faouzi Atig, Benedikt Bollig, and Peter Habermehl. Emptiness of ordered multipushdown automata is 2ETIME-complete. International Journal of Foundations of Computer Science (IJFCS), 28(8):945–??, December 2017. CODEN IFC-SEN. ISSN 0129-0541.

Anceaume:2011:DEC

[ABL^{+}11]  Emmanuelle  Anceaume,

Francisco Brasileiro, Romaric Ludinard, Bruno Sericola, and Frédéric Tronel. Dependability evaluation of cluster-based distributed systems. International Journal of Foundations of Computer Science (IJFCS), 22 (5):1123–1142, August 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Altundag:2016:NRD

[ABT16] Ferhan Nihan Altundag and Goksen Bacak-Turan. Neighbor rupture degree of Harary graphs. International Journal of Foundations of Computer Science (IJFCS), 27(6):739–??, September 2016. CODEN IFCSEN. ISSN 0129-0541.

Alba:2005:BPG

[AC05] E. Alba and F. Chicano. On the behavior of parallel genetic algorithms for optimal placement of antennae in telecommunications. International Journal of Foundations of Computer Science (IJFCS), 16(2):343– ??, April 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Alazemi:2011:CSU

[AČ11] Hamed M. K. Alazemi and Anton Černý. Counting subwords using a trie automaton. International Journal of Foundations of Computer Science (IJFCS), 22 (6):1457–1469, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Aydinian:2018:CMT

[ACDL18] Harout Aydinian, Ferdinando Cicalese, Christian Deppe, and Vladimir Lebedev. A combinatorial model of two-sided search. International Journal of Foundations of Computer Science (IJFCS), 29(4):481– 504, June 2018. CODEN IFCSEN. ISSN 0129-0541.

Andre:2009:IMP

[ACFE09] Étienne André, Thomas Chatain, Laurent Fribourg, and Emmanuelle Encrenaz. An inverse method for parametric timed automata. International Journal of Foundations of Computer Science (IJFCS), 20(5):819– 836, October 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Allauzen:2011:DCD

[ACM11] Cyril Allauzen, Corinna Cortes, and Mehryar Mohri. A dual coordinate descent algorithm for SVMs combined with rational kernels. International Journal of Foundations of Computer Science (IJFCS), 22 (8):1761–1779, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Almagor:2020:WCR

[ACMP20] Shaull Almagor, Michaël Cadilhac, Filip Mazowiecki, and Guillermo A. Pérez. Weak cost register automata are still powerful. International Journal of Foundations of Computer Science (IJFCS), 31(06):689– 709, September 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120410026.

Abdulla:2013:MAP

[ACV13] Parosh Aziz Abdulla, Jonathan Cederberg, and Tomáš Vojnar. Monotonic abstraction for programs with multiplylinked structures. International Journal of Foundations of Computer Science (IJFCS), 24(2):187–??, February 2013. CODEN IFCSEN. ISSN 0129-0541.

Arrighi:2012:PCT

[AD12] Pablo Arrighi and Gilles Dowek. The physical Church–Turing thesis and the principles of quantum theory. International Journal of Foundations of Computer Science (IJFCS), 23 (5):1131–??, August 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ada:2010:NDC

[Ada10] Anil Ada. On the nondeterministic communication complexity of regular languages. International Journal of Foundations of Computer Science (IJFCS), 21(4):479–493, August 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Angelini:2018:ABD

[ADD^{+}18] Patrizio Angelini, Giordano Da Lozzo, Marco Di Bartolomeo, Valentino Di Donato, Maurizio Patrignani, Vincenzo Roselli, and Ioannis G. Tollis. Algorithms and bounds for L-drawings of directed graphs. International Journal of Foundations of Computer Science (IJFCS), 29(4):461– 480, June 2018. CODEN IFCSEN. ISSN 0129-0541.

Abdulla:2009:MAE

[ADHR09] Parosh Aziz Abdulla, Giorgio Delzanno, Noomene Ben Henda, and Ahmed Rezine. Monotonic abstraction: on efficient verification of parameterized systems. International Journal of Foundations of Computer Science (IJFCS), 20(5):779– 801, October 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Abdulla:2011:AVD

[ADR11] Parosh Aziz Abdulla, Giorgio Delzanno, and Ahmed Rezine. Automatic verification of directory-based consistency protocols with graph constraints. International Journal of Foundations of Computer Science (IJFCS), 22(4):761– 782, June 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Azizoglu:1999:IND

[AE99] M. C. Azizoglu and Ö. Eğecioğlu. The isoperimetric number of d-dimensional k-ary arrays. International Journal of Foundations of Computer Science (IJFCS), 10 (3):289–300, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Arslan:2002:AAL

[AE02] Abdullah N. Arslan and Ömer Eğecioğlu. Approximation algorithms for local alignment with length constraints. International Journal of Foundations of Computer Science (IJFCS), 13 (5):751–??, October 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Arslan:2004:DLW

[AE04] Abdullah N. Arslan and Ömer Eğecioğlu. Dictionary look-up within small edit distance. International Journal of Foundations of Computer Science (IJFCS), 15(1):57–??, February 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Arslan:2005:ACL

[AE05] Abdullah N. Arslan and Ömer Eğecioğlu. Algorithms for the constrained longest common subsequence problems. International Journal of Foundations of Computer Science (IJFCS), 16(6):1099– ??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ammar:2021:TBV

[AEMY21] Ikhlass Ammar, Yamen El Touati, John Mullins, and Moez Yeddes. Timed bounded verification of inclusion based on timed bounded discretized language. International Journal of Foundations of Computer Science (IJFCS), 32(02):175–202, February 2021. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054121500106.

Abdollahyan:2018:IPR

[AES18] Maryam Abdollahyan, Greg Elgar, and Fabrizio Smeraldi. Identifying potential regulatory elements by transcription factor binding site alignment using partial order graphs. International Journal of Foundations of Computer Science (IJFCS), 29(8):1345–1354, December 2018. ISSN 0129-0541.

Alba:2005:GEN

[AETZ05] Enrique Alba, Fikret Ercal, El-Ghazali Talbi, and Albert Y. Zomaya. Guest editorial: Nature-inspired distributed computing. International Journal of Foundations of Computer Science (IJFCS), 16(2):239– ??, April 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ando:2020:TBU

[AF20] Hironori Ando and Satoshi Fujita. Tight bounds on the upload capacity to enable two-hop delivery in peer-to-peer video streaming systems. International Journal of Foundations of Computer Science (IJFCS), 31(03): 341–354, April 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500124.

Agarwala:1996:SAP

[AFB96] R. Agarwala and D. Fernandez-Baca. Simple-algorithms for perfect phylogeny and triangulating colored graphs. International Journal of Foundations of Computer Science (IJFCS), 7(1):11–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Alhazov:2022:TPS

[AFIV22] Artiom Alhazov, Rudolf Freund, Sergiu Ivanov, and Sergey Verlan. Tissue P systems with vesicles of multisets. International Journal of Foundations of Computer Science (IJFCS), 33(3–4):179–202, April–June 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122410015.

Alhazov:2006:CSC

[AFO06] Artiom Alhazov, Rudolf Freund, and Marion Oswald. Cell/symbol complexity of tissue P systems with symport/antiport rules. International Journal of Foundations of Computer Science (IJFCS), 17(1):3–??, February 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Abramov:2001:SNS

[AG01] Sergei Abramov and Robert Glück. From standard to non-standard semantics by semantics modifiers. International Journal of Foundations of Computer Science (IJFCS), 12(2):171–211 (or 171–212??), 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anselmo:2014:PPC

[AGM14] Marcella Anselmo, Dora Giammarresi, and Maria Madonia. Prefix picture codes: a decidable class of two-dimensional codes. International Journal of Foundations of Computer Science (IJFCS), 25(8):1017– ??, December 2014. CODEN IFCSEN. ISSN 0129-0541.

Anselmo:2019:SPA

[AGM19] Marcella Anselmo, Dora Giammarresi, and Maria Madonia. Sets of pictures avoiding overlaps. International Journal of Foundations of Computer Science (IJFCS), 30(6– 7):875–898, September– November 2019. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054119400215.

Aziz:2007:MAP

[AH07] Benjamin Aziz and Geoff Hamilton. Modelling and analysis of PKI-based systems using process calculi. International Journal of Foundations of Computer Science (IJFCS), 18(3):593– 618, June 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Atig:2011:YPL

[AH11] Mohamed Faouzi Atig and Peter Habermehl. On Yen’s path logic for Petri nets. International Journal of Foundations of Computer Science (IJFCS), 22(4):783– 799, June 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Abdulla:2007:BMT

[AHK07] Parosh Aziz Abdulla, Johanna Högberg, and Lisa Kaati. Bisimulation minimization of tree automata. International Journal of Foundations of Computer Science (IJFCS), 18(4):699– 713, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Axelsen:2017:DID

[AHK17] Holger Bock Axelsen, Markus Holzer, and Martin Kutrib. The degree of irreversibility in deterministic finite automata. International Journal of Foundations of Computer Science (IJFCS), 28 (5):503–??, August 2017. CODEN IFCSEN. ISSN 0129-0541.

Augustine:2013:TAS

[AHL^{+}13] John Augustine, Qi Han, Philip Loden, Sachin Lodha, and Sasanka Roy. Tight analysis of shortest path convergecast in wireless sensor networks. International Journal of Foundations of Computer Science (IJFCS), 24(1):31–??, January 2013. CODEN IFCSEN. ISSN 0129-0541.

Anceaume:2002:NDI

[AHR02]  Emmanuelle  Anceaume,

Jean-Michel Helary, and Michel Raynal. A note on the determination of the immediate predecessors in a distributed computation. International Journal of Foundations of Computer Science (IJFCS), 13(6):865– ??, December 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Alatabbi:2016:ALC

[AILR16] Ali Alatabbi, Costas S. Iliopoulos, Alessio Langiu, and M. Sohel Rahman. Algorithms for longest common Abelian factors. International Journal of Foundations of Computer Science (IJFCS), 27(5):529–??, August 2016. CODEN IFC-SEN. ISSN 0129-0541.

Asahiro:2021:GOE

[AJM^{+}21] Yuichi Asahiro, Jesper Jansson, Eiji Miyano, Hirotaka Ono, and T. P. Sandhya. Graph orientation with edge modifications. International Journal of Foundations of Computer Science (IJFCS), 32(02):209– 233, February 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412150012X.

Asahiro:2011:GOM

[AJMO11] Yuichi Asahiro, Jesper Jansson, Eiji Miyano, and Hirotaka Ono. Graph orientation to maximize the minimum weighted outdegree. International Journal of Foundations of Computer Science (IJFCS), 22(3):583– 601, April 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Afonin:2006:MFP

[AK06] Sergey Afonin and Elena Khazova. Membership and finiteness problems for rational sets of regular languages. International Journal of Foundations of Computer Science (IJFCS), 17(3):493– ??, June 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Afonin:2010:SFG

[AK10] Sergey Afonin and Elena Khazova. On the structure of finitely generated semigroups of unary regular languages. International Journal of Foundations of Computer Science (IJFCS), 21 (5):689–704, October 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Aizikowitz:2014:LCG

[AK14] Tamar Aizikowitz and Michael Kaminski. Linear conjunctive grammars and oneturn synchronized alternating pushdown automata. International Journal of Foundations of Computer Science (IJFCS), 25(6):781–??, September 2014. CODEN IFCSEN. ISSN 0129-0541.

Almagor:2019:SCM

[AKK19] Shaull Almagor, Denis Kuperberg, and Orna Kupferman. Sensing as a complexity measure. International Journal of Foundations of Computer Science (IJFCS), 30(6–7):831– 873, September–November 2019. ISSN 0129-0541. URL

                                        https://www.worldscientific.
                                        com/doi/10.1142/S0129054119400203.

Alhazov:2011:SUT

[AKM^{+}11] Artiom Alhazov, Marian Kogler, Maurice Margenstern, Yurii Rogozhin, and Sergey Verlan. Small universal TVDH and test tube systems. International Journal of Foundations of Computer Science (IJFCS), 22(1):143– 154, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Axelsen:2020:BRP

[AKMW20] Holger Bock Axelsen, Martin Kutrib, Andreas Malcher, and Matthias Wendlandt. Boosting reversible pushdown and queue machines by preprocessing. International Journal of Foundations of Computer Science (IJFCS), 31(08):1021– 1049, December 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120420022.

Arvind:1995:HIP

[AKS95] V. Arvind, J. Koebler, and R. Schuler. On helping and interactive proof systems. International Journal of Foundations of Computer Science (IJFCS), 6 (2):137–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Atig:2014:AOM

[AKS14] Mohamed Faouzi Atig, K. Narayan Kumar, and Prakash Saivasan. Adjacent ordered multi-pushdown systems. International Journal of Foundations of Computer Science (IJFCS), 25(8):1083–??, December 2014. CODEN IFCSEN. ISSN 0129-0541.

Akutsu:2006:APS

[Aku06]  Tatsuya Akutsu.  Algo-

rithms for point set matching with k-differences. International Journal of Foundations of Computer Science (IJFCS), 17(4):903– ??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Alirezazadeh:2016:PFA

[Ali16] Saeid Alirezazadeh. On pseudovarieties of forest algebras. International Journal of Foundations of Computer Science (IJFCS), 27(8):909–942, December 2016. CODEN IFCSEN. ISSN 0129-0541.

Alberich:2004:SPT

[ALR04] Ricardo Alberich, Mercè Llabrés, and Francesc Rosselló. Single-pushout transformation of total algebras. International Journal of Foundations of Computer Science (IJFCS), 15(1):205–??, February 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Allauzen:2003:FST

[AM03] Cyril Allauzen and Mehryar Mohri. Finitely subsequential transducers. International Journal of Foundations of Computer Science (IJFCS), 14(6):983–??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Allauzen:2009:WCW

[AM09] Cyril Allauzen and Mehryar Mohri. N-way composition of weighted finite-state transducers. International Journal of Foundations of Computer Science (IJFCS), 20(4):613–627, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Amir:2005:TIS

[Ami05] Amihood Amir. Theoretical issues of searching aerial photographs: a bird’s eye view. International Journal of Foundations of Computer Science (IJFCS), 16(6): 1075–??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Asahiro:2007:GOA

[AMOZ07] Yuichi Asahiro, Eiji Miyano, Hirotaka Ono, and Kouhei Zenmyo. Graph orientation algorithms to minimize the maximum outdegree. International Journal of Foundations of Computer Science (IJFCS), 18(2):197– 215, April 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Allauzen:2005:DPA

[AMR05] Cyril Allauzen, Mehryar Mohri, and Brian Roark. The design principles and algorithms of a weighted grammar library. International Journal of Foundations of Computer Science (IJFCS), 16(3):403– ??, June 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Almeida:2008:EGM

[AMR08] Marco Almeida, Nelma Moreira, and Rogério Reis. Exact generation of minimal acyclic deterministic finite automata. International Journal of Foundations of Computer Science (IJFCS), 19(4):751–765, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Almeida:2009:AMR

[AMR09] Marco Almeida, Nelma Moreira, and Rogério Reis. Antimirov and Mosses’s rewrite system revisited. International Journal of Foundations of Computer Science (IJFCS), 20(4):669– 684, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Allauzen:2011:GAT

[AMR11] Cyril Allauzen, Mehryar Mohri, and Ashish Rastogi. General algorithms for testing the ambiguity of finite automata and the doubletape ambiguity of finitestate transducers. International Journal of Foundations of Computer Science (IJFCS), 22(4):883– 904, June 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Amorim:2015:NLF

[AMR15] Ivone Amorim, António Machiavelo, and Rogério Reis. On the number of linear finite transducers. International Journal of Foundations of Computer Science (IJFCS), 26(7):873–??, November 2015. CODEN IFCSEN. ISSN 0129-0541.

Akutsu:2020:IHM

[AMT20] Tatsuya Akutsu, Avraham A. Melkman, and Takeyuki Tamura. Improved hardness of maximum common subgraph problems on labeled graphs of bounded treewidth and bounded degree. International Journal of Foundations of Computer Science (IJFCS), 31(2):253– 273, February 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500069.

Attou:2020:BPT

[AMZ20] Samira Attou, Ludovic Mignot, and Djelloul Ziadi. The bottom-up position tree automaton and the father automaton. International Journal of Foundations of Computer Science (IJFCS), 31(08):1051– 1068, December 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120420034.

Anceaume:2002:ESU

[Anc02] Emmanuelle Anceaume. Efficient solution to uniform atomic broadcast. International Journal of Foundations of Computer Science (IJFCS), 13(5):695– ??, October 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ahrabian:2009:CTA

[ANDZM09] H. Ahrabian, A. Nowzari-Dalini, and F. Zare-Mirakabad. A constant time algorithm for DNA add. International Journal of Foundations of Computer Science (IJFCS), 20(3):549–558, June 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:1997:AIV

[Ano97] Anonymous. Author index volume 8 (1997). International Journal of Foundations of Computer Science (IJFCS), 8(4):469–??, December 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:1998:AIV

[Ano98] Anonymous. Author index volume 9 (1998). International Journal of Foundations of Computer Science (IJFCS), 9(4):455–??, December 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:1999:AI

[Ano99] Anonymous. Author index. International Journal of Foundations of Computer Science (IJFCS), 10 (4):545–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2000:AI

[Ano00] Anonymous. Author index. International Journal of Foundations of Computer Science (IJFCS), 11 (4):651–??, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2001:AIV

[Ano01a] Anonymous. Author index volume 12 (2001). International Journal of Foundations of Computer Science (IJFCS), 12(6):849–??, December 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2001:P

[Ano01b] Anonymous. Preface. International Journal of Foundations of Computer Science (IJFCS), 12(1):??, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2001:SIF

[Ano01c] Anonymous. Special issue on functional and logic programming — part 1. International Journal of Foundations of Computer Science (IJFCS), 12(1):1–??, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2002:AIV

[Ano02] Anonymous. Author index volume 13 (2002). International Journal of Foundations of Computer Science (IJFCS), 13(6):937–??, December 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2003:AIV

[Ano03a] Anonymous. Author index volume 14 (2003). International Journal of Foundations of Computer Science (IJFCS), 14(6):1183– ??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2003:Pa

[Ano03b] Anonymous. Preface. International Journal of Foundations of Computer Science (IJFCS), 14(2):167– ??, April 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2003:Pb

[Ano03c] Anonymous. Preface. International Journal of Foundations of Computer Science (IJFCS), 14(3):335– ??, June 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2003:Pc

[Ano03d] Anonymous. Preface. International Journal of Foundations of Computer Science (IJFCS), 14(4):525– ??, August 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2003:Pd

[Ano03e] Anonymous. Preface. International Journal of Foundations of Computer Science (IJFCS), 14(5):711– ??, October 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2004:AIV

[Ano04a] Anonymous. Author index volume 15 (2004). International Journal of Foundations of Computer Science (IJFCS), 15(6):911–??, December 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2004:P

[Ano04b] Anonymous. Preface. International Journal of Foundations of Computer Science (IJFCS), 15(2):223– ??, April 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2005:AIV

[Ano05a] Anonymous. Author index volume 16 (2005). International Journal of Foundations of Computer Science (IJFCS), 16(6):1309– ??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2005:P

[Ano05b] Anonymous. Preface. International Journal of Foundations of Computer Science (IJFCS), 16(5):829– ??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2006:AIV

[Ano06] Anonymous. Author index volume 17 (2006). International Journal of Foundations of Computer Science (IJFCS), 17(6):1485–1490, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2007:AIV

[Ano07] Anonymous. Author index volume 18 (2007). International Journal of Foundations of Computer Science (IJFCS), 18(6):1453–1459, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2008:AIV

[Ano08] Anonymous. Author index volume 19 (2008). International Journal of Foundations of Computer Science (IJFCS), 19(6):1479–1485, December 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2009:AIV

[Ano09] Anonymous. Author index volume 20 (2009). International Journal of Foundations of Computer Science (IJFCS), 20(6):1161–1166, December 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2011:AIV

[Ano11] Anonymous. Author index volume 22 (2011). International Journal of Foundations of Computer Science (IJFCS), 22(8):1995–2004, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Anonymous:2012:AIV

[Ano12] Anonymous. Author index volume 23 (2012). International Journal of Foundations of Computer Science (IJFCS), 23(8):1767–??, December 2012. CODEN IFC-SEN. ISSN 0129-0541.

Anonymous:2013:AIV

[Ano13] Anonymous. Author index volume 24 (2013). International Journal of Foundations of Computer Science (IJFCS), 24(8):1355–??, December 2013. CODEN IFC-SEN. ISSN 0129-0541.

Anonymous:2014:AIV

[Ano14] Anonymous. Author index: Volume 25 (2014). International Journal of Foundations of Computer Science (IJFCS), 25(8):1177–??, December 2014. CODEN IFC-SEN. ISSN 0129-0541.

Anonymous:2015:AIV

[Ano15] Anonymous. Author index volume 26 (2015). International Journal of Foundations of Computer Science (IJFCS), 26(8):1191–??, December 2015. CODEN IFC-SEN. ISSN 0129-0541.

Anonymous:2016:AIV

[Ano16] Anonymous. Author index volume 27 (2016). International Journal of Foundations of Computer Science (IJFCS), 27(8):1005–??, December 2016. CODEN IFC-SEN. ISSN 0129-0541.

Anonymous:2017:AIV

[Ano17] Anonymous. Author index volume 28 (2017). International Journal of Foundations of Computer Science (IJFCS), 28(8):1073–??, December 2017. CODEN IFC-SEN. ISSN 0129-0541.

Anonymous:2018:AIV

[Ano18] Anonymous. Author index volume 29 (2018). International Journal of Foundations of Computer Science (IJFCS), 29(8):1373– 1379, December 2018. ISSN 0129-0541.

Anonymous:2019:AIV

[Ano19] Anonymous. Author index volume 30 (2019). International Journal of Foundations of Computer Science (IJFCS), 30(8):1381– 1387, December 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119990016.

Anonymous:2020:AIV

[Ano20] Anonymous. Author index volume 31 (2020). International Journal of Foundations of Computer Science (IJFCS), 31(08):1199– 1204, December 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120990014.

Anonymous:2021:AIV

[Ano21] Anonymous. Author index volume 32 (2021). International Journal of Foundations of Computer Science (IJFCS), 32(08):991– 995, December 2021. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S012905412199001X.

Anonymous:2022:AIV

[Ano22] Anonymous. Author index volume 33 (2022). International Journal of Foundations of Computer Science (IJFCS), 33(08):1053– 1058, December 2022. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122990015.

Aytac:2010:CRD

[AO10] Aysun Aytac and Zeynep Nihan Odabas. Computing the rupture degree in composite graphs. International Journal of Foundations of Computer Science (IJFCS), 21(3):311–319, June 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Aytac:2011:RCW

[AO11] Aysun Aytaç and Zeynep Nihan Odabaş. Residual closeness of wheels and related networks. International Journal of Foundations of Computer Science (IJFCS), 22(5):1229–1240, August 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ando:2010:SCL

[AOSY10] Ei Ando, Hirotaka Ono, Kunihiko Sadakane, and Masafumi Yamashita. The space complexity of leader election in anonymous networks. International Journal of Foundations of Computer Science (IJFCS), 21(3):427– 440, June 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ausiello:1990:LPA

[AP90] G. Ausiello and M. Protasi.

Limiting polynomial approximation of complexity classes. International Journal of Foundations of Computer Science (IJFCS), 1(2): 111–??, June 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Agostino:1992:PCO

[AP92a] S. D. Agostino and R. Petreschi. On PVsub chunk operations and matrogenic graphs. International Journal of Foundations of Computer Science (IJFCS), 3(1): 11–20, March 1992. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Antonelli:1992:CMP

[AP92b] S. Antonelli and S. Pelagatti. On the complexity of the mapping problem for massively parallel architectures. International Journal of Foundations of Computer Science (IJFCS), 3 (3):379–??, September 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Azimi:2017:MSL

[APMP17] Sepinoud Azimi, Charmi Panchal, Andrzej Mizera, and Ion Petre. Multistability, limit cycles, and period-doubling bifurcation with reaction systems. International Journal of Foundations of Computer Science (IJFCS), 28(8):1007– ??, December 2017. CODEN IFCSEN. ISSN 0129-0541.

Abadi:1991:FIM

[APP91] Martin Abadi, Benjamin Pierce, and Gordon Plotkin. Faithful ideal models for recursive polymorphic types. International Journal of Foundations of Computer Science (IJFCS), 2(1):1– ??, March 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ananichev:2006:CWP

[APV06] D. S. Ananichev, I. V. Petrov, and M. V. Volkov. Collapsing words: a progress report. International Journal of Foundations of Computer Science (IJFCS), 17 (3):507–??, June 2006. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Almeida:2016:SSA

[AR16] Jorge Almeida and Emanuele Rodaro. Semisimple synchronizing automata and the Wedderburn–Artin theory. International Journal of Foundations of Computer Science (IJFCS), 27(2):127– ??, February 2016. CODEN IFCSEN. ISSN 0129-0541.

Arnold:2017:IGR

[Arn17] Stefan Arnold. Identifying generalized Reed–Muller codewords by quantum queries. International Journal of Foundations of Computer Science (IJFCS), 28 (2):185–194, February 2017. CODEN IFCSEN. ISSN 0129-0541.

Allauzen:2011:FBA

[ARS11] Cyril Allauzen, Michael Riley, and Johan Schalkwyk. A filter-based algorithm for efficient composition of finite-state transducers. International Journal of Foundations of Computer Science (IJFCS), 22(8):1781– 1795, December 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Arslan:2015:FAL

[Ars15] Abdullah N. Arslan. Fast algorithms for local similarity queries in two sequences. International Journal of Foundations of Computer Science (IJFCS), 26(5):625–??, August 2015. CODEN IFC-SEN. ISSN 0129-0541.

Arvind:1997:CMP

[Arv97] V. Arvind. Constructivizing membership proofs in complexity classes. International Journal of Foundations of Computer Science (IJFCS), 8(4):433–??, December 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Alhazov:2007:MCS

[ARV07] Artiom Alhazov, Yurii Rogozhin, and Sergey Verlan. Minimal cooperation in symport/antiport tissue P systems. International Journal of Foundations of Computer Science (IJFCS), 18 (1):163–179, February 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Alhazov:2012:SUS

[ARV12] Artiom Alhazov, Yurii Rogozhin, and Sergey Verlan. On small universal splicing systems. International Journal of Foundations of Computer Science (IJFCS), 23(7):1423–??, November 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Alnasir:2018:TQN

[AS18] Jamie J. Alnasir and Hugh P.

Shanahan. Transcriptomics: Quantifying nonuniform read distribution using MapReduce. International Journal of Foundations of Computer Science (IJFCS), 29(8):1355–1372, December 2018. ISSN 0129-0541.

Asano:2023:TPA

[Asa23] Tetsuo Asano. Transportation problem allowing sending and bringing back. International Journal of Foundations of Computer Science (IJFCS), 34(05):487– 505, August 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500289.

Aslan:2016:WRD

[Asl16] Ersin Aslan. Weak-rupture degree of graphs. International Journal of Foundations of Computer Science (IJFCS), 27(6):725–??, September 2016. CODEN IFCSEN. ISSN 0129-0541.

Alba:2012:P

[ASTZ12] Enrique Alba, Franciszek Seredynski, El-Ghazali Talbi, and Albert Y. Zomaya. Preface. International Journal of Foundations of Computer Science (IJFCS), 23 (2):403–405, February 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Asveld:2007:GAC

[Asv07] Peter R. J. Asveld. Generating all circular shifts by context-free grammars in Greibach Normal Form. International Journal of Foundations of Computer Science (IJFCS), 18(6):1139– 1149, December 2007. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Aytac:2011:VVP

[AT11] Aysun Aytac and Tufan Turaci. Vertex vulnerability parameter of gear graphs. International Journal of Foundations of Computer Science (IJFCS), 22 (5):1187–1195, August 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Auger:2012:FDP

[AT12] David Auger and Olivier Teytaud. The frontier of decidability in partially observable recursive games. International Journal of Foundations of Computer Science (IJFCS), 23(7): 1439–??, November 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Aytac:2015:VMT

[AT15] Aysun Aytaç and Tufan Turaci. Vulnerability measures of transformation graph G^{xy+}. International Journal of Foundations of Computer Science (IJFCS), 26(6):667– ??, September 2015. CO-DEN IFCSEN. ISSN 0129-0541.

Atanasiu:2016:NOP

[AT16] Adrian Atanasiu and Wen Chean Teh. A new operator over Parikh languages. International Journal of Foundations of Computer Science (IJFCS), 27(6):757–??, September 2016. CODEN IFCSEN. ISSN 0129-0541.

Aytac:2023:AVS

[AT23] Vecdi Aytaç and Tufan Turacı. Analysis of vulnerability of some transformation networks. International Journal of Foundations of Computer Science (IJFCS), 34(01):11– 24, January 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500162.

Atanasiu:2007:BAW

[Ata07] Adrian Atanasiu. Binary amiable words. International Journal of Foundations of Computer Science (IJFCS), 18(2):387– 400, April 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Atanasiu:2011:EPM

[Ata11] Radu-Florian Atanasiu. Erratum: Parikh Matrix Mapping and Languages. International Journal of Foundations of Computer Science (IJFCS), 22(1):273, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Abderrahim:2012:HGQ

[ATK12] Allani Abderrahim, El-Ghazali Talbi, and Mellouli Khaled. Hybridization of genetic and quantum algorithm for gene selection and classification of microarray data. International Journal of Foundations of Computer Science (IJFCS), 23 (2):431–444, February 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Agrawal:1996:ICD

[AV96] M. Agrawal and S. Venkatesh. On the isomorphism conjecture for 2-DFA reductions. International Journal of Foundations of Computer Science (IJFCS), 7 (4):339–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Alzoubi:2003:MIS

[AWF03] Khaled M. Alzoubi, Peng-Jun Wan, and Ophir Frieder. Maximal independent set, weakly-connected dominating set, and induced spanners in wireless ad hoc networks. International Journal of Foundations of Computer Science (IJFCS), 14(2):287–??, April 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Andresen:1999:P

[AY99] D. Andresen and T. Yang.

Preface. International Journal of Foundations of Computer Science (IJFCS), 10 (4):373–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Berberler:2024:NLV

[BA24] Zeynep Nihan Berberler and Aysun Aytaç. Node and link vulnerability in complete multipartite networks. International Journal of Foundations of Computer Science (IJFCS), 35 (04):375–385, June 2024. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123500077.

Badr:2009:HM

[Bad09]  Andrew Badr.  Hyper-

minimization in O(n^{2}). International Journal of Foundations of Computer Science (IJFCS), 20(4):735– 746, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ben-Amram:2012:EDC

[BAK12] Amir M. Ben-Amram and Lars Kristiansen. On the edge of decidability in complexity analysis of loop programs. International Journal of Foundations of Computer Science (IJFCS), 23(7):1451–??, November 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Barbanera:1990:CTR

[Bar90] Franco Barbanera. Combining term rewriting and type assignment systems. International Journal of Foundations of Computer Science (IJFCS), 1(3):165–??, September 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Basten:1997:PPO

[Bas97] Twan Basten. Parsing partially ordered multisets. International Journal of Foundations of Computer Science (IJFCS), 8(4):379–??, December 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Beigel:1999:NPR

[BB99] R. Beigel and A. Bernasconi.

A note on the polynomial representation of Boolean functions over GF(2). International Journal of Foundations of Computer Science (IJFCS), 10(4):535– 542, 1999. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bartzis:2003:ESR

[BB03a] Constantinos Bartzis and Tevfik Bultan. Efficient symbolic representations for arithmetic constraints in verification. International Journal of Foundations of Computer Science (IJFCS), 14(4):605–??, August 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bouda:2003:EQI

[BB03b] Jan Bouda and Vladimír R.

Bužek. Encryption of quantum information. International Journal of Foundations of Computer Science (IJFCS), 14(5):741– ??, October 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Blin:2004:FAD

[BB04] Lélia Blin and Franck Butelle. The first approximated distributed algorithm for the minimum degree spanning tree problem on general graphs. International Journal of Foundations of Computer Science (IJFCS), 15(3):507– ??, June 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bianchi:2018:OMS

[BBB^{+}18] Maria Paola Bianchi, Hans-Joachim Böckenhauer, Tatjana Brülisauer, Dennis Komm, and Beatrice Palano. Online minimum spanning tree with advice. International Journal of Foundations of Computer Science (IJFCS), 29(4):505– 527, June 2018. CODEN IFCSEN. ISSN 0129-0541.

Berry:2000:GAM

[BBC00] Anne Berry, Jean-Paul Bordat, and Olivier Cogis. Generating all the minimal separators of a graph. International Journal of Foundations of Computer Science (IJFCS), 11(3):397– 404, 2000. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bjorklund:2024:TBG

[BBE24] Henrik Björklund, Johanna Björklund, , and Petter Ericson. Tree-based generation of restricted graph languages. International Journal of Foundations of Computer Science (IJFCS), 35(1–2):215–243, January– February 2024. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123480106.

Babvey:2006:SEI

[BBFZM06] Sharareh Babvey, Anu G.

Bourgeois, José Alberto Fernández-Zepeda, and Steven W. Mclaughlin. Scalable and efficient implementations of the LDPC decoder using reconfigurable models. International Journal of Foundations of Computer Science (IJFCS), 17(2):303– ??, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bensch:2017:DST

[BBK17] Suna Bensch, Johanna Björklund, and Martin Kutrib. Deterministic stack transducers. International Journal of Foundations of Computer Science (IJFCS), 28(5):583–??, August 2017. CODEN IFCSEN. ISSN 0129-0541.

Benattar:2012:CSF

[BBL^{+}12] Gilles Benattar, Béatrice Bérard, Didier Lime, John Mullins, Olivier H. Roux, and Mathieu Sassolas. Channel synthesis for finite transducers. International Journal of Foundations of Computer Science (IJFCS), 23(6):1241–??, September 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bertoni:1990:I

[BBM90] A. Bertoni, C. Bohm, and P. Miglioli. Introduction. International Journal of Foundations of Computer Science (IJFCS), 1(3):??, September 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Boukerche:2012:EPA

[BBM^{+}12] Azzedine Boukerche, Rodolfo Bezerra Batista, Alba Cristina Magalhaes Alves De Melo, Felipe Brandt Scarel, and Lavir Antonio Bahia Carvalho De Souza. Exact parallel alignment of megabase genomic sequences with tunable work distribution. International Journal of Foundations of Computer Science (IJFCS), 23(2):407– 429, February 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Beal:2011:QUB

[BBP11] Marie-Pierre Béal, Mikhail V. Berlinkov, and Dominique Perrin. A quadratic upper bound on the size of a synchronizing word in onecluster automata. International Journal of Foundations of Computer Science (IJFCS), 22(2):277– 288, February 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bes:2006:KTL

[BC06] Alexis Bès and Olivier Carton. A Kleene theorem for languages of words indexed by linear orderings. International Journal of Foundations of Computer Science (IJFCS), 17(3):519– ??, June 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Brim:2012:USI

[BC12] Luboš Brim and Jakub Chaloupka. Using strategy improvement to stay alive. International Journal of Foundations of Computer Science (IJFCS), 23(3):585– ??, April 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Beal:2014:P

[BC14] Marie-Pierre Béal and Olivier Carton. Preface. International Journal of Foundations of Computer Science (IJFCS), 25(8):933–??, December 2014. CODEN IFC-SEN. ISSN 0129-0541.

Bennoui:2012:SAI

[BCB12] Hammadi Bennoui, Allaoua Chaoui, and Kamel Barkaoui. On structural analysis of interacting behavioral Petri nets for distributed causal model-based diagnosis. International Journal of Foundations of Computer Science (IJFCS), 23(7):1523–??, November 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Breveglieri:1996:MPL

[BCC^{+}96] L. Breveglieri, A. Cherubini, C. Citrini, et al. Multi-push-down languages and grammars. International Journal of Foundations of Computer Science (IJFCS), 7 (3):253–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Besozzi:2011:MDS

[BCC^{+}11] Daniela Besozzi, Paolo Cazzaniga, Stefania Cocolo, Giancarlo Mauri, and Dario Pescini. Modeling diffusion in a signal transduction pathway: the use of virtual volumes in P systems. International Journal of Foundations of Computer Science (IJFCS), 22(1):89– 96, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Burgin:2013:ICM

[BCC13] Mark Burgin, Cristian S. Calude, and Elena Calude. Inductive complexity measures for mathematical problems. International Journal of Foundations of Computer Science (IJFCS), 24(4):487–??, June 2013. CODEN IFCSEN. ISSN 0129-0541.

Bertoni:2014:DIP

[BCD14] Alberto Bertoni, Christian Choffrut, and Flavio D’Alessandro. On the decidability of the intersection problem for quantum automata and context-free languages. International Journal of Foundations of Computer Science (IJFCS), 25(8):1065–??, December 2014. CODEN IFCSEN. ISSN 0129-0541.

Bhagat:2023:OGW

[BCDM23] Subhash Bhagat, Abhinav Chakraborty, Bibhuti Das, and Krishnendu Mukhopadhyaya. Optimal gathering over Weber meeting nodes in infinite grid. International Journal of Foundations of Computer Science (IJFCS), 34(01):25– 49, January 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500174.

Berardi:2008:ASC

[BCDP08] Daniela Berardi, Fahima Cheikh, Giuseppe De Giacomo, and Fabio Patrizi. Automatic service composition via simulation. International Journal of Foundations of Computer Science (IJFCS), 19(2):429– 451, April 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Burcsi:2012:AJP

[BCFL12] Péter Burcsi, Ferdinando Cicalese, Gabriele Fici, and Zsuzsanna Lipták. Algorithms for jumbled pattern matching in strings. International Journal of Foundations of Computer Science (IJFCS), 23(2):357– 374, February 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bastien:2007:RSG

[BCFR07] Cédric Bastien, Jurek Czyzowicz, Wojciech Fraczak, and Wojciech Rytter. Reducing simple grammars: Exponential against highlypolynomial time in practice. International Journal of Foundations of Computer Science (IJFCS), 18(4):715– 725, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Boichut:2009:HNL

[BCHK09] Yohan Boichut, Romeo Courbis, Pierre-Cyrille Heam, and Olga Kouchnarenko. Handling non left-linear rules when completing tree automata. International Journal of Foundations of Computer Science (IJFCS), 20(5):837– 849, October 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bera:2020:PQM

[BCMS20] Somnath Bera, Rodica Ceterchi, Kalpana Mahalingam, and K. G. Subramanian. Parikh q-matrices and q-ambiguous words. International Journal of Foundations of Computer Science (IJFCS), 31(1):23– 36, January 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412040002X.

Bischoff:2012:UPI

[BCN12] Bastian Bischoff, James Currie, and Dirk Nowotka. Unary patterns with involution. International Journal of Foundations of Computer Science (IJFCS), 23(8): 1641–??, December 2012. CODEN IFCSEN. ISSN 0129-0541.

BoavidaDeBrito:2022:PMC

[BCP22] Vasco Boavida De Brito, José Félix Costa, and Diogo Poças. The power of machines that control experiments. International Journal of Foundations of Computer Science (IJFCS), 33 (02):91–118, February 2022. CODEN IFCSEN. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122500010.

Bloem:2007:SIA

[BCPR07] Roderick Bloem, Alessandro Cimatti, Ingo Pill, and Marco Roveri. Symbolic implementation of alternating automata. International Journal of Foundations of Computer Science (IJFCS), 18(4):727–743, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bertoni:2011:IPC

[BCR11] Alberto Bertoni, Christian Choffrut, and Roberto Radicioni. The inclusion problem of context-free languages: Some tractable cases. International Journal of Foundations of Computer Science (IJFCS), 22(2):289– 299, February 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bondar:2023:CRA

[BCV23] Eugenija A. Bondar, David Casas, , and Mikhail V. Volkov. Completely reachable automata: an interplay between automata, graphs, and trees. International Journal of Foundations of Computer Science (IJFCS), 34(06):655– 690, September 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123450053.

Beek:2007:CCG

[BCVVH07] Maurice H. Ter Beek, Erzsébet Csuhaj-Varjú, György Vaszil, and Markus Holzer. On competence in CD grammar systems with parallel rewriting. International Journal of Foundations of Computer Science (IJFCS), 18(6):1425–1439, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Brzozowski:2019:MCN

[BD19] Janusz A. Brzozowski and Sylvie Davies. Most complex non-returning regular languages. International Journal of Foundations of Computer Science (IJFCS), 30(6–7):921–957, September–November 2019. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054119400239.

Bovet:1990:DPC

[BDC90] Daniel P. Bovet, Miriam Di Ianni, and Pierluigi Crescenzi. Deadlock prediction in the case of dynamic routing. International Journal of Foundations of Computer Science (IJFCS), 1 (3):185–??, September 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Berthe:2018:RSD

[BDD^{+}18] V. Berthé, F. Dolce, F. Durand, J. Leroy, and D. Perrin. Rigidity and substitutive dendric words. International Journal of Foundations of Computer Science (IJFCS), 29(5):705– 720, August 2018. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054118420017.

Bui:2001:RMA

[BDDN01] Marc Bui, Sajal K. Das, Ajoy K. Datta, and D. T. Nguyen. Randomized mobile agent based routing in wireless networks. International Journal of Foundations of Computer Science (IJFCS), 12(3):365– 384, 2001. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bougeret:2011:ADR

[BDG^{+}11] Marin Bougeret, Pierre-François Dutot, Alfredo Goldman, Yanik Ngoko, and Denis Trystram. Approximating the discrete resource sharing scheduling problem. International Journal of Foundations of Computer Science (IJFCS), 22(3):639– 656, April 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bujtas:2011:GPP

[BDI^{+}11] Csilla Bujtás, György Dósa, Csanád Imreh, Judit Nagy-György, and Zsolt Tuza. The graph-bin packing problem. International Journal of Foundations of Computer Science (IJFCS), 22(8):1971–1993, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Biegler:2008:CAM

[BDL08] Franziska Biegler, Mark Daley, and M. Elizabeth O. Locke. Computation by annotation: Modelling epigenetic regulation. International Journal of Foundations of Computer Science (IJFCS), 19(5):1087–1098, October 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bertolotti:2006:EST

[BDSV06] Ivan Cibrario Bertolotti, Luca Durante, Riccardo Sisto, and Adriano Valenzano. Exploiting symmetries for testing equivalence verification in the Spi calculus. International Journal of Foundations of Computer Science (IJFCS), 17(4):815– ??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bloom:1992:IA

[BE92] S. L. Bloom and Z. Esik. Iteration algebras. International Journal of Foundations of Computer Science (IJFCS), 3(3):245–302, September 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bloom:1993:IA

[BE93] S. L. Bloom and Z. Esik. Iteration algebras. International Journal of Foundations of Computer Science (IJFCS), 4(1):99–??, March 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bloom:1995:SEL

[BE95] S. L. Bloom and Z. Esik. Some equational laws of initiality in 2CCC’s. International Journal of Foundations of Computer Science (IJFCS), 6(2):95–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bloom:2011:ALO

[BÉ11] S. L. Bloom and Z. Ésik. Algebraic linear orderings. International Journal of Foundations of Computer Science (IJFCS), 22(2):491– 515, February 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Boyar:2019:TBR

[BE19] Joan Boyar and Faith Ellen.

Tight bounds for restricted grid scheduling. International Journal of Foundations of Computer Science (IJFCS), 30(3):375– 405, 2019. ISSN 0129-0541.

Bedon:2018:CBA

[Bed18] Nicolas Bedon. Complementation of branching automata for scattered and countable N-free posets. International Journal of Foundations of Computer Science (IJFCS), 29(5):769– 799, August 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118420042.

Beeson:1995:UNA

[Bee95] M. Beeson. Using nonstandard analysis to ensure the correctness of symbolic computations. International Journal of Foundations of Computer Science (IJFCS), 6(3):299–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Brijder:2011:TRS

[BEMR11] Robert Brijder, Andrzej Ehrenfeucht, Michael Main, and Grzegorz Rozenberg. A tour of reaction systems. International Journal of Foundations of Computer Science (IJFCS), 22(7):1499– 1517, November 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Berlinkov:2011:CCD

[Ber11] Mikhail V. Berlinkov. On a conjecture by Carpi and D’Alessandro. International Journal of Foundations of Computer Science (IJFCS), 22(7):1565–1576, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Berlinkov:2013:SQE

[Ber13] Mikhail V. Berlinkov. Synchronizing quasi-Eulerian and quasi-one-cluster automata. International Journal of Foundations of Computer Science (IJFCS), 24 (6):729–??, September 2013. CODEN IFCSEN. ISSN 0129-0541.

Bhika:2007:TDC

[BESW07] Charita Bhika, Sigrid Ewert, Ryan Schwartz, and Mutahi Waruhiu. Tabledriven context-free picture grammars. International Journal of Foundations of Computer Science (IJFCS), 18(6):1151–1160, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bouajjani:2003:GAS

[BET03] Ahmed Bouajjani, Javier Esparza, and Tayssir Touili. A generic approach to the static analysis of concurrent programs with procedures. International Journal of Foundations of Computer Science (IJFCS), 14(4):551– ??, August 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Berthome:1997:CIP

[BF97] Pascal Berthomé and Afonso Ferreira. Communication issues in parallel systems with optical interconnections. International Journal of Foundations of Computer Science (IJFCS), 8(2): 143–??, June 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Becha:2007:OCS

[BF07] Hanane Becha and Paola Flocchini. Optimal construction of sense of direction in a torus by a mobile agent. International Journal of Foundations of Computer Science (IJFCS), 18(3):529– 546, June 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bilbao:2002:CCV

[BFL02] J. M. Bilbao, J. R. Fernández, and J. J. López. On the complexity of computing values of restricted games. International Journal of Foundations of Computer Science (IJFCS), 13(5):633– ??, October 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bianco:2006:SRM

[BFM06] Luca Bianco, Federico Fontana, and Vincenzo Manca. P systems with reaction maps. International Journal of Foundations of Computer Science (IJFCS), 17(1):27–??, February 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Barriere:2011:USA

[BFMBS11] Lali Barrière, Paola Flocchini, Eduardo Mesa-Barrameda, and Nicola Santoro. Uniform scattering of autonomous mobile robots in a grid. International Journal of Foundations of Computer Science (IJFCS), 22(3):679– 697, April 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bordim:2010:P

[BFN10] Jacir L. Bordim, Akihiro Fujiwara, and Koji Nakano. Preface. International Journal of Foundations of Computer Science (IJFCS), 21 (3):385–386, June 2010. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bordim:2012:P

[BFN12] Jacir L. Bordim, Akihiro Fujiwara, and Koji Nakano. Preface. International Journal of Foundations of Computer Science (IJFCS), 23 (4):761–??, June 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bannai:2018:DPF

[BGI^{+}18] Hideo Bannai, Travis Gagie, Shunsuke Inenaga, Juha Kärkkäinen, Dominik Kempa, Marcin Pia̧tkowski, and Shiho Sugimoto. Diverse palindromic factorization is NP-complete. International Journal of Foundations of Computer Science (IJFCS), 29(2):143–??, February 2018. CODEN IFCSEN. ISSN 0129-0541.

Bartholdi:2020:NHA

[BGK^{+}20] Laurent Bartholdi, Thibault Godin, Ines Klimann, Camille Noûs, and Matthieu Picantin. A new hierarchy for automaton semigroups. International Journal of Foundations of Computer Science (IJFCS), 31(08):1069– 1089, December 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120420046.

Belazzougui:2018:BVO

[BGM^{+}18] Djamal Belazzougui, Travis Gagie, Veli Mäkinen, Marco Previtali, and Simon J.

Puglisi. Bidirectional variable-order de Bruijn graphs. International Journal of Foundations of Computer Science (IJFCS), 29 (8):1279–1295, December 2018. ISSN 0129-0541.

Bernardini:2008:HSA

[BGMV08] Francesco Bernardini, Marian Gheorghe, Maurice Margenstern, and Sergey Verlan. How to synchronize the activity of all components of a P system? International Journal of Foundations of Computer Science (IJFCS), 19(5):1183–1198, October 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bassino:2010:ASC

[BGN10] Frédérique Bassino, Laura Giambruno, and Cyril Nicaud. The average state complexity of rational operations on finite languages. International Journal of Foundations of Computer Science (IJFCS), 21(4):495– 516, August 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bednarova:2016:NRM

[BGRY16] Zuzana Bednárová, Viliam Geffert, Klaus Reinhardt, and Abuzer Yakaryilmaz. New results on the minimum amount of useful space. International Journal of Foundations of Computer Science (IJFCS), 27(2):259–??, February 2016. CODEN IFCSEN. ISSN 0129-0541.

Brzozowski:2011:CFL

[BGS11] Janusz Brzozowski, Elyot Grant, and Jeffrey Shallit. Closures in formal languages and Kuratowski’s theorem. International Journal of Foundations of Computer Science (IJFCS), 22(2):301– 321, February 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bergeron:2002:VAA

[BH02] A. Bergeron and S. Hamel.

Vector algorithms for approximate string matching. International Journal of Foundations of Computer Science (IJFCS), 13 (1):53–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Budaghyan:2011:ICP

[BH11] Lilya Budaghyan and Tor Helleseth. On isotopisms of commutative presemifields and CCZ-equivalence of functions. International Journal of Foundations of Computer Science (IJFCS), 22(6):1243–1258, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Beier:2020:DRO

[BH20] Simon Beier and Markus Holzer. Decidability of right one-way jumping finite automata. International Journal of Foundations of Computer Science (IJFCS), 31(06):805– 825, September 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120410063.

Bordihn:2005:ULO

[BHK05] Henning Bordihn, Markus Holzer, and Martin Kutrib. Unsolvability levels of operation problems for subclasses of context-free languages. International Journal of Foundations of Computer Science (IJFCS), 16 (3):423–??, June 2005. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bordihn:2007:HEF

[BHK07] Henning Bordihn, Markus Holzer, and Martin Kutrib. Hybrid extended finite automata. International Journal of Foundations of Computer Science (IJFCS), 18 (4):745–760, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bianchi:2018:STW

[BHK18a] Maria Paola Bianchi, Juraj Hromkovič, and Ivan Kováč.

On the size of two-way reasonable automata for the liveness problem. International Journal of Foundations of Computer Science (IJFCS), 29(2):187–??, February 2018. CODEN IFCSEN. ISSN 0129-0541.

Burjons:2018:OGC

[BHK^{+}18b] Elisabet Burjons, Juraj Hromkovič, Rastislav Královič, Richard Královič, Xavier Muñoz, and Walter Unger. Online graph coloring against a randomized adversary. International Journal of Foundations of Computer Science (IJFCS), 29(4):551– 569, June 2018. CODEN IFCSEN. ISSN 0129-0541.

Beier:2019:OSC

[BHK19] Simon Beier, Markus Holzer, and Martin Kutrib. Operational state complexity and decidability of jumping finite automata. International Journal of Foundations of Computer Science (IJFCS), 30(1):5– 27, January 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905411940001X.

Bermond:1997:NDM

[BHL^{+}97] Jean-Claude Bermond, H. A.

Harutyunyan, A. L. Liestman, et al. A note on the dimensionality of modified Knödel graphs. International Journal of Foundations of Computer Science (IJFCS), 8(2):109–??, June 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Brlek:2004:PCI

[BHNR04] S. Brlek, S. Hamel, M. Nivat, and C. Reutenauer. On the palindromic complexity of infinite words. International Journal of Foundations of Computer Science (IJFCS), 15(2):293– ??, April 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Brijder:2009:RGO

[BHR09] Robert Brijder, Hendrik Jan Hoogeboom, and Grzegorz Rozenberg. Reduction graphs from overlap graphs for gene assembly in ciliates. International Journal of Foundations of Computer Science (IJFCS), 20(2):271– 291, April 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bordim:2004:ISS

[BIIN04] Jacir L. Bordim, Oscar H. Ibarra, Yasuaki Ito, and Koji Nakano. Instancespecific solutions for accelerating the CKY parsing of large context-free grammars. International Journal of Foundations of Computer Science (IJFCS), 15(2):403– ??, April 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Birget:2011:CSI

[Bir11] J. C. Birget. On the circuitsize of inverses. International Journal of Foundations of Computer Science (IJFCS), 22(8):1925–1938, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Brzozowski:2005:RSC

[BJ05] Janusz Brzozowski and Helmut Jürgensen. Representation of semiautomata by canonical words and equivalences. International Journal of Foundations of Computer Science (IJFCS), 16 (5):831–??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic). See errata [BJ06].

Brzozowski:2006:ERS

[BJ06] Janusz Brzozowski and Helmut Jürgensen. Errata: Representation of Semiautomata by Canonical Words and Equivalences. International Journal of Foundations of Computer Science (IJFCS), 17(5):1231– ??, October 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic). See [BJ05].

Boldt:2007:SLN

[BJ07a] Oliver Boldt and Helmut Jürgensen. Soliton languages are nearly an anti-AFL. International Journal of Foundations of Computer Science (IJFCS), 18 (6):1161–1165, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Brzozowski:2007:RSC

[BJ07b] Janusz Brzozowski and Helmut Jürgensen. Representation of semiautomata by canonical words and equivalences, Part II: Specification of software modules. International Journal of Foundations of Computer Science (IJFCS), 18(5):1065–1087, October 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Basappa:2020:CKK

[BJD20] Manjanna Basappa, Ramesh K. Jallu, and Gautam K. Das. Constrained k-center problem on a convex polygon. International Journal of Foundations of Computer Science (IJFCS), 31(2):275– 291, February 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120500070.

Bruschi:1990:SSB

[BJY90] Danilo Bruschi, Deborah Joseph, and Paul Young.

Strong separations for the Boolean hierarchy over RP. International Journal of Foundations of Computer Science (IJFCS), 1(3):201– ??, September 1990. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Besnard:1995:ETB

[BK95] P. Besnard and J. Kohlas.

Evidence theory based on general consequence relations. International Journal of Foundations of Computer Science (IJFCS), 6 (2):119–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Berdinsky:2016:CAR

[BK16] Dmitry Berdinsky and Bakhadyr Khoussainov. Cayley automatic representations of wreath products. International Journal of Foundations of Computer Science (IJFCS), 27(2):147–??, February 2016. CODEN IFCSEN. ISSN 0129-0541.

Beaur:2024:EPG

[BK24] Pierre Béaur and Jarkko Kari. Effective projections on group shifts to decide properties of group cellular automata. International Journal of Foundations of Computer Science (IJFCS), 35(1–2):77–100, January– February 2024. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054123480040.

Borovansky:2001:RSE

[BKKR01] Peter Borovanský, Claude Kirchner, Hélène Kirchner, and C. Ringeissen. Rewriting with strategies in ELAN: a functional semantics. International Journal of Foundations of Computer Science (IJFCS), 12(1):69–95 (or 69–96??), 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Brzozowski:2020:SCO

[BKLS20] Janusz A. Brzozowski, Lila Kari, Bai Li, and Marek Szykuła. State complexity of overlap assembly. International Journal of Foundations of Computer Science (IJFCS), 31(08):1113– 1132, December 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S012905412042006X.

Bordihn:2011:UHR

[BKM11] Henning Bordihn, Martin Kutrib, and Andreas Malcher. Undecidability and hierarchy results for parallel communicating finite automata. International Journal of Foundations of Computer Science (IJFCS), 22(7):1577–1592, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bordihn:2012:CCP

[BKM12] Henning Bordihn, Martin Kutrib, and Andreas Malcher. On the computational capacity of parallel communicating finite automata. International Journal of Foundations of Computer Science (IJFCS), 23(3):713–??, April 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bordihn:2015:RPC

[BKM15] Henning Bordihn, Martin Kutrib, and Andreas Malcher. Returning parallel communicating finite automata with communication bounds: Hierarchies, decidabilities, and undecidabilities. International Journal of Foundations of Computer Science (IJFCS), 26(8):1101–??, December 2015. CODEN IFCSEN. ISSN 0129-0541.

Bournez:2018:RPO

[BKP18] Olivier Bournez, Oleksiy Kurganskyy, and Igor Potapov. Reachability problems for one-dimensional piecewise affine maps. International Journal of Foundations of Computer Science (IJFCS), 29(4):529– 549, June 2018. CODEN IFCSEN. ISSN 0129-0541.

Bui:2012:ARW

[BKS12] Alain Bui, Abdurusul Kudireti, and Devan Sohier. An adaptive random walk based distributed clustering algorithm. International Journal of Foundations of Computer Science (IJFCS), 23(4):803–??, June 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bibak:2018:AUH

[BKST18] Khodakhast Bibak, Bruce M. Kapron, Venkatesh Srinivasan, and László Tóth. On an almost-universal hash function family with applications to authentication and secrecy codes. International Journal of Foundations of Computer Science (IJFCS), 29(3):357– 375, April 2018. CODEN IFCSEN. ISSN 0129-0541.

Bruggemann-Klein:2002:RTW

[BKW02] A. Brüggemann-Klein and D. Wood. The regularity of two-way nondeterministic tree automata languages. International Journal of Foundations of Computer Science (IJFCS), 13 (1):67–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Blokh:2001:MTS

[BL01] David Blokh and E. Levner. The maximum traveling salesman problem on banded matrices. International Journal of Foundations of Computer Science (IJFCS), 12(6):809– 820, December 2001. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Brzozowski:2012:QCS

[BL12] Janusz Brzozowski and Bo Liu. Quotient complexity of star-free languages. International Journal of Foundations of Computer Science (IJFCS), 23(6): 1261–??, September 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Brzozowski:2014:SCS

[BL14] Janusz Brzozowski and Baiyu Li. Syntactic complexity of R- and J -trivial regular languages. International Journal of Foundations of Computer Science (IJFCS), 25(7):807–??, November 2014. CODEN IFCSEN. ISSN 0129-0541.

Bleak:2021:NAS

[Ble21] Collin Bleak. Normalish amenable subgroups of the R. Thompson groups. International Journal of Foundations of Computer Science (IJFCS), 32(06):785– 800, September 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121420089.

Brlek:2006:PCP

[BLL06] S. Brlek, G. Labelle, and A. Lacasse. Properties of the contour path of discrete sets. International Journal of Foundations of Computer Science (IJFCS), 17(3):543– ??, June 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bein:2003:BSH

[BLLS03] Wolfgang W. Bein, Lawrence L. Larmore, Shahram Latifi, and I. Hal Sudborough. Block sorting is hard. International Journal of Foundations of Computer Science (IJFCS), 14(3):425– ??, June 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Brandstadt:2004:GCG

[BLM04] Andreas Brandstädt, Hoàng-Oanh Le, and Raffaele Mosca. Gem- and co-gem-free graphs have bounded clique-width. International Journal of Foundations of Computer Science (IJFCS), 15(1):163–??, February 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Boyar:2015:FIP

[BLM15] Joan Boyar, Kim S. Larsen, and Abyayananda Maiti.

The frequent items problem in online streaming under various performance measures. International Journal of Foundations of Computer Science (IJFCS), 26 (4):413–??, June 2015. CO-DEN IFCSEN. ISSN 0129-0541.

Beaumont:2005:SSS

[BLMR05] Olivier Beaumont, Arnaud Legrand, Loris Marchal, and Yves Robert. Steadystate scheduling on heterogeneous clusters. International Journal of Foundations of Computer Science (IJFCS), 16(2):163– ??, April 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Barton:2018:FAC

[BLP18] Carl Barton, Chang Liu, and Solon P. Pissis. Fast average-case pattern matching on weighted sequences. International Journal of Foundations of Computer Science (IJFCS), 29(8): 1331–1343, December 2018. ISSN 0129-0541.

Bein:2009:KSC

[BLR09] Wolfgang Bein, Lawrence L.

Larmore, and Rüdiger Reischuk. Knowledge states for the caching problem in shared memory multiprocessor systems. International Journal of Foundations of Computer Science (IJFCS), 20(1):167– 183, February 2009. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bottoni:2020:NRS

[BLR20] Paolo Bottoni, Anna Labella, and Grzegorz Rozenberg. Networks of reaction systems. International Journal of Foundations of Computer Science (IJFCS), 31(1):53– 71, January 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120400043.

Borchert:2005:DDP

[BLS^{+}05] Bernd Borchert, Klaus-Jörn Lange, Frank Stephan, Pascal Tesson, and Denis Thérien. The dot-depth and the polynomial hierarchies correspond on the delta levels. International Journal of Foundations of Computer Science (IJFCS), 16(4):625– ??, August 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bell:2020:ANT

[BLS20] Jason Bell, Thomas F. Lidbetter, and Jeffrey Shallit. Additive number theory via approximation by regular languages. International Journal of Foundations of Computer Science (IJFCS), 31(06):667– 687, September 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120410014.

Brzozowski:2012:CET

[BLY12] Janusz Brzozowski, Baiyu Li, and Yuli Ye. On the complexity of the evaluation of transient extensions of Boolean functions. International Journal of Foundations of Computer Science (IJFCS), 23(1):21–35, January 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Barbuti:1990:RNF

[BM90] Roberto Barbuti and Maurizio Martelli. Recognizing non-floundering logic programs and goals. International Journal of Foundations of Computer Science (IJFCS), 1(2):151–??, June 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bera:2016:SAA

[BM16] Somnath Bera and Kalpana Mahalingam. Some algebraic aspects of Parikh q-matrices. International Journal of Foundations of Computer Science (IJFCS), 27(4):479–??, June 2016. CODEN IFCSEN. ISSN 0129-0541.

Blattmann:2023:CCW

[BM23] Malte Blattmann and Andreas Maletti. Compositions with constant weighted extended tree transducers. International Journal of Foundations of Computer Science (IJFCS), 34(06):539– 558, September 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122450010.

Broda:2011:ASC

[BMMR11] Sabine Broda, António Machiavelo, Nelma Moreira, and Rogério Reis. On the average state complexity of partial derivative automata: an analytic combinatorics approach. International Journal of Foundations of Computer Science (IJFCS), 22(7):1593–1606, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Broda:2012:ASG

[BMMR12] Sabine Broda, António Machiavelo, Nelma Moreira, and Rogério Reis. On the average size of Glushkov and partial derivative automata. International Journal of Foundations of Computer Science (IJFCS), 23 (5):969–??, August 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Broda:2019:ABR

[BMMR19] Sabine Broda, António Machiavelo, Nelma Moreira, and Rogério Reis. On average behaviour of regular expressions in strong star normal form. International Journal of Foundations of Computer Science (IJFCS), 30(6–7):899– 920, September–November 2019. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054119400227.

Bertoni:2003:GRD

[BMP03] Alberto Bertoni, Carlo Mereghetti, and Beatrice Palano. Golomb rulers and difference sets for succinct quantum automata. International Journal of Foundations of Computer Science (IJFCS), 14(5):871– ??, October 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bianchi:2015:POW

[BMP15] Maria Paola Bianchi, Carlo Mereghetti, and Beatrice Palano. On the power of one-way automata with quantum and classical states. International Journal of Foundations of Computer Science (IJFCS), 26 (7):895–??, November 2015. CODEN IFCSEN. ISSN 0129-0541.

Bonomo:2014:SCS

[BMR^{+}14] Silvia Bonomo, Sabrina Mantaci, Antonio Restivo, Giovanna Rosone, and Marinella Sciortino. Sorting conjugates and suffixes of words in a multiset. International Journal of Foundations of Computer Science (IJFCS), 25(8):1161–??, December 2014. CODEN IFC-SEN. ISSN 0129-0541.

Bertoni:1992:HGF

[BMS92] A. Bertoni, P. Massazza, and N. Sabadini. Holonomic generating functions and context free languages. International Journal of Foundations of Computer Science (IJFCS), 3(2):181– 192, June 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bonifacio:2012:MPC

[BMS12] Adilson Luiz Bonifacio, Arnaldo Vieira Moura, and Adenilso Simao. Model partitions and compact test case suites. International Journal of Foundations of Computer Science (IJFCS), 23(1):147– 172, January 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bera:2018:PPM

[BMS18] Somnath Bera, Kalpana Mahalingam, and K. G.

Subramanian. Properties of Parikh matrices of binary words obtained by an extension of a restricted shuffle operator. International Journal of Foundations of Computer Science (IJFCS), 29(3):403–413, April 2018. CODEN IFCSEN. ISSN 0129-0541.

Barbuti:2011:OOS

[BMSMT11] Roberto Barbuti, Andrea Maggiolo-Schettini, Paolo Milazzo, and Simone Tini. An overview on operational semantics in membrane computing. International Journal of Foundations of Computer Science (IJFCS), 22(1):119– 131, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bergstra:1991:UAS

[BMW91] J. A. Bergstra, S. Mauw, and F. Wiedijk. Uniform algebraic specifications of finite sets with equality. International Journal of Foundations of Computer Science (IJFCS), 2(1):43– ??, March 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Belovs:2017:CCC

[BMY17] Aleksandrs Belovs, J. Andres Montoya, and Abuzer Yakaryılmaz. On a conjecture by Christian Choffrut. International Journal of Foundations of Computer Science (IJFCS), 28(5):483– ??, August 2017. CODEN IFCSEN. ISSN 0129-0541.

Bordim:2007:P

[BN07] Jacir L. Bordim and Koji Nakano. Preface. International Journal of Foundations of Computer Science (IJFCS), 18(3):433– 434, June 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bordim:2008:P

[BN08] Jacir L. Bordim and Koji Nakano. Preface. International Journal of Foundations of Computer Science (IJFCS), 19(6):1277–1278, December 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Berlinkov:2020:SAG

[BN20] Mikhail V. Berlinkov and Cyril Nicaud. Synchronizing almost-group automata. International Journal of Foundations of Computer Science (IJFCS), 31(08):1091– 1112, December 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120420058.

Bahig:2020:UMP

[BNBN20] Hatem M. Bahig, Dieaa I. Nassr, Ashraf Bhery, and Abderrahmane Nitaj. A unified method for private exponent attacks on RSA using lattices. International Journal of Foundations of Computer Science (IJFCS), 31(2):207– 231, February 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500045.

Bordim:2011:P

[BNF11] Jacir L. Bordim, Koji Nakano, and Akihiro Fujiwara. Preface. International Journal of Foundations of Computer Science (IJFCS), 22(3):517–518, April 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Boeres:1999:CBT

[BNR99] C. Boeres, A. Nascimento, and V. E. F. Rebello. Cluster-based task scheduling for the LogP model. International Journal of Foundations of Computer Science (IJFCS), 10(4):405– 424, 1999. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bordim:2005:Fa

[BNR05a] Jacir L. Bordim, Koji Nakano, and Arnold L. Rosenberg. Foreword. International Journal of Foundations of Computer Science (IJFCS), 16(1):1–??, February 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bordim:2005:Fb

[BNR05b] Jacir L. Bordim, Koji Nakano, and Arnold L. Rosenberg. Foreword. International Journal of Foundations of Computer Science (IJFCS), 16(2):143– ??, April 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bordim:2003:SSC

[BNS03] Jacir Luiz Bordim, Koji Nakano, and Hong Shen. Sorting on single-channel wireless sensor networks. International Journal of Foundations of Computer Science (IJFCS), 14(3):391– ??, June 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Baumslag:1997:ISG

[BO97] Marc Baumslag and Bojana Obrenic. Index-shuffle graphs. International Journal of Foundations of Computer Science (IJFCS), 8 (3):289–??, September 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bodlaender:1991:CSC

[Bod91] Hans L. Bodlaender. On the complexity of some coloring games. International Journal of Foundations of Computer Science (IJFCS), 2(2):133–148, June 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Belohlavek:2008:FFS

[BOV08] Radim Bělohlávek, Jan Outrata, and Vilem Vychodil. Fast factorization by similarity of fuzzy concept lattices with hedges. International Journal of Foundations of Computer Science (IJFCS), 19(2):255– 269, April 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bournez:2011:P

[BP11] Olivier Bournez and Igor Potapov. Preface. International Journal of Foundations of Computer Science (IJFCS), 22(4):757– 760, June 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Baturo:2009:UDA

[BPR09] Paweł Baturo, Marcin Piatkowski, and Wojciech Rytter. Usefulness of directed acyclic subword graphs in problems related to standard Sturmian words. International Journal of Foundations of Computer Science (IJFCS), 20 (6):1005–1023, December 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Borozky:2006:PCN

[BPT06] Károly J. Börözky, János Pach, and Géza Tóth. Planar crossing numbers of graphs embeddable in another surface. International Journal of Foundations of Computer Science (IJFCS), 17(5):1005– ??, October 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bertier:2016:CCM

[BPT16] Marin Bertier, Matthieu Perrin, and Cédric Tedeschi. On the complexity of concurrent multiset rewriting. International Journal of Foundations of Computer Science (IJFCS), 27(1):67– ??, January 2016. CODEN IFCSEN. ISSN 0129-0541.

Balaban:2007:MRA

[BPZ07] Ittai Balaban, Amir Pnueli, and Lenore D. Zuck. Modular ranking abstraction. International Journal of Foundations of Computer Science (IJFCS), 18(1):5–44, February 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bertoni:2008:AMS

[BR08] Alberto Bertoni and Roberto Radicioni. Approximating the mean speedup in trace monoids. International Journal of Foundations of Computer Science (IJFCS), 19(3):497–511, June 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bucci:2018:GPL

[BR18] Michelangelo Bucci and Gwenaël Richomme. Greedy palindromic lengths. International Journal of Foundations of Computer Science (IJFCS), 29(3):331– 356, April 2018. CODEN IFCSEN. ISSN 0129-0541.

Baudru:2020:TWT

[BR20] Nicolas Baudru and Pierre-Alain Reynier. From twoway transducers to regular function expressions. International Journal of Foundations of Computer Science (IJFCS), 31(06):843– 873, September 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120410087.

Balla:2007:EAD

[BRM07] Sudha Balla, Sanguthevar Rajasekaran, and Ion I. Mandoiu. Efficient algorithms for degenerate primer search. International Journal of Foundations of Computer Science (IJFCS), 18(4):899–910, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Benoit:2011:RAS

[BRSRC11] Anne Benoit, Veronika Rehn-Sonigo, Yves Robert, and Henri Casanova. Resource allocation strategies for constructive in-network stream processing. International Journal of Foundations of Computer Science (IJFCS), 22(3):621– 638, April 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bordihn:2007:PAL

[BRST07] Henning Bordihn, Bernd Reichel, Ralf Stiebe, and Bianca Truthe. Preface: Aspects in language and automata theory: Special issue dedicated to Jürgen Dassow. International Journal of Foundations of Computer Science (IJFCS), 18 (6):1137–1138, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Blin:2013:MMI

[BRSV13] Guillaume Blin, Romeo Rizzi, Florian Sikora, and Stephane Vialette. Minimum Mosaic inference of a set of recombinants. International Journal of Foundations of Computer Science (IJFCS), 24(1):51–??, January 2013. CODEN IFC-SEN. ISSN 0129-0541.

Brzozowski:2013:SMC

[Brz13]  Janusz Brzozowski.  In

search of most complex regular languages. International Journal of Foundations of Computer Science (IJFCS), 24(6):691–??, September 2013. CODEN IFCSEN. ISSN 0129-0541.

Blanchet-Sadri:1992:DDG

[BS92] F. Blanchet-Sadri. The dotdepth of a generating class of aperiodic monoids is computable. International Journal of Foundations of Computer Science (IJFCS), 3(4): 419–442, December 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

BarbosaDaSilva:2001:EPJ

[BS01] Fabricio Alves Barbosa Da Silva and Isaac D. Scherson. Efficient parallel job scheduling using gang service. International Journal of Foundations of Computer Science (IJFCS), 12 (3):265–284, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Blanchet-Sadri:2012:ACP

[BS12] F. Blanchet-Sadri. Algorithmic combinatorics on partial words. International Journal of Foundations of Computer Science (IJFCS), 23(6):1189–??, September 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Brzozowski:2015:LAS

[BS15] Janusz Brzozowski and Marek Szykuła. Large aperiodic semigroups. International Journal of Foundations of Computer Science (IJFCS), 26(7):913–??, November 2015. CODEN IFCSEN. ISSN 0129-0541.

Bonnet:2016:NEI

[BS16] Édouard Bonnet and Florian Sikora. A note on edge isoperimetric numbers and regular graphs. International Journal of Foundations of Computer Science (IJFCS), 27(6):771–??, September 2016. CODEN IFCSEN. ISSN 0129-0541.

Blanchet-Sadri:2008:RTN

[BSBZ08] F. Blanchet-Sadri, L. Bromberg, and K. Zipple. Remarks on two nonstandard versions of periodicity in words. International Journal of Foundations of Computer Science (IJFCS), 19(6):1439–1448, December 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Blanchet-Sadri:2022:DWL

[BSCH22] Francine Blanchet-Sadri, Kun Chen, and Kenneth Hawes. Dyck words, lattice paths, and Abelian borders. International Journal of Foundations of Computer Science (IJFCS), 33(3–4):203–226, April– June 2022. ISSN 0129-

0541. URL https://www.
worldscientific.com/doi/
10.1142/S0129054122410027.

Bansal:2003:MHM

[BSG03] S. Bansal, S. Sreekanth, and P. Gupta. M-heap: a modified heap data structure. International Journal of Foundations of Computer Science (IJFCS), 14(3):491– ??, June 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Blanchet-Sadri:2010:FWT

[BSOR10] F. Blanchet-Sadri, Taktin Oey, and Timothy D. Rankin. Fine and Wilf’s theorem for partial words with arbitrarily many weak periods. International Journal of Foundations of Computer Science (IJFCS), 21(5):705– 722, October 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bresolin:2012:BMB

[BSS12] Davide Bresolin, Pietro Sala, and Guido Sciavicco. On begins, meets and before. International Journal of Foundations of Computer Science (IJFCS), 23(3):559– ??, April 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bourgeois:2000:RTD

[BT00] A. G. Bourgeois and J. L. Trahan. Relating twodimensional reconfigurable meshes with optically pipelined buses. International Journal of Foundations of Computer Science (IJFCS), 11 (4):553–572, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bae:2007:ADM

[BT07] Sung Eun Bae and Tadao Takaoka. Algorithms for K-disjoint maximum subarrays. International Journal of Foundations of Computer Science (IJFCS), 18(2):319– 339, April 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Brzozowski:2013:CAR

[BT13] Janusz Brzozowski and Hellis Tamm. Complexity of atoms of regular languages. International Journal of Foundations of Computer Science (IJFCS), 24(7):1009–??, November 2013. CODEN IFCSEN. ISSN 0129-0541.

Bramas:2017:RBC

[BT17] Quentin Bramas and Sébastien Tixeuil. The random bit complexity of mobile robots scattering. International Journal of Foundations of Computer Science (IJFCS), 28(2):111– 133, February 2017. CO-DEN IFCSEN. ISSN 0129-0541.

Bacak-Turan:2013:NIT

[BTK13] Goksen Bacak-Turan and Alpay Kirlangic. Neighbor integrity of transformation graphs. International Journal of Foundations of Computer Science (IJFCS), 24 (3):303–??, April 2013. CO-DEN IFCSEN. ISSN 0129-0541.

Bacak-Turan:2017:NRD

[BTO17] Goksen Bacak-Turan and Ekrem Oz. Neighbor rupture degree of transformation graphs G^{xy−}. International Journal of Foundations of Computer Science (IJFCS), 28(4):335–??, June 2017. CODEN IFCSEN. ISSN 0129-0541.

Burderi:2012:FMM

[Bur12a] Fabio Burderi. Full monoids and maximal codes. International Journal of Foundations of Computer Science (IJFCS), 23(8):1677–??, December 2012. CODEN IFC-SEN. ISSN 0129-0541.

Burgin:2012:DUA

[Bur12b] Mark Burgin. Decidability and universality in the axiomatic theory of computability and algorithms. International Journal of Foundations of Computer Science (IJFCS), 23(7): 1465–??, November 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Banerjee:1998:DSC

[BV98a] Ayan Banerjee and Emmanouel (Manos) Varvarigos. A dynamic scheduling communication protocol and its analysis for hypercube networks. International Journal of Foundations of Computer Science (IJFCS), 9(1):39–??, March 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Burtschick:1998:LQL

[BV98b] Hans-Jörg Burtschick and H. Vollmer. Lindström quantifiers and leaf language definability. International Journal of Foundations of Computer Science (IJFCS), 9(3):277–??, September 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Belohlavek:2008:BAA

[BV08] Radim Bělohlávek and Vilem Vychodil. Basic algorithm for attribute implications and functional dependencies in graded setting. International Journal of Foundations of Computer Science (IJFCS), 19(2):297– 317, April 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bordihn:2020:DLS

[BV20] Henning Bordihn and György Vaszil. Deterministic Lindenmayer systems with dynamic control of parallelism. International Journal of Foundations of Computer Science (IJFCS), 31(1):37– 51, January 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120400031.

Berwanger:2018:CGA

[BvdB18] Dietmar Berwanger and Marie van den Bogaard. Consensus game acceptors and iterated transductions. International Journal of Foundations of Computer Science (IJFCS), 29(2):165– ??, February 2018. CODEN IFCSEN. ISSN 0129-0541.

Bonizzoni:2000:AMI

[BVM00] P. Bonizzoni, G. D. Vedova, and G. Mauri. Approximating the maximum isomorphic agreement subtree is hard. International Journal of Foundations of Computer Science (IJFCS), 11 (4):579–590, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Boykett:2014:RAN

[BW14] Tim Boykett and Gerhard Wendt. I_{2} radical in automata near rings. International Journal of Foundations of Computer Science (IJFCS), 25(5):585–??, August 2014. CODEN IFC-SEN. ISSN 0129-0541.

Berberler:2018:LVN

[BY18] Zeynep Nihan Berberler and Esin Yigit. Link vulnerability in networks. International Journal of Foundations of Computer Science (IJFCS), 29(3):447– 456, April 2018. CODEN IFCSEN. ISSN 0129-0541.

Berberler:2021:ABN

[BYİT21] Zeynep Nihan Berberler, Halil İbrahim Yildirim, Tolga İltüzer, and İzzet Tunç. Agglomeration-based node importance analysis in wheel-type networks. International Journal of Foundations of Computer Science (IJFCS), 32(03):269– 288, April 2021. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054121500210.

Baeza-Yates:1995:HOA

[BYP95] R. A. Baeza-Yates and P. V. Poblete. Higher-order analysis of 2-3 trees. International Journal of Foundations of Computer Science (IJFCS), 6(1):1–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bruda:2010:CHP

[BZ10] Stefan D. Bruda and Yuanqiao Zhang. Collapsing the hierarchy of parallel computational models. International Journal of Foundations of Computer Science (IJFCS), 21(3):441– 457, June 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Bienkowski:2013:HCM

[BZ13] Marcin Bienkowski and Paweł Zalewski. (1, 2)- Hamiltonian completion on a matching. International Journal of Foundations of Computer Science (IJFCS), 24(1):95–??, January 2013. CODEN IFCSEN. ISSN 0129-0541.

Ciftci:2018:EIN

[ÇA18] Canan Çiftçi and Aysun Aytaç. Exponential independence number of some graphs. International Journal of Foundations of Computer Science (IJFCS), 29(7):1151– 1164, November 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118500260.

Cai:1994:CJN

[Cai94] J.-Y. Cai. Computing Jordan Normal Forms exactly for commuting matrices in polynomial time. International Journal of Foundations of Computer Science (IJFCS), 5(3/4):??, 1994. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Calude:2005:P

[Cal05] Cristian S. Calude. Preface. International Journal of Foundations of Computer Science (IJFCS), 16(4):623– ??, August 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Calamoneri:2015:OJK

[Cal15] Tiziana Calamoneri. Optimal L(j, k)-edge-labeling of regular grids. International Journal of Foundations of Computer Science (IJFCS), 26(4):523–??, June 2015. CODEN IFCSEN. ISSN 0129-0541.

Campeanu:2014:DCE

[Câm14] Cezar Câmpeanu. Descriptional complexity in encoded Blum static complexity spaces. International Journal of Foundations of Computer Science (IJFCS), 25(7):917–??, November 2014. CODEN IFCSEN. ISSN 0129-0541.

Campeanu:2020:IAA

[Câm20] Cezar Câmpeanu. Implementations and applications of Automata 2018: Preface. International Journal of Foundations of Computer Science (IJFCS), 31(08): 979–982, December 2020. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120020037.

Caporaso:1996:STM

[Cap96] S. Caporaso. Safe Turing machines, Grzegorczyk classes and polytime. International Journal of Foundations of Computer Science (IJFCS), 7(3):241–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Carlet:2011:MVB

[Car11] Claude Carlet. More vectorial Boolean functions with unbounded nonlinearity profile. International Journal of Foundations of Computer Science (IJFCS), 22(6):1259–1269, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Castellani:1995:ODP

[Cas95] I. Castellani. Observing distribution in processes: Static and dynamic localities. International Journal of Foundations of Computer Science (IJFCS), 6 (4):353–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Casanova:2005:NMI

[Cas05] Henri Casanova. Network modeling issues for Grid application scheduling. International Journal of Foundations of Computer Science (IJFCS), 16(2):145– ??, April 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Christodoulakis:2009:EDC

[CB09] Manolis Christodoulakis and Gerhard Brey. Edit distance with combinations and splits and its applications in OCR name matching. International Journal of Foundations of Computer Science (IJFCS), 20 (6):1047–1068, December 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chiang:1998:TPS

[CC98] Wei-Kuo Chiang and Rong-Jaye Chen. Topological properties of the (n, k)-star graph. International Journal of Foundations of Computer Science (IJFCS), 9(2): 235–??, June 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Champarnaud:2005:ENA

[CC05] J.-M. Champarnaud and F. Coulon. Enumerating nondeterministic automata for a given language without constructing the canonical automaton. International Journal of Foundations of Computer Science (IJFCS), 16(6):1253–??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chai:2007:EIB

[CCD07] Zhenchuan Chai, Zhenfu Cao, and Xiaolei Dong. Efficient ID-based multireceiver threshold decryption. International Journal of Foundations of Computer Science (IJFCS), 18 (5):987–1004, October 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cantone:2008:SCP

[CCF08] Domenico Cantone, Salvatore Cristofaro, and Simone Faro. On some combinatorial problems concerning the harmonic structure of musical chord sequences. International Journal of Foundations of Computer Science (IJFCS), 19 (1):103–124, February 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cantone:2009:NEB

[CCF09] Domenico Cantone, Salvatore Cristofaro, and Simone Faro. New efficient bit-parallel algorithms for the (δ, α)-matching problem with applications in music information retrieval. International Journal of Foundations of Computer Science (IJFCS), 20(6):1087–1108, December 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Campanelli:2012:PMS

[CCFG12] Matteo Campanelli, Domenico Cantone, Simone Faro, and Emanuele Giaquinta. Pattern matching with swaps in practice. International Journal of Foundations of Computer Science (IJFCS), 23(2):323–342, February 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cienciala:2007:PDS

[CCFS07] Luděk Cienciala, Lucie Ciencialová, Pierluigi Frisco, and Petr Sosík. On the power of deterministic and sequential communicating P systems. International Journal of Foundations of Computer Science (IJFCS), 18(2):415–431, April 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Czumaj:2011:AAB

[CCG^{+}11] Artur Czumaj, Jurek Czyzowicz, Leszek Ga̧sieniec, Jesper Jansson, Andrzej Lingas, and Pawel Zylinski. Approximation algorithms for buy-at-bulk geometric network design. International Journal of Foundations of Computer Science (IJFCS), 22(8):1949–1969, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Christou:2012:IAA

[CCI12] Michalis Christou, Maxime Crochemore, and Costas S. Iliopoulos. Identifying all Abelian periods of a string in quadratic time and relevant problems. International Journal of Foundations of Computer Science (IJFCS), 23(6):1371– ??, September 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Calvin:1997:MOC

[CCM97] C. Calvin, L. Colombet, and P. Michallon. Methods to overlap communications in parallel numerical algorithms. International Journal of Foundations of Computer Science (IJFCS), 8(2): 211–??, June 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Calude:2011:OBQ

[CCM11] Cristian S. Calude, Matteo Cavaliere, and Radu Mardare. An observerbased de-quantisation of Deutsch’s Algorithm. International Journal of Foundations of Computer Science (IJFCS), 22(1):191– 201, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Champarnaud:2005:BFD

[CCP05] J.-M. Champarnaud, F. Coulon, and T. Paranthoën. Brute force determinization of NFAs by means of state covers. International Journal of Foundations of Computer Science (IJFCS), 16(3):441– ??, June 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chauve:2018:CGA

[CCP18] Cedric Chauve, Julien Courtiel, and Yann Ponty. Counting, generating, analyzing and sampling tree alignments. International Journal of Foundations of Computer Science (IJFCS), 29(5):741–767, August 2018. ISSN 0129-0541. URL

                                        https://www.worldscientific.
                                        com/doi/10.1142/S0129054118420030.

Cai:2004:HAM

[CCPS04] Jin-Yi Cai, Denis Charles, A. Pavan, and Samik Sengupta. On higher Arthur– Merlin classes. International Journal of Foundations of Computer Science (IJFCS), 15(1):3–??, February 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cheng:2024:LCL

[CCQ24] Xiangdong Cheng, Xiwang Cao, , and Liqin Qian. Linear codes and linear complementary pairs of codes over a non-chain ring. International Journal of Foundations of Computer Science (IJFCS), 35(03):297– 311, April 2024. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412350003X.

Cherubini:1990:BDG

[CCR^{+}90] Alessandra Cherubini, Claudio Citrini, Stefano Crespi Reghizzi, et al. Breadth and depth grammars and deque automata. International Journal of Foundations of Computer Science (IJFCS), 1(3):219–??, September 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Clementi:1995:OSP

[CD95] A. Clementi and M. Di Ianni. Optimum schedule problems in store and forward networks. International Journal of Foundations of Computer Science (IJFCS), 6(2):155–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cheng:2002:SI

[CD02] S.-W. Cheng and T. Dey. Special issue. International Journal of Foundations of Computer Science (IJFCS), 13(2):161–??, 2002. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cavaliere:2006:FRT

[CD06] Matteo Cavaliere and Vincenzo Deufemia. Further results on time-free P systems. International Journal of Foundations of Computer Science (IJFCS), 17(1):69– ??, February 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cvetkovic:2009:MIN

[CD09] Dragoš Cvetković and Tatjana Davidović. Multiprocessor interconnection networks with small tightness. International Journal of Foundations of Computer Science (IJFCS), 20(5):941– 963, October 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Calude:2015:AAN

[CD15] Cristian S. Calude and Damien Desfontaines. Anytime algorithms for nonending computations. International Journal of Foundations of Computer Science (IJFCS), 26(4):465–??, June 2015. CODEN IFCSEN. ISSN 0129-0541.

Chen:2020:SAS

[CD20] Ke Chen and Adrian Dumitrescu. Selection algorithms with small groups. International Journal of Foundations of Computer Science (IJFCS), 31(03): 355–369, April 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500136.

Carpi:2021:CEA

[CD21] Arturo Carpi and Flavio D’Alessandro. On the commutative equivalence of algebraic formal series and languages. International Journal of Foundations of Computer Science (IJFCS), 32(04):341–367, June 2021. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500192.

Cambie:2023:EBP

[CdBD23] Stijn Cambie, Michiel de Bondt, and Henk Don. Extremal binary PFAs with small number of states. International Journal of Foundations of Computer Science (IJFCS), 34(2–3):85–115, February–April 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122440038.

Chimani:2019:ATC

[CDFK19] Markus Chimani, Giuseppe Di Battista, Fabrizio Frati, and Karsten Klein. Advances on testing c-planarity of embedded flat clustered graphs. International Journal of Foundations of Computer Science (IJFCS), 30 (2):197–230, February 2019. CODEN IFCSEN. ISSN 0129-0541. Champarnaud:2009:EAT

[CDJ09] Jean-Marc Champarnaud, Jean Philippe Dubernard, and Hadrien Jeanne. An efficient algorithm to test whether a binary and prolongeable regular language is geometrical. International Journal of Foundations of Computer Science (IJFCS), 20(4):763–774, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Carpi:2004:RFU

[CdL04] Arturo Carpi and Aldo de Luca. Repetitions, fullness, and uniformity in twodimensional words. International Journal of Foundations of Computer Science (IJFCS), 15(2):355– ??, April 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cinque:2005:SLC

[CDLW05] Luigi Cinque, Sergio De Agostino, Franco Liberati, and Bart Westgeest. A simple lossless compression heuristic for grey scale images. International Journal of Foundations of Computer Science (IJFCS), 16(6): 1111–??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chatterjee:2013:CC

[CDM13] Krishnendu Chatterjee, Luca De Alfaro, and Rupak Majumdar. The complexity of coverage. International Journal of Foundations of Computer Science (IJFCS), 24(2):165–??, February 2013. CODEN IFCSEN. ISSN 0129-0541.

Carrier:2011:AOD

[CDPR11] Fabienne Carrier, Stéphane Devismes, Franck Petit, and Yvan Rivierre. Asymptotically optimal deterministic rendezvous. International Journal of Foundations of Computer Science (IJFCS), 22(5):1143– 1159, August 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Caron:2016:SSP

[CDPT16] Eddy Caron, Ajoy K. Datta, Franck Petit, and Cédric Tedeschi. Self-stabilizing prefix tree based overlay networks. International Journal of Foundations of Computer Science (IJFCS), 27(5):607–??, August 2016. CODEN IFCSEN. ISSN 0129-0541.

Cao:2021:KMB

[CDX21] Mengyue Cao, Tongtong Ding, and Min Xu. The (n, k)-modified-bubble-sort graph: a generalized modifiedbubble-sort graph. International Journal of Foundations of Computer Science (IJFCS), 32(07):849– 860, November 2021. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054121500258.

Cappello:1998:PLB

[CE98] Peter Cappello and Ömer Egecioğlu. Processor lower bound formulas for array computations and parametric Diophantine systems. International Journal of Foundations of Computer Science (IJFCS), 9(4):351–??, December 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cerny:2008:SSW

[Čer08] Anton Černý. On subword symmetry of words. International Journal of Foundations of Computer Science (IJFCS), 19(1):243– 250, February 2008. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cevik:2020:PCC

[Çev20] Ahmet Çevik. Palindromic characteristic of committed graphs and some model theoretic properties. International Journal of Foundations of Computer Science (IJFCS), 31(04):483– 498, June 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500203.

Cantone:2006:SEB

[CF06] Domenico Cantone and Simone Faro. A space efficient bit-parallel algorithm for the multiple string matching problem. International Journal of Foundations of Computer Science (IJFCS), 17(6):1235–1251, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cantone:2012:ABM

[CFG12] Domenico Cantone, Simone Faro, and Emanuele Giaquinta. Adapting Boyer– Moore-like algorithms for searching Huffman encoded texts. International Journal of Foundations of Computer Science (IJFCS), 23 (2):343–356, February 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Clarke:2003:ACG

[CFH^{+}03] Edmund Clarke, Ansgar Fehnker, Zhi Han, Bruce Krogh, Joël Ouaknine, Olaf Stursberg, and Michael Theobald. Abstraction and counterexample-guided refinement in model checking of hybrid systems. International Journal of Foundations of Computer Science (IJFCS), 14(4):583– ??, August 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Crochemore:2010:NOP

[CFIJ10] Maxime Crochemore, Szilárd Zsolt Fazekas, Costas S. Iliopoulos, and Inuka Jayasekera. Number of occurrences of powers in strings. International Journal of Foundations of Computer Science (IJFCS), 21(4):535– 547, August 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cadilhac:2012:BPA

[CFM12] Michaël Cadilhac, Alain Finkel, and Pierre Mckenzie. Bounded Parikh automata. International Journal of Foundations of Computer Science (IJFCS), 23(8):

1691–??, December 2012. CODEN IFCSEN. ISSN 0129-0541.

Cadilhac:2013:UCA

[CFM13] Michaël Cadilhac, Alain Finkel, and Pierre Mckenzie. Unambiguous constrained automata. International Journal of Foundations of Computer Science (IJFCS), 24(7):1099– ??, November 2013. CO-DEN IFCSEN. ISSN 0129-0541.

Cappello:2005:AMH

[CFMR05] Franck Cappello, Pierre Fraigniaud, Bernard Mans, and Arnold L. Rosenberg. An algorithmic model for heterogeneous hyperclusters: Rationale and experience. International Journal of Foundations of Computer Science (IJFCS), 16(2):195–??, April 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Casteigts:2015:SFF

[CFMS15] Arnaud Casteigts, Paola Flocchini, Bernard Mans, and Nicola Santoro. Shortest, fastest, and foremost broadcast in dynamic networks. International Journal of Foundations of Computer Science (IJFCS), 26 (4):499–??, June 2015. CO-DEN IFCSEN. ISSN 0129-0541.

Czyzowicz:2003:LTP

[CFPR03] Jurek Czyzowicz, Wojciech Fraczak, Andrzej Pelc, and Wojciech Rytter. Lineartime prime decomposition of regular prefix codes. International Journal of Foundations of Computer Science (IJFCS), 14(6):1019– ??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cellier:2008:PAE

[CFRD08] Peggy Cellier, Sébastien Ferré, Olivier Ridoux, and Mireille Ducassé. A parameterized algorithm to explore formal contexts with a taxonomy. International Journal of Foundations of Computer Science (IJFCS), 19(2):319–343, April 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chigahara:2016:OWJ

[CFY16] Hiroyuki Chigahara, Szilárd Zsolt Fazekas, and Akihiro Yamamura. One-way jumping finite automata. International Journal of Foundations of Computer Science (IJFCS), 27(3):391–??, April 2016. CODEN IFC-SEN. ISSN 0129-0541.

Ciobanu:2006:MMA

[CG06] Gabriel Ciobanu and Mihai Gontineac. Mealy multiset automata. International Journal of Foundations of Computer Science (IJFCS), 17(1):111–??, February 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ciobanu:2009:EM

[CG09] Gabriel Ciobanu and Mihai Gontineac. Encodings of multisets. International Journal of Foundations of Computer Science (IJFCS), 20(3):381–393, June 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Champarnaud:2005:CTF

[CGH05] Jean-Marc Champarnaud, Franck Guingne, and Georges Hansel. Cover transducers for functions with finite domain. International Journal of Foundations of Computer Science (IJFCS), 16(5):851– ??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Czyzowicz:2021:SLB

[CGK^{+}21] Jurek Czyzowicz, Konstantinos Georgiou, Evangelos Kranakis, Danny Krizanc, Lata Narayanan, Jaroslav Opatrny, and Sunil Shende. Search on a line by Byzantine robots. International Journal of Foundations of Computer Science (IJFCS), 32(04):369– 387, June 2021. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054121500209.

Champarnaud:2008:AJA

[CGKN08] Jean-Marc Champarnaud, Franck Guingne, André Kempe, and Florent Nicart. Algorithms for the join and auto-intersection of multitape weighted finite-state machines. International Journal of Foundations of Computer Science (IJFCS), 19(2):453–476, April 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cui:2011:SCT

[CGKY11] Bo Cui, Yuan Gao, Lila Kari, and Sheng Yu. State complexity of two combined operations: Catenationunion and catenation-intersection. International Journal of Foundations of Computer Science (IJFCS), 22(8): 1797–1812, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cui:2012:SCT

[CGKY12] Bo Cui, Yuan Gao, Lila Kari, and Sheng Yu. State complexity of two combined operations: Catenation-star and catenation-reversal. International Journal of Foundations of Computer Science (IJFCS), 23(1):51–66, January 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Crochemore:2012:LCS

[CGL12] Maxime Crochemore, Laura Giambruno, and Alessio Langiu. On-line construction of a small automaton for a finite set of words. International Journal of Foundations of Computer Science (IJFCS), 23 (2):281–301, February 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Carioni:2013:ATM

[CGR13] Alessandro Carioni, Silvio Ghilardi, and Silvio Ranise. Automated termination in model-checking modulo theories. International Journal of Foundations of Computer Science (IJFCS), 24(2):211– ??, February 2013. CODEN IFCSEN. ISSN 0129-0541.

Cao:2015:TBF

[CH15] Xiwang Cao and Lei Hu.

Two Boolean functions with five-valued Walsh spectra and high nonlinearity. International Journal of Foundations of Computer Science (IJFCS), 26(5):537–??, August 2015. CODEN IFC-SEN. ISSN 0129-0541.

Cheng:1992:RAD

[CHA^{+}92] Y. Cheng, F. K. Hwang, I. F. Akyildiz, et al. Routing algorithms for double loop networks. International Journal of Foundations of Computer Science (IJFCS), 3(3):323–332, September 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Changizi:1997:LNI

[Cha97] Mark A. Changizi. Learning with natural imprecision. International Journal of Foundations of Computer Science (IJFCS), 8(4):409– ??, December 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Champarnaud:2002:ETI

[Cha02] J.-M. Champarnaud. Evaluation of three implicit structures to implement nondeterministic automata from regular expressions. International Journal of Foundations of Computer Science (IJFCS), 13(1):99–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Champarnaud:2003:P

[Cha03] Jean-Marc Champarnaud. Preface. International Journal of Foundations of Computer Science (IJFCS), 14 (6):953–??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chen:2022:CSI

[Che22a]  Yen Hung Chen.  The

clustered selected-internal Steiner tree problem. International Journal of Foundations of Computer Science (IJFCS), 33(01):55– 66, January 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500362.

Cheng:2022:ECS

[Che22b] Dongqin Cheng. Extra connectivity and structure connectivity of 2-dimensional torus networks. International Journal of Foundations of Computer Science (IJFCS), 33(02):155– 173, February 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500058.

Czeizler:2007:ISW

[CHKL07] Elena Czeizler, Štěpán Holub, Juhani Karhumäki, and Markku Laine. Intricacies of simple word equations: an example. International Journal of Foundations of Computer Science (IJFCS), 18(6):1167–1175, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chen:2009:GAC

[CHWX09] Danny Z. Chen, Mark A.

Healy, Chao Wang, and Bin Xu. Geometric algorithms for the constrained 1-D Kmeans clustering problems and IMRT applications. International Journal of Foundations of Computer Science (IJFCS), 20(2):361– 377, April 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chiu:2014:AHA

[CHYT14] Chiao-Wei Chiu, Kuo-Si Huang, Chang-Biau Yang, and Chiou-Ting Tseng. An adaptive heuristic algorithm with the probabilistic safety vector for fault-tolerant routing on the (n, k)-star graph. International Journal of Foundations of Computer Science (IJFCS), 25 (6):723–??, September 2014. CODEN IFCSEN. ISSN 0129-0541.

Cleophas:2006:TRA

[CHZ06] Loek Cleophas, Kees Hemerik, and Gerard Zwaan. Two related algorithms for root-to-frontier tree pattern matching. International Journal of Foundations of Computer Science (IJFCS), 17 (6):1253–1272, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ciglaric:2004:CND

[Cig04] Mojca Ciglarič. Content networks: Distributed routing decisions in presence of repeated queries. International Journal of Foundations of Computer Science (IJFCS), 15(3):555– ??, June 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Christodoulakis:2008:IRM

[CIRS08] Manolis Christodoulakis, Costas S. Iliopoulos, M. Sohel Rahman, and William F. Smyth. Identifying rhythms in musical texts. International Journal of Foundations of Computer Science (IJFCS), 19(1):37–51, February 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cartigny:2003:RRS

[CIS03] Julien Cartigny, François Ingelrest, and David Simplot. RNG relay subset flooding protocols in mobile adhoc networks. International Journal of Foundations of Computer Science (IJFCS), 14(2):253–??, April 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chin:2016:A

[CIS16] Francis Chin, Oscar H.

Ibarra, and Sartaj K. Sahni. Announcement. International Journal of Foundations of Computer Science (IJFCS), 27(8):895– 896, December 2016. CO-DEN IFCSEN. ISSN 0129-0541.

Crochemore:2007:SFO

[CISH07] Maxime Crochemore, Lucian Ilie, and Emine Seid-Hilmi. The structure of factor oracles. International Journal of Foundations of Computer Science (IJFCS), 18(4):781–797, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chin:2012:SY

[CISS12] F. Chin, O. Ibarra, S. Sahni, and A. Salomaa. Sheng yu. International Journal of Foundations of Computer Science (IJFCS), 23 (2):243–246, February 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Calude:2001:CMP

[CIY01] Christian S. Calude, Hajime Ishihara, and Takeshi Yamaguchi. Coding with minimal programs. International Journal of Foundations of Computer Science (IJFCS), 12(4):479–489 (or 479–490??), 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Catalano:2020:SPF

[CJ20] Costanza Catalano and Raphaël M. Jungers. The synchronizing probability function for primitive sets of matrices. International Journal of Foundations of Computer Science (IJFCS), 31(06):777– 803, September 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120410051.

Case:1992:LLP

[CJS92] J. Case, S. Jain, and A. Sharma. On learning limiting programs. International Journal of Foundations of Computer Science (IJFCS), 3(1):93–??, March 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Czeizler:2007:NPS

[CK07] Elena Czeizler and Juhani Karhumäki. On nonperiodic solutions of independent systems of word equations over three unknowns. International Journal of Foundations of Computer Science (IJFCS), 18 (4):873–897, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Campeanu:2008:SCS

[CK08a] Cezar Câmpeanu and Stavros Konstantinidis. State complexity of the subword closure operation with applications to DNA coding. International Journal of Foundations of Computer Science (IJFCS), 19(5):1099–1112, October 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Conley:2008:UAM

[CK08b] Ehud S. Conley and Shmuel T. Klein. Using alignment for multilingual text compression. International Journal of Foundations of Computer Science (IJFCS), 19 (1):89–101, February 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cassaigne:2016:ACF

[CK16] Julien Cassaigne and Idrissa Kaboré. Abelian complexity and frequencies of letters in infinite words. International Journal of Foundations of Computer Science (IJFCS), 27(5):631–??, August 2016. CODEN IFC-SEN. ISSN 0129-0541.

Chung:2018:ERN

[CK18] Heewon Chung and Myungsun Kim. Encoding of rational numbers and their homomorphic computations for FHE-based applications. International Journal of Foundations of Computer Science (IJFCS), 29(6):1023–1044, September 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118500193.

Culik:2002:NSA

[CKK02] K. Culik II, J. Karhumäki, and J. Kari. A note on synchronized automata and road coloring problem. International Journal of Foundations of Computer Science (IJFCS), 13(3):459–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Christensen:2015:SHT

[CKL15] Jan Christensen, Anders Nicolai Knudsen, and Kim S. Larsen. Soccer is harder than football. International Journal of Foundations of Computer Science (IJFCS), 26(4):477–??, June 2015. CODEN IFCSEN. ISSN 0129-0541.

Coetser:2009:REH

[CKW09] Wikus Coetser, Derrick G.

Kourie, and Bruce W. Watson. On regular expression hashing to reduce FA size. International Journal of Foundations of Computer Science (IJFCS), 20 (6):1069–1086, December 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chang:2017:SER

[CKZ17] Zuling Chang, Pinhui Ke, and Yongcheng Zhao. Some enumeration results on binary 2^{n}-periodic sequences. International Journal of Foundations of Computer Science (IJFCS), 28(2):171– 184, February 2017. CO-DEN IFCSEN. ISSN 0129-0541.

Chen:1998:SRN

[CL98] Guihai Chen and Francis C. M. Lau. Shuffle-ring: a new constant-degree network. International Journal of Foundations of Computer Science (IJFCS), 9(1): 77–??, March 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chen:2003:ZAC

[CL03] Yuanzhu Peter Chen and Arthur L. Liestman. A zonal algorithm for clustering an hoc networks. International Journal of Foundations of Computer Science (IJFCS), 14(2):305–??, April 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cheng:2007:FRC

[CL07a] Eddie Cheng and László Lipták. Fault resiliency of Cayley graphs generated by transpositions. International Journal of Foundations of Computer Science (IJFCS), 18(5):1005–1022, October 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Couceiro:2007:EVI

[CL07b] Miguel Couceiro and Erkko Lehtonen. On the effect of variable identification on the essential arity of functions on finite sets. International Journal of Foundations of Computer Science (IJFCS), 18(5):975– 986, October 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chang:2010:ETF

[CL10] Ching-Lueh Chang and Yuh-Dauh Lyuu. Efficient testing of forecasts. International Journal of Foundations of Computer Science (IJFCS), 21(1):61–72, February 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Carnino:2014:FUA

[CL14] Vincent Carnino and Sylvain Lombardy. Factorizations and universal automaton of Omega languages. International Journal of Foundations of Computer Science (IJFCS), 25(8):1111– ??, December 2014. CODEN IFCSEN. ISSN 0129-0541.

Carnino:2015:DUW

[CL15] Vincent Carnino and Sylvain Lombardy. On determinism and unambiguity of weighted two-way automata. International Journal of Foundations of Computer Science (IJFCS), 26(8):1127–??, December 2015. CODEN IFCSEN. ISSN 0129-0541.

Cheng:2008:MPA

[CLLL08] Eddie Cheng, Linda Lesniak, Marc J. Lipman, and László Lipták. Matching preclusion for alternating group graphs and their generalizations. International Journal of Foundations of Computer Science (IJFCS), 19(6):1413–1437, December 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Caron:2016:SCC

[CLMP16] Pascal Caron, Jean-Gabriel Luque, Ludovic Mignot, and Bruno Patrou. State complexity of catenation combined with a Boolean operation: A unified approach. International Journal of Foundations of Computer Science (IJFCS), 27(6):675– ??, September 2016. CO-DEN IFCSEN. ISSN 0129-0541.

Champarnaud:2004:RWE

[CLOZ04] Jean-Marc Champarnaud, Éric Laugerotte, Faissal Ouardi, and Djelloul Ziadi. From regular weighted expressions to finite automata. International Journal of Foundations of Computer Science (IJFCS), 15(5):687– ??, October 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Charlier:2019:P

[CLR19] Émilie Charlier, Julien Leroy, and Michel Rigo. Preface. International Journal of Foundations of Computer Science (IJFCS), 30 (1):1–4, January 2019. CO-DEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119020015.

Chang:2009:TEB

[CLT09] Ching-Lueh Chang, Yuh-Dauh Lyuu, and Yen-Wu Ti. Testing embeddability between metric spaces. International Journal of Foundations of Computer Science (IJFCS), 20(2):313– 329, April 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chen:2014:TRA

[CLT14] Jheng-Cheng Chen, Chia-Jui Lai, and Chang-Hsiung Tsai. A three-round adaptive diagnostic algorithm in a distributed system modeled by dual-cubes. International Journal of Foundations of Computer Science (IJFCS), 25(2):125–??, February 2014. CODEN IFCSEN. ISSN 0129-0541.

Cantin:2009:CCH

[CLW09] François Cantin, Axel Legay, and Pierre Wolper. Computing convex hulls by automata iteration. International Journal of Foundations of Computer Science (IJFCS), 20(4):647– 667, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Campadelli:1992:LCL

[CM92] P. Campadelli and A. Morpurgo. Learning classes of linearly separable Boolean functions from positive examples. International Journal of Foundations of Computer Science (IJFCS), 3(1): 41–54, March 1992. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chatterjee:2012:DAG

[CM12] Krishnendu Chatterjee and Rupak Majumdar. Discounting and averaging in games across time scales. International Journal of Foundations of Computer Science (IJFCS), 23(3):609– ??, April 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Casasnovas:2004:AMC

[CMMR04] Jaume Casasnovas, Joe Miró, Manuel Moyà, and Francesc Rosselló. An approach to membrane computing under inexactitude. International Journal of Foundations of Computer Science (IJFCS), 15(6):841– ??, December 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cortes:2007:DEP

[CMR07] Corinna Cortes, Mehryar Mohri, and Ashish Rastogi. L_{p} distance and equivalence of probabilistic automata. International Journal of Foundations of Computer Science (IJFCS), 18 (4):761–779, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cortes:2008:CRE

[CMRR08] Corinna Cortes, Mehryar Mohri, Ashish Rastogi, and Michael Riley. On the computation of the relative entropy of probabilistic automata. International Journal of Foundations of Computer Science (IJFCS), 19 (1):219–242, February 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Couceiro:2019:CVS

[CMWZ19] Miguel Couceiro, Miklós Maróti, Tamás Waldhauser, and László Zádori. Computing version spaces in the qualitative approach to multicriteria decision aid. International Journal of Foundations of Computer Science (IJFCS), 30(2):333– 353, February 2019. ISSN 0129-0541.

Costa:2022:OGB

[CNT22] José Carlos Costa, Conceição Nogueira, and Maria Lurdes Teixeira. The overlap gap between left-infinite and right-infinite words. International Journal of Foundations of Computer Science (IJFCS), 33(01):45– 53, January 2022. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054121500350.

Comon:1990:SSO

[Com90]  Hubert Comon.  Solv-

ing symbolic ordering constraints. International Journal of Foundations of Computer Science (IJFCS), 1 (4):387–??, December 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Coons:2017:RSJ

[Coo17]  Michael Coons.  Regu-

lar sequences and the joint spectral radius. International Journal of Foundations of Computer Science (IJFCS), 28(2):135– 140, February 2017. CO-DEN IFCSEN. ISSN 0129-0541.

Costantini:1990:SMP

[Cos90] Stefania Costantini. Semantics of a metalogic programming language. International Journal of Foundations of Computer Science (IJFCS), 1(3):233–??, September 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chaturvedi:2012:LVO

[COT12] Namit Chaturvedi, Jörg Olschewski, and Wolfgang Thomas. Languages versus ω-languages in regular infinite games. International Journal of Foundations of Computer Science (IJFCS), 23(5):985–??, August 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Calamoneri:1999:OLT

[CP99] T. Calamoneri and R. Petreschi. Optimal layout of trivalent Cayley interconnection networks. International Journal of Foundations of Computer Science (IJFCS), 10(3):277– 288, 1999. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Campeanu:2003:CNM

[CP03] Cezar Câmpeanu and Andrei Păun. Counting the number of minimal DFCA obtained by merging states. International Journal of Foundations of Computer Science (IJFCS), 14(6):995– ??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cardelli:2006:URM

[CP06] Luca Cardelli and Gheorghe Păun. An universality result for a (mem)brane calculus based on mate/drip operations. International Journal of Foundations of Computer Science (IJFCS), 17(1):49– ??, February 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Caissy:2016:EFH

[CP16] David Caissy and Andrzej Pelc. Exploration of faulty Hamiltonian graphs. International Journal of Foundations of Computer Science (IJFCS), 27(7):809–??, November 2016. CODEN IFCSEN. ISSN 0129-0541.

Campeanu:2019:P

[CP19] Cezar Câmpeanu and Giovanni Pighizzini. Preface. International Journal of Foundations of Computer Science (IJFCS), 30 (6–7):827–829, September– November 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119020039.

Chung:1999:PMM

[CPC99] Y. Chung, K. Park, and Y. Cho. Parallel maximum matching algorithms in interval graphs. International Journal of Foundations of Computer Science (IJFCS), 10(1):47–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ceterchi:2006:SCP

[CPJ06] Rodica Ceterchi and Mario J.

Pérez-Jiménez. On simulating a class of parallel architectures. International Journal of Foundations of Computer Science (IJFCS), 17(1):91–??, February 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Campeanu:2002:EAC

[CPY02] C. Campeanu, A. Păun, and S. Yu. An efficient algorithm for constructing minimal cover automata for finite languages. International Journal of Foundations of Computer Science (IJFCS), 13(1):83–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cordasco:2014:SSP

[CR14] Gennaro Cordasco and Arnold L. Rosenberg. On scheduling series–parallel DAGs to maximize area. International Journal of Foundations of Computer Science (IJFCS), 25(5):597–??, August 2014. CODEN IFC-SEN. ISSN 0129-0541.

Corson:2015:ACR

[CR15] Jon M. Corson and Lance L. Ross. Automata with counters that recognize word problems of free products. International Journal of Foundations of Computer Science (IJFCS), 26(1):79– ??, January 2015. CODEN IFCSEN. ISSN 0129-0541.

Charlier:2012:EDP

[CRS12] Émilie Charlier, Narad Rampersad, and Jeffrey Shallit. Enumeration and decidable properties of automatic sequences. International Journal of Foundations of Computer Science (IJFCS), 23(5):1035– ??, August 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cassaigne:2011:AAP

[CRSZ11] Julien Cassaigne, Gwénaël Richomme, Kalle Saari, and Luca Q. Zamboni. Avoiding Abelian powers in binary words with bounded Abelian complexity. International Journal of Foundations of Computer Science (IJFCS), 22(4):905– 920, June 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Crescenzi:1993:AMD

[CS93] P. Crescenzi and R. Silvestri. Average measure, descriptive complexity and approximation of maximization problems. International Journal of Foundations of Computer Science (IJFCS), 4(1):15–30, March 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chen:1996:OOR

[CS96] L. Chen and R. Schott. Optimal operations on redblack trees. International Journal of Foundations of Computer Science (IJFCS), 7(3):227–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cho:1999:MDE

[CS99] S. Cho and S. Sahni. Mergeable double-ended priority queues. International Journal of Foundations of Computer Science (IJFCS), 10 (1):1–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cho:2000:NWB

[CS00a] S. Cho and S. Sahni. A new weight balanced binary search tree. International Journal of Foundations of Computer Science (IJFCS), 11(3):485–514, 2000. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cho:2000:PRP

[CS00b] S. Cho and S. Sahni. Part 2 (regular papers). International Journal of Foundations of Computer Science (IJFCS), 11(3):485–??, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cheng:2002:RP

[CS02] E. Y. C. Cheng and S. Sahni. Regular papers. International Journal of Foundations of Computer Science (IJFCS), 13 (3):405–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Charlier:2018:PNB

[CS18] Émilie Charlier and Wolfgang Steiner. Permutations and negative beta-shifts. International Journal of Foundations of Computer Science (IJFCS), 29(5):721– 740, August 2018. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054118420029.

Chern:2020:NSP

[CSAT20] Zi Jing Chern, K. G. Subramanian, Azhana Ahmad, and Wen Chean Teh. A new study of Parikh matrices restricted to terms. International Journal of Foundations of Computer Science (IJFCS), 31(05):621– 638, August 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500306.

Camargo:2021:RKC

[CSN21] Priscila P. Camargo, Uéverton S.

Souza, and Julliano R. Nascimento. Remarks on k-clique, k-independent set and 2-contamination in complementary prisms. International Journal of Foundations of Computer Science (IJFCS), 32(01):37– 52, January 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500027.

Calude:2012:SSH

[CSR12] Cristian S. Calude, Kai Salomaa, and Tania K. Roblot. State-size hierarchy for finite-state complexity. International Journal of Foundations of Computer Science (IJFCS), 23(1):37– 50, January 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chen:2017:SIE

[CST^{+}17] Chin-Ling Chen, Jungpil Shin, Yu-Ting Tsai, Aniello Castiglione, and Francesco Palmieri. Securing information exchange in VANETs by using pairing-based cryptography. International Journal of Foundations of Computer Science (IJFCS), 28(6):781–??, September 2017. CODEN IFCSEN. ISSN 0129-0541.

Campeanu:2002:SDR

[CSV02] C. Câmpeanu, K. Salomaa, and S. Vágvölgyi. Shuffle decompositions of regular languages. International Journal of Foundations of Computer Science (IJFCS), 13(6):799–??, December 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Campeanu:2003:FSP

[CSY03] Cezar Câmpeanu, Kai Salomaa, and Sheng Yu. A formal study of practical regular expressions. International Journal of Foundations of Computer Science (IJFCS), 14(6):1007– ??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Czyba:2018:FAI

[CTS18] Christopher Czyba, Wolfgang Thomas, and Christopher Spinrath. Finite automata over infinite alphabets: Two models with transitions for local change. International Journal of Foundations of Computer Science (IJFCS), 29(2):213–??, February 2018. CODEN IFCSEN. ISSN 0129-0541.

Chantrapornchai:2001:REA

[CTZ01] Chantana Chantrapornchai, Sissades Tongsima, and Albert Zomaya. Resource estimation algorithm under impreciseness using inclusion scheduling. International Journal of Foundations of Computer Science (IJFCS), 12(5):581– 598, October 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Csuhaj-Varju:2013:PDA

[CV13] Erzsébet Csuhaj-Varjú. P and DP automata: unconventional versus classical automata. International Journal of Foundations of Computer Science (IJFCS), 24 (7):995–??, November 2013. CODEN IFCSEN. ISSN 0129-0541.

Cao:2014:SEC

[CV14] Xiang-Lan Cao and Elkin Vumar. Super edge connectivity of Kronecker products of graphs. International Journal of Foundations of Computer Science (IJFCS), 25(1):59–??, January 2014. CODEN IFC-SEN. ISSN 0129-0541.

Csuhaj-Varju:2010:VCB

[CVDV10] Erzsébet Csuhaj-Varjú, Jürgen Dassow, and György Vaszil.

Variants of competencebased derivations in CD grammar systems. International Journal of Foundations of Computer Science (IJFCS), 21(4):549– 569, August 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Csuhaj-Varju:2022:P

[CVDV22] Erzsébet Csuhaj-Varjú, Pál Dömösi, and György Vaszil. Preface. International Journal of Foundations of Computer Science (IJFCS), 33(3–4):175–178, April– June 2022. CODEN IFCSEN. ISSN 0129-

0541. URL https://www.
worldscientific.com/doi/
10.1142/S0129054122020014.

Csuhaj-Varju:2010:P

[CVÉ10] Erzsébet Csuhaj-Varjú and Zoltán Ésik. Preface. International Journal of Foundations of Computer Science (IJFCS), 21(5):687– 688, October 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Csuhaj-Varju:2020:PSI

[CVM20] Erzsébet Csuhaj-Varjú and Florin Manea. Preface — special issue: A collection of papers in honour of the 60th birthday of Victor Mitrana. International Journal of Foundations of Computer Science (IJFCS), 31(1):1–6, January 2020. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120020013.

Csuhaj-Varju:2000:PCP

[CVMVMV00] E. Csuhaj-Varjú, C. Martín-Vide, V. Mitrana, and G. Vaszil. Parallel communicating pushdown automata systems. International Journal of Foundations of Computer Science (IJFCS), 11 (4):633–650, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Csuhaj-Varju:2011:PGS

[CVOV11] Erzsébet Csuhaj-Varjú, Marion Oswald, and György Vaszil. PC grammar systems with clusters of components. International Journal of Foundations of Computer Science (IJFCS), 22(1):203– 212, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Csuhaj-Varju:2008:TLS

[CVPV08] Erzsébet Csuhaj-Varjú, Gheorghe Păun, and György Vaszil. Tissue-like P systems with dynamically emerging requests. International Journal of Foundations of Computer Science (IJFCS), 19(3):729– 745, June 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Csuhaj-Varju:2008:P

[CVV08] Erzsébet Csuhaj-Varjú and György Vaszil. Preface. International Journal of Foundations of Computer Science (IJFCS), 19(5):1181– 1182, October 2008. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chen:2011:POI

[CW11] Danny Z. Chen and Haitao Wang. Processing an offline insertion-query sequence with applications. International Journal of Foundations of Computer Science (IJFCS), 22(6):1439–1456, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Cortina:1998:CRI

[CX98] Thomas J. Cortina and Zhi-Wei Xu. The cube-of-rings interconnection network. International Journal of Foundations of Computer Science (IJFCS), 9(1): 25–??, March 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chen:2012:MLD

[CYS^{+}12] Guaning Chen, Chih-Wei Yi, Min-Te Sun, Fang-Chu Liu, and Wei-Chi Lan. Minimum local disk cover sets for broadcasting in heterogeneous multihop wireless networks. International Journal of Foundations of Computer Science (IJFCS), 23 (5):1147–??, August 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Chen:2014:OSM

[CYZ14] Lin Chen, Deshi Ye, and Guochuan Zhang. Online scheduling of mixed CPU–GPU jobs. International Journal of Foundations of Computer Science (IJFCS), 25(6):745–??, September 2014. CODEN IFCSEN. ISSN 0129-0541.

Collins:2011:CSC

[CZ11] Pieter Collins and Ivan S.

Zapreev. Computable semantics for Ctl* on discretetime and continuous-space dynamic systems. International Journal of Foundations of Computer Science (IJFCS), 22(4):801– 821, June 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Calvo-Zaragoza:2017:CEE

[CZOdlH17] Jorge Calvo-Zaragoza, Jose Oncina, and Colin de la Higuera. Computing the expected edit distance from a string to a probabilistic finite-state automaton. International Journal of Foundations of Computer Science (IJFCS), 28(5):603–??, August 2017. CODEN IFC-SEN. ISSN 0129-0541.

Cai:2013:NOF

[CZTH13] Han Cai, Xiangyong Zeng, Xiaohu Tang, and Lei Hu. New optimal frequency hopping sequence sets from balanced nested difference packings of partition-type. International Journal of Foundations of Computer Science (IJFCS), 24(4):533– ??, June 2013. CODEN IFCSEN. ISSN 0129-0541.

Dai:1997:CDS

[Dai97] H. K. Dai. The complexity of deciding strictly nonblocking concentration and generalized-concentration properties. International Journal of Foundations of Computer Science (IJFCS), 8(3):237–??, September 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dangalchev:2011:RCG

[Dan11] Chavdar Dangalchev. Residual closeness and generalized closeness. International Journal of Foundations of Computer Science (IJFCS), 22(8):1939–1948, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dangalchev:2022:ACC

[Dan22] Chavdar Dangalchev. Additional closeness of cycle graphs. International Journal of Foundations of Computer Science (IJFCS), 33(08):1033– 1052, December 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500149.

Darmann:2013:PST

[Dar13] Andreas Darmann. Popular spanning trees. International Journal of Foundations of Computer Science (IJFCS), 24(5):655–??, August 2013. CODEN IFC-SEN. ISSN 0129-0541.

Dassow:2004:DCL

[Das04] Jürgen Dassow. On the descriptional complexity of Lindenmayer systems. International Journal of Foundations of Computer Science (IJFCS), 15(4):663– ??, August 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dassow:2019:OAS

[Das19] Jürgen Dassow. Operational accepting state complexity: The unary and finite case. International Journal of Foundations of Computer Science (IJFCS), 30 (6–7):959–978, September– November 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400240.

Dassow:2021:FRO

[Das21] Jürgen Dassow. Further remarks on the operational nonterminal complexity. International Journal of Foundations of Computer Science (IJFCS), 32(05):439– 453, August 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121410021.

deBondt:2019:LBS

[dBDZ19] Michiel de Bondt, Henk Don, and Hans Zantema. Lower bounds for synchronizing word lengths in partial automata. International Journal of Foundations of Computer Science (IJFCS), 30(1):29– 60, January 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400021.

Das:2013:NEA

[DCS13] Ashok Kumar Das, Santanu Chatterjee, and Jamuna Kanta Sing. A novel efficient access control scheme for large-scale distributed wireless sensor networks. International Journal of Foundations of Computer Science (IJFCS), 24 (5):625–??, August 2013. CODEN IFCSEN. ISSN 0129-0541.

DeLuca:2006:CPS

[DD06] Aldo De Luca and Alessandro De Luca. Combinatorial properties of Sturmian palindromes. International Journal of Foundations of Computer Science (IJFCS), 17(3):557–??, June 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Daykin:2008:PCU

[DD08] David E. Daykin and Jacqueline W. Daykin. Properties and construction of unique maximal factorization families for strings. International Journal of Foundations of Computer Science (IJFCS), 19(4):1073– 1084, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dominguez:2012:PPF

[DD12] César Domínguez and Dominique Duval. A parameterization process: from a functorial point of view. International Journal of Foundations of Computer Science (IJFCS), 23(1):225– 242, January 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Deorowicz:2013:BPA

[DD13] Sebastian Deorowicz and Agnieszka Danek. Bitparallel algorithms for the merged longest common subsequence problem. International Journal of Foundations of Computer Science (IJFCS), 24(8):1281–??, December 2013. CODEN IFC-SEN. ISSN 0129-0541.

Darmann:2018:MSP

[DDD18] Andreas Darmann, Janosch Döcker, and Britta Dorn. The monotone satisfiability problem with bounded variable appearances. International Journal of Foundations of Computer Science (IJFCS), 29(6):979– 993, September 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118500168.

Datta:2011:SSE

[DDHL11] Ajoy K. Datta, Stéphane Devismes, Florian Horn, and Lawrence L. Larmore. Self-stabilizing k-out-of-ℓ exclusion in tree networks. International Journal of Foundations of Computer Science (IJFCS), 22(3):657– 677, April 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Daley:2007:IMT

[DDM07] Mark Daley, Michael Domaratzki, and Alexis Morris. Intra-molecular templateguided recombination. International Journal of Foundations of Computer Science (IJFCS), 18(6):1177– 1186, December 2007. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dieudonne:2019:ECA

[DDPS19] Yoann Dieudonné, Shlomi Dolev, Franck Petit, and Michael Segal. Explicit communication among stigmergic robots. International Journal of Foundations of Computer Science (IJFCS), 30(2):315– 332, February 2019. ISSN 0129-0541.

DeAgostino:2006:BSD

[De 06] Sergio De Agostino. Bounded size dictionary compression: Relaxing the LRU deletion heuristic. International Journal of Foundations of Computer Science (IJFCS), 17(6):1273–1280, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ducrou:2008:IUI

[DE08] Jon Ducrou and Peter Eklund. An intelligent user interface for browsing and searching MPEG-7 images using concept lattices. International Journal of Foundations of Computer Science (IJFCS), 19(2):359– 381, April 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Domosi:2012:P

[DÉ12] Pál Dömösi and Zoltán Ésik. Preface. International Journal of Foundations of Computer Science (IJFCS), 23(6):1185–??, September 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Droste:2022:TPC

[DÉK22] Manfred Droste, Zoltán Ésik, and Werner Kuich. The triple-pair construction for weighted ω-pushdown automata. International Journal of Foundations of Computer Science (IJFCS), 33(3–4):227–246, April– June 2022. ISSN 0129-

                                        0541. URL https://www.
                                        worldscientific.com/doi/
                                        10.1142/S0129054122410039.

Duncan:2006:DFE

[DEKW06] Christian A. Duncan, Alon Efrat, Stephen Kobourov, and Carola Wenk. Drawing with fat edges. International Journal of Foundations of Computer Science (IJFCS), 17(5):1143– ??, October 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dolev:2011:TTU

[DEKZ11] Shlomi Dolev, Yuval Elovici, Alex Kesselman, and Polina Zilberman. Trawling traffic under attack overcoming DDOS attacks by targetcontrolled traffic filtering. International Journal of Foundations of Computer Science (IJFCS), 22(5): 1073–1098, August 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dutot:2005:SLS

[DEMT05] Pierre-François Dutot, Lionel Eyraud, Grégory Mounié, and Denis Trystram. Scheduling on large scale distributed platforms: From models to implementations. International Journal of Foundations of Computer Science (IJFCS), 16(2):217– ??, April 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Deng:2002:P

[Den02] Xiaotie Deng. Preface. International Journal of Foundations of Computer Science (IJFCS), 13(5):629– ??, October 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dixon:2009:ABS

[DES09] Ryan Dixon, Ömer Eḡecioḡlu, and Timothy Sherwood. Analysis of bit-split languages for packet scanning and experiments with wildcard matching. International Journal of Foundations of Computer Science (IJFCS), 20(4):597– 612, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Domaratzki:2005:NUR

[DESW05] Michael Domaratzki, Keith Ellul, Jeffrey Shallit, and Ming-Wei Wang. Nonuniqueness and radius of cyclic unary NFAs. International Journal of Foundations of Computer Science (IJFCS), 16(5):883– ??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Devillers:2002:DH

[Dev02] O. Devillers. The Delaunay hierarchy. International Journal of Foundations of Computer Science (IJFCS), 13(2):163–??, 2002. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Diessel:2001:DTS

[DEZ01] Oliver Diessel, Hossam Elgindy, and Albert Zomaya. On dynamic task scheduling for FPGA-based systems. International Journal of Foundations of Computer Science (IJFCS), 12 (5):645–669 (or 645–670??), October 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Droste:2023:DBF

[DFK23] Manfred Droste, Zoltán Fülöp, , and Dávid Kószó. Decidability boundaries for the finite-image property of weighted finite automata. International Journal of Foundations of Computer Science (IJFCS), 34(06):633–653, September 2023. ISSN 0129-

0541. URL https://www.
worldscientific.com/doi/
10.1142/S0129054123450041.

Deng:2002:PMT

[DFLL02] Xiaotie Deng, Haodi Feng, Guojun Li, and Guizhen Liu. A PTAS for minimizing total completion time of bounded batch scheduling. International Journal of Foundations of Computer Science (IJFCS), 13(6):817– ??, December 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dassow:1999:LDM

[DFP99] J. Dassow, H. Fernau, and G. Păun. On the leftmost derivation in matrix grammars. International Journal of Foundations of Computer Science (IJFCS), 10 (1):61–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Droste:1990:UIS

[DG90] Manfred Droste and Rudiger Gobel. Universal information systems. International Journal of Foundations of Computer Science (IJFCS), 1(4):413–??, December 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Diderich:1998:EDO

[DG98] Claude G. Diderich and Marc Gengler. An extended dimension order token distribution algorithm on kary d-cubes and its complexity. International Journal of Foundations of Computer Science (IJFCS), 9(2): 213–??, June 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dawson:2009:TAR

[DG09] Jeremy E. Dawson and Rajeev Goré. Termination of abstract reduction systems. International Journal of Foundations of Computer Science (IJFCS), 20(1):57– 82, February 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Diekert:2008:SSF

[DGK08] Volker Diekert, Paul Gastin, and Manfred Kufleitner. A survey on small fragments of first-order logic over finite words. International Journal of Foundations of Computer Science (IJFCS), 19(3):513–548, June 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Diks:1993:PAF

[DGL93] K. Diks, O. Garrido, and A. Lingas. Parallel algorithms for finding maximal k-dependent sets and maximal f-matchings. International Journal of Foundations of Computer Science (IJFCS), 4(2):179–??, June 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dumitran:2015:BPS

[DGMM15] Marius Dumitran, Javier Gil, Florin Manea, and Victor Mitrana. Bounded prefix-suffix duplication: Language theoretic and algorithmic results. International Journal of Foundations of Computer Science (IJFCS), 26(7):933–??, November 2015. CODEN IFCSEN. ISSN 0129-0541.

Das:2007:IAM

[DGN07] Gautam K. Das, Sasthi C.

Ghosh, and Subhas C. Nandy. Improved algorithm for minimum cost range assignment problem for linear radio networks. International Journal of Foundations of Computer Science (IJFCS), 18(3):619– 635, June 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Droste:2024:WLD

[DGR24] Manfred Droste, Gustav Grabolle, , and George Rahonis. Weighted linear dynamic logic. International Journal of Foundations of Computer Science (IJFCS), 35(1–2):145–177, January– February 2024. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054123480088.

Das:1996:CDS

[DH96] G. Das and P. J. Heffernan. Constructing degree-3 spanners with other sparseness properties. International Journal of Foundations of Computer Science (IJFCS), 7(2):121–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dassow:2005:LFD

[DH05] Jürgen Dassow and Markus Holzer. Language families defined by a ciliate biooperation: Hierarchies and decision problems. International Journal of Foundations of Computer Science (IJFCS), 16(4):645– ??, August 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Derakhshan:2018:OBD

[DH18] Parisa Derakhshan and Walter Hussak. Optimal bounds for disjoint Hamilton cycles in star graphs. International Journal of Foundations of Computer Science (IJFCS), 29(3):377–389, April 2018. CODEN IFCSEN. ISSN 0129-0541.

Das:1997:ECF

[DHIÖ97] Sajal K. Das, Dirk H. Hohndel, Maximilian Ibel, and Sabine R. Öhring. Efficient communication in folded Petersen networks. International Journal of Foundations of Computer Science (IJFCS), 8(2):163–??, June 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Doyen:2008:ELM

[DHR08] Laurent Doyen, Thomas A.

Henzinger, and Jean-François Raskin. Equivalence of labeled Markov chains. International Journal of Foundations of Computer Science (IJFCS), 19(3):549– 563, June 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dang:2002:ECT

[DI02] Zhe Dang and Oscar H.

Ibarra. The existence of ωchains for transitive mixed linear relations and its applications. International Journal of Foundations of Computer Science (IJFCS), 13 (6):911–??, December 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dang:2005:OMS

[DI05] Zhe Dang and Oscar H. Ibarra. On one-membrane P systems operating in sequential mode. International Journal of Foundations of Computer Science (IJFCS), 16(5):867– ??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dickerson:1993:GPD

[Dic93] M. T. Dickerson. General polynomial decomposition and the s-1-decomposition are NP-hard. International Journal of Foundations of Computer Science (IJFCS), 4(2):147–156, June 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dolzhenko:2012:TDL

[DJ12] Egor Dolzhenko and Nataša Jonoska. Two-dimensional languages and cellular automata. International Journal of Foundations of Computer Science (IJFCS), 23 (1):185–206, January 2012.

CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dessmark:2007:AMM

[DJL^{+}07] Anders Dessmark, Jesper Jansson, Andrzej Lingas, Eva-Marta Lundell, and Mia Persson. On the approximability of maximum and minimum edge clique partition problems. International Journal of Foundations of Computer Science (IJFCS), 18(2):217–226, April 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dartois:2018:AST

[DJR18] Luc Dartois, Ismaël Jecker, and Pierre-Alain Reynier. Aperiodic string transducers. International Journal of Foundations of Computer Science (IJFCS), 29(5):801– 824, August 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118420054.

Droste:1998:RLD

[DK98] Manfred Droste and Dietrich Kuske. Recognizable and logically definable languages of infinite computations in concurrent automata. International Journal of Foundations of Computer Science (IJFCS), 9 (3):295–??, September 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Diekert:2011:INC

[DK11] Volker Diekert and Steffen Kopecki. It is NL-complete to decide whether a hairpin completion of regular languages is regular. International Journal of Foundations of Computer Science (IJFCS), 22(8):1813–1828, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dudzinski:2012:FDC

[DK12] Krystian Dudzinski and Stavros Konstantinidis. Formal descriptions of code properties: Decidability, complexity, implementation. International Journal of Foundations of Computer Science (IJFCS), 23(1):67– 85, January 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Daley:2011:OST

[DKSS11] Mark Daley, Lila Kari, Shinnosuke Seki, and Petr Sosìk. Orthogonal shuffle on trajectories. International Journal of Foundations of Computer Science (IJFCS), 22(1):213– 222, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

DAgostino:2012:MCF

[DL12] Giovanna D’Agostino and Giacomo Lenzi. On modal μ-calculus over finite graphs with small components or small tree width. International Journal of Foundations of Computer Science (IJFCS), 23(3):627– ??, April 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ding:2014:OMM

[DLC^{+}14] Ning Ding, Yan Lan, Xin Chen, György Dósa, He Guo, and Xin Han. Online minimum makespan scheduling with a buffer. International Journal of Foundations of Computer Science (IJFCS), 25(5):525–??, August 2014. CODEN IFC-SEN. ISSN 0129-0541.

Durand-Lose:2012:P

[DLMS12] Jérôme Durand-Lose, Maurice Margenstern, and Klaus Sutner. Preface. International Journal of Foundations of Computer Science (IJFCS), 23(7):1419– ??, November 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

DiGiacomo:2006:EGT

[DLT06] Emilio Di Giacomo, Giuseppe Liotta, and Francesco Trotta. On embedding a graph on two sets of points. International Journal of Foundations of Computer Science (IJFCS), 17(5):1071– ??, October 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Deng:2002:CCA

[DLW02] Xiaotie Deng, Zhong-Fei Li, and Shou-Yang Wang. Computational complexity of arbitrage in frictional security market. International Journal of Foundations of Computer Science (IJFCS), 13 (5):681–??, October 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Daley:2005:FMV

[DM05] Mark Daley and Ian Mcquillan. Formal modelling of viral gene compression. International Journal of Foundations of Computer Science (IJFCS), 16(3):453– ??, June 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dragoi:2008:DCA

[DM08] Cezara Drăgoi and Florin Manea. On the descriptional complexity of accepting networks of evolutionary processors with filtered connections. International Journal of Foundations of Computer Science (IJFCS), 19(5):1113–1132, October 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Droste:2011:WAR

[DM11] Manfred Droste and Ingmar Meinecke. Weighted automata and regular expressions over valuation monoids. International Journal of Foundations of Computer Science (IJFCS), 22(8):1829–1844, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Diekert:2012:GEI

[DM12] Volker Diekert and Alexei Myasnikov. Group extensions over infinite words. International Journal of Foundations of Computer Science (IJFCS), 23(5):1001– ??, August 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lopez:2020:SIM

[dMLBPP20] Luis Fernando de Mingo López, Nuria Gómez Blas, Angel Luis Castellanos Peñuela, and Juan Bautista Castellanos Peñuela. Swarm intelligence models: Ant colony systems applied to BNF grammars rule derivation. International Journal of Foundations of Computer Science (IJFCS), 31(1):103– 116, January 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120400079.

Dassow:2014:IPC

[DMMM14] Jürgen Dassow, Florin Manea, Robert Mercaş, and Mike Müller. Inner palindromic closure. International Journal of Foundations of Computer Science (IJFCS), 25(8):1049–??, December 2014. CODEN IFC-SEN. ISSN 0129-0541.

Du:2016:DAF

[DMSS16] Chen Fei Du, Hamoon Mousavi, Luke Schaeffer, and Jeffrey Shallit. Decision algorithms for Fibonacciautomatic words, III: Enumeration and Abelian properties. International Journal of Foundations of Computer Science (IJFCS), 27(8):943– 964, December 2016. CO-DEN IFCSEN. ISSN 0129-0541.

Demri:2007:RAT

[DN07] Stéphane Demri and David Nowak. Reasoning about transfinite sequences. International Journal of Foundations of Computer Science (IJFCS), 18(1):87–112, February 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Diekert:2011:P

[DN11] Volker Diekert and Dirk Nowotka. Preface. International Journal of Foundations of Computer Science (IJFCS), 22(2):275– 276, February 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

DeFelice:2016:ACA

[DN16] Sven De Felice and Cyril Nicaud. Average case analysis of Brzozowski’s algorithm. International Journal of Foundations of Computer Science (IJFCS), 27 (2):109–??, February 2016. CODEN IFCSEN. ISSN 0129-0541.

Domaratzki:2004:IBN

[Dom04] Michael Domaratzki. Improved bounds on the number of automata accepting finite languages. International Journal of Foundations of Computer Science (IJFCS), 15(1):143–??, February 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dombek:2012:SIA

[Dom12] Daniel Dombek. Substitutions over infinite alphabet generating (−β)-integers. International Journal of Foundations of Computer Science (IJFCS), 23(8): 1627–??, December 2012.

CODEN IFCSEN. ISSN 0129-0541.

DeFraysseix:2006:TTP

[DOR06] Hubert De Fraysseix, Patrice Ossona De Mendez, and Pierre Rosenstiehl. Trémaux trees and planarity. International Journal of Foundations of Computer Science (IJFCS), 17(5):1017– ??, October 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

DeAgostino:1990:PRA

[DP90] Sergio De Agostino and Rossella Petreschi. Parallel recognition algorithms for graphs with restricted neighbourhoods. International Journal of Foundations of Computer Science (IJFCS), 1(2):123–??, June 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Delzanno:2013:P

[DP13] Giorgio Delzanno and Igor Potapov. Preface. International Journal of Foundations of Computer Science (IJFCS), 24(2):161–??, February 2013. CODEN IFCSEN. ISSN 0129-0541.

Droste:2014:WNW

[DP14] Manfred Droste and Bundit Pibaljommee. Weighted nested word automata and logics over strong bimonoids.

International Journal of Foundations of Computer Science (IJFCS), 25(5):641– ??, August 2014. CODEN IFCSEN. ISSN 0129-0541.

Damm:2007:GTV

[DPR07] Werner Damm, Guilherme Pinto, and Stefan Ratschan. Guaranteed termination in the verification of LTL properties of non-linear robust discrete time hybrid systems. International Journal of Foundations of Computer Science (IJFCS), 18(1):63– 86, February 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dongarra:2008:RMP

[DPR^{+}08] Jack Dongarra, Jean-François Pineau, Yves Robert, Zhiao Shi, and Frédéric Vivien. Revisiting matrix product on master-worker platforms. International Journal of Foundations of Computer Science (IJFCS), 19(6): 1317–1336, December 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dassow:1993:GBP

[DPS93] J. Dassow, G. Păun, and A. Salomaa. Grammars based on patterns. International Journal of Foundations of Computer Science (IJFCS), 4(1):1–14, March 1993. CODEN IFCSEN.

ISSN 0129-0541 (print), 1793-6373 (electronic).

Dumitrescu:1997:PLV

[DPS97] Sorina Dumitrescu, Georghe Păun, and Arto Salomaa. Pattern languages versus parallel communicating grammar systems. International Journal of Foundations of Computer Science (IJFCS), 8(1):67–??, March 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

DeQueirosVieiraMartins:1999:DAR

[DPS99] E. De Queiros Vieira Martins, M. M. B. Pascoal, and J. L. E. Dos Santos. Deviation algorithms for ranking shortest paths. International Journal of Foundations of Computer Science (IJFCS), 10(3):247– 262, 1999. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Devillers:2002:WT

[DPT02] O. Devillers, S. Pion, and M. Teillaud. Walking in a triangulation. International Journal of Foundations of Computer Science (IJFCS), 13(2):181–??, 2002. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dong:2012:NCV

[DQFL12] Deshuai Dong, Longjiang Qu, Shaojing Fu, and Chao Li. New constructions of vectorial Boolean functions with good cryptographic properties. International Journal of Foundations of Computer Science (IJFCS), 23(3):749–??, April 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Duris:1994:CDR

[DR94] P. Duris and J. D. P. Rolim.

Conjunctive and disjunctive reducibilities to sparse and tally sets revisited. International Journal of Foundations of Computer Science (IJFCS), 5(3/4):??, 1994. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Davila:2005:RSP

[DR05] Jaime Davila and Sanguthevar Rajasekaran. Randomized sorting on the POPS network. International Journal of Foundations of Computer Science (IJFCS), 16(1):105–??, February 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

DeFelice:2006:P

[DR06] Clelia De Felice and Antonio Restivo. Preface. International Journal of Foundations of Computer Science (IJFCS), 17(3):489– ??, June 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Domaratzki:2012:APW

[DR12] Michael Domaratzki and Narad Rampersad. Abelian primitive words. International Journal of Foundations of Computer Science (IJFCS), 23(5):1021– ??, August 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Das:2008:VBS

[DRDN08] Gautam K. Das, Sasanka Roy, Sandip Das, and Subhas C. Nandy. Variations of base-station placement problem on the boundary of a convex region. International Journal of Foundations of Computer Science (IJFCS), 19(2):405– 427, April 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Drewes:2007:L

[Dre07] Frank Drewes. Links. International Journal of Foundations of Computer Science (IJFCS), 18(6):1187–1196, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Droste:1992:CAD

[Dro92] M. Droste. Concurrent automata and domains. International Journal of Foundations of Computer Science (IJFCS), 3(4):389–418, December 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Day:2014:DPC

[DRS14] Joel D. Day, Daniel Reidenbach, and Johannes C. Schneider. On the dual post correspondence problem. International Journal of Foundations of Computer Science (IJFCS), 25(8): 1033–??, December 2014. CODEN IFCSEN. ISSN 0129-0541.

Droste:2023:SII

[DRS23] Manfred Droste, George Rahonis, , and Arto Salomaa. Special issue: International Colloquium: Recent Advances of Quantitative Models in Computer Science (RAQM 2021) — Preface. International Journal of Foundations of Computer Science (IJFCS), 34(06): 537–538, September 2023. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123020033.

Duval:1996:CCS

[DS96] A. M. Duval and W. F. Smyth. Covering a circular string with substrings of fixed length. International Journal of Foundations of Computer Science (IJFCS), 7(1):87–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

DeFrancesco:2002:FDM

[DS02] Nicoletta De Francesco and Antonella Santone. A formula-driven modular attack on state explosion. International Journal of Foundations of Computer Science (IJFCS), 13(5):719– ??, October 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dsouza:2003:LCE

[D’s03] Deepak D’souza. A logical characterisation of event clock automata. International Journal of Foundations of Computer Science (IJFCS), 14(4):625– ??, August 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Domaratzki:2005:RST

[DS05] Michael Domaratzki and Kai Salomaa. Restricted sets of trajectories and decidability of shuffle decompositions. International Journal of Foundations of Computer Science (IJFCS), 16(5):897–??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Domaratzki:2008:P

[DS08] Michael Domaratzki and Kai Salomaa. Preface. International Journal of Foundations of Computer Science (IJFCS), 19(5):1085– 1086, October 2008. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Domaratzki:2011:P

[DS11] Michael Domaratzki and Kai Salomaa. Preface. International Journal of Foundations of Computer Science (IJFCS), 22(8):1759– 1760, December 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ozelim:2018:IDF

[dSMOC18] Luan Carlos de Sena Monteiro Ozelim and Andre Luis Brasil Cavalcante. The iota-delta function as an alternative to Boolean formalism. International Journal of Foundations of Computer Science (IJFCS), 29(3):415– 423, April 2018. CODEN IFCSEN. ISSN 0129-0541.

Dobrev:2008:USM

[DSS08] Stefan Dobrev, Nicola Santoro, and Wei Shi. Using scattered mobile agents to locate a black hole in an un-oriented ring with tokens. International Journal of Foundations of Computer Science (IJFCS), 19 (6):1355–1372, December 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Du:2015:OBS

[DSS15] Chen Fei Du, Jeffrey Shallit, and Arseny M. Shur. Optimal bounds for the similarity density of the Thue– Morse word with overlapfree and 7/3-power-free infinite binary words. International Journal of Foundations of Computer Science (IJFCS), 26(8):1147–??, December 2015. CODEN IFC-SEN. ISSN 0129-0541.

Dassow:2010:GCS

[DST10] Jürgen Dassow, Ralf Stiebe, and Bianca Truthe. Generative capacity of subregularly tree controlled grammars. International Journal of Foundations of Computer Science (IJFCS), 21(5):723– 740, October 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dassow:2020:NEP

[DT20] Jürgen Dassow and Bianca Truthe. Networks with evolutionary processors and ideals and codes as filters. International Journal of Foundations of Computer Science (IJFCS), 31(1):73– 89, January 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120400055.

Devismes:2015:WVS

[DTY15] Stéphane Devismes, Sébastien Tixeuil, and Masafumi Yamashita. Weak vs. self vs. probabilistic stabilization. International Journal of Foundations of Computer Science (IJFCS), 26(3):293– ??, April 2015. CODEN IFCSEN. ISSN 0129-0541.

Dubacq:1995:HST

[Dub95] J.-C. Dubacq. How to simulate Turing machines by invertible one-dimensional cellular automata. International Journal of Foundations of Computer Science (IJFCS), 6(4):395–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Durand:2013:DUR

[Dur13] Fabien Durand. Decidability of uniform recurrence of morphic sequences. International Journal of Foundations of Computer Science (IJFCS), 24(1):123–??, January 2013. CODEN IFC-SEN. ISSN 0129-0541.

Dassow:2011:NAS

[DV11] Jürgen Dassow and György Vaszil. On the number of active symbols in Lindenmayer systems. International Journal of Foundations of Computer Science (IJFCS), 22 (1):223–235, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Droste:2014:CST

[DV14] Manfred Droste and Heiko Vogler. The Chomsky– Schützenberger theorem for quantitative contextfree languages. International Journal of Foundations of Computer Science (IJFCS), 25(8):955–??, December 2014. CODEN IFC-SEN. ISSN 0129-0541.

Dold:2003:CVR

[DVG03] Axel Dold, Friedrich Von Henke, and Wolfgang Goerigk. A completely verified realistic bootstrap compiler. International Journal of Foundations of Computer Science (IJFCS), 14(4):659– ??, August 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dantsin:2003:RDC

[DW03] Evgeny Dantsin and Alexander Wolpert. A robust DNA computation model that captures Pspace. International Journal of Foundations of Computer Science (IJFCS), 14(5):933– ??, October 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Desmedt:2004:AVC

[DW04] Yvo Desmedt and Yongge Wang. Analyzing vulnerabilities of critical infrastructures using flows and critical vertices in And/Or graphs. International Journal of Foundations of Computer Science (IJFCS), 15 (1):107–??, February 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ding:2011:P

[DW11] Cunsheng Ding and Qi Wang.

Preface. International Journal of Foundations of Computer Science (IJFCS), 22 (6):1241, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ding:2015:NSS

[DWS15] Yihua Ding, James Z. Wang, and Pradip K. Srimani. New self-stabilizing algorithms for minimal weakly connected dominating sets. International Journal of Foundations of Computer Science (IJFCS), 26(2):229–??, February 2015. CODEN IFCSEN. ISSN 0129-0541.

Ding:2020:NID

[DXZ20] Tongtong Ding, Min Xu, and Qiang Zhu. The non-inclusive diagnosability of hypercubes under the MM* model. International Journal of Foundations of Computer Science (IJFCS), 31(07):929– 940, November 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500355.

Dimitrijevs:2019:URP

[DY19] Maksims Dimitrijevs and Abuzer Yakaryılmaz. Uncountable realtime probabilistic classes. International Journal of Foundations of Computer Science (IJFCS), 30(8):1317– 1333, December 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905411950028X.

Dong:2000:SAA

[DZ00] G. Dong and L. Zhang. Separating auxiliary arity hierarchy of first-order incremental evaluation systems using (3k + 1)-ary input relations. International Journal of Foundations of Computer Science (IJFCS), 11 (4):573–578, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Deng:2016:ECP

[DZH16] Lunzhi Deng, Jiwen Zeng, and Huawei Huang. Efficient certificateless proxy signature scheme. International Journal of Foundations of Computer Science (IJFCS), 27(1):85–??, January 2016. CODEN IFC-SEN. ISSN 0129-0541. Etherington:2016:PAC

[EAB^{+}16] Carole J. Etherington, Matthew W. Anderson, Eric Bach, Jon T. Butler, and Pantelimon Stănică. A parallel approach in computing correlation immunity up to six variables. International Journal of Foundations of Computer Science (IJFCS), 27(4):511–??, June 2016. CODEN IFCSEN. ISSN 0129-0541.

Estivill-Castro:2002:CWV

[ECY02] V. Estivill-Castro and J. Yang.

Clustering Web visitors by fast, robust and convergent algorithms. International Journal of Foundations of Computer Science (IJFCS), 13(4):497–??, 2002. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Edelsbrunner:2002:SIM

[EG02] H. Edelsbrunner and D. Guoy.

Sink insertion for mesh improvement. International Journal of Foundations of Computer Science (IJFCS), 13(2):223–??, 2002. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Elkind:2010:QDO

[EGPS10] Edith Elkind, Blaise Genest, Doron Peled, and Paola Spoletini. Quantifying the discord: Order discrepancies in message sequence charts. International Journal of Foundations of Computer Science (IJFCS), 21 (2):211–233, April 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ediger:2012:EAT

[EH12] Patrick Ediger and Rolf Hoffmann. Efficiency analysis of the time-shuffling method for the evolution of agent behavior. International Journal of Foundations of Computer Science (IJFCS), 23(2):523– 542, February 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Eom:2015:SCB

[EH15] Hae-Sung Eom and Yo-Sub Han. State complexity of boundary of prefix-free regular languages. International Journal of Foundations of Computer Science (IJFCS), 26(6):697–??, September 2015. CODEN IFCSEN. ISSN 0129-0541.

Emerson:2006:MMC

[EHK06] E. Allen Emerson, Kristina D. Hager, and Jay H. Konieczka. Molecular model checking. International Journal of Foundations of Computer Science (IJFCS), 17(4):733– ??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Eom:2015:SCK

[EHS15] Hae-Sung Eom, Yo-Sub Han, and Kai Salomaa. State complexity of k-union and k-intersection for prefixfree regular languages. International Journal of Foundations of Computer Science (IJFCS), 26(2):211–??, February 2015. CODEN IFCSEN. ISSN 0129-0541.

Esik:2014:OCS

[ÉI14] Zoltán Ésik and Szabolcs Iván. Operational characterization of scattered MCFLs. International Journal of Foundations of Computer Science (IJFCS), 25(8):1001–??, December 2014. CODEN IFCSEN. ISSN 0129-0541.

Eremondi:2018:DCF

[EIM18] Joey Eremondi, Oscar H.

Ibarra, and Ian McQuillan. On the density of contextfree and counter languages. International Journal of Foundations of Computer Science (IJFCS), 29(2):233– ??, February 2018. CODEN IFCSEN. ISSN 0129-0541.

Edixhoven:2023:BCR

[EJ23] Luc Edixhoven and Sung-Shik Jongmans. Balancedby-construction regular and ω-regular languages. International Journal of Foundations of Computer Science (IJFCS), 34(2–3):117–144, February–April 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122440026.

Esik:2007:BFS

[ÉK07] Zoltán Ésik and Werner Kuich. Boolean fuzzy sets. International Journal of Foundations of Computer Science (IJFCS), 18 (6):1197–1207, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Endrullis:2018:DIW

[EKKS18] Jörg Endrullis, Juhani Karhumäki, Jan Willem Klop, and Aleksi Saarela. Degrees of infinite words, polynomials and atoms. International Journal of Foundations of Computer Science (IJFCS), 29(5):825– 843, August 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118420066.

Ehmsen:2013:TEC

[EL13] Martin R. Ehmsen and Kim S. Larsen. A technique for exact computation of precoloring extension on interval graphs. International Journal of Foundations of Computer Science (IJFCS), 24(1):109–??, January 2013. CODEN IFC-SEN. ISSN 0129-0541.

Elbl:2001:PRP

[Elb01] B. Elbl. Part 2 (regular papers). International Journal of Foundations of Computer Science (IJFCS), 12 (3):385–??, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Elmasry:2006:PQW

[Elm06] Amr Elmasry. A priority queue with the workingset property. International Journal of Foundations of Computer Science (IJFCS), 17(6):1455–1465, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Esparza:2015:FGS

[ELS15] Javier Esparza, Michael Luttenberger, and Maximilian Schlund. FPSOLVE: A generic solver for fixpoint equations over semirings. International Journal of Foundations of Computer Science (IJFCS), 26(7):805– ??, November 2015. CO-DEN IFCSEN. ISSN 0129-0541.

Esik:2011:CSW

[ÉM11] Zoltán Ésik and Andreas Maletti. The category of simulations for weighted tree automata. International Journal of Foundations of Computer Science (IJFCS), 22(8):1845–1859, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ehrenfeucht:2010:CLD

[EMR10] Andrzej Ehrenfeucht, Michael Main, and Grzegorz Rozenberg. Combinatorics of life and death for reaction systems. International Journal of Foundations of Computer Science (IJFCS), 21(3):345– 356, June 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ehrenfeucht:2011:FDR

[EMR11] Andrzej Ehrenfeucht, Michael Main, and Grzegorz Rozenberg. Functions defined by reaction systems. International Journal of Foundations of Computer Science (IJFCS), 22(1):167– 178, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ehrenfeucht:2012:SCR

[EMRB12] Andrzej Ehrenfeucht, Michael Main, Grzegorz Rozenberg, and Allison Thompson Brown. Stability and chaos in reaction systems. International Journal of Foundations of Computer Science (IJFCS), 23(5):1173– ??, August 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Emerson:2003:RAR

[EN03] E. Allen Emerson and Kedar S. Namjoshi. On reasoning about rings. International Journal of Foundations of Computer Science (IJFCS), 14(4):527– ??, August 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Esik:2013:CFL

[ÉO13] Zoltán Ésik and Satoshi Okawa. On context-free languages of scattered words. International Journal of Foundations of Computer Science (IJFCS), 24(7): 1029–??, November 2013. CODEN IFCSEN. ISSN 0129-0541.

Elouasbi:2017:DRD

[EP17] Samir Elouasbi and Andrzej Pelc. Deterministic rendezvous with detection using beeps. International Journal of Foundations of Computer Science (IJFCS), 28(1):77–97, January 2017. CODEN IFCSEN. ISSN 0129-0541.

Ehrenfeucht:2006:CT

[ER06] A. Ehrenfeucht and G. Rozenberg. Covers from templates. International Journal of Foundations of Computer Science (IJFCS), 17 (2):475–??, April 2006. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ehrenfeucht:2014:ZSR

[ER14] Andrzej Ehrenfeucht and Grzegorz Rozenberg. Zoom structures and reaction systems yield exploration systems. International Journal of Foundations of Computer Science (IJFCS), 25 (3):275–??, April 2014. CO-DEN IFCSEN. ISSN 0129-0541.

Egecioglu:2004:CGW

[ERW04] Ömer Eğecioğlu, Jeffrey B. Remmel, and S. Gill Williamson. A class of graphs which has efficient ranking and unranking algorithms for spanning trees and forests. International Journal of Foundations of Computer Science (IJFCS), 15(4):619– ??, August 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Elbl:2001:NDR

[ES01] Birgit Elbl and Jiawen Su. A non-definability result for a predicational language with the usual control. International Journal of Foundations of Computer Science (IJFCS), 12(3):385– 396, 2001. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Esik:2012:OAC

[Ési12] Zoltán Ésik. Ordinal automata and Cantor normal form. International Journal of Foundations of Computer Science (IJFCS), 23(1):87– 98, January 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Esik:2015:P

[Ési15] Zoltán Ésik. Preface. International Journal of Foundations of Computer Science (IJFCS), 26(8):1007–??, December 2015. CODEN IFC-SEN. ISSN 0129-0541.

Egecioglu:2020:KFC

[ESS20] Ömer Eğecioğlu, Elif Saygı, and Zülfükar Saygı. k-Fibonacci cubes: A family of subgraphs of Fibonacci cubes. International Journal of Foundations of Computer Science (IJFCS), 31(05):639– 661, August 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500318.

Egecioglu:2021:ALC

[ESS21] Ömer Eğecioğlu, Elif Saygı, and Zülfükar Saygı. Alternate Lucas cubes. International Journal of Foundations of Computer Science (IJFCS), 32(07):871– 899, November 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500271.

Elmasry:2014:FST

[ET14] Amr Elmasry and Yung H. Tsin. On finding sparse three-edge-connected and three-vertex-connected spanning subgraphs. International Journal of Foundations of Computer Science (IJFCS), 25(3):355–??, April 2014. CODEN IFC-SEN. ISSN 0129-0541.

Eshaghian:2001:MAH

[EZ01] Mary Mehrnoosh Eshaghian and Albert Zomaya. Mapping arbitrary heterogeneous task graphs onto arbitrary heterogeneous system graphs. International Journal of Foundations of Computer Science (IJFCS), 12 (5):599–628, October 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fukagawa:2006:FAC

[FA06] Daiji Fukagawa and Tatsuya Akutsu. Fast algorithms for comparison of similar unordered trees. International Journal of Foundations of Computer Science (IJFCS), 17(3):703–??, June 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Faal:2019:MVE

[Faa19] Hossein Teimoori Faal. A multiset version of even-odd permutations identity. International Journal of Foundations of Computer Science (IJFCS), 30(5):683– 691, August 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119500163.

Farhadian:2020:AEV

[Far20] Ameneh Farhadian. Almost every n-vertex graph is determined by its 3 log_{2} nvertex subgraphs. International Journal of Foundations of Computer Science (IJFCS), 31(05):611– 619, August 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412050029X.

Fazekas:2008:IBS

[Faz08] Szilárd Zsolt Fazekas. On inequalities between subword histories. International Journal of Foundations of Computer Science (IJFCS), 19(4):1039–1047, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fazekas:2011:PRL

[Faz11] Szilárd Zsolt Fazekas. Powers of regular languages. International Journal of Foundations of Computer Science (IJFCS), 22(2):323– 330, February 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fujimoto:2001:MPT

[FBHH01] Noriyuki Fujimoto, Tomoki Baba, Takashi Hashimoto, and K. Hagihara. On message packaging in task scheduling for distributed memory parallel machines. International Journal of Foundations of Computer Science (IJFCS), 12(3):285– 306, 2001. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fontaine:2005:BBA

[FBK05] Marc Fontaine, Stefan Burkhardt, and Juha Kärkkäinen. BDD-based analysis of gapped q-gram filters. International Journal of Foundations of Computer Science (IJFCS), 16(6):1121– ??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fung:2005:OSU

[FCS05] Stanley P. Y. Fung, Francis Y. L. Chin, and Hong Shen. Online scheduling of unit jobs with bounded importance ratio. International Journal of Foundations of Computer Science (IJFCS), 16(3):581–??, June 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fajardo-Delgado:2012:RSS

[FDFZB12] Daniel Fajardo-Delgado, José Alberto Fernández-Zepeda, and Anu G. Bourgeois. Randomized selfstabilizing leader election in preference-based anonymous trees. International Journal of Foundations of Computer Science (IJFCS), 23(4):853–??, June 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Feng:2022:SPK

[Fen22]  Kai Feng.  Subnetwork

preclusion of (n, k)-star networks. International Journal of Foundations of Computer Science (IJFCS), 33 (05):439–451, August 2022. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054122500095.

Fernau:2007:PGR

[Fer07]  Henning Fernau.  Pro-

grammed grammars with rule queues. International Journal of Foundations of Computer Science (IJFCS), 18(6):1209–1213, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fernau:2015:FIR

[FFH15] Henning Fernau, Rudolf Freund, and Markus Holzer.

The finite index restriction meets hybrid modes in cooperating distributed grammar systems. International Journal of Foundations of Computer Science (IJFCS), 26(8):1167–??, December 2015. CODEN IFC-SEN. ISSN 0129-0541.

Fernandes:2019:PAS

[FFMW19] Cristina G. Fernandes, Carlos E. Ferreira, Flávio K. Miyazawa, and Yoshiko Wakabayashi. Prices of anarchy of selfish 2D bin packing games. International Journal of Foundations of Computer Science (IJFCS), 30(3):355–374, 2019. CO-DEN IFCSEN. ISSN 0129-0541.

Fredriksson:2008:EAM

[FG08] Kimmo Fredriksson and Szymon Grabowski. Efficient algorithms for (δ, γ, α) and (δ, κ_{δ}, α)-matching. International Journal of Foundations of Computer Science (IJFCS), 19(1):163– 183, February 2008. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fix:2007:VVL

[FGH^{+}07] Limor Fix, Orna Grumberg, Amnon Heyman, Tamir Heyman, and Assaf Schuster. Verifying very large industrial circuits using 100 processes and beyond. International Journal of Foundations of Computer Science (IJFCS), 18(1):45–61, February 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Falaschi:1990:NGH

[FGL^{+}90] Moreno Falaschi, Maurizio Gabbrielli, Giorgio Levi, et al. Nested guarded Horn clauses. International Journal of Foundations of Computer Science (IJFCS), 1 (3):249–??, September 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Freund:2011:P

[FGM^{+}11] Rudolf Freund, Marian Gheorghe, Solomon Marcus, Victor Mitrana, and Mario J. Pérez-Jiménez. Preface. International Journal of Foundations of Computer Science (IJFCS), 22 (1):1–6, January 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fachini:1990:SST

[FGS^{+}90] Emanuela Fachini, Jozef Gruska, Andrea Maggiolo Schettini, et al. Simulation of systolic tree automata on trellis automata. International Journal of Foundations of Computer Science (IJFCS), 1(2):87–??, June 1990. CODEN IFCSEN.

ISSN 0129-0541 (print), 1793-6373 (electronic).

Fouquet:1999:BGT

[FGV99] J. L. Fouquet, V. Giakoumakis, and J. M. Vanherpe. Bipartite graphs totally decomposable by canonical decomposition. International Journal of Foundations of Computer Science (IJFCS), 10(4):513– 534, 1999. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Faliszewski:2005:ASS

[FH05] Piotr Faliszewski and Lane A. Hemaspaandra. Advice for semifeasible sets and the complexity-theoretic cost(lessness) of algebraic properties. International Journal of Foundations of Computer Science (IJFCS), 16(5):913–??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Frisco:2011:STS

[FH11] Pierluigi Frisco and Hendrik Jan Hoogeboom. P systems and topology: Some suggestions for research. International Journal of Foundations of Computer Science (IJFCS), 22(1):179– 190, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Frei:2023:TWN

[FHKK23] Fabian Frei, Juraj Hromkovič, Rastislav Královič, and Richard Královič. Twoway non-uniform finite automata. International Journal of Foundations of Computer Science (IJFCS), 34 (2–3):145–162, February– April 2023. ISSN 0129-

0541. URL https://www.
worldscientific.com/doi/
10.1142/S0129054122440063.

Flocchini:2007:DCR

[FHL07] Paola Flocchini, Miao Jun Huang, and Flaminia L. Luccio. Decontaminating chordal rings and tori using mobile agents. International Journal of Foundations of Computer Science (IJFCS), 18(3):547–563, June 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fialik:2008:SBC

[Fia08] Ivan Fialík. Separation between classical and quantum winning strategies for the matching game. International Journal of Foundations of Computer Science (IJFCS), 19(6):1449–1459, December 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Finkel:2004:RLI

[Fin04] Olivier Finkel. On recognizable languages of infinite pictures. International Journal of Foundations of Computer Science (IJFCS), 15(6):823–??, December 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Finkel:2012:TAR

[Fin12] Olivier Finkel. Three applications to rational relations of the high undecidability of the infinite post correspondence problem in a regular ω-language. International Journal of Foundations of Computer Science (IJFCS), 23(7):1481–??, November 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Finkel:2019:ITL

[Fin19] Olivier Finkel. Incompleteness theorems, large cardinals, and automata over finite words. International Journal of Foundations of Computer Science (IJFCS), 30(3):449–467, 2019. ISSN 0129-0541.

Finkel:2021:TEP

[Fin21] Olivier Finkel. Two effective properties of ωrational functions. International Journal of Foundations of Computer Science (IJFCS), 32(07):901– 920, November 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500283.

Freund:2008:ESN

[FIO08] Rudolf Freund, Mihai Ionescu, and Marion Oswald. Extended spiking neural P systems with decaying spikes and/or total spiking. International Journal of Foundations of Computer Science (IJFCS), 19(5):1223–1234, October 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Franek:2012:CRA

[FJ12] Frantisek Franek and Mei Jiang. Crochemore’s repetitions algorithm revisited: Computing runs. International Journal of Foundations of Computer Science (IJFCS), 23(2):389– 401, February 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Forsyth:2016:RPW

[FJPS16] Michael Forsyth, Amlesh Jayakumar, Jarkko Peltomäki, and Jeffrey Shallit. Remarks on privileged words. International Journal of Foundations of Computer Science (IJFCS), 27 (4):431–??, June 2016. CO-DEN IFCSEN. ISSN 0129-0541.

Fujiwara:2005:PMI

[FK05] Akihiro Fujiwara and Satoshi Kamio. Procedures for multiple input functions with DNA molecules. International Journal of Foundations of Computer Science (IJFCS), 16(1):37–??, February 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fehnker:2006:HSV

[FK06] Ansgar Fehnker and Bruce Krogh. Hybrid system verification is not a sinecure — the electronic throttle control case study. International Journal of Foundations of Computer Science (IJFCS), 17(4):885– ??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fribourg:2013:PVT

[FK13] Laurent Fribourg and Ulrich Kühne. Parametric verification and test coverage for hybrid automata using the inverse method. International Journal of Foundations of Computer Science (IJFCS), 24(2):233–??, February 2013. CODEN IFCSEN. ISSN 0129-0541.

Faran:2019:PAA

[FK19] Rachel Faran and Orna Kupferman. A parametrized analysis of algorithms on hierarchical graphs. International Journal of Foundations of Computer Science (IJFCS), 30(6–7):979– 1003, September–November 2019. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054119400252.

Fazekas:2021:LBS

[FKM^{+}21] Szilárd Zsolt Fazekas, Hwee Kim, Ryuichi Matsuoka, Reoto Morita, and Shinnosuke Seki. Linear bounds on the size of conformations in greedy deterministic oritatami. International Journal of Foundations of Computer Science (IJFCS), 32(05):575– 596, August 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121410082.

Fujiwara:2011:P

[FKN11] Akihiro Fujiwara, Hirotsugu Kakugawa, and Koji Nakano. Preface. International Journal of Foundations of Computer Science (IJFCS), 22(5):999– 1000, August 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fujiwara:2007:PCM

[FKT07] Akihiro Fujiwara, Satoshi Kamio, and Akiko Takehara. Procedures for computing the maximum with DNA. International Journal of Foundations of Computer Science (IJFCS), 18(3):475– 493, June 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Friedgut:2006:BCM

[FKV06] Ehud Friedgut, Orna Kupferman, and Moshe Y. Vardi. Büchi complementation made tighter. International Journal of Foundations of Computer Science (IJFCS), 17 (4):851–??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Finta:1997:CTG

[FL97] Lucian Finta and Zhen Liu. Complexity of task graph scheduling with fixed communication capacity. International Journal of Foundations of Computer Science (IJFCS), 8(1):43–??, March 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Faro:2009:EVB

[FL09] Simone Faro and Thierry Lecroq. Efficient variants of the backward-oracle-matching algorithm. International Journal of Foundations of Computer Science (IJFCS), 20(6):967– 984, December 2009. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Friedmann:2012:TLS

[FL12] Oliver Friedmann and Martin Lange. Two local strategy iteration schemes for parity game solving. International Journal of Foundations of Computer Science (IJFCS), 23(3):669– ??, April 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fleischer:1996:SBS

[Fle96] R. Fleischer. A simple balanced search tree with O(1) worst-case update time. International Journal of Foundations of Computer Science (IJFCS), 7(2):137–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fleischer:2020:IPF

[Fle20] Lukas Fleischer. The intersection problem for finite semigroups. International Journal of Foundations of Computer Science (IJFCS), 31(06):827– 842, September 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120410075.

Fu:2019:CBO

[FLFR19] Qiang Fu, Ruihu Li, Fangwei Fu, and Yi Rao. On the construction of binary optimal LCD codes with short length. International Journal of Foundations of Computer Science (IJFCS), 30(8):1237– 1245, December 2019. CO-DEN IFCSEN. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054119500242.

Fleischmann:2021:RWR

[FLM^{+}21] Pamela Fleischmann, Marie Lejeune, Florin Manea, Dirk Nowotka, and Michel Rigo. Reconstructing words from right-bounded-block words. International Journal of Foundations of Computer Science (IJFCS), 32(06):619– 640, September 2021. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054121420016.

Floderus:2013:TME

[FLP13] Peter Floderus, Andrzej Lingas, and Mia Persson. Towards more efficient infection and fire fighting. International Journal of Foundations of Computer Science (IJFCS), 24(1):3–??, January 2013. CODEN IFC-SEN. ISSN 0129-0541.

Fazekas:2012:NPP

[FLST12] Szilárd Zsolt Fazekas, Peter Leupold, and Kayoko Shikishima-Tsuji. On non-primitive palindromic context-free languages. International Journal of Foundations of Computer Science (IJFCS), 23(6):1277– ??, September 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fuerer:1996:AAE

[FM96] M. Fuerer and W. Miller.

Alignment-to-alignment editing with “move gap” operations. International Journal of Foundations of Computer Science (IJFCS), 7 (1):23–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fimmel:2001:OSP

[FM01] Dirk Fimmel and J. Muller.

Optimal software pipelining under resource constraints. International Journal of Foundations of Computer Science (IJFCS), 12(6):697– 718, December 2001. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fazekas:2013:NDS

[FM13] Szilárd Zsolt Fazekas and Robert Mercaş. A note on the decidability of subword inequalities. International Journal of Foundations of Computer Science (IJFCS), 24(4):445–??, June 2013. CODEN IFCSEN. ISSN 0129-0541.

Fujiwara:2004:PLA

[FMC04] Akihiro Fujiwara, Ken’Ichi Matsumoto, and Wei Chen. Procedures for logic and arithmetic operations with DNA molecules. International Journal of Foundations of Computer Science (IJFCS), 15(3):461– ??, June 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fredriksson:2006:FMR

[FMN06] Kimmo Fredriksson, Veli Mäkinen, and Gonzalo Navarro. Flexible music retrieval in sublinear time. International Journal of Foundations of Computer Science (IJFCS), 17(6):1345– 1364, December 2006. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fazekas:2020:PSD

[FMR20a] Szilárd Zsolt Fazekas, Robert Mercaş, and Daniel Reidenbach. On the prefix– suffix duplication reduction. International Journal of Foundations of Computer Science (IJFCS), 31(1):91– 102, January 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120400067.

Filiot:2020:PLA

[FMR20b] Emmanuel Filiot, Nicolas Mazzocchi, and Jean-François Raskin. A pattern logic for automata with outputs. International Journal of Foundations of Computer Science (IJFCS), 31(06):711– 748, September 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120410038.

Fominykh:2013:PAS

[FMV13] F. M. Fominykh, P. V. Martyugin, and M. V. Volkov. P(l)aying for synchronization. International Journal of Foundations of Computer Science (IJFCS), 24(6):765– ??, September 2013. CO-DEN IFCSEN. ISSN 0129-0541.

Fujita:2016:FSC

[FNI16] Toru Fujita, Koji Nakano, and Yasuaki Ito. Fast simulation of Conway’s Game of Life using bitwise parallel bulk computation on a GPU. International Journal of Foundations of Computer Science (IJFCS), 27(8):981– 1004, December 2016. CO-DEN IFCSEN. ISSN 0129-0541.

Freund:2007:PHS

[FO07] Rudolf Freund and Marion Oswald. Partial halting in P systems. International Journal of Foundations of Computer Science (IJFCS), 18(6):1215–1225, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Freund:2008:CGS

[FO08] Rudolf Freund and Marion Oswald. Cd grammar systems with regular start conditions. International Journal of Foundations of Computer Science (IJFCS), 19 (4):767–779, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Freund:2005:ORC

[FOP05] Rudolf Freund, Marion Oswald, and Andrei Păun. Optimal results for the computational completeness of gemmating (tissue) P systems. International Journal of Foundations of Computer Science (IJFCS), 16(5):929– ??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Forsell:2010:PCS

[For10] Martti Forsell. On the performance and cost of some PRAM models on CMP hardware. International Journal of Foundations of Computer Science (IJFCS), 21(3):387–404, June 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ferrante:2004:VCP

[FP04] Alessandro Ferrante and Mimmo Parente. On the vertex-connectivity problem for graphs with sharpened triangle inequality. International Journal of Foundations of Computer Science (IJFCS), 15(5):701– ??, October 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fantozzi:2003:GPS

[FPP03] Carlo Fantozzi, Andrea Pietracaprina, and Geppino Pucci. A general PRAM simulation scheme for clustered machines. International Journal of Foundations of Computer Science (IJFCS), 14(6):1147–??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ferro:2003:FCM

[FPPS03] A. Ferro, G. Pigola, A. Pulvirenti, and D. Shasha. Fast clustering and minimum weight matching algorithms for very large mobile backbone wireless networks. International Journal of Foundations of Computer Science (IJFCS), 14(2):223– ??, April 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Faloutsos:2002:EAL

[FPS02] Michalis Faloutsos, Rajesh Pankaj, and Kenneth C. Sevcik. The effect of asymmetry on the on-line multicast routing problem. International Journal of Foundations of Computer Science (IJFCS), 13(6):889–??, December 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Frey:2002:BTA

[Fre02] W. H. Frey. Boundary triangulations approximating developable surfaces that interpolate a close space curve. International Journal of Foundations of Computer Science (IJFCS), 13 (2):285–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Freund:2005:SWS

[Fre05] Rudolf Freund. P systems working in the sequential mode on arrays and strings. International Journal of Foundations of Computer Science (IJFCS), 16 (4):663–??, August 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Freivalds:2008:NCM

[Fre08]  Rūsiņš Freivalds.  Non-

constructive methods for finite probabilistic automata. International Journal of Foundations of Computer Science (IJFCS), 19(3):565– 580, June 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Friedmann:2010:SSA

[Fri10]  Oliver Friedmann.  The

stevens-stirling-algorithm for solving parity games locally requires exponential time. International Journal of Foundations of Computer Science (IJFCS), 21(3):277– 287, June 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Freydenberger:2006:UMI

[FRS06] Dominik D. Freydenberger, Daniel Reidenbach, and Johannes C. Schneider. Unambiguous morphic images of strings. International Journal of Foundations of Computer Science (IJFCS), 17(3):601–??, June 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fulop:2024:SIA

[FRS24] Zoltán Fülöp, George Rahonis, , and Kai Salomaa. Special issue: Articles dedicated to the memory of Magnus Steinby — preface. International Journal of Foundations of Computer Science (IJFCS), 35(1–2): ??, January–February 2024. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123020057.

Freund:2019:VNE

[FRV19] Rudolf Freund, Vladimir Rogojin, and Sergey Verlan. Variants of networks of evolutionary processors with polarizations and a small number of processors. International Journal of Foundations of Computer Science (IJFCS), 30(6– 7):1005–1027, September– November 2019. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054119400264.

Flocchini:1998:TCS

[FS98] Paola Flocchini and Nicola Santoro. Topological constraints for sense of direction. International Journal of Foundations of Computer Science (IJFCS), 9(2): 179–??, June 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Franek:2005:SST

[FS05] Frantisek Franek and W. F. Smyth. Sorting suffixes of two-pattern strings. International Journal of Foundations of Computer Science (IJFCS), 16(6):1135– ??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Franek:2006:RSA

[FS06] Frantisek Franek and William F. Smyth. Reconstructing a suffix array. International Journal of Foundations of Computer Science (IJFCS), 17(6):1281–1295, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fleischer:2021:RLS

[FS21] Lukas Fleischer and Jeffrey Shallit. Recognizing lexicographically smallest words and computing successors in regular languages. International Journal of Foundations of Computer Science (IJFCS), 32(06):641– 662, September 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121420028.

Friese:2011:ENF

[FSM11] Sylvia Friese, Helmut Seidl, and Sebastian Maneth. Earliest normal form and minimization for bottom-up tree transducers. International Journal of Foundations of Computer Science (IJFCS), 22(7):1607–1623, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fazekas:2016:P

[FSTY16] Szilárd Zsolt Fazekas, Kayoko Shikishima-Tsuji, and Akihiro Yamamura. Preface. International Journal of Foundations of Computer Science (IJFCS), 27(3):301– ??, April 2016. CODEN IFCSEN. ISSN 0129-0541.

Feng:2011:GDA

[FSWF11] Xiutao Feng, Zhenqing Shi, Chuankun Wu, and Dengguo Feng. On guess and determine analysis of Rabbit. International Journal of Foundations of Computer Science (IJFCS), 22 (6):1283–1296, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Felscher:2009:CRC

[FT09] Ingo Felscher and Wolfgang Thomas. Compositionality and reachability with conditions on path lengths. International Journal of Foundations of Computer Science (IJFCS), 20(5):851– 868, October 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fujiwara:2011:LAO

[FT11] Akihiro Fujiwara and Takeshi Tateishi. Logic and arithmetic operations with a constant number of steps in membrane computing. International Journal of Foundations of Computer Science (IJFCS), 22(3):547– 564, April 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Filiot:2010:TAG

[FTT10] Emmanuel Filiot, Jean-Marc Talbot, and Sophie Tison. Tree automata with global constraints. International Journal of Foundations of Computer Science (IJFCS), 21(4):571– 596, August 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fujita:2016:PLG

[Fuj16] Satoshi Fujita. On the power of lookahead in greedy scheme for finding a minimum CDS for unit disk graphs. International Journal of Foundations of Computer Science (IJFCS), 27 (7):829–??, November 2016. CODEN IFCSEN. ISSN 0129-0541.

Fujiyoshi:2017:PAU

[Fuj17] Akio Fujiyoshi. A practical algorithm for the uniform membership problem of labeled multidigraphs of tree-width 2 for spanning tree automata. International Journal of Foundations of Computer Science (IJFCS), 28(5):563–??, August 2017. CODEN IFC-SEN. ISSN 0129-0541.

Fulop:2017:MZE

[Fül17] Zoltán Fülöp. In memoriam Zoltán Ésik (1951– 2016). International Journal of Foundations of Computer Science (IJFCS), 28(5):441– ??, August 2017. CODEN IFCSEN. ISSN 0129-0541.

Fulop:2024:CSR

[FV24] Zoltán Fülöp and Heiko Vogler. A comparison of sets of recognizable weighted tree languages over specific sets of bounded lattices. International Journal of Foundations of Computer Science (IJFCS), 35(1–2):51–76, January– February 2024. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123480039.

Farmer:1990:RCT

[FW90] William M. Farmer and Ronald J. Watro. Redex capturing in term graph rewriting. International Journal of Foundations of Computer Science (IJFCS), 1(4):369–??, December 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Filiot:2024:SCF

[FW24] Emmanuel Filiot and Sarah Winter. Synthesizing computable functions from rational specifications over infinite words. International Journal of Foundations of Computer Science (IJFCS), 35(1–2):179–214, January– February 2024. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412348009X.

Feng:2015:LFT

[FWZ15] Kai Feng, Shiying Wang, and Guozhen Zhang. Link failure tolerance in the arrangement graphs. International Journal of Foundations of Computer Science (IJFCS), 26(2):241–??, February 2015. CODEN IFCSEN. ISSN 0129-0541.

Franek:2008:ALB

[FY08] Frantisek Franek and Qian Yang. An asymptotic lower bound for the maximal number of runs in a string. International Journal of Foundations of Computer Science (IJFCS), 19(1):195– 203, February 2008. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Feng:2011:VBF

[FY11] Keqin Feng and Jing Yang.

Vectorial Boolean functions with good cryptographic properties. International Journal of Foundations of Computer Science (IJFCS), 22(6):1271–1282, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fang:2002:LIP

[FZ02] Qizhi Fang and Shanfeng Zhu. Linear and integer programming techniques for cooperative games. International Journal of Foundations of Computer Science (IJFCS), 13(5):653– ??, October 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fan:2003:OCN

[FZ03] Guangbin Fan and Jingyuan Zhang. Optimal cellular network deployment reusing existing base stations. International Journal of Foundations of Computer Science (IJFCS), 14(2):169– ??, April 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fearnley:2012:PMG

[FZ12] John Fearnley and Martin Zimmermann. Playing Muller games in a hurry. International Journal of Foundations of Computer Science (IJFCS), 23(3):649– ??, April 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fenner:2013:CHS

[FZ13] Stephen Fenner and Yong Zhang. On the complexity of the hidden subgroup problem. International Journal of Foundations of Computer Science (IJFCS), 24(8):1221–??, December 2013. CODEN IFC-SEN. ISSN 0129-0541.

Fenner:2015:QAS

[FZ15] Stephen Fenner and Yong Zhang. Quantum algorithms for a set of group theoretic problems. International Journal of Foundations of Computer Science (IJFCS), 26(2):255–??, February 2015. CODEN IFCSEN. ISSN 0129-0541.

Fernandez-Zepeda:2008:AAE

[FZAM08] José Alberto Fernández-Zepeda and Juan Paulo Alvarado-Magaña. Analysis of the average execution time for a selfstabilizing leader election algorithm. International Journal of Foundations of Computer Science (IJFCS), 19(6):1387–1402, December 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fernandez-Zepeda:2008:SML

[FZCFB08] José Alberto Fernández-Zepeda, Carlos Alberto Córdova-Flores, and Anu G. Bourgeois. Simulating an R-mesh on an LR-mesh in constant time. International Journal of Foundations of Computer Science (IJFCS), 19(6):1337–1354, December 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fernandez-Zepeda:2005:DFT

[FZEBB05] José Alberto Fernández-Zepeda, Alejandro Estrella-Balderrama, and Anu G. Bourgeois. Designing fault tolerant algorithms for reconfigurable meshes. International Journal of Foundations of Computer Science (IJFCS), 16(1):71–??, February 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fernandez-Zepeda:2005:ESA

[FZFDCHB05] José Alberto Fernández-Zepeda, Daniel Fajardo-Delgado, José Antonio Cárdenas-Haro, and Anu G. Bourgeois. Efficient simulation of an acyclic directed reconfigurable model on an undirected reconfigurable model. International Journal of Foundations of Computer Science (IJFCS), 16(1):55–??, February 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Fan:2014:NCI

[FZT14] Jia Fan, Yuliang Zheng, and Xiaohu Tang. A new construction of identity-based signcryption without random oracles. International Journal of Foundations of Computer Science (IJFCS), 25(1):1–??, January 2014. CODEN IFCSEN. ISSN 0129-0541.

Gazdag:2006:DSP

[Gaz06] Zsolt Gazdag. Decidability of the shape preserving property of bottom-up tree transducers. International Journal of Foundations of Computer Science (IJFCS), 17(2):395–??, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gheorghiu:2003:SFF

[GB03] Mihaela Gheorghiu and Janusz Brzozowski. Simulation of feedback-free circuits in the algebra of transients. International Journal of Foundations of Computer Science (IJFCS), 14(6):1033–??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gandhi:2015:AAS

[GC15] Ratnik Gandhi and Samaresh Chatterji. Applications of algebra for some game theoretic problems. International Journal of Foundations of Computer Science (IJFCS), 26(1):51–??, January 2015. CODEN IFC-SEN. ISSN 0129-0541.

Guellouma:2018:TAR

[GC18] Younes Guellouma and Hadda Cherroun. From tree automata to rational tree expressions. International Journal of Foundations of Computer Science (IJFCS), 29(6):1045– 1062, September 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905411850020X.

Gu:2020:CCH

[GCH20] Mei-Mei Gu, Jou-Ming Chang, and Rong-Xia Hao. On component connectivity of hierarchical star networks. International Journal of Foundations of Computer Science (IJFCS), 31 (03):313–326, April 2020. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054120500100.

Gupta:2008:MPC

[GCK08] Sunil Kumar Gupta, R. K. Chauhan, and Parveen Kumar. A minimum-process coordinated checkpointing protocol for mobile computing systems. International Journal of Foundations of Computer Science (IJFCS), 19(4):1015–1038, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Garcia:1998:SOR

[GD98] Roman Garcia and Jose Duato. Suboptimal-optimal routing for LAN internetworking using transparent bridges. International Journal of Foundations of Computer Science (IJFCS), 9(2): 139–??, June 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gudys:2012:PAC

[GD12] Adam Gudyś and Sebastian Deorowicz. A parallel algorithm for the constrained multiple sequence alignment problem designed for GPUs. International Journal of Foundations of Computer Science (IJFCS), 23(4):877–??, June 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gecseg:2007:CTL

[Géc07] Ferenc Gécseg. Classes of tree languages determined by classes of monoids. International Journal of Foundations of Computer Science (IJFCS), 18(6):1237–1246, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Grammatikakis:1998:CRP

[GFK98] Miltos D. Grammatikakis, Eric Fleury, and Miro Kraetzl. Continuous routing in packet switches. International Journal of Foundations of Computer Science (IJFCS), 9(2):121–??, June 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gonze:2019:IBC

[GGJ^{+}19] François Gonze, Vladimir V. Gusev, Raphaël M. Jungers, Balázs Gerencsér, and Mikhail V. Volkov. On the interplay between Černý and Babai’s conjectures. International Journal of Foundations of Computer Science (IJFCS), 30(1):93– 114, January 2019. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054119400057.

Ghasemi:2014:AFS

[GGR14] Taha Ghasemi, Hossein Ghasemalizadeh, and Mohammadreza Razzazi. An algorithmic framework for solving geometric covering problems — with applications. International Journal of Foundations of Computer Science (IJFCS), 25(5):623– ??, August 2014. CODEN IFCSEN. ISSN 0129-0541.

Gao:2007:SSP

[GH07] Yan Gao and Hendrik Jan Hoogeboom. P systems with single passenger carriers. International Journal of Foundations of Computer Science (IJFCS), 18 (6):1227–1235, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gudmundsson:2009:P

[GH09] Joachim Gudmundsson and James Harland. Preface. International Journal of Foundations of Computer Science (IJFCS), 20(1):1–2, February 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gruber:2013:PSR

[GH13] Hermann Gruber and Markus Holzer. Provably shorter regular expressions from finite automata. International Journal of Foundations of Computer Science (IJFCS), 24(8):1255–??, December 2013. CODEN IFC-SEN. ISSN 0129-0541.

Gruber:2015:FAR

[GH15] Hermann Gruber and Markus Holzer. From finite automata to regular expressions and back — a summary on descriptional complexity. International Journal of Foundations of Computer Science (IJFCS), 26(8):1009–??, December 2015. CODEN IFCSEN. ISSN 0129-0541.

Goddard:2005:SSA

[GHJS05] Wayne Goddard, Stephen T.

Hedetniemi, David P. Jacobs, and Pradip K. Srimani. Self-stabilizing algorithms for orderings and colorings. International Journal of Foundations of Computer Science (IJFCS), 16 (1):19–??, February 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Goc:2013:ATP

[GHS13] Daniel Goč, Dane Henshall, and Jeffrey Shallit. Automatic theorem-proving in combinatorics on words. International Journal of Foundations of Computer Science (IJFCS), 24(6):781–??, September 2013. CODEN IFCSEN. ISSN 0129-0541.

Geser:2005:FFA

[GHWZ05] Alfons Geser, Dieter Hofbauer, Johannes Waldmann, and Hans Zantema. Finding finite automata that certify termination of string rewriting systems. International Journal of Foundations of Computer Science (IJFCS), 16(3):471–??, June 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gelle:2019:RUF

[GI19] Kitti Gelle and Szabolcs Iván. Recognizing union– find trees is NP-Complete, even without rank info. International Journal of Foundations of Computer Science (IJFCS), 30(6– 7):1029–1045, September– November 2019. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054119400276.

Gelle:2022:DCR

[GI22] Kitti Gelle and Szabolcs Iván. Descriptive complexity of reversible languages having finitely many reduced automata. International Journal of Foundations of Computer Science (IJFCS), 33(3–4):247– 262, April–June 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122410040.

Giammarresi:2011:EIT

[Gia11] Dora Giammarresi. Exploring inside tiling recognizable picture languages to find deterministic subclasses. International Journal of Foundations of Computer Science (IJFCS), 22(7):1519– 1532, November 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gudmundsson:2007:P

[GJ07] Joachim Gudmundsson and Barry Jay. Preface. International Journal of Foundations of Computer Science (IJFCS), 18(2):195– 196, April 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gehrke:2018:PSU

[GJKS18] Jan Clemens Gehrke, Klaus Jansen, Stefan E. J. Kraft, and Jakob Schikowski. A PTAS for scheduling unrelated machines of few different types. International Journal of Foundations of Computer Science (IJFCS), 29(4):591–621, June 2018. CODEN IFCSEN. ISSN 0129-0541.

Gravier:2006:QOG

[GJMP06] Sylvain Gravier, Philippe Jorrand, Mehdi Mhalla, and Charles Payan. Quantum octal games. International Journal of Foundations of Computer Science (IJFCS), 17(4):919–??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ganzinger:2000:PA

[GJV00a] H. Ganzinger, F. Jacquemard, and M. Veanes. Part 1 (Asian ’98). International Journal of Foundations of Computer Science (IJFCS), 11(1):3–??, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ganzinger:2000:RRN

[GJV00b] H. Ganzinger, F. Jacquemard, and M. Veanes. Rigid reachability, the nonsymmetric form of rigid Eunification. International Journal of Foundations of Computer Science (IJFCS), 11(1):3–28, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Goresky:2011:SPA

[GK11] Mark Goresky and Andrew Klapper. Statistical properties of the arithmetic correlation of sequences. International Journal of Foundations of Computer Science (IJFCS), 22(6):1297–1315, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gasieniec:1999:MCL

[GKKP99] L. Ga̧sieniec, E. Kranakis, D. Krizanc, and A. Pelc. Minimizing congestion of layouts for ATM networks with faulty links. International Journal of Foundations of Computer Science (IJFCS), 10(4):503– 512, 1999. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gawrychowski:2010:FGR

[GKRS10] Paweł Gawrychowski, Dalia Krieger, Narad Rampersad, and Jeffrey Shallit. Finding the growth rate of a regular or context-free language in polynomial time. International Journal of Foundations of Computer Science (IJFCS), 21(4):597– 618, August 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Grigoriev:2017:YMP

[GKS17] Dima Grigoriev, Laszlo B. Kish, and Vladimir Shpilrain. Yao’s millionaires’ problem and public-key encryption without computational assumptions. International Journal of Foundations of Computer Science (IJFCS), 28(4):379–??, June 2017. CODEN IFCSEN. ISSN 0129-0541.

Grosse:2019:FAD

[GKS^{+}19] Ulrike Große, Christian Knauer, Fabian Stehn, Joachim Gudmundsson, and Michiel Smid. Fast algorithms for diameteroptimally augmenting paths and trees. International Journal of Foundations of Computer Science (IJFCS), 30(2):293–313, February 2019. ISSN 0129-0541.

Gao:2019:GAC

[GKSZ19] Alice L. L. Gao, Sergey Kitaev, Wolfgang Steiner, and Philip B. Zhang. On a greedy algorithm to construct universal cycles for permutations. International Journal of Foundations of Computer Science (IJFCS), 30(1):61– 72, January 2019. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054119400033.

Glockler:2007:FAU

[Glö07] Jens Glöckler. Forgetting automata and unary languages. International Journal of Foundations of Computer Science (IJFCS), 18 (4):813–827, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Glockler:2010:TDF

[Glö10] Jens Glöckler. A taxonomy of deterministic forgetting automata. International Journal of Foundations of Computer Science (IJFCS), 21(4):619–631, August 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gruska:2007:FSS

[GLP07] Jozef Gruska, Salvatore La Torre, and Mimmo Parente. The firing squad synchronization problem on squares, toruses and rings. International Journal of Foundations of Computer Science (IJFCS), 18(3):637– 654, June 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Guillon:2022:WSI

[GLPP22] Bruno Guillon, Giovanna J.

Lavado, Giovanni Pighizzini, and Luca Prigioniero. Weakly and strongly irreversible regular languages. International Journal of Foundations of Computer Science (IJFCS), 33(3–4):263– 284, April–June 2022. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122410052.

Garcia:2014:EDF

[GLV14] Pedro García, Damián López, and Manuel Vázquez De Parga. Efficient deterministic finite automata split-minimization derived from Brzozowski’s algorithm. International Journal of Foundations of Computer Science (IJFCS), 25 (6):679–??, September 2014. CODEN IFCSEN. ISSN 0129-0541.

Gorrieri:1990:THD

[GM90] Roberto Gorrieri and Ugo Montanari. Towards hierarchical description of systems: a proof system for strong prefixing. International Journal of Foundations of Computer Science (IJFCS), 1(3):277–??, September 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gorain:2019:AAB

[GM19] Barun Gorain and Partha Sarathi Mandal. Approximation algorithms for barrier sweep coverage. International Journal of Foundations of Computer Science (IJFCS), 30(3):425–448, 2019. ISSN 0129-0541.

Giambruno:2015:GGB

[GMNS15] Laura Giambruno, Sabrina Mantaci, Jean Néraud, and Carla Selmi. A generalization of Girod’s bidirectional decoding method to codes with a finite deciphering delay. International Journal of Foundations of Computer Science (IJFCS), 26(6):733– ??, September 2015. CO-DEN IFCSEN. ISSN 0129-0541.

Grigoriev:2015:NMS

[GMU15] Alexander Grigoriev, Bert Marchal, and Natalya Usotskaya. A note on the minimum H-subgraph edge deletion. International Journal of Foundations of Computer Science (IJFCS), 26(3):399– ??, April 2015. CODEN IFCSEN. ISSN 0129-0541.

Gergatsoulis:2004:PPT

[GN04] Manolis Gergatsoulis and Christos Nomikos. A proof procedure for temporal logic programming. International Journal of Foundations of Computer Science (IJFCS), 15(2):417–??, April 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gazdag:2011:KTB

[GN11] Zsolt Gazdag and Zoltán L.

Németh. A Kleene theorem for bisemigroup and binoid languages. International Journal of Foundations of Computer Science (IJFCS), 22(2):427– 446, February 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Guingne:2003:VOV

[GNC^{+}03] Franck Guingne, Florent Nicart, Jean-Marc Champarnaud, Lauri Karttunen, Tamás Gaál, and André Kempe. Virtual operations on virtual networks: the priority union. International Journal of Foundations of Computer Science (IJFCS), 14(6):1055– ??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Grabowski:2006:SAI

[GNP^{+}06] Szymon Grabowski, Gonzalo Navarro, RafaŁ Przywarski, Alejandro Salinger, and Veli Mäkinen. A simple alphabet-independent FMindex. International Journal of Foundations of Computer Science (IJFCS), 17 (6):1365–1384, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gebauer:2009:FET

[GO09] Heidi Gebauer and Yoshio Okamoto. Fast exponentialtime algorithms for the forest counting and the Tutte polynomial computation in graph classes. International Journal of Foundations of Computer Science (IJFCS), 20(1):25–44, February 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Goldwurm:1990:SLD

[Gol90] Massimiliano Goldwurm.

Some limit distributions in analysis of algorithms for problems on trace languages. International Journal of Foundations of Computer Science (IJFCS), 1 (3):265–??, September 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Golovnev:2014:AAT

[Gol14] Alexander Golovnev. Approximating asymmetric TSP in exponential time. International Journal of Foundations of Computer Science (IJFCS), 25(1):89– ??, January 2014. CODEN IFCSEN. ISSN 0129-0541.

Gonzalez:2001:SMM

[Gon01] Teofilo F. Gonzalez. On solving multimessage multicasting problems. International Journal of Foundations of Computer Science (IJFCS), 12(6):791– 808, December 2001. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Geffert:2008:P

[GP08] Viliam Geffert and Giovanni Pighizzini. Preface. International Journal of Foundations of Computer Science (IJFCS), 19(4):747– 749, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Groote:2009:SG

[GP09] Jan Friso Groote and Bas Ploeger. Switching graphs. International Journal of Foundations of Computer Science (IJFCS), 20(5):869– 886, October 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Geffert:2013:UCN

[GP13] Viliam Geffert and Dana Pardubská. Unary coded NP-complete languages in ASPACE(log log n). International Journal of Foundations of Computer Science (IJFCS), 24(7):1167– ??, November 2013. CO-DEN IFCSEN. ISSN 0129-0541.

Gusev:2015:RTA

[GP15] Vladimir V. Gusev and Elena V. Pribavkina. Reset thresholds of automata with two cycle lengths. International Journal of Foundations of Computer Science (IJFCS), 26(7):953–??, November 2015. CODEN IFCSEN. ISSN 0129-0541.

Gurski:2017:IRS

[GP17] Frank Gurski and Patrick Gwydion Poullie. Interval routing schemes for circulararc graphs. International Journal of Foundations of Computer Science (IJFCS), 28(1):30–60, January 2017. CODEN IFCSEN. ISSN 0129-0541.

Galle:2009:PUS

[GPC09] Matthias Gallé, Pierre Peterlongo, and François Coste. In-place update of suffix array while recoding words. International Journal of Foundations of Computer Science (IJFCS), 20(6):1025–1045, December 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Guillon:2020:NSE

[GPP20] Bruno Guillon, Giovanni Pighizzini, and Luca Prigioniero. Non-self-embedding grammars, constant-height pushdown automata, and limited automata. International Journal of Foundations of Computer Science (IJFCS), 31(08):1133– 1157, December 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120420071.

Gheorghe:2013:RFM

[GPPJR13] Marian Gheorghe, Gheorghe Păun, Mario J. Pérez-Jiménez, and Grzegorz Rozenberg. Research frontiers of membrane computing: Open problems and research topics. International Journal of Foundations of Computer Science (IJFCS), 24(5):547–??, August 2013. CODEN IFCSEN. ISSN 0129-0541.

Goc:2014:NSC

[GPS14] Daniel Goč, Alexandros Palioudakis, and Kai Salomaa. Nondeterministic state complexity of proportional removals. International Journal of Foundations of Computer Science (IJFCS), 25(7):823–??, November 2014. CODEN IFCSEN. ISSN 0129-0541.

Gruska:2015:PQF

[GQZ15] Jozef Gruska, Daowen Qiu, and Shenggen Zheng. Potential of quantum finite automata with exact acceptance. International Journal of Foundations of Computer Science (IJFCS), 26 (3):381–??, April 2015. CO-DEN IFCSEN. ISSN 0129-0541.

Golumbic:2000:CWS

[GR00] M. C. Golumbic and U. Rotics. On the clique-width of some perfect graph classes. International Journal of Foundations of Computer Science (IJFCS), 11(3):423– 444, 2000. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Grimmell:2003:SBR

[GR03] William C. Grimmell and Nageswara S. V. Rao. On source-based route computation for quickest paths under dynamic bandwidth constraints. International Journal of Foundations of Computer Science (IJFCS), 14(3):503–??, June 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gradel:1990:NLT

[Gra90] Erich Gradel. On the notion of linear time computability. International Journal of Foundations of Computer Science (IJFCS), 1 (3):295–??, September 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Grassl:2003:EQC

[GRB03] Markus Grassl, Martin Rötteler, and Thomas Beth. Efficient quantum circuits for non-qubit quantum error-correcting codes. International Journal of Foundations of Computer Science (IJFCS), 14(5):757– ??, October 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Greenlaw:1996:SID

[Gre96] R. Greenlaw. Subtree isomorphism is in DLOG for nested trees. International Journal of Foundations of Computer Science (IJFCS), 7(2):161–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Greeni:2024:EAC

[GRI24] A. Berin Greeni, R. Sundara Rajan, , and Paul Immanuel. Embedding augmented cube into certain trees and windmill graphs. International Journal of Foundations of Computer Science (IJFCS), 35(04): 409–434, June 2024. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054123500090.

Grover:2003:IQS

[Gro03] Lov K. Grover. An improved quantum scheduling algorithm. International Journal of Foundations of Computer Science (IJFCS), 14 (5):715–??, October 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Goc:2014:NAB

[GRRS14] Daniel Goč, Narad Rampersad, Michel Rigo, and Pavel Salimov. On the number of Abelian bordered words (with an example of automatic theorem-proving). International Journal of Foundations of Computer Science (IJFCS), 25(8):1097– ??, December 2014. CODEN IFCSEN. ISSN 0129-0541.

Gabric:2021:INP

[GRS21] Daniel Gabric, Narad Rampersad, and Jeffrey Shallit. An inequality for the number of periods in a word. International Journal of Foundations of Computer Science (IJFCS), 32(05):597– 614, August 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121410094.

Geeraerts:2010:ECM

[GRV10] Gilles Geeraerts, Jean-François Raskin, and Laurent Van Begin. On the efficient computation of the minimal coverability set of Petri nets. International Journal of Foundations of Computer Science (IJFCS), 21(2):135–165, April 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gudmundsson:2009:SGG

[GS09] Joachim Gudmundsson and Michiel Smid. On spanners of geometric graphs. International Journal of Foundations of Computer Science (IJFCS), 20(1):135– 149, February 2009. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Goddard:2012:SSM

[GS12a] Wayne Goddard and Pradip K. Srimani. Self-stabilizing master-slave token circulation and efficient sizecomputation in a unidirectional ring of arbitrary size. International Journal of Foundations of Computer Science (IJFCS), 23(4):763– ??, June 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gorbunova:2012:PWA

[GS12b] Irina A. Gorbunova and Arseny M. Shur. On Pansiot words avoiding 3repetitions. International Journal of Foundations of Computer Science (IJFCS), 23(8):1583–??, December 2012. CODEN IFCSEN. ISSN 0129-0541.

Gasnikov:2018:SFP

[GS18] Daniil Gasnikov and Arseny M. Shur. Squarefree partial words with many wildcards. International Journal of Foundations of Computer Science (IJFCS), 29(5):846– 860, August 2018. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054118420078.

Ghosh:2003:NAE

[GSD03] Sasthi C. Ghosh, Bhabani P.

Sinha, and Nabanita Das. A new approach to efficient channel assignment for hexagonal cellular networks. International Journal of Foundations of Computer Science (IJFCS), 14(3):439– ??, June 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Greco:1999:GAO

[GSZ99] S. Greco, D. Saccà, and C. Zaniolo. Grammars and automata to optimize chain logic queries. International Journal of Foundations of Computer Science (IJFCS), 10(3):349–372, 1999. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Glasser:2009:ICC

[GSZ09] Christian Glaßer, Alan L.

Selman, and Liyu Zhang. The informational content of canonical disjoint NPpairs. International Journal of Foundations of Computer Science (IJFCS), 20(3):501– 522, June 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gazdag:2019:PLP

[GTCV19] Zsolt Gazdag, Krisztián Tichler, and Erzsébet Csuhaj-Varjú. A pumping lemma for permitting semi-conditional languages. International Journal of Foundations of Computer Science (IJFCS), 30(1):73–92, January 2019. ISSN 0129-0541. URL

                                        https://www.worldscientific.
                                        com/doi/10.1142/S0129054119400045.

Guan:2021:LKA

[Gua21] Albert Guan. A lightweight key agreement protocol with authentication capability. International Journal of Foundations of Computer Science (IJFCS), 32(04): 389–404, June 2021. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054121500222.

Gurevich:2016:PP

[Gur16] Yuri Gurevich. Past present. International Journal of Foundations of Computer Science (IJFCS), 27(2):103– ??, February 2016. CODEN IFCSEN. ISSN 0129-0541.

Gusev:2013:LBL

[Gus13] Vladimir V. Gusev. Lower bounds for the length of reset words in Eulerian automata. International Journal of Foundations of Computer Science (IJFCS), 24 (2):251–??, February 2013. CODEN IFCSEN. ISSN 0129-0541.

Giakoumakis:2003:LTR

[GV03] V. Giakoumakis and J. M. Vanherpe. Linear time recognition and optimizations for weak-bisplit graphs, bi-cographs and bipartite P_{6}-free graphs. International Journal of Foundations of Computer Science (IJFCS), 14(1):107–??, February 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gazdag:2023:CHC

[GV23] Zsolt Gazdag and Sándor Vágvölgyi. The component hierarchy of chain-free cooperating distributed regular tree grammars revisited. International Journal of Foundations of Computer Science (IJFCS), 34(04): 411–427, June 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412250023X.

Giraud:2007:PED

[GVL07] Mathieu Giraud, Phillipe Veber, and Dominique Lavenier. Path-equivalent developments in acyclic weighted automata. International Journal of Foundations of Computer Science (IJFCS), 18(4):799–811, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gazda:2018:CYO

[GW18] M. W. Gazda and T. A. C.

Willemse. Cooking your own parity game preorders through matching plays. International Journal of Foundations of Computer Science (IJFCS), 29(4):571– 590, June 2018. CODEN IFCSEN. ISSN 0129-0541.

Guo:2024:BEV

[GW24] Fei Guo and Zilong Wang. Balanced even-variable rotation symmetric Boolean functions with optimal algebraic immunity, maximum algebraic degree and higher nonlinearity. International Journal of Foundations of Computer Science (IJFCS), 35(03):245–270, April 2024. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123500016.

Guo:2024:EHC

[GWF^{+}24] Ruyan Guo, Yan Wang, Jianxi Fan, , and Weibei Fan. Embedding hierarchical cubic networks into krooted complete binary trees for minimum wirelength. International Journal of Foundations of Computer Science (IJFCS), 35(03):327– 352, April 2024. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123500053.

Gao:2002:DRS

[GWL02] W. Gao, S. Wang, and B. Liu. A dynamic recommendation system based on log mining. International Journal of Foundations of Computer Science (IJFCS), 13(4):521–??, 2002. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Guo:2017:FFU

[GWL^{+}17] Lili Guo, Xi Wang, Cheng-Kuan Lin, Jingya Zhou, and Jianxi Fan. A fault-free unicast algorithm in the generalized hypercube with restricted faulty vertices. International Journal of Foundations of Computer Science (IJFCS), 28(7):915–??, November 2017. CODEN IFCSEN. ISSN 0129-0541.

Gao:2012:SCA

[GY12] Yuan Gao and Sheng Yu.

State complexity and approximation. International Journal of Foundations of Computer Science (IJFCS), 23(5):1085–??, August 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Gimbert:2012:BOS

[GZ12] Hugo Gimbert and Wiesław Zielonka. Blackwell optimal strategies in priority meanpayoff games. International Journal of Foundations of Computer Science (IJFCS), 23(3):687–??, April 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ge:2024:SFT

[GZY24] Huifen Ge, Shumin Zhang, , and Chengfu Ye. The structure fault tolerance of alternating group networks. International Journal of Foundations of Computer Science (IJFCS), 35(04):483– 500, June 2024. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054123500120.

Guo:2021:REE

[GZZX21] Litao Guo, Mingzu Zhang, Shaohui Zhai, and Liqiong Xu. Relation of extra edge connectivity and component edge connectivity for regular networks. International Journal of Foundations of Computer Science (IJFCS), 32(02):137– 149, February 2021. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054121500076.

Han:2013:IPF

[Han13]  Yo-Sub Han.  An im-

proved prefix-free regularexpression matching. International Journal of Foundations of Computer Science (IJFCS), 24(5):679–??, August 2013. CODEN IFC-SEN. ISSN 0129-0541.

Harju:2024:NSB

[Har24] Tero Harju. A note on squares in binary words. International Journal of Foundations of Computer Science (IJFCS), 35(1–2): 101–106, January–February 2024. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054123480052.

Hashimoto:2000:FCC

[Has00] M. Hashimoto. First-class contexts in ML. International Journal of Foundations of Computer Science (IJFCS), 11(1):65–88, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hakem:2006:CPS

[HB06] Mourad Hakem and Franck Butelle. Critical path scheduling parallel programs on an unbounded number of processors. International Journal of Foundations of Computer Science (IJFCS), 17(2):287– ??, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Higa:2008:RST

[HBIT08] Yasuto Higa, Hideo Bannai, Shunsuke Inenaga, and Masayuki Takeda. Reachability on suffix tree graphs. International Journal of Foundations of Computer Science (IJFCS), 19(1):147– 162, February 2008. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hamrouni:2008:SMG

[HBN08] Tarek Hamrouni, Sadok Ben Yahia, and Engelbert Mephu Nguifo. Succinct minimal generators: Theoretical foundations and applications. International Journal of Foundations of Computer Science (IJFCS), 19(2):271– 296, April 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hernandez-Castro:2012:MTA

[HCETPL^{+}12] Julio Cesar Hernandez-Castro, Juan Manuel Estevez-Tapiador, Pedro Peris-Lopez, John A. Clark, and El-Ghazali Talbi. Metaheuristic traceability attack against SLMAP, an RFID lightweight authentication protocol. International Journal of Foundations of Computer Science (IJFCS), 23(2):543– 553, February 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hanks:1996:FTV

[HCG96] L. Hanks, R. K. Cytron, and W. Gillett. On finding topologically valid matchings in restriction-fragment maps. International Journal of Foundations of Computer Science (IJFCS), 7 (1):59–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Head:2011:CLT

[Hea11]  Tom Head.  Computing

with light: Toward parallel Boolean algebra. International Journal of Foundations of Computer Science (IJFCS), 22(7):1625–1637, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Heirich:1997:SDA

[Hei97] A. Heirich. A scalable diffusion algorithm for dynamic mapping and load balancing on networks of arbitrary topology. International Journal of Foundations of Computer Science (IJFCS), 8(3):329–??, September 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Henriksen:2002:EET

[Hen02] J. G. Henriksen. An expressive extension of TLC. International Journal of Foundations of Computer Science (IJFCS), 13(3):341–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hinze:2009:RMC

[HFLD09] Thomas Hinze, Raffael Fassler, Thorsten Lenser, and Peter Dittrich. Register machine computations on binary numbers by oscillating and catalytic chemical reactions modelled using mass-action kinetics. International Journal of Foundations of Computer Science (IJFCS), 20(3):411– 426, June 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hu:2011:PTK

[HG11] Honggang Hu and Guang Gong. Periods on two kinds of nonlinear feedback shift registers with time varying feedback functions. International Journal of Foundations of Computer Science (IJFCS), 22(6):1317–1329, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Halava:2011:RTB

[HH11] Vesa Halava and Štěpán Holub. Reduction tree of the binary generalized post correspondence problem. International Journal of Foundations of Computer Science (IJFCS), 22 (2):473–490, February 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hadravova:2012:LSB

[HH12] Jana Hadravová and Štĕpán Holub. Large simple binary equality words. International Journal of Foundations of Computer Science (IJFCS), 23(6):1385– ??, September 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hospodar:2020:RAS

[HH20] Michal Hospodár and Markus Holzer. The ranges of accepting state complexities of languages resulting from some operations. International Journal of Foundations of Computer Science (IJFCS), 31(08):1159– 1177, December 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120420083.

Han:2022:MEP

[HH22] Yijie Han and Xin He. More efficient parallel integer sorting. International Journal of Foundations of Computer Science (IJFCS), 33(05): 411–427, August 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500071.

Halava:2024:BGP

[HH24] Vesa Halava and Štěpán Holub. Binary generalized PCP for two periodic morphisms is decidable in polynomial time. International Journal of Foundations of Computer Science (IJFCS), 35(1–2):129–144, January– February 2024. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054123480076.

Halava:2007:UBI

[HHH07] Vesa Halava, Tero Harju, and Mika Hirvensalo. Undecidability bounds for integer matrices using Claus instances. International Journal of Foundations of Computer Science (IJFCS), 18 (5):931–948, October 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hemaspaandra:1995:NSS

[HHN^{+}95] L. A. Hemaspaandra, A. Hoene, A. V. Naik, et al. Nondeterministically selective sets. International Journal of Foundations of Computer Science (IJFCS), 6 (4):403–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Halava:2023:IWA

[HHNP23] Vesa Halava, Tero Harju, Reino Niskanen, and Igor Potapov. Integer weighted automata on infinite words. International Journal of Foundations of Computer Science (IJFCS), 34(2–3): 163–182, February–April 2023. ISSN 0129-0541. URL

                                        https://www.worldscientific.
                                        com/doi/10.1142/S0129054122440014.

Huang:2017:BBC

[HHP17] Jianye Huang, Qiong Huang, and Chunhua Pan. A black-box construction of strongly unforgeable signature scheme in the leakage setting. International Journal of Foundations of Computer Science (IJFCS), 28 (6):761–??, September 2017. CODEN IFCSEN. ISSN 0129-0541.

Hemaspaandra:1999:SSM

[HHW99] L. A. Hemaspaandra, H. Hempel, and G. Wechsung. Selfspecifying machines. International Journal of Foundations of Computer Science (IJFCS), 10(3):263– 276, 1999. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hyyro:2018:DRC

[HI18] Heikki Hyyrö and Shunsuke Inenaga. Dynamic RLEcompressed edit distance tables under general weighted cost functions. International Journal of Foundations of Computer Science (IJFCS), 29(4):623–645, June 2018. CODEN IFCSEN. ISSN 0129-0541.

Harvath:2001:CPP

[HIIW01] Géza Harváth, Katsushi Inoue, Akira Ito, and Y. Wang. Closure property of probabilistic Turing machines and alternating Turing machines with sublogarithmic spaces. International Journal of Foundations of Computer Science (IJFCS), 12(3):397– 409, 2001. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hinze:2001:SIF

[Hin01] R. Hinze. Special issue on functional and logic programming — part 2. International Journal of Foundations of Computer Science (IJFCS), 12(2):125–??, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hirokawa:1991:PTA

[Hir91]  Sachio Hirokawa.  Prin-

cipal type assignment to lambda terms. International Journal of Foundations of Computer Science (IJFCS), 2(2):149–162, June 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hromkovic:1992:POW

[HIR^{+}92] J. Hromkovic, K. Inoue, B. Rovan, et al. On the power of one-way synchronized alternating machines with small space. International Journal of Foundations of Computer Science (IJFCS), 3(1):65–80, March 1992. CODEN IFCSEN.

ISSN 0129-0541 (print), 1793-6373 (electronic).

Hittmeir:2020:RIF

[Hit20] Markus Hittmeir. A reduction of integer factorization to modular tetration. International Journal of Foundations of Computer Science (IJFCS), 31(04): 461–481, June 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500197.

Hemachandra:1991:LLR

[HJ91] Lane A. Hemachandra and Sanjay Jain. On the limitations of locally robust positive reductions. International Journal of Foundations of Computer Science (IJFCS), 2(3):237–256, September 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hemaspaandra:1997:LRM

[HJ97] Lane A. Hemaspaandra and Zhigen Jiang. Logspace reducibility: Models and equivalences. International Journal of Foundations of Computer Science (IJFCS), 8(1):95–??, March 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Holzer:2013:EAE

[HJ13] Markus Holzer and Sebastian Jakobi. From equivalence to almost-equivalence, and beyond: Minimizing automata with errors. International Journal of Foundations of Computer Science (IJFCS), 24(7):1083– ??, November 2013. CO-DEN IFCSEN. ISSN 0129-0541.

Holzer:2014:NBT

[HJ14] Markus Holzer and Sebastian Jakobi. Nondeterministic biautomata and their descriptional complexity. International Journal of Foundations of Computer Science (IJFCS), 25(7):837–??, November 2014. CODEN IFCSEN. ISSN 0129-0541.

Holzer:2016:MHM

[HJ16] Markus Holzer and Sebastian Jakobi. Minimal and hyper-minimal biautomata. International Journal of Foundations of Computer Science (IJFCS), 27(2):161– ??, February 2016. CODEN IFCSEN. ISSN 0129-0541.

Holzer:2017:MMF

[HJ17] Markus Holzer and Sebastian Jakobi. More on minimizing finite automata with errors — nondeterministic machines. International Journal of Foundations of Computer Science (IJFCS), 28(3):229–??, April 2017. CODEN IFC-SEN. ISSN 0129-0541.

Holzer:2012:MNP

[HJK12] Markus Holzer, Sebastian Jakobi, and Martin Kutrib. The magic number problem for subregular language families. International Journal of Foundations of Computer Science (IJFCS), 23(1):115– 131, January 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Holzer:2018:MRD

[HJK18] Markus Holzer, Sebastian Jakobi, and Martin Kutrib. Minimal reversible deterministic finite automata. International Journal of Foundations of Computer Science (IJFCS), 29(2):251–??, February 2018. CODEN IFCSEN. ISSN 0129-0541.

Hospodar:2019:DCF

[HJM19] Michal Hospodár, Galina Jirásková, and Peter Mlynárčik. Descriptional complexity of the forever operator. International Journal of Foundations of Computer Science (IJFCS), 30(1):115– 134, January 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400069.

Harren:2013:TOT

[HJP^{+}13] Rolf Harren, Klaus Jansen, Lars Prädel, Ulrich M. Schwarz, and Rob Van Stee. Two for one: Tight approximation of 2D bin packing. International Journal of Foundations of Computer Science (IJFCS), 24(8): 1299–??, December 2013. CODEN IFCSEN. ISSN 0129-0541.

Hemaspaandra:1993:BRT

[HJV93] L. A. Hemaspaandra, S. Jain, and N. K. Vereshchagin. Banishing robust Turing completeness. International Journal of Foundations of Computer Science (IJFCS), 4(3):245–266, September 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Henzinger:2011:FSM

[HJW11] Thomas Henzinger, Barbara Jobstmann, and Verena Wolf. Formalisms for specifying Markovian population models. International Journal of Foundations of Computer Science (IJFCS), 22(4):823–841, June 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hayashi:1995:NFF

[HK95] S. Hayashi and S. Kobayashi. A new formalization of Feferman’s system of functions and classes and its relation to Frege structure. International Journal of Foundations of Computer Science (IJFCS), 6(3):187–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Harel:2002:SSB

[HK02] D. Harel and H. Kugler.

Synthesizing state-based object systems from LSC specifications. International Journal of Foundations of Computer Science (IJFCS), 13(1):5–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Holzer:2003:NDC

[HK03] Markus Holzer and Martin Kutrib. Nondeterministic descriptional complexity of regular languages. International Journal of Foundations of Computer Science (IJFCS), 14(6):1087– ??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Harju:2008:P

[HK08] Tero Harju and Juhani Karhumäki. Preface. International Journal of Foundations of Computer Science (IJFCS), 19(3):495– 496, June 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hempel:2009:APC

[HK09a] Harald Hempel and Madlen Kimmritz. Aspects of persistent computations. International Journal of Foundations of Computer Science (IJFCS), 20(4):701– 715, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Holzer:2009:NFA

[HK09b] Markus Holzer and Martin Kutrib. Nondeterministic finite automata — recent results on the descriptional and computational complexity. International Journal of Foundations of Computer Science (IJFCS), 20(4):563– 580, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Holzer:2011:CRL

[HK11] Markus Holzer and Martin Kutrib. The complexity of regular(-like) expressions. International Journal of Foundations of Computer Science (IJFCS), 22 (7):1533–1548, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Holzer:2015:P

[HK15] Markus Holzer and Martin Kutrib. Preface. International Journal of Foundations of Computer Science (IJFCS), 26(7):803–??, November 2015. CODEN IFCSEN. ISSN 0129-0541.

Holzer:2019:OTN

[HK19] Markus Holzer and Martin Kutrib. One-time nondeterministic computations. International Journal of Foundations of Computer Science (IJFCS), 30(6– 7):1069–1089, September– November 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905411940029X.

Holzer:2021:P

[HK21] Markus Holzer and Martin Kutrib. Preface. International Journal of Foundations of Computer Science (IJFCS), 32(05):417– 418, August 2021. CO-DEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121020019.

Henriksson:2023:FPF

[HK23] Viktor Henriksson and Manfred Kufleitner. Forbidden patterns for FO^{2} alternation over finite and infinite words. International Journal of Foundations of Computer Science (IJFCS), 34(2–3):183–224, February–April 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123440021.

Hromkovic:2013:DVN

[HKKŠ13] Juraj Hromkovič, Rastislav Královič, Richard Královič, and Richard Štefanec. Determinism vs. nondeterminism for two-way automata: Representing the meaning of states by logical formulæ. International Journal of Foundations of Computer Science (IJFCS), 24(7):955– ??, November 2013. CO-DEN IFCSEN. ISSN 0129-0541.

Holzer:2022:IDD

[HKMW22] Markus Holzer, Martin Kutrib, Andreas Malcher, and Matthias Wendlandt. Input-driven double-head pushdown automata. International Journal of Foundations of Computer Science (IJFCS), 33(3–4):285– 311, April–June 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122410064.

Han:2016:SCI

[HKNS16] Yo-Sub Han, Sang-Ki Ko, Timothy Ng, and Kai Salomaa. State complexity of insertion. International Journal of Foundations of Computer Science (IJFCS), 27 (7):863–??, November 2016. CODEN IFCSEN. ISSN 0129-0541.

Han:2019:SAR

[HKRS19] Yo-Sub Han, Hwee Kim, Trent A. Rogers, and Shinnosuke Seki. Self-attraction removal from Oritatami systems. International Journal of Foundations of Computer Science (IJFCS), 30(6– 7):1047–1067, September– November 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400288.

Han:2013:EDB

[HKS13] Yo-Sub Han, Sang-Ki Ko, and Kai Salomaa. The editdistance between a regular language and a contextfree language. International Journal of Foundations of Computer Science (IJFCS), 24(7):1067–??, November 2013. CODEN IFCSEN. ISSN 0129-0541. Halldorsson:2000:MID

[HKT00] M. M. Halldórsson, J. Kratochvíl, and J. A. Telle. Mod-2 independence and domination in graphs. International Journal of Foundations of Computer Science (IJFCS), 11(3):355– 364, 2000. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hadid:2017:SAF

[HKV17] Rachid Hadid, Mehmet Hakan Karaata, and Vincent Villain. A stabilizing algorithm for finding two nodedisjoint paths in arbitrary networks. International Journal of Foundations of Computer Science (IJFCS), 28(4):411–??, June 2017. CODEN IFCSEN. ISSN 0129-0541.

Hon:2001:ANN

[HL01] Wing-Kai Hon and Tak-Wah Lam. Approximating the nearest neighbor interchange distance for nonuniform-degree evolutionary trees. International Journal of Foundations of Computer Science (IJFCS), 12 (4):533–550, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ho:2004:DCP

[HL04] Kevin I.-J. Ho and Joseph Y.-T. Leung. A dual criteria preemptive scheduling problem for minimax error of imprecise computation tasks. International Journal of Foundations of Computer Science (IJFCS), 15(5):717– ??, October 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Healy:2006:TFP

[HL06] Patrick Healy and Karol Lynch. Two fixed-parameter tractable algorithms for testing upward planarity. International Journal of Foundations of Computer Science (IJFCS), 17(5): 1095–??, October 2006. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hu:2019:AAA

[HLC^{+}19] Yupu Hu, Zhizhu Lian, Jiangshan Chen, Baocang Wang, and Shanshan Zhang. Algebraic attacks against several weak variants of GVW 13 ABE. International Journal of Foundations of Computer Science (IJFCS), 30(4):607– 618, June 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905411940015X.

Hwang:2019:ELS

[HLH19] Min-Shiang Hwang, Cheng-Chi Lee, and Shih-Ting Hsu. An ElGamal-like secure channel free public key encryption with keyword search scheme. International Journal of Foundations of Computer Science (IJFCS), 30(2):255– 273, February 2019. ISSN 0129-0541.

Hsu:2006:SCS

[HLHH06] Hong-Chun Hsu, Cheng-Kuan Lin, Hua-Min Hung, and Lih-Hsing Hsu. The spanning connectivity of the (n, k)-star graphs. International Journal of Foundations of Computer Science (IJFCS), 17(2):415– ??, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Huo:2009:PSA

[HLW09] Yumei Huo, Joseph Y.-T. Leung, and Xin Wang. Preemptive scheduling algorithms with nested processing set restriction. International Journal of Foundations of Computer Science (IJFCS), 20(6):1147–1160, December 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hon:2004:STD

[HLY^{+}04] Wing-Kai Hon, Tak-Wah Lam, Siu-Ming Yiu, Ming-Yang Kao, and Wing-Kin Sung. Subtree transfer distance for degree-D phylogenies. International Journal of Foundations of Computer Science (IJFCS), 15(6):893– ??, December 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hong:2004:AWS

[HM04] Dawei Hong and Shushuang Man. Analysis of Web search algorithm hits. International Journal of Foundations of Computer Science (IJFCS), 15(4):649– ??, August 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Han:2005:AAM

[HMZ05] Yonghua Han, Bin Ma, and Kaizhong Zhang. An automata approach to match gapped sequence tags against protein database. International Journal of Foundations of Computer Science (IJFCS), 16(3):487– ??, June 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Harju:2004:MDE

[HN04] Tero Harju and Dirk Nowotka.

Minimal Duval extensions. International Journal of Foundations of Computer Science (IJFCS), 15(2):349– ??, April 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hyyro:2006:BPC

[HN06] Heikki Hyyrö and Gonzalo Navarro. Bit-parallel computation of local similarity score matrices with unitary weights. International Journal of Foundations of Computer Science (IJFCS), 17(6):1325–1344, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Holub:2010:RBP

[HN10] Štěpán Holub and Dirk Nowotka. On the relation between periodicity and unbordered factors of finite words. International Journal of Foundations of Computer Science (IJFCS), 21 (4):633–645, August 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Holzrichter:1999:GBD

[HO99] M. Holzrichter and S. Oliveira. A graph based Davidson algorithm for the graph partitioning problem. International Journal of Foundations of Computer Science (IJFCS), 10(2):223–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hsiang:2000:SIA

[HO00] J. Hsiang and A. Ohori. Special issue on Advances in Computing Science — Asian ’98. International Journal of Foundations of Computer Science (IJFCS), 11(1):1–??, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Holub:2005:F

[Hol05] Jan Holub. Foreword. International Journal of Foundations of Computer Science (IJFCS), 16(6):1071– ??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Holub:2006:F

[Hol06] Jan Holub. Foreword. International Journal of Foundations of Computer Science (IJFCS), 17(6):1233–1234, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Holub:2008:F

[Hol08] Jan Holub. Foreword. International Journal of Foundations of Computer Science (IJFCS), 19(1):1–3, February 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Holub:2009:F

[Hol09] Jan Holub. Foreword. International Journal of Foundations of Computer Science (IJFCS), 20(6):965– 966, December 2009. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Holub:2011:BMS

[Hol11] Štěpán Holub. Binary morphisms with stable suffix complexity. International Journal of Foundations of Computer Science (IJFCS), 22(3):699–712, April 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Holub:2012:P

[Hol12] Jan Holub. Preface. International Journal of Foundations of Computer Science (IJFCS), 23(2):247– 248, February 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Honkala:2002:RCD

[Hon02] Juha Honkala. Remarks concerning the D0L ωequivalence problem. International Journal of Foundations of Computer Science (IJFCS), 13(5):769– ??, October 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Honkala:2006:BPD

[Hon06] Juha Honkala. The base problem for D0L Parikh sets. International Journal of Foundations of Computer Science (IJFCS), 17(2):465– ??, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Honkala:2007:DEP

[Hon07] Juha Honkala. The D0L ωequivalence problem. International Journal of Foundations of Computer Science (IJFCS), 18(1):181– 194, February 2007. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Honkala:2012:ESM

[Hon12]  Juha Honkala.  Equal-

ity sets of morphic word sequences. International Journal of Foundations of Computer Science (IJFCS), 23(8):1749–??, December 2012. CODEN IFCSEN. ISSN 0129-0541.

Halava:2008:P

[HP08] Vesa Halava and Igor Potapov. Preface. International Journal of Foundations of Computer Science (IJFCS), 19(4):915– 917, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Halava:2009:P

[HP09a] Vesa Halava and Igor Potapov. Preface. International Journal of Foundations of Computer Science (IJFCS), 20(5):775– 777, October 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Heuberger:2009:ACM

[HP09b] Clemens Heuberger and Helmut Prodinger. Analysis of complements in multiexponentiation algorithms using signed digit representations. International Journal of Foundations of Computer Science (IJFCS), 20(3):443–453, June 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Herley:1999:DBB

[HPP99] K. T. Herley, A. Pietracaprina, and G. Pucci. Deterministic branch-and-bound on distributed memory machines. International Journal of Foundations of Computer Science (IJFCS), 10 (4):391–404, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Habib:1999:PRT

[HPV99] M. Habib, C. Paul, and L. Viennot. Partition refinement techniques: An interesting algorithmic tool kit. International Journal of Foundations of Computer Science (IJFCS), 10 (2):147–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Han:2017:Pb

[HRS17] Jinguang Han, Yogachandran Rahulamathavan, and Willy Susilo. Preface. International Journal of Foundations of Computer Science (IJFCS), 28(6):641–??, September 2017. CODEN IFCSEN. ISSN 0129-0541.

Hintikka:1995:WLP

[HS95] J. Hintikka and G. Sandu. What is the logic of parallel processing? International Journal of Foundations of Computer Science (IJFCS), 6(1):27–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Han:2008:SCU

[HS08] Yo-Sub Han and Kai Salomaa. State complexity of union and intersection of finite languages. International Journal of Foundations of Computer Science (IJFCS), 19(3):581– 595, June 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Han:2011:OFL

[HS11] Yo-Sub Han and Kai Salomaa. Overlap-free languages and solid codes. International Journal of Foundations of Computer Science (IJFCS), 22(5):1197– 1209, August 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Han:2017:Pa

[HS17] Yo-Sub Han and Kai Salomaa. Preface. International Journal of Foundations of Computer Science (IJFCS), 28(5):437–??, August 2017. CODEN IFC-SEN. ISSN 0129-0541.

Herlihy:2007:KBE

[HSS07] Brian Herlihy, Peter Schachte, and Harald Søndergaard. Un-Kleene Boolean equation solving. International Journal of Foundations of Computer Science (IJFCS), 18(2):227–250, April 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Huang:2019:DCS

[HSS19] Daitao Huang, Minjia Shi, and Patrick Solé. Double circulant self-dual and LCD codes over Z_{p^{2}}. International Journal of Foundations of Computer Science (IJFCS), 30(3):407– 416, 2019. ISSN 0129-0541.

Hinze:2001:PCC

[HST01] Ralf Hinze, M. Sato, and Y. Toyama. Prolog’s control constructs in a functional setting — axioms and implementation. International Journal of Foundations of Computer Science (IJFCS), 12(2):125–170, 2001. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hsu:1998:SII

[Hsu98] D. Frank Hsu. Special issue on interconnection networks — Editors’ foreword. International Journal of Foundations of Computer Science (IJFCS), 9(1):1–??, March 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Haralambides:1995:BOS

[HT95] J. Haralambides and S. Tragoudas.

Bipartitioning into overlapping sets. International Journal of Foundations of Computer Science (IJFCS), 6(1):67–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hoang:2004:EMP

[HT04a] Thanh Minh Hoang and Thomas Thierauf. Erratum: on the minimal polynomial of a matrix. International Journal of Foundations of Computer Science (IJFCS), 15(4):685– ??, August 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic). See [HT04b].

Hoang:2004:MPM

[HT04b] Thanh Minh Hoang and Thomas Thierauf. On the minimal polynomial of a matrix. International Journal of Foundations of Computer Science (IJFCS), 15 (1):89–??, February 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic). See erratum [HT04a].

Hung:2009:NOB

[HT09] Regant Y. S. Hung and H. F.

Ting. A near-optimal broadcasting protocol for mobile video-on-demand. International Journal of Foundations of Computer Science (IJFCS), 20(1):45–55, February 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Harbich:2012:CDC

[HT12] Ronny Harbich and Bianca Truthe. A comparison of the descriptional complexity of classes of limited Lindenmayer systems: Part I. International Journal of Foundations of Computer Science (IJFCS), 23(1):99–114, January 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hutter:2002:FSA

[Hut02] M. Hutter. The fastest and shortest algorithm for all well-defined problems. International Journal of Foundations of Computer Science (IJFCS), 13(3):431–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Huynh:1991:EDC

[Huy91] Dung T. Huynh. Efficient detectors and constructors for simple languages. International Journal of Foundations of Computer Science (IJFCS), 2(3):183–206, September 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hartog:2002:VPP

[HV02] J. I. D. Hartog and E. P. D.

Vink. Verifying probabilistic programs using a Hoare like logic. International Journal of Foundations of Computer Science (IJFCS), 13 (3):315–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hempel:2000:OMM

[HW00] H. Hempel and G. Wechsung. The operators min and max on the polynomial hierarchy. International Journal of Foundations of Computer Science (IJFCS), 11 (2):315–342, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Han:2005:GGA

[HW05] Yo-Sub Han and Derick Wood. The generalization of generalized automata: Expression automata. International Journal of Foundations of Computer Science (IJFCS), 16(3):499– ??, June 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Huang:2010:CSB

[HW10] Geng-Dian Huang and Bow-Yaw Wang. Complete SATbased model checking for context-free processes. International Journal of Foundations of Computer Science (IJFCS), 21(2):115– 134, April 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hong:2017:IAA

[HW17] Wenyi Hong and Zhenbo Wang. Improved approximation algorithm for the combination of parallel machine scheduling and vertex cover. International Journal of Foundations of Computer Science (IJFCS), 28(8):977– ??, December 2017. CODEN IFCSEN. ISSN 0129-0541.

Han:2006:IFR

[HWW06] Yo-Sub Han, Yajun Wang, and Derick Wood. Infixfree regular expressions and languages. International Journal of Foundations of Computer Science (IJFCS), 17(2):379–??, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hu:1997:FTS

[HY97] Shuo-Cheng Hu and Chang-Biau Yang. Fault tolerance on star graphs. International Journal of Foundations of Computer Science (IJFCS), 8(2):127–??, June 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Hong:2006:P

[HY06] Seok-Hee Hong and Hsu-Chun Yen. Preface. International Journal of Foundations of Computer Science (IJFCS), 17(5):1003– ??, October 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Han:2020:CPE

[HYLF20] Yuejuan Han, Lantao You, Cheng-Kuan Lin, and Jianxi Fan. Communication performance evaluation of the locally twisted cube. International Journal of Foundations of Computer Science (IJFCS), 31(2):233– 252, February 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500057.

Ho:2015:SYE

[HYT15] Tzu-Hsin Ho, Li-Hsing Yen, and Chien-Chao Tseng. Simple-yet-efficient construction and revocation of group signatures. International Journal of Foundations of Computer Science (IJFCS), 26(5):611–??, August 2015. CODEN IFC-SEN. ISSN 0129-0541.

Hu:2012:LGG

[HZZT12] Liang Hu, Meng Zhang, Yi Zhang, and Jijun Tang. Label-guided graph exploration with adjustable ratio of labels. International Journal of Foundations of Computer Science (IJFCS), 23(4):903–??, June 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Inenaga:2012:FCS

[IB12] Shunsuke Inenaga and Hideo Bannai. Finding characteristic substrings from compressed texts. International Journal of Foundations of Computer Science (IJFCS), 23(2):261– 280, February 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibarra:2002:VQC

[Iba02] O. H. Ibarra. Verification in queue-connected multicounter machines. International Journal of Foundations of Computer Science (IJFCS), 13(1):115–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibarra:2011:SRS

[Iba11] Oscar H. Ibarra. On strong reversibility in P systems and related problems. International Journal of Foundations of Computer Science (IJFCS), 22(1):7–14, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibarra:2015:AFV

[Iba15]  Oscar H. Ibarra.  On

the ambiguity and finitevaluedness problems in acceptors and transducers. International Journal of Foundations of Computer Science (IJFCS), 26(7):967–??, November 2015. CODEN IFCSEN. ISSN 0129-0541.

Ibrahim:2022:MMC

[Ibr22] Salisu Ibrahim. Mathematical modelling and computational analysis of Covid-19 epidemic in Erbil Kurdistan using modified Lagrange interpolating polynomial. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):513–529, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420023.

Ibarra:2001:RSI

[IBS01] Oscar H. Ibarra, T. Bultan, and J. Su. On reachability and safety in infinitestate systems. International Journal of Foundations of Computer Science (IJFCS), 12(6):821–836, December 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibarra:1997:CCA

[IDR97] Oscar H. Ibarra, Pedro C.

Diniz, and Martin C. Rinard. On the complexity of commutativity analysis. International Journal of Foundations of Computer Science (IJFCS), 8(1):81– ??, March 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibarra:2008:CMR

[IDY08] Oscar H. Ibarra, Zhe Dang, and Linmin Yang. On counter machines, reachability problems, and Diophantine equations. International Journal of Foundations of Computer Science (IJFCS), 19(4):919– 934, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Inoue:2004:NRT

[IIK^{+}04] Katsushi Inoue, Akira Ito, Takashi Kamiura, Holger Petersen, and Lan Zhang. A note on rebound Turing machines. International Journal of Foundations of Computer Science (IJFCS), 15 (5):791–??, October 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Inoue:1991:ATM

[IIT91] K. Inoue, A. Ito, and I. Takanami. Alternating Turing machines with modified accepting structure. International Journal of Foundations of Computer Science (IJFCS), 2(4):401–418, December 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibarra:2021:SCS

[IJMP21] Oscar H. Ibarra, Jozef Jirásek, Jr., Ian McQuillan, and Luca Prigioniero. Space complexity of stack automata models. International Journal of Foundations of Computer Science (IJFCS), 32(06):801– 823, September 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121420090.

Ibarra:1993:ETW

[IJT^{+}93] O. H. Ibarra, T. Jiang, N. Tran, et al. On the equivalence of two-way pushdown automata and counter machines over bounded languages. International Journal of Foundations of Computer Science (IJFCS), 4(2): 135–146, June 1993. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibrahimov:2021:EFA

[IKPY21] Rishat Ibrahimov, Kamil Khadiev, Krišjānis Prūsis, and Abuzer Yakaryılmaz. Error-free affine, unitary, and probabilistic OBDDs. International Journal of Foundations of Computer Science (IJFCS), 32(07):

827–847, November 2021. ISSN 0129-0541. URL

                                        https://www.worldscientific.
                                        com/doi/10.1142/S0129054121500246.

Ipate:2011:FVS

[ILT11] Florentin Ipate, Raluca Lefticaru, and Cristina Tudose. Formal verification of P systems using spin. International Journal of Foundations of Computer Science (IJFCS), 22(1):133– 142, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibarra:2020:SFL

[IM20] Oscar H. Ibarra and Ian Mc-Quillan. Semilinearity of families of languages. International Journal of Foundations of Computer Science (IJFCS), 31(08):1179– 1198, December 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120420095.

Ibarra:2021:GCS

[IM21] Oscar H. Ibarra and Ian McQuillan. Generalizations of checking stack automata: Characterizations and hierarchies. International Journal of Foundations of Computer Science (IJFCS), 32(05):481– 508, August 2021. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054121410045.

Idwan:2004:FPM

[IML04] Sahar Idwan, Dinesh P.

Mehta, and Mario A. Lopez. Fast pursuit of mobile nodes using TPR trees. International Journal of Foundations of Computer Science (IJFCS), 15(5):753– ??, October 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Iliopoulos:2005:FAF

[IMP^{+}05] Costas S. Iliopoulos, James Mchugh, Pierre Peterlongo, Nadia Pisanti, Wojciech Rytter, and Marie-France Sagot. A first approach to finding common motifs with gaps. International Journal of Foundations of Computer Science (IJFCS), 16(6):1145–??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Iliopoulos:2012:PAM

[IMP12] Costas S. Iliopoulos, Mirka Miller, and Solon P. Pissis. Parallel algorithms for mapping short degenerate and weighted DNA sequences to a reference genome. International Journal of Foundations of Computer Science (IJFCS), 23 (2):249–259, February 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ivanyos:2003:EQA

[IMS03] Gábor Ivanyos, Frédéric Magniez, and Miklos Santha. Efficient quantum algorithms for some instances of the non-Abelian hidden subgroup problem. International Journal of Foundations of Computer Science (IJFCS), 14(5):723– ??, October 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ito:2005:FSL

[IN05] Yasuaki Ito and Koji Nakano. FM screening by the local exhaustive search, with hardware acceleration. International Journal of Foundations of Computer Science (IJFCS), 16(1):89– ??, February 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ito:2008:NFS

[IN08] Yasuaki Ito and Koji Nakano. A new FM screening method to generate cluster-dot binary images using the local exhaustive search with FPGA acceleration. International Journal of Foundations of Computer Science (IJFCS), 19 (6):1373–1386, December 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ito:2010:LLC

[IN10] Yasuaki Ito and Koji Nakano. Low-latency connected component labeling using an FPGA. International Journal of Foundations of Computer Science (IJFCS), 21(3):405– 425, June 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Iwama:2013:RSO

[IN13] Kazuo Iwama and Harumichi Nishimura. Recovering strings in oracles: Quantum and classic. International Journal of Foundations of Computer Science (IJFCS), 24(7):979–??, November 2013. CODEN IFCSEN. ISSN 0129-0541.

Ito:2007:EHA

[INY07] Yasuaki Ito, Koji Nakano, and Youhei Yamagishi. Efficient hardware algorithms for n choose k counters using the bitonic merger. International Journal of Foundations of Computer Science (IJFCS), 18(3):517– 528, June 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ishdorj:2008:GAM

[IP08] Tseren-Onolt Ishdorj and Ion Petre. Gene assembly models and Boolean circuits. International Journal of Foundations of Computer Science (IJFCS), 19(5): 1133–1145, October 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ishdorj:2007:CPI

[IPR07] Tseren-Onolt Ishdorj, Ion Petre, and Vladimir Rogojin. Computational power of intramolecular gene assembly. International Journal of Foundations of Computer Science (IJFCS), 18(5): 1123–1136, October 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibarra:2009:P

[IR09] Oscar H. Ibarra and Bala Ravikumar. Preface. International Journal of Foundations of Computer Science (IJFCS), 20(4):559– 561, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibarra:2014:SDQ

[IR14] Oscar H. Ibarra and Bala Ravikumar. Some decision questions concerning the time complexity of language acceptors. International Journal of Foundations of Computer Science (IJFCS), 25(8):1127–??, December 2014. CODEN IFC-SEN. ISSN 0129-0541.

Ibarra:2012:CBS

[IS12] Oscar H. Ibarra and Shinnosuke Seki. Characterizations of bounded semilinear languages by oneway and two-way deterministic machines. International Journal of Foundations of Computer Science (IJFCS), 23(6):1291–??, September 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibarra:2021:A

[IS21] Oscar H. Ibarra and Sartaj K. Sahni. Announcement. International Journal of Foundations of Computer Science (IJFCS), 32 (07):??, November 2021. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121010012.

Imani:2008:ICM

[ISAZ08] N. Imani, H. Sarbazi-Azad, and A. Zomaya. Intruder capturing in mesh and torus networks. International Journal of Foundations of Computer Science (IJFCS), 19(4):1049–1071, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Inenaga:2005:FCP

[IST05] Shunsuke Inenaga, Ayumi Shinohara, and Masayuki Takeda. A fully compressed pattern matching algorithm for simple collage systems. International Journal of Foundations of Computer Science (IJFCS), 16(6):1155–??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibarra:2013:HSH

[IT13] Oscar H. Ibarra and Nicholas Q. Tran. How to synchronize the heads of a multitape automaton. International Journal of Foundations of Computer Science (IJFCS), 24(6):799–??, September 2013. CODEN IFCSEN. ISSN 0129-0541.

Ito:2010:P

[Ito10] Masami Ito. Preface. International Journal of Foundations of Computer Science (IJFCS), 21(4):477– 478, August 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Iliopoulos:2018:P

[IV18] Costas S. Iliopoulos and Fatima Vayani. Preface. International Journal of Foundations of Computer Science (IJFCS), 29(8):1247– 1248, December 2018. CO-DEN IFCSEN. ISSN 0129-0541.

Ibarra:2007:CRL

[IW07] Oscar H. Ibarra and Sara Woodworth. Characterizing regular languages by spiking neural P systems. International Journal of Foundations of Computer Science (IJFCS), 18(6):1247–1256, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibarra:2007:P

[IY07] Oscar H. Ibarra and Hsu-Chun Yen. Preface. International Journal of Foundations of Computer Science (IJFCS), 18(4):667– 668, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibarra:2005:VNP

[IYD05] Oscar H. Ibarra, Hsu-Chun Yen, and Zhe Dang. On various notions of parallelism in P systems. International Journal of Foundations of Computer Science (IJFCS), 16(4):683–??, August 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ilie:2004:WCR

[IYZ04] Lucian Ilie, Sheng Yu, and Kaizhong Zhang. Word complexity and repetitions in words. International Journal of Foundations of Computer Science (IJFCS), 15(1):41–??, February 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ibarra:2004:CCC

[IZ04] Oscar H. Ibarra and Louxin Zhang. Computing and combinatorics conference — COCOON’02. International Journal of Foundations of Computer Science (IJFCS), 15(1):1–??, February 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Isobe:1999:PTA

[IZN99] S. Isobe, X. Zhou, and T. Nishizeki. A polynomialtime algorithm for finding total colorings of partial ktrees. International Journal of Foundations of Computer Science (IJFCS), 10 (2):171–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ito:2005:PTS

[IZN05] Takehiro Ito, Xiao Zhou, and Takao Nishizeki. Partitioning trees of supply and demand. International Journal of Foundations of Computer Science (IJFCS), 16(4):803–??, August 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jain:1995:ICF

[Jai95] S. Jain. An infinite class of functions identifiable using minimal programs in all Kolmogorov numberings. International Journal of Foundations of Computer Science (IJFCS), 6 (1):89–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jain:1998:MCI

[Jai98] Sanjay Jain. Minimal concept identification and reliability. International Journal of Foundations of Computer Science (IJFCS), 9 (3):315–??, September 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jansen:1993:SIJ

[Jan93] K. Jansen. Scheduling of incompatible jobs on unrelated machines. International Journal of Foundations of Computer Science (IJFCS), 4(4):275–292, December 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jiao:2003:CLM

[JC03] Li Jiao and To-Yat Cheung. Characterizing liveness monotonicity for weighted petri nets in terms of siphonbased properties. International Journal of Foundations of Computer Science (IJFCS), 14(4):641– ??, August 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jez:2008:CGG

[Jeż08] Artur Jeż. Conjunctive grammars generate nonregular unary languages. International Journal of Foundations of Computer Science (IJFCS), 19(3):597– 615, June 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jecker:2018:MSW

[JF18] Ismaël Jecker and Emmanuel Filiot. Multisequential word relations. International Journal of Foundations of Computer Science (IJFCS), 29(2):271– ??, February 2018. CODEN IFCSEN. ISSN 0129-0541.

Jendrsczok:2008:IHP

[JHK08] Johannes Jendrsczok, Rolf Hoffmann, and Jörg Keller. Implementing Hirschberg’s PRAM-algorithm for connected components on a global cellular automaton. International Journal of Foundations of Computer Science (IJFCS), 19(6): 1299–1316, December 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jiraskova:2011:MNT

[Jir11] Galina Jirásková. Magic numbers and ternary alphabet. International Journal of Foundations of Computer Science (IJFCS), 22 (2):331–344, February 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jiraskova:2014:RSC

[Jir14]  Galina Jirásková.  The

ranges of state complexities for complement, star, and reversal of regular languages. International Journal of Foundations of Computer Science (IJFCS), 25 (1):101–??, January 2014. CODEN IFCSEN. ISSN 0129-0541.

Jirasek:2005:SCC

[JJS05] Jozef Jirásek, Galina Jirásková, and Alexander Szabari. State complexity of concatenation and complementation. International Journal of Foundations of Computer Science (IJFCS), 16(3):511– ??, June 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jirasek:2008:DBU

[JJS08] Jozef Jirásek, Galina Jirásková, and Alexander Szabari. Deterministic blow-ups of minimal nondeterministic finite automata over a fixed alphabet. International Journal of Foundations of Computer Science (IJFCS), 19(3):617– 631, June 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jirasek:2018:OUF

[JJŠ18] Jozef Jirásek, Jr., Galina Jirásková, and Juraj Šebej. Operations on unambiguous finite automata. International Journal of Foundations of Computer Science (IJFCS), 29(5):861– 876, August 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905411842008X.

Jurgensen:2007:SAB

[JK07] Helmut Jürgensen and Pauline Kraak. Soliton automata based on trees. International Journal of Foundations of Computer Science (IJFCS), 18(6):1257– 1270, December 2007. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jonoska:2014:ATSa

[JK14a] Nataša Jonoska and Daria Karpenko. Active tile selfassembly, part 1: Universality at temperature 1. International Journal of Foundations of Computer Science (IJFCS), 25(2):141–??, February 2014. CODEN IFCSEN. ISSN 0129-0541.

Jonoska:2014:ATSb

[JK14b] Nataša Jonoska and Daria Karpenko. Active tile selfassembly, part 2: Selfsimilar structures and structural recursion. International Journal of Foundations of Computer Science (IJFCS), 25(2):165–??, February 2014. CODEN IFCSEN. ISSN 0129-0541.

Jirskov:2019:SDA

[JK19] Galina Jirásková and Ivana Krajňáková. Square on deterministic, alternating, and Boolean finite automata. International Journal of Foundations of Computer Science (IJFCS), 30(6– 7):1117–1134, September– November 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400318.

Jacobsen:2001:VTR

[JL01] Lars Jacobsen and Kim S.

Larsen. Variants of (A, B)- trees with relaxed balance. International Journal of Foundations of Computer Science (IJFCS), 12 (4):455–478, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jansson:2023:OAN

[JLL23] Jesper Jansson, Christos Levcopoulos, , and Andrzej Lingas. Online and approximate network construction from bounded connectivity constraints. International Journal of Foundations of Computer Science (IJFCS), 34(05):453– 468, August 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500265.

Jorrand:2003:SPQ

[JM03] Philippe Jorrand and Mehdi Mhalla. Separability of pure N-qubit states: Two characterizations. International Journal of Foundations of Computer Science (IJFCS), 14(5):797–??, October 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jiraskova:2011:CUF

[JM11] Galina Jirásková and Tomáš Masopust. Complexity in union-free regular languages. International Journal of Foundations of Computer Science (IJFCS), 22 (7):1639–1653, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jez:2013:HMD

[JM13] Artur Jeż and Andreas Maletti. Hyper-minimization for deterministic tree automata. International Journal of Foundations of Computer Science (IJFCS), 24 (6):815–??, September 2013.

CODEN IFCSEN. ISSN 0129-0541.

Jiang:1991:SCM

[JMR91] Tao Jiang, Edward McDowell, and B. Ravikumar. The structure and complexity of minimal NFA’s over a unary alphabet. International Journal of Foundations of Computer Science (IJFCS), 2(2):163–??, June 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jansen:2005:AAF

[JMSO05] Klaus Jansen, Monaldo Mastrolilli, and Roberto Solis-Oba. Approximation algorithms for flexible job shop problems. International Journal of Foundations of Computer Science (IJFCS), 16(2):361– ??, April 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jurdzinski:2007:SRA

[JO07] Tomasz Jurdziński and Friedrich Otto. Shrinking restarting automata. International Journal of Foundations of Computer Science (IJFCS), 18(2):361– 385, April 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Johansson:2000:NDP

[Joh00]  Ö. Johansson.  NLC_{2}-

decomposition in polynomial time. International Journal of Foundations of Computer Science (IJFCS), 11(3):373–396, 2000. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Justin:2004:EWS

[JP04] Jacques Justin and Giuseppe Pirillo. Episturmian words: Shifts, morphisms and numeration systems. International Journal of Foundations of Computer Science (IJFCS), 15(2):329– ??, April 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jonoska:2006:TTD

[JP06] Nataša Jonoska and Joni Burnette Pirnot. Transitivity in two-dimensional local languages defined by dot systems. International Journal of Foundations of Computer Science (IJFCS), 17(2):435– ??, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jansson:2007:ODR

[JP07] Jesper Jansson and Zeshan Peng. Online and dynamic recognition of squarefree strings. International Journal of Foundations of Computer Science (IJFCS), 18(2):401–414, April 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jalsenius:2008:SSC

[JP08] Markus Jalsenius and Kasper Pedersen. A systematic scan for 7-colourings of the grid. International Journal of Foundations of Computer Science (IJFCS), 19 (6):1461–1477, December 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jirsek:2019:KAG

[JPŠ19] Jozef Jirásek, Jr., Matú Palmovský, and Juraj Šebej. Kuratowski algebras generated by prefix-, suffix-, factor-, and subword-free languages under star and complementation. International Journal of Foundations of Computer Science (IJFCS), 30(6–7): 1091–1115, September–November 2019. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054119400306.

Jurgensen:2014:P

[JR14] Helmut Jürgensen and Rogério Reis. Preface. International Journal of Foundations of Computer Science (IJFCS), 25(7):803–??, November 2014. CODEN IFCSEN. ISSN 0129-0541.

Jack:2008:DNM

[JRPIP08] John Jack, Alfonso Rodríguez-Patón, Oscar H. Ibarra, and Andrei Păun. Discrete nondeterministic modeling of the FAS pathway. International Journal of Foundations of Computer Science (IJFCS), 19(5):1147–1162, October 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jia:1997:TLN

[JS97] Xingde Jia and Weidong Su. Triple loop networks with minimal transmission delay. International Journal of Foundations of Computer Science (IJFCS), 8 (3):305–??, September 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jia:2002:CCH

[JS02] Weijia Jia and Zhibin Sun. On computational complexity of hierarchical optimization. International Journal of Foundations of Computer Science (IJFCS), 13(5):667– ??, October 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jung:2003:SBS

[JS03] Haejae Jung and Sartaj Sahni. Supernode binary search trees. International Journal of Foundations of Computer Science (IJFCS), 14(3):465–??, June 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jonoska:2021:P

[JS21] Nataša Jonoska and Dmytro Savchuk. Preface. International Journal of Foundations of Computer Science (IJFCS), 32(06):615– 617, September 2021. CO-DEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121020020.

Jonoska:2020:SSD

[JSKM20] Nataša Jonoska, Masahico Saito, Hwee Kim, and Brad Mostowski. Symbol separation in double occurrence words. International Journal of Foundations of Computer Science (IJFCS), 31(07):915– 928, November 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500343.

Jansen:2010:ASS

[JSO10] Klaus Jansen and Roberto Solis-Oba. Approximation schemes for scheduling jobs with chain precedence constraints. International Journal of Foundations of Computer Science (IJFCS), 21 (1):27–49, February 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jain:2003:PPH

[JSPD03] Anuj Jain, Sartaj Sahni, Jatinder Palta, and James Dempsey. Partitioning 3D phantoms into homogeneous cuboids. International Journal of Foundations of Computer Science (IJFCS), 14 (5):905–??, October 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jung:2014:SAV

[Jun14] Haejae Jung. A simple array version of M-heap. International Journal of Foundations of Computer Science (IJFCS), 25(1):67–??, January 2014. CODEN IFC-SEN. ISSN 0129-0541.

Jurgensen:2008:CIE

[Jür08] H. Jürgensen. Complexity, information, energy. International Journal of Foundations of Computer Science (IJFCS), 19(4):781– 793, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jiang:2008:ASP

[JW08] Zhen Jiang and Jie Wu. On achieving the shortest-path routing in 2-D meshes. International Journal of Foundations of Computer Science (IJFCS), 19(6):1279– 1297, December 2008. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Janzing:2003:CPP

[JWB03] Dominik Janzing, Paweł Wocjan, and Thomas Beth. On the computational power of physical interactions: Bounds on the number of time steps for simulating arbitrary interaction graphs. International Journal of Foundations of Computer Science (IJFCS), 14(5):889– ??, October 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Jin-Yi:1991:PSC

[JYF91] Cai Jin-Yi and Merrick Furst. PSPACE survives constant-width bottlenecks. International Journal of Foundations of Computer Science (IJFCS), 2(1):67– ??, March 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Januszewski:2016:IOA

[JZ16] Janusz Januszewski and Łukasz Zielonka. Improved online algorithms for 2-space bounded 2dimensional bin packing. International Journal of Foundations of Computer Science (IJFCS), 27(4):407–??, June 2016. CODEN IFC-SEN. ISSN 0129-0541.

Kurkcu:2018:CBE

[KA18] Ömür Kıvanç Kürkçü and Ersin Aslan. A comparison between edge neighbor rupture degree and edge scattering number in graphs. International Journal of Foundations of Computer Science (IJFCS), 29(7):1119– 1142, November 2018. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054118500247.

Kameyama:1995:TFT

[Kam95] Y. Kameyama. A type-free theory of half-monotone inductive definitions. International Journal of Foundations of Computer Science (IJFCS), 6(3):203–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kamareddine:1998:SES

[Kam98] Fairouz Kamareddine. The soundness of explicit substitution with nameless variables. International Journal of Foundations of Computer Science (IJFCS), 9 (3):321–??, September 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kantabutra:2015:FSP

[Kan15] Sanpawat Kantabutra. Fast sequential and parallel vertex relabelings of K_{m,m}. International Journal of Foundations of Computer Science (IJFCS), 26(1):33–??, January 2015. CODEN IFC-SEN. ISSN 0129-0541.

Kapoutsis:2005:NRT

[Kap05] Christos Kapoutsis. Nonrecursive trade-offs for twoway machines. International Journal of Foundations of Computer Science (IJFCS), 16(5):943–??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Katangur:2005:ROM

[KAPF05] Ajay K. Katangur, Somasheker Akkaladevi, Yi Pan, and Martin D. Fraser. Routing in optical multistage networks with limited crosstalk using ant colony optimization. International Journal of Foundations of Computer Science (IJFCS), 16(2):301–??, April 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Karaata:1999:SSA

[Kar99] M. H. Karaata. A selfstabilizing algorithm for finding articulation points. International Journal of Foundations of Computer Science (IJFCS), 10(1):33– ??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Karhumaki:2009:PCM

[Kar09] Juhani Karhumäki. On the power of cooperating morphisms via reachability problems. International Journal of Foundations of Computer Science (IJFCS), 20(5):803– 818, October 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kunt:2020:EIN

[KB20] Aysun Asena Kunt and Zeynep Nihan Berberler. Efficient identification of node importance based on agglomeration in cyclerelated networks. International Journal of Foundations of Computer Science (IJFCS), 31(07):969– 978, November 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120500379.

Krithivasan:1999:DPA

[KBH99a] K. Krithivasan, M. S. Balan, and P. Harsha. Distributed processing in automata. International Journal of Foundations of Computer Science (IJFCS), 10(4):443– 464, 1999. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Krithivasan:1999:PRP

[KBH99b] K. Krithivasan, N. S. Balan, and P. Harsha. Part 2 (regular papers). International Journal of Foundations of Computer Science (IJFCS), 10(4):443–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Krings:1999:RTD

[KD99] A. W. Krings and M. Dror.

Real-time dispatching: Scheduling stability and precedence. International Journal of Foundations of Computer Science (IJFCS), 10 (3):313–328, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ko:2016:OSC

[KEH16] Sang-Ki Ko, Hae-Sung Eom, and Yo-Sub Han. Operational state complexity of subtree-free regular tree languages. International Journal of Foundations of Computer Science (IJFCS), 27 (6):705–??, September 2016. CODEN IFCSEN. ISSN 0129-0541.

Karmakar:2011:ADM

[KG11] Sushanta Karmakar and Arobinda Gupta. Adaptive distributed mutual exclusion by dynamic topology switching. International Journal of Foundations of Computer Science (IJFCS), 22(3):713–737, April 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ko:2021:LBT

[KH21] Sang-Ki Ko and Yo-Sub Han. Left is better than right for reducing nondeterminism of NFAs. International Journal of Foundations of Computer Science (IJFCS), 32(05):531– 550, August 2021. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054121410069.

Kuo:2012:MFA

[KHLC12] Chi-Jung Kuo, Chiun-Chieh Hsu, Hon-Ren Lin, and Da-Ren Chen. Minimum feedback arc sets in rotator and incomplete rotator graphs. International Journal of Foundations of Computer Science (IJFCS), 23(4):931–??, June 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ko:2021:GCL

[KHS21] Sang-Ki Ko, Yo-Sub Han, and Kai Salomaa. Generalizations of code languages with marginal errors. International Journal of Foundations of Computer Science (IJFCS), 32(05):509– 529, August 2021. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054121410057.

Korah:1990:DOB

[KK90] A. P. Korah and M. R. Kaimal. Dynamic optimal binary search tree. International Journal of Foundations of Computer Science (IJFCS), 1(4):449–??, December 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kaufman:2005:SLT

[KK05] Yair Kaufman and Shmuel T.

Klein. Semi-lossless text compression. International Journal of Foundations of Computer Science (IJFCS), 16(6):1167–??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Klein:2007:CFG

[KK07] Andreas Klein and Martin Kutrib. Context-free grammars with linked nonterminals. International Journal of Foundations of Computer Science (IJFCS), 18(6):1271–1282, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kamei:2010:SSD

[KK10] Sayaka Kamei and Hirotsugu Kakugawa. A selfstabilizing distributed approximation algorithm for the minimum connected dominating set. International Journal of Foundations of Computer Science (IJFCS), 21(3):459– 476, June 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kumari:2019:ALC

[KK19] Priti Kumari and Pramod Kumar Kewat. 2-adic and linear complexities of a class of Whiteman’s generalized cyclotomic sequences of order four. International Journal of Foundations of Computer Science (IJFCS), 30(5):759– 779, August 2019. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054119500205.

Keshavarz-Kohjerdi:2024:LPP

[KKB24] Fatemeh Keshavarz-Kohjerdi and Alireza Bagheri. The longest path problem in odd-sized O-shaped grid graphs. International Journal of Foundations of Computer Science (IJFCS), 35 (03):353–374, April 2024. ISSN 0129-0541. URL

                                        https://www.worldscientific.
                                        com/doi/10.1142/S0129054123500065.

Keqin:1990:GFF

[KKH90] Li Keqin and Cheng Kam-Hoi. Generalized first-fit algorithms in two and three dimensions. International Journal of Foundations of Computer Science (IJFCS), 1(2):131–??, June 1990. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kranakis:1997:HCT

[KKP97] Evangelos Kranakis, Danny Krizanc, and Andrzej Pelc. Hop-congestion trade-offs for high-speed networks. International Journal of Foundations of Computer Science (IJFCS), 8(2):117–??, June 1997. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kari:2005:BFL

[KKS05a] Lila Kari, Stavros Konstantinidis, and Petr Sosík. Bond-free languages: Formalizations, maximality and construction methods. International Journal of Foundations of Computer Science (IJFCS), 16(5):1039– ??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kari:2005:OTA

[KKS05b] Lila Kari, Stavros Konstantinidis, and Petr Sosík. Operations on trajectories with applications to coding and bioinformatics. International Journal of Foundations of Computer Science (IJFCS), 16(3):531– ??, June 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kobler:2000:OSE

[KL00] J. Köbler and W. Lindner. Oracles in Σ_{2}^{p} are sufficient for exact learning. International Journal of Foundations of Computer Science (IJFCS), 11 (4):615–632, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Karhumaki:2003:EPF

[KL03] J. Karhumäki and L. P. Lisovik. The equivalence problem of finite substitutions on ab ∗ c, with applications. International Journal of Foundations of Computer Science (IJFCS), 14(4):699– ??, August 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kohrt:2005:LSR

[KL05] Jens S. Kohrt and Kim S. Larsen. On-line seat reservations via off-line seating arrangements. International Journal of Foundations of Computer Science (IJFCS), 16(2):381–??, April 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kupferman:2010:LSR

[KL10] Orna Kupferman and Yoad Lustig. Latticed simulation relations and games. International Journal of Foundations of Computer Science (IJFCS), 21(2):167– 189, April 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kufleitner:2011:POT

[KL11] Manfred Kufleitner and Alexander Lauser. Partially ordered two-way Büchi automata. International Journal of Foundations of Computer Science (IJFCS), 22 (8):1861–1876, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kufleitner:2012:ADD

[KL12] Manfred Kufleitner and Alexander Lauser. Around dot-depth one. International Journal of Foundations of Computer Science (IJFCS), 23(6):1323– ??, September 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kalampakas:2013:MPD

[KLB13] Antonios Kalampakas and Olympia Louscou-Bozapalidou. Minimization of planar directed acyclic graph algebras. International Journal of Foundations of Computer Science (IJFCS), 24 (4):519–??, June 2013. CO-DEN IFCSEN. ISSN 0129-0541.

Ko:2016:SCR

[KLH16] Sang-Ki Ko, Ha-Rim Lee, and Yo-Sub Han. State complexity of regular tree languages for tree matching. International Journal of Foundations of Computer Science (IJFCS), 27(8):965– 980, December 2016. CO-DEN IFCSEN. ISSN 0129-0541.

Kloks:1996:TCG

[Klo96a] T. Kloks. Treewidth of circle graphs. International Journal of Foundations of Computer Science (IJFCS), 7(2):111–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Klostermeyer:1996:STS

[Klo96b] W. F. Klostermeyer. Scheduling two salesmen in a network. International Journal of Foundations of Computer Science (IJFCS), 7 (4):353–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kim:2020:TRO

[KLP20] Eunkyung Kim, Hyang-Sook Lee, and Jeongeun Park. Towards roundoptimal secure multiparty computations: Multikey FHE without a CRS. International Journal of Foundations of Computer Science (IJFCS), 31(2):157– 174, February 2020. CO-DEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412050001X.

Krawetz:2005:SCM

[KLS05] Bryan Krawetz, John Lawrence, and Jeffrey Shallit. State complexity and the monoid of transformations of a finite set. International Journal of Foundations of Computer Science (IJFCS), 16(3):547– ??, June 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Klein:2019:FFP

[KLS^{+}19] Rolf Klein, Elmar Langetepe, Barbara Schwarzwald, Christos Levcopoulos, and Andrzej Lingas. On a fire fighter’s problem. International Journal of Foundations of Computer Science (IJFCS), 30(2):231– 246, February 2019. ISSN 0129-0541.

Katajainen:1990:TCO

[KM90] Jyrki Katajainen and Erkki Makinen. Tree compression and optimization with applications. International Journal of Foundations of Computer Science (IJFCS), 1 (4):425–??, December 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kamareddine:2002:EAT

[KM02] F. Kamareddine and F. Monin. An extension of an automated termination method of recursive functions. International Journal of Foundations of Computer Science (IJFCS), 13(3):361–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kari:2007:IBW

[KM07a] Lila Kari and Kalpana Mahalingam. Involutively bordered words. International Journal of Foundations of Computer Science (IJFCS), 18(5):1089–1106, October 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kutrib:2007:WCR

[KM07b] Martin Kutrib and Andreas Malcher. When Church– Rosser becomes context free. International Journal of Foundations of Computer Science (IJFCS), 18 (6):1293–1302, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kari:2008:WCB

[KM08] Lila Kari and Kalpana Mahalingam. Watson–Crick bordered words and their syntactic monoid. International Journal of Foundations of Computer Science (IJFCS), 19(5):1163–1179, October 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Krivka:2015:JG

[KM15] Zbyněk Křivka and Alexander Meduna. Jumping grammars. International Journal of Foundations of Computer Science (IJFCS), 26(6):709–??, September 2015. CODEN IFCSEN. ISSN 0129-0541.

Kapoutsis:2017:LCS

[KM17] Christos A. Kapoutsis and Lamana Mulaffer. A logical characterization of small 2NFAs. International Journal of Foundations of Computer Science (IJFCS), 28 (5):445–??, August 2017. CODEN IFCSEN. ISSN 0129-0541.

Kavand:2018:TST

[KM18] Pardis Kavand and Ali Mohades. A time-space tradeoff for the shortest path tree in a simple polygon. International Journal of Foundations of Computer Science (IJFCS), 29(3):391– 402, April 2018. CODEN IFCSEN. ISSN 0129-0541.

Kunc:2019:GRT

[KM19] Michal Kunc and Jan Meitner. The generalized rank of trace languages. International Journal of Foundations of Computer Science (IJFCS), 30(1):135– 169, January 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400070.

Kirgiz:2022:NBS

[KM22] Havva Kırgız and A. Dilek Maden. New bounds for some topological indices. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):953–965, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420230.

Kutrib:2023:RTS

[KM23] Martin Kutrib and Uwe Meyer. Reversible top-down syntax analysis. International Journal of Foundations of Computer Science (IJFCS), 34(2–3):225–251, February–April 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412244004X.

Krithivasan:2011:SLG

[KMG11] Kamala Krithivasan, Venkata Padmavati Metta, and Deepak Garg. On string languages generated by spiking neural P systems with antispikes. International Journal of Foundations of Computer Science (IJFCS), 22 (1):15–27, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kusano:2009:AVS

[KMIS09] Kazuhiko Kusano, Wataru Matsubara, Akira Ishino, and Ayumi Shinohara. Average value of sum of exponents of runs in a string. International Journal of Foundations of Computer Science (IJFCS), 20(6):1135– 1146, December 2009. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kuppusamy:2011:AIS

[KMK11] Lakshmanan Kuppusamy, Anand Mahendran, and Kamala Krithivasan. On the ambiguity of insertion systems. International Journal of Foundations of Computer Science (IJFCS), 22 (7):1747–1758, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kupferman:2006:TRA

[KMM06] Orna Kupferman, Gila Morgenstern, and Aniello Murano. Typeness for ω-regular automata. International Journal of Foundations of Computer Science (IJFCS), 17(4):869–??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kutrib:2010:STP

[KMO10] Martin Kutrib, Hartmut Messerschmidt, and Friedrich Otto. On stateless twopushdown automata and restarting automata. International Journal of Foundations of Computer Science (IJFCS), 21(5):781– 798, October 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Konstantinidis:2020:RET

[KMRY20] Stavros Konstantinidis, Nelma Moreira, Rogério Reis, and Joshua Young. Regular expressions and transducers over alphabet-invariant and user-defined labels. International Journal of Foundations of Computer Science (IJFCS), 31(08):983– 1019, December 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120420010.

Klarlund:2002:RP

[KMS02] N. Klarlund, A. Møller, and M. I. Schwatzbach. Regular papers. International Journal of Foundations of Computer Science (IJFCS), 13 (4):571–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Krivka:2006:GLR

[KMS06] Zbyněk Křivka, Alexander Meduna, and Rudolf Schönecker. Generation of languages by rewriting systems that resemble automata. International Journal of Foundations of Computer Science (IJFCS), 17 (5):1223–??, October 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kari:2011:PPP

[KMS11] Lila Kari, Benoît Masson, and Shinnosuke Seki. Properties of pseudo-primitive words and their applications. International Journal of Foundations of Computer Science (IJFCS), 22 (2):447–471, February 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Konstantinidis:2021:ZAT

[KMŠ21] Stavros Konstantinidis, Mitja Mastnak, and Juraj Šebej. Zero-avoiding transducers, length separable relations, and the rational asymmetric partition problem. International Journal of Foundations of Computer Science (IJFCS), 32(05):455– 480, August 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121410033.

Kappes:2012:MCK

[KMW12] Martin Kappes, Andreas Malcher, and Detlef Wotschke. In memoriam: Chandra Kintala. International Journal of Foundations of Computer Science (IJFCS), 23 (1):5–19, January 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kutrib:2014:SUO

[KMW14a] Martin Kutrib, Andreas Malcher, and Matthias Wendlandt. Simulations of unary one-way multi-head finite automata. International Journal of Foundations of Computer Science (IJFCS), 25(7):877–??, November 2014. CODEN IFCSEN. ISSN 0129-0541.

Kutrib:2014:SOW

[KMW14b] Martin Kutrib, Andreas Malcher, and Matthias Wendlandt. Stateless oneway multi-head finite automata with pebbles. International Journal of Foundations of Computer Science (IJFCS), 25(8):1141–??, December 2014. CODEN IFC-SEN. ISSN 0129-0541.

Kutrib:2016:SA

[KMW16] Martin Kutrib, Andreas Malcher, and Matthias Wendlandt. Set automata. International Journal of Foundations of Computer Science (IJFCS), 27(2):187– ??, February 2016. CODEN IFCSEN. ISSN 0129-0541.

Kang:2019:NBK

[KMZS19] Burong Kang, Xinyu Meng, Lei Zhang, and Yinxia Sun. Nonce-based key agreement protocol against bad randomness. International Journal of Foundations of Computer Science (IJFCS), 30(4):619– 633, June 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400161.

Kamareddine:1993:SES

[KN93] F. Kamareddine and R. Nederpelt. On stepwise explicit substitution. International Journal of Foundations of Computer Science (IJFCS), 4(3):197–240, September 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kari:2021:DCS

[KN21] Lila Kari and Timothy Ng. Descriptional complexity of semi-simple splicing systems. International Journal of Foundations of Computer Science (IJFCS), 32(06):685– 711, September 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121420041.

Kravchenko:2018:GSF

[KNR18] Dmitry Kravchenko, Nikolajs Nahimovs, and Alexander Rivosh. Grover’s search with faults on some marked elements. International Journal of Foundations of Computer Science (IJFCS), 29(4):647–662, June 2018.

CODEN IFCSEN. ISSN

0129-0541.

Koechlin:2021:SUE

[KNR21] Florent Koechlin, Cyril Nicaud, and Pablo Rotondo. Simplifications of uniform expressions specified by systems. International Journal of Foundations of Computer Science (IJFCS), 32(06):733– 760, September 2021. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054121420065.

Kutrib:2013:DCW

[KO13] Martin Kutrib and Friedrich Otto. On the descriptional complexity of the window size for deleting restarting automata. International Journal of Foundations of Computer Science (IJFCS), 24(6):831–??, September 2013. CODEN IFCSEN. ISSN 0129-0541.

Kari:2018:P

[KO18a] Jarkko Kari and Alexander Okhotin. Preface. International Journal of Foundations of Computer Science (IJFCS), 29(4):457– 459, June 2018. CODEN IFCSEN. ISSN 0129-0541.

Kwee:2018:NOR

[KO18b] Kent Kwee and Friedrich Otto. Nondeterministic ordered restarting automata. International Journal of Foundations of Computer Science (IJFCS), 29(4):663– 685, June 2018. CODEN IFCSEN. ISSN 0129-0541.

Koga:2018:CFP

[Kog18] Toshihiro Koga. Contextfreeness of parsing expression languages is undecidable. International Journal of Foundations of Computer Science (IJFCS), 29(7):1203–1213, November 2018. ISSN 0129-

0541. URL https://www.
worldscientific.com/doi/
10.1142/S0129054118500296.

Koga:2021:PPT

[Kog21] Toshihiro Koga. A proof of Parikh’s theorem via Dickson’s Lemma. International Journal of Foundations of Computer Science (IJFCS), 32(02):163– 173, February 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412150009X.

Kopra:2021:IDT

[Kop21]  Johan Kopra.  On the

interplay of direct topological factorizations and cellular automata dynamics on beta-shifts. International Journal of Foundations of Computer Science (IJFCS), 32(06):663– 683, September 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412142003X.

Korner:2003:TSE

[Kör03] Heiko Körner. A time and space efficient algorithm for minimizing cover automata for finite languages. International Journal of Foundations of Computer Science (IJFCS), 14(6):1071– ??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Klima:2010:HPT

[KP10a] Ondřej Klíma and Libor Polák. Hierarchies of piecewise testable languages. International Journal of Foundations of Computer Science (IJFCS), 21(4):517– 533, August 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Klima:2010:LIL

[KP10b] Ondřej Klíma and Libor Polák. Literally idempotent languages and their varieties — two letter case. International Journal of Foundations of Computer Science (IJFCS), 21(5):761– 780, October 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kouri:2015:RMA

[KPM15] Tina M. Kouri, Daniel Pascua, and Dinesh P. Mehta. Random models and analyses for chemical graphs. International Journal of Foundations of Computer Science (IJFCS), 26(2):269–??, February 2015. CODEN IFCSEN. ISSN 0129-0541.

Kirschenhofer:1993:MDS

[KPS93] P. Kirschenhofer, H. Prodinger, and W. Szpankowski. Multidimensional digital searching and some new parameters in tries. International Journal of Foundations of Computer Science (IJFCS), 4(1):69–84, March 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Karhumaki:2013:FWT

[KPS13] Juhani Karhumäki, Svetlana Puzynina, and Aleksi Saarela. Fine and Wilf’s theorem for k-Abelian periods. International Journal of Foundations of Computer Science (IJFCS), 24(7): 1135–??, November 2013. CODEN IFCSEN. ISSN 0129-0541.

Klimann:2018:CSR

[KPS18] Ines Klimann, Matthieu Picantin, and Dmytro Savchuk. A connected 3-state reversible Mealy automaton cannot generate an infinite Burnside group. International Journal of Foundations of Computer Science (IJFCS), 29(2):297–??, February 2018. CODEN IFCSEN. ISSN 0129-0541.

Kurganskyy:2008:RPL

[KPSC08] Oleksiy Kurganskyy, Igor Potapov, and Fernando Sancho-Caparrini. Reachability problems in lowdimensional iterative maps. International Journal of Foundations of Computer Science (IJFCS), 19(4):935– 951, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Keum:1997:DAS

[KR97] Young Wook Keum and Hwakyung Rim. Design and analysis of the Symmetric Banyan Network (SBN): a min with high performance and high fault tolerance. International Journal of Foundations of Computer Science (IJFCS), 8(3):253–??, September 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Klappenecker:2003:QSR

[KR03] Andreas Klappenecker and Martin Rötteler. Quantum software reusability. International Journal of Foundations of Computer Science (IJFCS), 14(5):777– ??, October 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kutrib:2008:OSW

[KR08] Martin Kutrib and Jens Reimann. Optimal simulations of weak restarting automata. International Journal of Foundations of Computer Science (IJFCS), 19 (4):795–811, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kostolanyi:2016:AAW

[KR16] Peter Kostolányi and Branislav Rovan. Automata with auxiliary weights. International Journal of Foundations of Computer Science (IJFCS), 27(7):787–??, November 2016. CODEN IFCSEN. ISSN 0129-0541.

Krebs:2021:MRP

[Kre21] Thijmen J. P. Krebs. A more reasonable proof of Cobham’s theorem. International Journal of Foundations of Computer Science (IJFCS), 32(02):203– 207, February 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500118.

Krishnan:1992:CTC

[Kri92] P. Krishnan. A calculus of timed communicating systems. International Journal of Foundations of Computer Science (IJFCS), 3(3): 303–322, September 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Krishnan:1997:PAA

[Kri97] Padmanabhan Krishnan. A process algebraic approach to time granularity semantics. International Journal of Foundations of Computer Science (IJFCS), 8(4):363– ??, December 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kuppusamy:2016:SDC

[KRK16] Lakshmanan Kuppusamy, Indhumathi Raman, and Kamala Krithivasan. On succinct description of certain context-free languages by ins-del and matrix ins-del systems. International Journal of Foundations of Computer Science (IJFCS), 27 (7):775–??, November 2016. CODEN IFCSEN. ISSN 0129-0541.

Klein:2006:CPM

[KS06] Shmuel T. Klein and Dana Shapira. Compressed pattern matching in JPEG images. International Journal of Foundations of Computer Science (IJFCS), 17 (6):1297–1306, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kellerer:2010:MTW

[KS10] Hans Kellerer and Vitaly A. Strusevich. Minimizing total weighted earliness-tardiness on a single machine around a small common due date: an FPTAS using quadratic knapsack. International Journal of Foundations of Computer Science (IJFCS), 21(3):357–383, June 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kari:2011:SPI

[KS11] Lila Kari and Shinnosuke Seki. Schema for parallel insertion and deletion: Revisited. International Journal of Foundations of Computer Science (IJFCS), 22(7):1655–1668, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Keeler:2019:BMN

[KS19] Chris Keeler and Kai Salomaa. Branching measures and nearly acyclic NFAs. International Journal of Foundations of Computer Science (IJFCS), 30(6– 7):1135–1155, September– November 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905411940032X.

Kaiser:2008:AAT

[KSJ08] Tim B. Kaiser, Stefan E.

Schmidt, and Cliff A. Joslyn. Adjusting annotated taxonomies. International Journal of Foundations of Computer Science (IJFCS), 19(2):345–358, April 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kumar:2022:NAT

[KSM22] Sunil Kumar, Harshdeep Singh, and Gaurav Mittal. A novel approach towards degree and Walsh-transform of Boolean functions. International Journal of Foundations of Computer Science (IJFCS), 33(05):453– 479, August 2022. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122500101.

Kedad-Sidhoum:2018:FSA

[KSMMT18] Safia Kedad-Sidhoum, Florence Monna, Grégory Mounié, and Denis Trystram. A family of scheduling algorithms for hybrid parallel platforms. International Journal of Foundations of Computer Science (IJFCS), 29(1):63–??, January 2018. CODEN IFCSEN. ISSN 0129-0541.

Klein:2008:MDE

[KSS08] Shmuel T. Klein, Tamar C. Serebro, and Dana Shapira. Modeling delta encoding of compressed files. International Journal of Foundations of Computer Science (IJFCS), 19(1):137– 146, February 2008. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kosub:2000:UCC

[KSV00] S. Kosub, H. Schmitz, and H. Vollmer. Uniform characterizations of complexity classes of functions. International Journal of Foundations of Computer Science (IJFCS), 11(4):525– 552, 2000. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Krithivasan:2003:DA

[KSV03] Kamala Krithivasan, K. Sharda, and Sandeep V. Varma. Distributed ω-automata. International Journal of Foundations of Computer Science (IJFCS), 14(4):681– ??, August 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kucuk:2014:FAA

[KSY14] Uğur Küçük, A. C. Cem Say, and Abuzer Yakaryilmaz. Finite automata with advice tapes. International Journal of Foundations of Computer Science (IJFCS), 25(8):987–??, December 2014. CODEN IFC-SEN. ISSN 0129-0541.

Kosmatopoulos:2020:SOH

[KTT20] Andreas Kosmatopoulos, Athanasios Tsakalidis, and Kostas Tsichlas. A spaceoptimal hidden surface removal algorithm for isooriented rectangles. International Journal of Foundations of Computer Science (IJFCS), 31(05):539– 549, August 2020. CO-DEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500240.

Kudlek:2007:SRQ

[Kud07] Manfred Kudlek. Some remarks on quantum automata. International Journal of Foundations of Computer Science (IJFCS), 18 (6):1283–1292, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kunimochi:2016:SPE

[Kun16] Yoshiyuki Kunimochi. Some properties of extractable codes and insertable codes. International Journal of Foundations of Computer Science (IJFCS), 27(3):327– ??, April 2016. CODEN IFCSEN. ISSN 0129-0541.

Kuroda:2020:MGA

[Kur20] Masamichi Kuroda. Monomial generalized almost perfect nonlinear functions. International Journal of Foundations of Computer Science (IJFCS), 31(03):411– 419, April 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500161.

Kutrib:2005:PNR

[Kut05] Martin Kutrib. The phenomenon of non-recursive trade-offs. International Journal of Foundations of Computer Science (IJFCS), 16(5):957–??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kari:2012:BSR

[KX12] Lila Kari and Zhi Xu.

De Bruijn sequences revisited. International Journal of Foundations of Computer Science (IJFCS), 23(6): 1307–??, September 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kenyon:1990:EBF

[KY90] Claire Kenyon and Andrew C. Yao. On evaluating Boolean functions with unreliable tests. International Journal of Foundations of Computer Science (IJFCS), 1(1):1–??, March 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Kobayashi:1996:FNL

[KY96] S. Kobayashi and T. Yokomori. Families of noncounting languages and their learnability from positive data. International Journal of Foundations of Computer Science (IJFCS), 7 (4):309–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ke:2017:AMB

[KYZS17] Pinhui Ke, Zhifan Ye, Zhengchun Zhou, and Jian Shen. Autocorrelation of the modified binary two-prime Sidelnikov sequence. International Journal of Foundations of Computer Science (IJFCS), 28(4):391–??, June 2017. CODEN IFCSEN. ISSN 0129-0541.

Kaminski:2010:FMA

[KZ10] Michael Kaminski and Daniel Zeitlin. Finitememory automata with non-deterministic reassignment. International Journal of Foundations of Computer Science (IJFCS), 21(5):741– 760, October 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lagogiannis:2014:PQD

[Lag14] George Lagogiannis. Parent queries over dynamic balanced parenthesis strings. International Journal of Foundations of Computer Science (IJFCS), 25(1):25– ??, January 2014. CODEN IFCSEN. ISSN 0129-0541.

Lagogiannis:2017:QOP

[Lag17] George Lagogiannis. Queryoptimal partially persistent B-trees with constant worstcase update time. International Journal of Foundations of Computer Science (IJFCS), 28(2):141– 169, February 2017. CO-DEN IFCSEN. ISSN 0129-0541.

Lam:2014:BSP

[LAHN14] Nhat Lam, Min Kyung An, Dung T. Huynh, and Trac Nguyen. Broadcast scheduling problem in SINR model. International Journal of Foundations of Computer Science (IJFCS), 25 (3):331–??, April 2014. CO-DEN IFCSEN. ISSN 0129-0541.

Larsen:1998:SOP

[Lar98]  Kim S. Larsen.  Sort

order problems in relational databases. International Journal of Foundations of Computer Science (IJFCS), 9(4):399–??, December 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Larsen:1999:GRA

[Lar99] K. S. Larsen. On grouping in relational algebra. International Journal of Foundations of Computer Science (IJFCS), 10(3):301– 312, 1999. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lazar:2013:BBS

[Láz13] Katalin Anna Lázár. A bridge between self-organizing networks and grammar systems theory. International Journal of Foundations of Computer Science (IJFCS), 24(4):501–??, June 2013. CODEN IFCSEN. ISSN 0129-0541.

Li:2004:QAU

[LB04] Xiaoyu Li and Howard Barnum. Quantum authentication using entangled states. International Journal of Foundations of Computer Science (IJFCS), 15(4):609– ??, August 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lipman:2003:NAA

[LBJ03] Justin Lipman, Paul Boustead, and John Judge. Neighbor Aware Adaptive Power flooding (NAAP) in mobile ad hoc networks. International Journal of Foundations of Computer Science (IJFCS), 14(2):237– ??, April 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lu:2006:CRC

[LBL06] Shiyong Lu, Arthur J. Bernstein, and Philip M. Lewis.

Completeness and realizability: Conditions for automatic generation of workflows. International Journal of Foundations of Computer Science (IJFCS), 17(1):223– ??, February 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ling:2002:SI

[LC02] C. X. Ling and N. Cercone. Special issue. International Journal of Foundations of Computer Science (IJFCS), 13(4):473–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Liu:2018:REC

[LC18] Yu-Liang Liu and Jou-Ming Chang. Realizing exchanged crossed cube communication patterns on linear array WDM optical networks. International Journal of Foundations of Computer Science (IJFCS), 29(6):1003– 1021, September 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118500181.

Li:2022:FCB

[LC22] Kun Li and Yanjun Chen.

Fuzzy clustering-based financial data mining system analysis and design. International Journal of Foundations of Computer Science (IJFCS), 33(6– 7):603–624, September– November 2022. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122420060.

Lu:2006:PFS

[LCL06] Shiyong Lu, Feng Cao, and Yi Lu. PAMA: a fast string matching algorithm. International Journal of Foundations of Computer Science (IJFCS), 17(2):357– 378, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lazar:2009:DFC

[LCVLV09] Katalin Anna Lázár, Erzsébet Csuhaj-Varjú, András Lőrincz, and György Vaszil. Dynamically formed clusters of agents in eco-grammar systems. International Journal of Foundations of Computer Science (IJFCS), 20(2):293– 311, April 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Liu:2019:TCP

[LCXS19] Xianping Liu, Yuan Chen, Yunge Xu, and Zhimin Sun. Triple-cycle permutations over finite fields of characteristic two. International Journal of Foundations of Computer Science (IJFCS), 30(2):272– 292, February 2019. ISSN 0129-0541.

Li:2012:MAR

[LCY12] Guoqiang Li, Xiaojuan Cai, and Shoji Yuen. Modeling and analysis of real-time systems with mutex components. International Journal of Foundations of Computer Science (IJFCS), 23(4):831– ??, June 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lederer:2001:ARV

[LD01] Edgar F. A. Lederer and Romeo A. Dumitrescu. Automatic result verification by complete run-time checking of computations. International Journal of Foundations of Computer Science (IJFCS), 12(1):97–124, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lai:2004:SGS

[LD04] Chun-Pong Lai and Cunsheng Ding. Several generalizations of Shamir’s secret sharing scheme. International Journal of Foundations of Computer Science (IJFCS), 15(2):445– ??, April 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Li:2017:ERD

[LDLW17] Yinkui Li, Mingzhe Du, Hongyan Li, and Xiaolin Wang. Edge rupture degree of graphs. International Journal of Foundations of Computer Science (IJFCS), 28(8):993–??, December 2017. CODEN IFC-SEN. ISSN 0129-0541.

Leopold:2003:CMA

[Leo03] Claudia Leopold. Cache miss analysis of 2D stencil codes with tiled time loop. International Journal of Foundations of Computer Science (IJFCS), 14(1):39– ??, February 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Leung:2004:ICA

[Leu04] Joseph Y.-T. Leung. Improved competitive algorithms for two-processor real-time systems. International Journal of Foundations of Computer Science (IJFCS), 15(5):733– ??, October 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Leung:2005:DCN

[Leu05] Hing Leung. Descriptional complexity of NFA of different ambiguity. International Journal of Foundations of Computer Science (IJFCS), 16(5):975–??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Leupold:2016:GIL

[Leu16] Peter Leupold. General idempotency languages over small alphabets. International Journal of Foundations of Computer Science (IJFCS), 27(3):343–??, April 2016. CODEN IFC-SEN. ISSN 0129-0541.

Larsen:1996:ERB

[LF96] K. S. Larsen and R. Fagerberg. Efficient rebalancing of B-trees with relaxed balance. International Journal of Foundations of Computer Science (IJFCS), 7 (2):169–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lin:2011:NIB

[LH11] Han-Yu Lin and Chien-Lung Hsu. A novel identitybased key-insulated convertible authenticated encryption scheme. International Journal of Foundations of Computer Science (IJFCS), 22(3):739–756, April 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Li:2024:GGN

[LHD^{+}24] Xianyong Li, Jiaming Huang, Yajun Du, Yongquan Fan, , and Xiaoliang Chen. The g-good-neighbor conditional diagnosabilities of hypermesh optical interconnection networks under the PMC and comparison models. International Journal of Foundations of Computer Science (IJFCS), 35(03): 313–325, April 2024. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123500041.

Li:2011:LBS

[LHG11] Xuelian Li, Yupu Hu, and Juntao Gao. Lower bounds on the second order nonlinearity of Boolean functions. International Journal of Foundations of Computer Science (IJFCS), 22 (6):1331–1349, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lee:2009:NCA

[LHT09] Cheng-Chi Lee, Min-Shiang Hwang, and Shiang-Feng Tzeng. A new convertible authenticated encryption scheme based on the ElGamal cryptosystem. International Journal of Foundations of Computer Science (IJFCS), 20(2):351– 359, April 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Li:2000:MEE

[Li00a] K. Li. A method for evaluating the expected load of dynamic tree embeddings in hypercubes. International Journal of Foundations of Computer Science (IJFCS), 11(2):207–230, 2000. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Li:2000:PRC

[Li00b] K. Li. Part 1 (randomized computing). International Journal of Foundations of Computer Science (IJFCS), 11(2):207–??, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Li:2001:EJS

[Li01]  Keqin Li.  An efficient

job scheduling algorithm in partitionable mesh connected systems. International Journal of Foundations of Computer Science (IJFCS), 12(6):763– 774, December 2001. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Li:2007:IDA

[Li07] Ming Li. Information distance and its applications. International Journal of Foundations of Computer Science (IJFCS), 18(4):669– 681, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Li:2012:PAE

[Li12a] Keqin Li. Performance analysis and evaluation of random walk algorithms on wireless networks. International Journal of Foundations of Computer Science (IJFCS), 23(4):779– ??, June 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Li:2012:PHC

[Li12b] Keqin Li. Probing highcapacity peers to reduce download times in P2P file sharing systems with stochastic service capacities. International Journal of Foundations of Computer Science (IJFCS), 23(6): 1341–??, September 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lin:2007:FSS

[Lin07] Cho-Chin Lin. A framework for solving sequence problem of multiple input streams. International Journal of Foundations of Computer Science (IJFCS), 18(5):1043–1064, October 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lindell:2008:NFF

[Lin08a] Steven Lindell. A normal form for first-order logic over doubly-linked data structures. International Journal of Foundations of Computer Science (IJFCS), 19(1):205– 217, February 2008. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Link:2008:IMD

[Lin08b] Sebastian Link. On the implication of multivalued dependencies in partial database relations. International Journal of Foundations of Computer Science (IJFCS), 19(3):691– 715, June 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lisitsa:1993:CUC

[Lis93] A. P. Lisitsa. Complexity of universal circumscription. International Journal of Foundations of Computer Science (IJFCS), 4(3): 241–244, September 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Li:2017:MMA

[LJ17] Rongjia Li and Chenhui Jin. Meet-in-the-middle attack on 11-round 3D block cipher. International Journal of Foundations of Computer Science (IJFCS), 28 (1):19–28, January 2017. CODEN IFCSEN. ISSN 0129-0541.

Lin:2009:VST

[LJA09] Jie Lin, Yue Jiang, and Don Adjeroh. The virtual suffix tree. International Journal of Foundations of Computer Science (IJFCS), 20(6):1109–1133, December 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Liu:2022:P

[LJF22] Jia-Bao Liu, Muhammad Javaid, and Mohammad Reza Farahani. Preface. International Journal of Foundations of Computer Science (IJFCS), 33(6–7):481–487, September–November 2022. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122020026.

Lin:2017:IID

[LJH^{+}17] Jie Lin, Yue Jiang, E. James Harner, Bing-Hua Jiang, and Don Adjeroh. IDPM: An improved degenerate pattern matching algorithm for biological sequences. International Journal of Foundations of Computer Science (IJFCS), 28(7):889–??, November 2017. CODEN IFCSEN. ISSN 0129-0541.

Langer:2011:PAE

[LK11] Miroslav Langer and Alica Kelemenová. Positioned agents in eco-grammar systems. International Journal of Foundations of Computer Science (IJFCS), 22(1):237– 246, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lorincz:2002:IHP

[LKM02] A. Lörincz, I. Kókai, and A. Meretei. Intelligent high-performance crawlers used to reveal topic-specific structure of WWW. International Journal of Foundations of Computer Science (IJFCS), 13(4):477–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Liu:2016:SNC

[LL16] Haibo Liu and Qunying Liao. Some new constructions for generalized zerodifference balanced functions. International Journal of Foundations of Computer Science (IJFCS), 27(8):897– 908, December 2016. CO-DEN IFCSEN. ISSN 0129-0541.

Landwehr:2020:PBT

[LL20] Patrick Landwehr and Christof Löding. Projection for Büchi tree automata with constraints between siblings. International Journal of Foundations of Computer Science (IJFCS), 31(06):749–775, September 2020. ISSN 0129-

0541. URL https://www.
worldscientific.com/doi/
10.1142/S012905412041004X.

Lin:2023:GBP

[LL23] Lan Lin and Yixun Lin. Graph bipartization problem with applications to via minimization in VLSI design. International Journal of Foundations of Computer Science (IJFCS), 34 (04):347–361, June 2023. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500198.

Li:2021:HCE

[LLL21] Chunfang Li, Shangwei Lin, and Shengjia Li. Hamiltonian cycle embeddings in faulty hypercubes under the forbidden faulty set model. International Journal of Foundations of Computer Science (IJFCS), 32(01):53– 72, January 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500039.

Li:2022:FEC

[LLL22] Pingshan Li, Rong Liu, and Xianglin Liu. The (E)FTSM-(edge) connectivity of Cayley graphs generated by transposition trees. International Journal of Foundations of Computer Science (IJFCS), 33(01):33– 43, January 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500349.

Li:2006:MTW

[LLQ06] Shuguang Li, Guojun Li, and Xingqin Qi. Minimizing total weighted completion time on identical parallel batch machines. International Journal of Foundations of Computer Science (IJFCS), 17(6):1441–1453, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lobel:2021:BTW

[LLS21] Raphaela Löbel, Michael Luttenberger, and Helmut Seidl. On the balancedness of tree-to-word transducers. International Journal of Foundations of Computer Science (IJFCS), 32(06):761– 783, September 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121420077.

Lin:2018:ELE

[LLW18] Nianfeng Lin, Damei Lü, and Jinhua Wang. L(2, 1)- edge-labelings of the edgepath-replacement of a graph. International Journal of Foundations of Computer Science (IJFCS), 29(1):91– ??, January 2018. CODEN IFCSEN. ISSN 0129-0541.

Lv:2021:EAC

[LLW21] Yali Lv, Cheng-Kuan Lin, and Guijuan Wang. An exchanged 3-ary n-cube interconnection network for parallel computation. International Journal of Foundations of Computer Science (IJFCS), 32(03):235– 252, April 2021. CO-DEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500131.

Lin:2022:OPD

[LLW^{+}22] Chih-Yuan Lin, Jia-Jie Liu, Yue-Li Wang, William Chung-Kung Yen, and Chiun-Chieh Hsu. The outer-paired domination of graphs. International Journal of Foundations of Computer Science (IJFCS), 33(02):141–148, February 2022. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054122500034.

Li:2013:SSC

[LLY13] Jing Li, Di Liu, and Jun Yuan. The super spanning connectivity and super spanning laceability of tori with faulty elements. International Journal of Foundations of Computer Science (IJFCS), 24(6):921–??, September 2013. CODEN IFCSEN. ISSN 0129-0541.

Leung:2007:STM

[LLZ07] Joseph Y.-T. Leung, Haibing Li, and Hairong Zhao. Scheduling two-machine flow shops with exact delays. International Journal of Foundations of Computer Science (IJFCS), 18(2):341– 359, April 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Li:2020:MUD

[LMG20] Juyan Li, Chunguang Ma, and Zhen Gu. Multi-use deterministic public key proxy re-encryption from lattices in the auxiliary-input setting. International Journal of Foundations of Computer Science (IJFCS), 31(05): 551–567, August 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500252.

Loukil:2012:PHG

[LMM^{+}12] Lakhdar Loukil, Malika Mehdi, Nouredine Melab, El-Ghazali Talbi, and Pascal Bouvry. Parallel hybrid genetic algorithms for solving Q3Ap on computational grid. International Journal of Foundations of Computer Science (IJFCS), 23 (2):483–500, February 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Loos:2008:DCS

[LMW08] Remco Loos, Andreas Malcher, and Detlef Wotschke. Descriptional complexity of splicing systems. International Journal of Foundations of Computer Science (IJFCS), 19(4):813– 826, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Li:2020:UMM

[LMZC20] Jing Li, Chris Melekian, Shurong Zuo, and Eddie Cheng. Unpaired many-to-many disjoint path covers on bipartite k-ary n-cube networks with faulty elements. International Journal of Foundations of Computer Science (IJFCS), 31(03): 371–383, April 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120500148.

Lisitsa:2008:RAV

[LN08] Alexei Lisitsa and Andrei P. Nemytykh. Reachability analysis in verification via supercompilation. International Journal of Foundations of Computer Science (IJFCS), 19(4):953– 969, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

LaTorre:2016:SBP

[LNP16] Salvatore La Torre, Margherita Napoli, and Gennaro Parlato. Scope-bounded pushdown languages. International Journal of Foundations of Computer Science (IJFCS), 27(2):215–??, February 2016. CODEN IFCSEN. ISSN 0129-0541.

Lehtinen:2010:BGG

[LO10] Tommi Lehtinen and Alexander Okhotin. Boolean grammars and GSM mappings. International Journal of Foundations of Computer Science (IJFCS), 21(5):799– 815, October 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lehtinen:2011:ESN

[LO11] Tommi Lehtinen and Alexander Okhotin. On equations over sets of numbers and their limitations. International Journal of Foundations of Computer Science (IJFCS), 22(2):377– 393, February 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lehtinen:2013:HPD

[LO13] Tommi Lehtinen and Alexander Okhotin. Homomorphisms preserving deterministic context-free languages. International Journal of Foundations of Computer Science (IJFCS), 24(7): 1049–??, November 2013. CODEN IFCSEN. ISSN 0129-0541.

Lauer:2007:UEDa

[LOD07a] Tobias Lauer, Thomas Ottmann, and Amitava Datta. Update-efficient data structures for dynamic IP router tables. International Journal of Foundations of Computer Science (IJFCS), 18(1):139– 161, February 2007. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lauer:2007:UEDb

[LOD07b] Tobias Lauer, Thomas Ottmann, and Amitava Datta. Update-efficient data structures for dynamic IP router tables. International Journal of Foundations of Computer Science (IJFCS), 18(2):295–317, April 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Loding:2015:SPD

[Löd15] Christof Löding. Simplification problems for deterministic pushdown automata on infinite words. International Journal of Foundations of Computer Science (IJFCS), 26(8):1041–??, December 2015. CODEN IFC-SEN. ISSN 0129-0541.

Lohrey:2005:DCA

[Loh05] Markus Lohrey. Decidability and complexity in automatic monoids. International Journal of Foundations of Computer Science (IJFCS), 16(4):707– ??, August 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lohrey:2010:CMP

[Loh10] Markus Lohrey. Compressed membership problems for regular expressions and hierarchical automata. International Journal of Foundations of Computer Science (IJFCS), 21(5):817– 841, October 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lopez-Ortiz:2018:AOC

[LOPR18] Alejandro López-Ortiz, Cynthia B. Perez, and Jazmín Romero. Arbitrary overlap constraints in graph packing problems. International Journal of Foundations of Computer Science (IJFCS), 29(1):101–??, January 2018. CODEN IFCSEN. ISSN 0129-0541.

Liestman:1998:NPD

[LOZ98] A. L. Liestman, J. Opatrny, and M. Zaragoza. Network properties of double and triple fixed step graphs. International Journal of Foundations of Computer Science (IJFCS), 9(1): 57–??, March 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Laine:2011:WEO

[LP11] Markku Laine and Wojciech Plandowski. Word equations with one unknown. International Journal of Foundations of Computer Science (IJFCS), 22 (2):345–375, February 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lavado:2019:CRR

[LP19] Giovanna J. Lavado and Luca Prigioniero. Concise representations of reversible automata. International Journal of Foundations of Computer Science (IJFCS), 30(6–7): 1157–1175, September–November 2019. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054119400331.

Li:2011:DPF

[LPC11] Yamin Li, Shietung Peng, and Wanming Chu. Disjointpaths and fault-tolerant routing on recursive dualnet. International Journal of Foundations of Computer Science (IJFCS), 22 (5):1001–1018, August 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Luccio:1992:AIP

[LPP92] F. Luccio, A. Pietracaprina, and G. Pucci. Analysis and implementation of parallel uniform hashing. International Journal of Foundations of Computer Science (IJFCS), 3(1):55–64, March 1992. CODEN IFCSEN.

ISSN 0129-0541 (print), 1793-6373 (electronic).

Luccio:2007:NDP

[LPS07] Fabrizio Luccio, Linda Pagli, and Nicola Santoro. Network decontamination in presence of local immunity. International Journal of Foundations of Computer Science (IJFCS), 18(3):457– 474, June 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Luo:2004:PDE

[LR04] Jun Luo and Sanguthevar Rajasekaran. Parallelizing 1-dimensional estuarine model. International Journal of Foundations of Computer Science (IJFCS), 15 (6):809–??, December 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lancia:2008:FLM

[LRR08] Giuseppe Lancia, Franca Rinaldi, and Romeo Rizzi. Flipping letters to minimize the support of a string. International Journal of Foundations of Computer Science (IJFCS), 19(1):5–17, February 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lodaya:1992:TLC

[LRT92] K. Lodaya, R. Ramanujam, and P. S. Thiagarajan. Temporal logics for communicating sequential agents: I. International Journal of Foundations of Computer Science (IJFCS), 3(2):117– 160, June 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

LeVerge:1998:NRC

[LS98] Hervé Le Verge and Yannic Saouter. New results on computability of recurrence equations. International Journal of Foundations of Computer Science (IJFCS), 9(3):249–??, 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lomuscio:2013:AGR

[LSWW13] Alessio Lomuscio, Ben Strulo, Nigel Walker, and Peng Wu. Assume-guarantee reasoning with local specifications. International Journal of Foundations of Computer Science (IJFCS), 24 (4):419–??, June 2013. CO-DEN IFCSEN. ISSN 0129-0541.

Levit:2021:RGS

[LT21] Vadim E. Levit and David Tankus. Recognizing generating subgraphs revisited. International Journal of Foundations of Computer Science (IJFCS), 32(01):93– 114, January 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500052.

Loding:2024:BCD

[LT24] Christof Löding and Wolfgang Thomas. On the Boolean closure of deterministic top-down tree automata. International Journal of Foundations of Computer Science (IJFCS), 35(1–2):11–22, January– February 2024. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123480015.

Lepere:2002:AAS

[LTW02] R. Lepère, D. Trystram, and G. J. Woeginger. Approximation algorithms for scheduling malleable tasks under precedence constraints. International Journal of Foundations of Computer Science (IJFCS), 13 (4):613–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lee:2012:PMF

[LTZ12] Young Choon Lee, Javid Taheri, and Albert Y. Zomaya. A parallel metaheuristic framework based on harmony search for scheduling in distributed computing systems. International Journal of Foundations of Computer Science (IJFCS), 23(2):445– 464, February 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lucanu:2009:RLB

[Luc09] Dorel Lucanu. Rewriting logic-based semantics of P systems and the maximal concurrency. International Journal of Foundations of Computer Science (IJFCS), 20(3):395–410, June 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Luck:2018:QQO

[Lüc18] Martin Lück. Quirky quantifiers: Optimal models and complexity of computation tree logic. International Journal of Foundations of Computer Science (IJFCS), 29(1):17–??, January 2018. CODEN IFCSEN. ISSN 0129-0541.

Lugiez:2011:FAD

[Lug11] Denis Lugiez. Forward analysis of dynamic network of pushdown systems is easier without order. International Journal of Foundations of Computer Science (IJFCS), 22(4):843–862, June 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lozin:2008:CWB

[LV08] Vadim V. Lozin and Jordan Volz. The clique-width of bipartite graphs in monogenic classes. International Journal of Foundations of Computer Science (IJFCS), 19(2):477–494, April 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lai:1993:TUA

[LW93] T. W. Lai and D. Wood.

A top-down updating algorithm for weight-balanced trees. International Journal of Foundations of Computer Science (IJFCS), 4(4): 309–324, December 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Laube:2005:CIS

[LW05] Uli Laube and Maik Weinard.

Conditional inequalities and the shortest common superstring problem. International Journal of Foundations of Computer Science (IJFCS), 16(6):1219– ??, December 2005. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic). See erratum [LW06a].

Laube:2006:ECI

[LW06a] U. Laube and M. Weinard.

Erratum: “Conditional Inequalities and the Shortest Common Superstring Problem”. International Journal of Foundations of Computer Science (IJFCS), 17(1):247– ??, February 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic). See [LW05].

Liu:2006:ESM

[LW06b] Joseph K. Liu and Duncan S. Wong. Enhanced security models and a generic construction approach for linkable ring signature. International Journal of Foundations of Computer Science (IJFCS), 17(6):1403– 1422, December 2006. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lu:2021:UMM

[LW21] Huazhong Lü and Tingzeng Wu. Unpaired many-to-many disjoint path cover of balanced hypercubest. International Journal of Foundations of Computer Science (IJFCS), 32(08):943– 956, December 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500301.

Little:2010:AMS

[LWJ^{+}10] Scott Little, David Walter, Kevin Jones, Chris Myers, and Alper Sen. Analog/ mixed-signal circuit verification using models generated from simulation traces. International Journal of Foundations of Computer Science (IJFCS), 21(2):191– 210, April 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Lv:2020:SAP

[LWS^{+}20] Jiaxian Lv, Yi Wang, Jinshu Su, Rongmao Chen, and Wenjun Wu. Security of auditing protocols against subversion attacks. International Journal of Foundations of Computer Science (IJFCS), 31(2):193– 206, February 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500033.

Liebers:2000:HRB

[LWW00] A. Liebers, D. Wagner, and K. Weihe. On the hardness of recognizing bundles in time table graphs. International Journal of Foundations of Computer Science (IJFCS), 11(3):467– 484, 2000. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Liu:2022:MMM

[LWW22] Changwei Liu, Kexin Wang, and Aman Wu. Management and monitoring of multi-behavior recommendation systems using graph convolutional neural networks. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):583–601, September– November 2022. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122420059.

Li:2014:TTG

[LWYL14] Yinkui Li, Zongtian Wei, Xiaokui Yue, and Erqiang Liu. Tenacity of total graphs. International Journal of Foundations of Computer Science (IJFCS), 25 (5):553–??, August 2014. CODEN IFCSEN. ISSN 0129-0541.

Leiberherr:1994:CAS

[LX94] K. L. Leiberherr and C. Xiao. Customizing adaptive software to object-oriented software using grammars. International Journal of Foundations of Computer Science (IJFCS), 5(3/4):??, 1994. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Li:2017:SSC

[LX17] Pingshan Li and Min Xu. The super spanning connectivity of arrangement graphs. International Journal of Foundations of Computer Science (IJFCS), 28(8):1047–??, December 2017. CODEN IFCSEN. ISSN 0129-0541.

Li:2019:FFH

[LX19] Pingshan Li and Min Xu.

Fault-free Hamiltonian cycles in balanced hypercubes with conditional edge faults. International Journal of Foundations of Computer Science (IJFCS), 30(5):693– 717, August 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119500175.

Leung:1994:HMN

[LY94] J. Y.-T. Leung and V. K. M.

Yu. Heuristic for minimizing the number of late jobs on two processors. International Journal of Foundations of Computer Science (IJFCS), 5(3/4):??, 1994. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Li:2017:HTN

[LYG17] Jing Li, Yuxing Yang, and Xiaohui Gao. Hamiltonicity of the torus network under the conditional fault model. International Journal of Foundations of Computer Science (IJFCS), 28 (3):211–??, April 2017. CO-DEN IFCSEN. ISSN 0129-0541.

Li:2015:CFT

[LYH^{+}15] Xianyong Li, Xiaofan Yang, Li He, Cui Yu, and Jing Zhang. Conditional fault tolerance of hypermesh optical interconnection networks. International Journal of Foundations of Computer Science (IJFCS), 26 (1):159–??, January 2015. CODEN IFCSEN. ISSN 0129-0541.

Liu:2019:SEC

[LYHW19] Xiangxin Liu, Xiaoyuan Yang, Yiliang Han, and Xu An Wang. A secure and efficient code-based signature scheme. International Journal of Foundations of Computer Science (IJFCS), 30(4):635– 645, June 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400173.

Lai:2019:NIB

[LYX^{+}19] Qiqi Lai, Bo Yang, Zhe Xia, Yannan Li, Yuan Chen, and Zhenlong Li. Novel identity-based hash proof system with compact master public key from lattices in the standard model. International Journal of Foundations of Computer Science (IJFCS), 30(4):589– 606, June 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400148.

Liu:2021:CCS

[LYY^{+}21] Jinhui Liu, Yong Yu, Bo Yang, Jianwei Jia, and Qiqi Lai. Cryptanalysis of Cramer–Shoup like cryptosystems based on index exchangeable family. International Journal of Foundations of Computer Science (IJFCS), 32(01):73– 91, January 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500040.

Lange:1993:LRL

[LZ93] S. Lange and T. Zeugmann. Learning recursive languages with a bounded number of mind changes. International Journal of Foundations of Computer Science (IJFCS), 4(2):157– 178, June 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Liao:2012:AST

[LZ12] Chung-Shou Liao and Louxin Zhang. Approximating the spanning k-tree forest problem. International Journal of Foundations of Computer Science (IJFCS), 23(7): 1543–??, November 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Liao:2015:NOC

[LZ15] Qunying Liao and Juan Zhu.

A note on optimal constant dimension codes. International Journal of Foundations of Computer Science (IJFCS), 26(1):143–??, January 2015. CODEN IFC-SEN. ISSN 0129-0541.

Lu:2016:MRL

[LZGF16] Jiyong Lu, Jun Zhang, Xuan Guang, and Fang-Wei Fu. Multiple repair localities with distinct erasure tolerance. International Journal of Foundations of Computer Science (IJFCS), 27(6):665– ??, September 2016. CO-DEN IFCSEN. ISSN 0129-0541.

Leporati:2006:SIB

[LZGN06] Alberto Leporati, Claudio Zandron, and Miguel A. Gutiérrez-Naranjo. P systems with input in binary form. International Journal of Foundations of Computer Science (IJFCS), 17(1):127– ??, February 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Liu:2022:CEC

[LZZN22] Xiaoqing Liu, Shuming Zhou, Hong Zhang, and Baohua Niu. The component (edge) connectivity of round matching composition networks. International Journal of Foundations of Computer Science (IJFCS), 33(08):1005– 1018, December 2022. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122500125.

Macarie:1996:NMF

[Mac96] I. I. Macarie. A note of multihead finite-state automata. International Journal of Foundations of Computer Science (IJFCS), 7 (4):329–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Madhu:2003:PRS

[Mad03] Mutyam Madhu. Probabilistic rewriting P systems. International Journal of Foundations of Computer Science (IJFCS), 14(1):157–??, February 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Morris:2009:USP

[MAG09] Peter Morris, Thorsten Altenkirch, and Neil Ghani. A universe of strictly positive families. International Journal of Foundations of Computer Science (IJFCS), 20 (1):83–107, February 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Maletti:2005:RTS

[Mal05] Andreas Maletti. Relating tree series transducers and weighted tree automata. International Journal of Foundations of Computer Science (IJFCS), 16(4):723– ??, August 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Maletti:2007:PS

[Mal07]  Andreas Maletti.  Pure

and O-substitution. International Journal of Foundations of Computer Science (IJFCS), 18(4):829– 845, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Maletti:2015:PWR

[Mal15] Andreas Maletti. The power of weighted regularitypreserving multi bottomup tree transducers. International Journal of Foundations of Computer Science (IJFCS), 26(7):987–??, November 2015. CODEN IFCSEN. ISSN 0129-0541.

Maletti:2018:CTT

[Mal18] Andreas Maletti. Compositions of tree-to-tree statistical machine translation models. International Journal of Foundations of Computer Science (IJFCS), 29(5):877– 892, August 2018. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054118420091.

Maletti:2024:CWE

[Mal24] Andreas Maletti. Compositions of weighted extended tree transducers

— the unambiguous case. International Journal of Foundations of Computer Science (IJFCS), 35(1–2):

107–127, January–February 2024. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054123480064.

Miura:2005:CDR

[MAN05] Kazuyuki Miura, Machiko Azuma, and Takao Nishizeki. Canonical decomposition, realizer, Schnyder labeling and orderly spanning trees of plane graphs. International Journal of Foundations of Computer Science (IJFCS), 16(1):117–??, February 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Miura:2006:CDP

[MAN06] Kazuyuki Miura, Machiko Azuma, and Takao Nishizeki. Convex drawings of plane graphs of minimum outer apices. International Journal of Foundations of Computer Science (IJFCS), 17 (5):1115–??, October 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Maneth:2015:SDE

[Man15] Sebastian Maneth. A survey on decidable equivalence problems for tree transducers. International Journal of Foundations of Computer Science (IJFCS), 26(8):1069–??, December 2015. CODEN IFCSEN. ISSN 0129-0541.

Marche:1992:WPA

[Mar92] C. Marche. The word problem of ACD-ground theories is undecidable. International Journal of Foundations of Computer Science (IJFCS), 3(1):81–92, March 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Martin:1997:ETA

[Mar97] Bruno Martin. Embedding torus automata into a ring of automata. International Journal of Foundations of Computer Science (IJFCS), 8(4):425–??, December 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Margenstern:2008:FTP

[Mar08a] Maurice Margenstern. The finite tiling problem is undecidable in the hyperbolic plane. International Journal of Foundations of Computer Science (IJFCS), 19(4):971– 982, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Margenstern:2008:CCA

[Mar08b] Maurice Margenstern. On a characterization of cellular automata in tilings of the hyperbolic plane. International Journal of Foundations of Computer Science (IJFCS), 19(5):1235–1257, October 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Martyugin:2009:LSR

[Mar09]  Pavel Martyugin.  The

length of subset reachability in nondeterministic automata. International Journal of Foundations of Computer Science (IJFCS), 20 (5):887–900, October 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mastrolilli:2004:SMM

[Mas04] Monaldo Mastrolilli. Scheduling to minimize max flow time: Off-line and on-line algorithms. International Journal of Foundations of Computer Science (IJFCS), 15(2):385–??, April 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Masopust:2009:TDM

[Mas09]  Tomáš Masopust.  On

the terminating derivation mode in cooperating distributed grammar systems with forbidding components. International Journal of Foundations of Computer Science (IJFCS), 20(2):331– 340, April 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Masopust:2013:NLP

[Mas13] Tomáš Masopust. A note on limited pushdown alphabets in stateless deterministic pushdown automata. International Journal of Foundations of Computer Science (IJFCS), 24(3):319–??, April 2013. CODEN IFC-SEN. ISSN 0129-0541.

Maslennikova:2019:RCI

[Mas19] Marina Maslennikova. Reset complexity of ideal languages over a binary alphabet. International Journal of Foundations of Computer Science (IJFCS), 30(6– 7):1177–1196, September– November 2019. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054119400343.

Matsumae:2004:SMS

[Mat04] Susumu Matsumae. Simulation of meshes with separable buses by meshes with multiple partitioned buses. International Journal of Foundations of Computer Science (IJFCS), 15(3):475– ??, June 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Martini:2003:DHM

[MB03] Paul M. Martini and Walter A. Burkhard. Double hashing with multiple passbits. International Journal of Foundations of Computer Science (IJFCS), 14(6): 1165–??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Muskulus:2006:CBC

[MB06] Michael Muskulus and Robert Brijder. Complexity of bio-computation: Symbolic dynamics in membrane systems. International Journal of Foundations of Computer Science (IJFCS), 17 (1):147–??, February 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Moslehi:2017:SBP

[MB17] Zahra Moslehi and Alireza Bagheri. Separating bichromatic point sets by minimal triangles with a fixed angle. International Journal of Foundations of Computer Science (IJFCS), 28 (4):309–??, June 2017. CO-DEN IFCSEN. ISSN 0129-0541.

Masse:2018:P

[MBR18] Alexandre Blondin Massé, Srečko Brlek, and Christophe Reutenauer. Preface. International Journal of Foundations of Computer Science (IJFCS), 29(5):701– 704, August 2018. CO-DEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118020033.

Mantler:2002:CRB

[MC02] Andrea Mantler and Helen Cameron. Constructing redblack tree shapes. International Journal of Foundations of Computer Science (IJFCS), 13(6):837–??, December 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mumme:2013:EFS

[MC13] Malcolm Mumme and Gianfranco Ciardo. An efficient fully symbolic bisimulation algorithm for nondeterministic systems. International Journal of Foundations of Computer Science (IJFCS), 24(2):263–??, February 2013. CODEN IFCSEN. ISSN 0129-0541.

Montoro:2011:FPN

[MCM^{+}11] Fernando Arroyo Montoro, Juan Castellanos, Victor Mitrana, Eugenio Santos, and Jose M. Sempere. Filter position in networks of substitution processors does not matter. International Journal of Foundations of Computer Science (IJFCS), 22 (1):155–165, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

McNaughton:1990:DFL

[McN90] Robert McNaughton. The development of formal language theory since 1956. International Journal of Foundations of Computer Science (IJFCS), 1(4):355–??, December 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mardare:2008:LCR

[MCS08] Radu Mardare, Matteo Cavaliere, and Sean Sedwards. A logical characterization of robustness, mutants and species in colonies of agents. International Journal of Foundations of Computer Science (IJFCS), 19(5): 1199–1221, October 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mahapatra:2000:RSG

[MD00] N. R. Mahapatra and S. Dutt. Random seeking: a general, efficient, and informed randomized scheme for dynamic load balancing. International Journal of Foundations of Computer Science (IJFCS), 11 (2):231–246, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Martinez-Del-Amor:2011:SAM

[MDAPHPJ^{+}11] M. A. Martínez-Del-Amor, I. Pérez-Hurtado, M. J. Pérez-Jiménez, A. Riscos-Núñez, and F. Sancho-Caparrini. A simulation algorithm for multienvironment probabilistic P systems: a formal verification. International Journal of Foundations of Computer Science (IJFCS), 22(1):107– 118, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Masse:2013:MW

[MDGH13] Alexandre Blondin Massé, Sarah Desmeules, Sébastien Gaboury, and Sylvain Hallé. Multipseudoperiodic words. International Journal of Foundations of Computer Science (IJFCS), 24(7): 1153–??, November 2013. CODEN IFCSEN. ISSN 0129-0541.

Monien:1997:CLS

[MDL97] Burkhard Monien, Ralf Diekmann, and Reinhard Lüling. The construction of large scale reconfigurable parallel computing systems (the architecture of the SC320). International Journal of Foundations of Computer Science (IJFCS), 8 (3):347–??, September 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Meer:2012:SIT

[Mee12] Klaus Meer. Some initial thoughts on bounded query computations over the reals. International Journal of Foundations of Computer Science (IJFCS), 23(7): 1499–??, November 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Melnikov:1993:ECI

[Mel93]  B. F. Melnikov.  The

equality condition for infinite catenations of two sets of finite words. International Journal of Foundations of Computer Science (IJFCS), 4(3):267–??, September 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mereghetti:2008:TDP

[Mer08] Carlo Mereghetti. Testing the descriptional power of small Turing machines on nonregular language acceptance. International Journal of Foundations of Computer Science (IJFCS), 19(4):827– 843, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mandal:2014:SAT

[MG14] Partha Sarathi Mandal and Anil K. Ghosh. A statistical approach towards secure location verification in noisy wireless channels. International Journal of Foundations of Computer Science (IJFCS), 25(5):563–??, August 2014. CODEN IFC-SEN. ISSN 0129-0541.

Mandal:2020:CTT

[MG20] Subhrangsu Mandal and Arobinda Gupta. Convergecast tree on temporal graphs. International Journal of Foundations of Computer Science (IJFCS), 31(03):385– 409, April 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412050015X.

Montoro:2020:LTS

[MGCVdlP20] Fernando Arroyo Montoro, Sandra Gómez-Canaval, Karina Jiménez Vega, and Alfonso Ortega de la Puente. A linear time solution for Nqueens problem using generalized networks of evolutionary polarized processors. International Journal of Foundations of Computer Science (IJFCS), 31(1):7– 21, January 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120400018.

Melodelima:2008:MAA

[MGGP08] Christelle Melodelima, Laurent Gueguen, Christian Gautier, and Didier Piau. A Markovian approach for the analysis of the gene structure. International Journal of Foundations of Computer Science (IJFCS), 19(1):19– 35, February 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Meng:2019:SDD

[MGJ19] Wenjuan Meng, Jianhua Ge, and Tao Jiang. Secure data deduplication with reliable data deletion in cloud. International Journal of Foundations of Computer Science (IJFCS), 30(4):551– 570, June 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400124.

Ma:2023:SMC

[MGL23] Meijie Ma, Chaoming Guo, , and Xiang-Jun Li. Strong Menger connectivity of folded hypercubes with faulty subcube. International Journal of Foundations of Computer Science (IJFCS), 34(05):443– 451, August 2023. CO-DEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500253.

Mang:2006:CCA

[MH06] Freddy Y. C. Mang and Pei-Hsin Ho. Controllability and cooperativeness analysis for automatic abstraction refinement. International Journal of Foundations of Computer Science (IJFCS), 17(4):763–??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Masakova:2012:P

[MH12] Zuzana Masáková and Štěpán Holub. Preface. International Journal of Foundations of Computer Science (IJFCS), 23(8):1579–??, December 2012. CODEN IFC-SEN. ISSN 0129-0541.

Matsumoto:2009:RTE

[MHT09] Tetsuya Matsumoto, Kazuhito Hagio, and Masayuki Takeda. A run-time efficient implementation of compressed pattern matching automata. International Journal of Foundations of Computer Science (IJFCS), 20(4):717– 733, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mignosi:1990:SWA

[Mig90]  Filippo Mignosi.  Stur-

mian words and ambiguous context-free languages. International Journal of Foundations of Computer Science (IJFCS), 1(3):309–??, September 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Man:2011:EPS

[MIN11] Duhu Man, Yasuaki Ito, and Koji Nakano. An efficient parallel sorting compatible with the standard qsort. International Journal of Foundations of Computer Science (IJFCS), 22(5):1057– 1071, August 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mondal:2011:MQS

[MKB^{+}11] Debashis Mondal, Abhay Kumar, Arijit Bishnu, Krishnendu Mukhopadhyaya, and Subhas C. Nandy. Measuring the quality of surveillance in a wireless sensor network. International Journal of Foundations of Computer Science (IJFCS), 22(4):983–998, June 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Maur:2012:P

[ML12] Giancarlo Maur and Alberto Leporati. Preface. International Journal of Foundations of Computer Science (IJFCS), 23(5):965– ??, August 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ma:2017:LBI

[MLO17] Chunguang Ma, Juyan Li, and Weiping Ouyang. Lattice-based identity-based homomorphic conditional proxy re-encryption for secure big data computing in cloud environment. International Journal of Foundations of Computer Science (IJFCS), 28(6):645–??, September 2017. CODEN IFCSEN. ISSN 0129-0541.

Manacher:1997:FMC

[MM97] Glenn K. Manacher and Terrance A. Mankus. Finding a maximum clique in a set of proper circular arcs in time O(n) with applications. International Journal of Foundations of Computer Science (IJFCS), 8(4):443– ??, December 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mancheron:2005:CCL

[MM05] Alban Mancheron and Christophe Moan. Combinatorial characterization of the language recognized by factor and suffix oracles. International Journal of Foundations of Computer Science (IJFCS), 16(6):1179–??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mandal:2007:MAB

[MM07] Partha Sarathi Mandal and Krishnendu Mukhopadhyaya. Mobile agent based checkpointing with concurrent initiations. International Journal of Foundations of Computer Science (IJFCS), 18(5):1107–1122, October 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Manca:2011:LGS

[MM11] Vincenzo Manca and Luca Marchetti. Log-gain stoichiometric stepwise regression for MP systems. International Journal of Foundations of Computer Science (IJFCS), 22(1):97–106, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mairesse:2017:USS

[MM17] Jean Mairesse and Irène Marcovici. Uniform sampling of subshifts of finite type on grids and trees. International Journal of Foundations of Computer Science (IJFCS), 28(3):263–??, April 2017. CODEN IFC-SEN. ISSN 0129-0541.

Martisko:2022:CGS

[MMK22] Jakub Martiško, Alexander Meduna, and Zbyněk Křivka. CD grammar systems with two propagating scattered context components characterize the family of context sensitive languages. International Journal of Foundations of Computer Science (IJFCS), 33(3–4):335–348, April–June 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122410088.

Malcher:2010:SSB

[MMP10] Andreas Malcher, Carlo Mereghetti, and Beatrice Palano. Sublinearly space bounded iterative arrays. International Journal of Foundations of Computer Science (IJFCS), 21(5):843– 858, October 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mantaci:2010:BPD

[MMR10] Roberto Mantaci, Sabrina Mantaci, and Antonio Restivo. Balance properties and distribution of squares in circular words. International Journal of Foundations of Computer Science (IJFCS), 21(4):647– 664, August 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mahalingam:2020:WCJ

[MMR20] Kalpana Mahalingam, Ujjwal Kumar Mishra, and Rama Raghavan. Watson–Crick jumping finite automata. International Journal of Foundations of Computer Science (IJFCS), 31(07): 891–913, November 2020. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054120500331.

Merkle:2005:DPC

[MMS05] Daniel Merkle, Martin Middendorf, and Alexander Scheidler. Decentralized packet clustering in routerbased networks. International Journal of Foundations of Computer Science (IJFCS), 16(2):321– ??, April 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mazumder:2017:PSK

[MMS17] Rashed Mazumder, Atsuko Miyaji, and Chunhua Su. Probably secure keyedfunction based authenticated encryption schemes for big data. International Journal of Foundations of Computer Science (IJFCS), 28(6):661–??, September 2017. CODEN IFCSEN. ISSN 0129-0541.

Marchal:2023:MCT

[MMSV23] Loris Marchal, Samuel Mc-Cauley, Bertrand Simon, and Frédéric Vivien. Minimizing I/Os in out-of-core task tree scheduling. International Journal of Foundations of Computer Science (IJFCS), 34(01):51– 80, January 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500186.

Manea:2010:SRH

[MMY10] Florin Manea, Victor Mitrana, and Takashi Yokomori. Some remarks on the hairpin completion. International Journal of Foundations of Computer Science (IJFCS), 21(5):859– 872, October 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Muller:2000:TIF

[MN00] M. Müller and S. Nishimura. Type inference for first-class messages with feature constraints. International Journal of Foundations of Computer Science (IJFCS), 11 (1):29–64, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Miura:2006:CGD

[MNN06] Kazuyuki Miura, Shin-Ichi Nakano, and Takao Nishizeki. Convex grid drawings of four-connected plane graphs. International Journal of Foundations of Computer Science (IJFCS), 17(5):1031– ??, October 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Montanari:2012:P

[MNP12] Angelo Montanari, Margherita Napoli, and Mimmo Parente. Preface. International Journal of Foundations of Computer Science (IJFCS), 23(3):555–??, April 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mizuki:2011:ASN

[MNS11] Takaaki Mizuki, Satoru Nakayama, and Hideaki Sone. An application of ST-numbering to secret key agreement. International Journal of Foundations of Computer Science (IJFCS), 22(5):1211–1227, August 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Manea:2018:CSR

[MNS18] Florin Manea, Dirk Nowotka, and Markus L. Schmid. On the complexity of solving restricted word equations. International Journal of Foundations of Computer Science (IJFCS), 29(5):893– 909, August 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118420108.

Maelbrancke:1994:DTR

[MO94] R. Maelbrancke and H. Olivie.

Dynamic tree rebalancing using recurrent rotations. International Journal of Foundations of Computer Science (IJFCS), 5(3/4):??, 1994. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Messerschmidt:2007:CDS

[MO07] Hartmut Messerschmidt and Friedrich Otto. Cooperating distributed systems of restarting automata. International Journal of Foundations of Computer Science (IJFCS), 18(6):1333–1342, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Messerschmidt:2009:DCS

[MO09] Hartmut Messerschmidt and Friedrich Otto. On deterministic CD-systems of restarting automata. International Journal of Foundations of Computer Science (IJFCS), 20(1):185– 209, February 2009. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Moriya:2010:APS

[MO10] Etsuro Moriya and Friedrich Otto. On alternating phrase-structure grammars. International Journal of Foundations of Computer Science (IJFCS), 21(1):1– 25, February 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mrykhin:2023:HLK

[MO23] Mikhail Mrykhin and Alexander Okhotin. The hardest LL(k) language. International Journal of Foundations of Computer Science (IJFCS), 34(2–3):289–319, February–April 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412344001X.

Modanese:2021:STL

[Mod21] Augusto Modanese. Sublineartime language recognition and decision by onedimensional cellular automata. International Journal of Foundations of Computer Science (IJFCS), 32(06):713–731, September 2021. ISSN 0129-

0541. URL https://www.
worldscientific.com/doi/
10.1142/S0129054121420053.

Mohri:2002:GRI

[Moh02]  M. Mohri.  Generic e-

removal and input e-normalization algorithms for weighted transducers. International Journal of Foundations of Computer Science (IJFCS), 13(1):129–??, 2002. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mohri:2003:EDW

[Moh03]  Mehryar Mohri.  Edit-

distance of weighted automata: General definitions and algorithms. International Journal of Foundations of Computer Science (IJFCS), 14(6):957–??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mohri:2013:DFA

[Moh13] Mehryar Mohri. On the disambiguation of finite automata and functional transducers. International Journal of Foundations of Computer Science (IJFCS), 24(6):847–??, September 2013. CODEN IFCSEN. ISSN 0129-0541.

Marti-Oliet:1991:PNL

[MOM91] N. Marti-Oliet and J. Meseguer. From Petri nets to linear logic through categories: a survey. International Journal of Foundations of Computer Science (IJFCS), 2(4): 297–400, December 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Morin:2010:USM

[Mor10] Rémi Morin. Unambiguous shared-memory systems. International Journal of Foundations of Computer Science (IJFCS), 21(4):665– 685, August 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mignot:2018:EAC

[MOSZ18] Ludovic Mignot, Nadia Ouali-Sebti, and Djelloul Ziadi. An efficient algorithm for the construction of the equation tree automaton. International Journal of Foundations of Computer Science (IJFCS), 29(6):

951–978, September 2018. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118500156.

Monti:1991:STB

[MP91] A. Monti and D. Parente.

Systolic tree with base automata. International Journal of Foundations of Computer Science (IJFCS), 2(3): 221–236, September 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Moffat:1993:HS

[MP93] A. Moffat and O. Petersson.

Historical searching. International Journal of Foundations of Computer Science (IJFCS), 4(1):85–98, March 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Magalini:2007:PCU

[MP07] Enzo Magalini and Giovanni Pighizzini. A pumping condition for ultralinear languages. International Journal of Foundations of Computer Science (IJFCS), 18(6):1303–1312, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mcquillan:2012:P

[MP12] Ian Mcquillan and Giovanni Pighizzini. Preface. International Journal of Foundations of Computer Science (IJFCS), 23(1):1–3, January 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mahalingam:2022:HPT

[MP22] Kalpana Mahalingam and Palak Pandoh. HVpalindromes in two-dimensional words. International Journal of Foundations of Computer Science (IJFCS), 33 (05):389–409, August 2022. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412250006X.

Mraz:2007:ARA

[MPJ07] František Mráz, Martin Plátek, and Tomasz Jurdziński. Ambiguity by restarting automata. International Journal of Foundations of Computer Science (IJFCS), 18(6):1343–1352, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Martins:1999:DAR

[MPS99] E. D. Q. V. Martins, M. M. B. Pascoal, and J. L. E. D. Santos. Deviation algorithms for ranking shortest paths. International Journal of Foundations of Computer Science (IJFCS), 10(3):247–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mkrtchyan:2024:AAP

[MPS24] Vahan Mkrtchyan, Ojas Parekh, , and K. Subramani. Approximation algorithms for partial vertex covers in trees. International Journal of Foundations of Computer Science (IJFCS), 35(04): 387–407, June 2024. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123500089.

Margolis:2004:WGM

[MPV04] S. W. Margolis, J.-E. Pin, and M. V. Volkov. Words guaranteeing minimum image. International Journal of Foundations of Computer Science (IJFCS), 15(2):259– ??, April 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Maletti:2011:OHM

[MQ11] Andreas Maletti and Daniel Quernheim. Optimal hyperminimization. International Journal of Foundations of Computer Science (IJFCS), 22(8):1877–1891, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Maletti:2012:UWH

[MQ12] Andreas Maletti and Daniel Quernheim. Unweighted and weighted hyper-minimization. International Journal of Foundations of Computer Science (IJFCS), 23(6): 1207–??, September 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Metivier:1991:SOF

[MR91] Yves Metivier and Brigitte Rozoy. On the star operation in free partially commutative monoids. International Journal of Foundations of Computer Science (IJFCS), 2(3):257–266, September 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Makowsky:1999:CWG

[MR99] J. A. Makowsky and U. Rotics. On the clique-width of graphs with few P_{4}’s. International Journal of Foundations of Computer Science (IJFCS), 10(3):329– 348, 1999. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Manuel:2011:CCA

[MR11] Amaldev Manuel and R. Ramanujam. Class counting automata on datawords. International Journal of Foundations of Computer Science (IJFCS), 22(4):863– 882, June 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Moreira:2013:P

[MR13] Nelma Moreira and Rogério Reis. Preface. International Journal of Foundations of Computer Science (IJFCS), 24(6):689–??, September 2013. CODEN IFCSEN. ISSN 0129-0541.

Moreira:2023:SII

[MR23] Nelma Moreira and Rogéxrio Reis. Special issue: 25th international conference on developments in language theory (DLT 2021) — preface. International Journal of Foundations of Computer Science (IJFCS), 34(2–3): 81–83, February–April 2023. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412302001X.

Mahajan:2006:ABS

[MRRV06] Meena Mahajan, Raghavan Rama, Venkatesh Raman, and S. Vijaykumar. Approximate block sorting. International Journal of Foundations of Computer Science (IJFCS), 17(2):337– ??, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mateescu:1997:GTL

[MRS97] Alexandru Mateescu, Grzegorz Rozenberg, and Arto Salomaa. Geometric transformations of language families: The power of symmetry. International Journal of Foundations of Computer Science (IJFCS), 8(1): 1–??, March 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mol:2019:CTA

[MRSS19] Lucas Mol, Narad Rampersad, Jeffrey Shallit, and Manon Stipulanti. Cobham’s theorem and automaticity. International Journal of Foundations of Computer Science (IJFCS), 30(8):1363– 1379, December 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119500308.

Monserrat:1995:WCM

[MRT95] M. Monserrat, F. Rossello, and J. Torrens. When is a category of many-sorted partial algebras Cartesianclosed? International Journal of Foundations of Computer Science (IJFCS), 6 (1):51–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mongelli:1999:PRM

[MS99a] H. Mongelli and S. W. Song. Parallel range minima on coarse grained multicomputers. International Journal of Foundations of Computer Science (IJFCS), 10 (4):375–390, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mongelli:1999:PI

[MS99b] H. Mongelli and S. W.

Song. Part 1 (Irregular ’99). International Journal of Foundations of Computer Science (IJFCS), 10 (4):375–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mateescu:2004:MIS

[MS04] Alexandru Mateescu and Arto Salomaa. Matrix indicators for subword occurrences and ambiguity. International Journal of Foundations of Computer Science (IJFCS), 15(2):277– ??, April 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Malcher:2007:MPC

[MS07] Andreas Malcher and Bettina Sunckel. On metalinear parallel communicating grammar systems. International Journal of Foundations of Computer Science (IJFCS), 18(6):1313–1322, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Mahalingam:2012:PPM

[MS12] Kalpana Mahalingam and K. G. Subramanian. Product of Parikh matrices and commutativity. International Journal of Foundations of Computer Science (IJFCS), 23(1):207– 223, January 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Meduna:2016:C

[MS16a] Alexander Meduna and Ondřej Soukup. Corrigendum: “Simple Matrix Grammars and Their Leftmost Variants [3]”. International Journal of Foundations of Computer Science (IJFCS), 27(5):651–??, August 2016. CODEN IFC-SEN. ISSN 0129-0541.

Meduna:2016:SMG

[MS16b] Alexander Meduna and Ondřej Soukup. Simple matrix grammars and their leftmost variants. International Journal of Foundations of Computer Science (IJFCS), 27(3):359–??, April 2016. CODEN IFCSEN. ISSN 0129-0541.

Michalewski:2018:SUT

[MS18] Henryk Michalewski and Michał Skrzypczak. On the strength of unambiguous tree automata. International Journal of Foundations of Computer Science (IJFCS), 29(5):911– 933, August 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905411842011X.

Mohanty:2019:IOE

[MS19] Sraban Kumar Mohanty and G. Sajith. An input/output efficient algorithm for Hessenberg reduction. International Journal of Foundations of Computer Science (IJFCS), 30(8):1279–1300, December 2019. ISSN 0129-

0541. URL https://www.
worldscientific.com/doi/
10.1142/S0129054119500266.

Madejski:2020:MPT

[MS20] Grzegorz Madejski and Andrzej Szepietowski. Membership problem for twodimensional general row jumping finite automata. International Journal of Foundations of Computer Science (IJFCS), 31(04): 527–538, June 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500239.

Mukhtar:2022:MPN

[MSMR22] Salman Mukhtar, Muhammad Salman, Ayse Dilek Maden, and Masood Ur Rehman. Metric properties of non-commuting graph associated to two groups. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):931–952, September– November 2022. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122420229.

Malik:2006:CFT

[MSR06] Robi Malik, David Streader, and Steve Reeves. Conflicts and fair testing. International Journal of Foundations of Computer Science (IJFCS), 17(4):797– ??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Maneth:2023:DRT

[MSV23] Sebastian Maneth, Helmut Seidl, and Martin Vu. Definability results for top-down tree transducers. International Journal of Foundations of Computer Science (IJFCS), 34(2–3):253–287, February–April 2023. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122440051.

Mason:1995:RAO

[MT95a] I. Mason and C. Talcott. Reasoning about object systems in VTLoE. International Journal of Foundations of Computer Science (IJFCS), 6(3):265–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Musikaev:1995:FBP

[MT95b] I. K. Musikaev and M. A.

Taitslin. Flat backtracking Prolog for databases: a formal semantics, the computational complexity and the expressibility. International Journal of Foundations of Computer Science (IJFCS), 6(1):11–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Maletti:2010:PQR

[MT10] Andreas Maletti and Cătălin Ionuţ Tîrnăucă. Properties of quasi-relabeling tree bimorphisms. International Journal of Foundations of Computer Science (IJFCS), 21 (3):257–276, June 2010. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Miura:1999:LTA

[MTNN99] K. Miura, D. Takahashi, S. I. Nakano, and T. Nishizeki. A linear-time algorithm to find four independent spanning trees in four connected planar graphs. International Journal of Foundations of Computer Science (IJFCS), 10(2):195–??, 1999. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Meier:2009:CSF

[MTVM09] Arne Meier, Michael Thomas, Heribert Vollmer, and Martin Mundhenk. The complexity of satisfiability for fragments of CT L and CT L∗. International Journal of Foundations of Computer Science (IJFCS), 20 (5):901–918, October 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Meier:2015:ECS

[MTVM15] Arne Meier, Michael Thomas, Heribert Vollmer, and Martin Mundhenk. Erratum: The Complexity of Satisfiability for Fragments of CTL and CTL^{⋆}. International Journal of Foundations of Computer Science (IJFCS), 26(8):1189–??, December 2015. CODEN IFC-SEN. ISSN 0129-0541.

Mukund:1992:PNS

[Muk92] M. Mukund. Petri nets and step transition systems. International Journal of Foundations of Computer Science (IJFCS), 3(4):443–478, December 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Miyazawa:2011:BCT

[MV11] Flávio K. Miyazawa and André L. Vignatti. Bounds on the convergence time of distributed selfish bin packing. International Journal of Foundations of Computer Science (IJFCS), 22(3):565– 582, April 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Martin-Vide:2007:DPP

[MVM07] Carlos Martin-Vide and Victor Mitrana. Decision problems on path-controlled grammars. International Journal of Foundations of Computer Science (IJFCS), 18(6):1323–1332, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Martin-Vide:2002:PFA

[MVMM02] Carlos Martín-Vide, Alexandru Mateescu, and Victor Mitrana. Parallel finite automata systems communicating by states. International Journal of Foundations of Computer Science (IJFCS), 13(5):733– ??, October 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Marais:2022:DCN

[MvZ22] Laurette Marais and Lynette van Zijl. Descriptional complexity of non-unary selfverifying symmetric difference automata. International Journal of Foundations of Computer Science (IJFCS), 33(3–4):313–333, April–June 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122410076.

Muscholl:2005:NCF

[MW05] Anca Muscholl and Igor Walukiewicz. An NPcomplete fragment of LTL. International Journal of Foundations of Computer Science (IJFCS), 16(4):743– ??, August 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Maza:2011:BDP

[MX11] Marc Moreno Maza and Yuzhen Xie. Balanced dense polynomial multiplication on multi-cores. International Journal of Foundations of Computer Science (IJFCS), 22(5):1035– 1055, August 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ma:2004:TSP

[MXY^{+}04] Weimin Ma, Yinfeng Xu, Jane You, James Liu, and Kanliang Wang. On the k-truck scheduling problem. International Journal of Foundations of Computer Science (IJFCS), 15(1):127– ??, February 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

MeyerAufmHofe:2001:SRT

[MZ01] Harald Meyer Auf’m Hofe and Albert Zomaya. Solving rostering tasks by generic methods for constraint optimization. International Journal of Foundations of Computer Science (IJFCS), 12(5):671– 693, October 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Meduna:2012:JFA

[MZ12] Alexander Meduna and Petr Zemek. Jumping finite automata. International Journal of Foundations of Computer Science (IJFCS), 23(7):1555–??, November 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nagamochi:2006:PSR

[Nag06] Hiroshi Nagamochi. Packing soft rectangles. International Journal of Foundations of Computer Science (IJFCS), 17(5):1165– ??, October 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nagy:2020:MPP

[Nag20] Benedek Nagy. On the membership problem of permutation grammars — a direct proof of NP-completeness. International Journal of Foundations of Computer Science (IJFCS), 31(04): 515–525, June 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120500227.

Nagy:2021:UFR

[Nag21]  Benedek Nagy.  Union-

freeness revisited — between deterministic and nondeterministic union-free languages. International Journal of Foundations of Computer Science (IJFCS), 32(05):551–573, August 2021. ISSN 0129-0541. URL

                                        https://www.worldscientific.
                                        com/doi/10.1142/S0129054121410070.

Nakano:2003:LLG

[Nak03] Koji Nakano. Linear layout of generalized hypercubes. International Journal of Foundations of Computer Science (IJFCS), 14 (1):137–??, February 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nakano:2004:TEO

[Nak04] Koji Nakano. Time and energy optimal list ranking algorithms on the kchannel broadcast communication model with no collision detection. International Journal of Foundations of Computer Science (IJFCS), 15(1):73–??, February 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nuida:2015:MPS

[NAK^{+}15] Koji Nuida, Takuro Abe, Shizuo Kaji, Toshiaki Maeno, and Yasuhide Numata. A mathematical problem for security analysis of hash functions and pseudorandom generators. International Journal of Foundations of Computer Science (IJFCS), 26(2):169–??, February 2015. CODEN IFCSEN. ISSN 0129-0541.

Nakanishi:2018:QPA

[Nak18] Masaki Nakanishi. Quantum pushdown automata with garbage tape. International Journal of Foundations of Computer Science (IJFCS), 29(3):425– 446, April 2018. CODEN IFCSEN. ISSN 0129-0541.

Nadeem:2022:LNB

[NAS22] Muhammad Faisal Nadeem, Waqar Ali, and Hafiz Muhammad Afzal Siddiqui. Locating number of biswapped networks. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):667–690, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420096.

Nakano:2006:P

[NB06] Koji Nakano and Jacir L.

Bordim. Preface. International Journal of Foundations of Computer Science (IJFCS), 17(2):249– ??, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nouri-Baygi:2018:CCA

[NB18] Mostafa Nouri-Baygi. The computational complexity of and approximation algorithms for variants of the component selection problem. International Journal of Foundations of Computer Science (IJFCS), 29(7):1231– 1245, November 2018. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054118500314.

Nicart:2007:LMT

[NCC^{+}07] Florent Nicart, Jean-Marc Champarnaud, Tibor Csáki, Tamás Gaál, and André Kempe. Labelling multitape automata with constrained symbol classes. International Journal of Foundations of Computer Science (IJFCS), 18(4):847– 858, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nedjah:2002:PMC

[ND02] Nadia Nedjah and Luiza De Macedo Mourelle. Pattern matching code minimization in rewriting-based programming languages. International Journal of Foundations of Computer Science (IJFCS), 13(6):873–??, December 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Neggazi:2015:ESS

[NGHK15] Brahim Neggazi, Nabil Guellati, Mohammed Haddad, and Hamamache Kheddouci. Efficient selfstabilizing algorithm for independent strong dominating sets in arbitrary graphs. International Journal of Foundations of Computer Science (IJFCS), 26(6):751– ??, September 2015. CO-DEN IFCSEN. ISSN 0129-0541.

Ng:2002:IAW

[NH02] V. Ng and M. K. Ho. An intelligent agent for Web advertisements. International Journal of Foundations of Computer Science (IJFCS), 13(4):531–??, 2002. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nishimura:2003:QCR

[Nis03] Harumichi Nishimura. Quantum computation with restricted amplitudes. International Journal of Foundations of Computer Science (IJFCS), 14(5):853– ??, October 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nishida:2007:MAB

[Nis07] Taishin Y. Nishida. Membrane algorithm with Brownian subalgorithm and genetic subalgorithm. International Journal of Foundations of Computer Science (IJFCS), 18(6):1353–1360, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nakanishi:2022:EAC

[NKP^{+}22] Masaki Nakanishi, Kamil Khadiev, Krisjanis Prusis, Jevgenijs Vihrovs, and Abuzer Yakaryılmaz. Exact affine counter automata. International Journal of Foundations of Computer Science (IJFCS), 33(3–4):349– 370, April–June 2022. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S012905412241009X.

Ngassam:2008:IPT

[NKW08] Ernest Ketcha Ngassam, Derrick G. Kourie, and Bruce W. Watson. On implementation and performance of table-driven DFAbased string processors. International Journal of Foundations of Computer Science (IJFCS), 19(1):53–70, February 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nakano:1993:SFT

[NN93] S.-I. Nakano and T. Nishizeki.

Scheduling file transfers under port and channel constraints. International Journal of Foundations of Computer Science (IJFCS), 4(2): 101–116, June 1993. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nakano:1999:GEI

[NO99] K. Nakano and S. Olariu.

Guest Editors’ introduction. International Journal of Foundations of Computer Science (IJFCS), 10 (2):123–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nochefranca:1998:DHC

[Noc98] Luz R. Nochefranca. The diameter and Hamiltonian cycle of the generalized de Bruijn graphs UG_{b}(n, n(n + 1)). International Journal of Foundations of Computer Science (IJFCS), 9(1):13– ??, March 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nicolas:2009:URM

[NP09] Francois Nicolas and Yuri Pritykin. On uniformly recurrent morphic sequences. International Journal of Foundations of Computer Science (IJFCS), 20(5):919– 940, October 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Navarro:2011:SQ

[NPPS11] Gonzalo Navarro, Rodrigo Paredes, Patricio V. Poblete, and Peter Sanders. Stronger quickheaps. International Journal of Foundations of Computer Science (IJFCS), 22(4):945–969, June 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nikoletseas:2000:CPR

[NPSY00] S. Nikoletseas, K. Palem, P. Spirakis, and M. Yung. Connectivity properties in random regular graphs with edge faults. International Journal of Foundations of Computer Science (IJFCS), 11(2):247–262, 2000. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nahimovs:2018:QWT

[NR18] Nikolajs Nahimovs and Alexander Rivosh. Quantum walks on two-dimensional grids with multiple marked locations. International Journal of Foundations of Computer Science (IJFCS), 29(4):687–700, June 2018. CODEN IFCSEN. ISSN 0129-0541.

Nicaud:2021:RRE

[NR21] Cyril Nicaud and Pablo Rotondo. Random regular expression over huge alphabets. International Journal of Foundations of Computer Science (IJFCS), 32(05): 419–438, August 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412141001X.

Ng:2018:SCN

[NRS18] Timothy Ng, David Rappaport, and Kai Salomaa. State complexity of neighbourhoods and approximate pattern matching. International Journal of Foundations of Computer Science (IJFCS), 29(2):315–??, February 2018. CODEN IFCSEN. ISSN 0129-0541.

Ng:2019:SCS

[NRS19] Timothy Ng, David Rappaport, and Kai Salomaa. State complexity of suffix distance. International Journal of Foundations of Computer Science (IJFCS), 30(6– 7):1197–1216, September– November 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400355.

Nishimura:2000:FSS

[NRT00] N. Nishimura, P. Ragde, and D. M. Thilikos. Finding smallest supertrees under minor containment. International Journal of Foundations of Computer Science (IJFCS), 11(3):445– 466, 2000. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nochefrance:1998:DHC

[NS98] L. R. Nochefrance and P. W. Sy. The diameter and Hamiltonian cycle of the generalized de Bruijn graphs UGB(n, n(n + 1)). International Journal of Foundations of Computer Science (IJFCS), 9(1):13–??, 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Narayanaswamy:2013:UFB

[NS13] N. S. Narayanaswamy and N. Sadagopan. A unified framework for bi(tri)connectivity and chordal augmentation. International Journal of Foundations of Computer Science (IJFCS), 24(1):67– ??, January 2013. CODEN IFCSEN. ISSN 0129-0541.

Nowotka:2018:OVW

[NS18] Dirk Nowotka and Aleksi Saarela. One-variable word equations and three-variable constant-free word equations. International Journal of Foundations of Computer Science (IJFCS), 29(5):935– 950, August 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118420121.

Niu:2012:TPS

[NSVA12] Yunyun Niu, K. G. Subramanian, Ibrahim Venkat, and Rosni Abdullah. A tissue P system based solution to quadratic assignment problem. International Journal of Foundations of Computer Science (IJFCS), 23(7):1511–??, November 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nakata:2006:TFE

[NTSH06] Akio Nakata, Tadaaki Tanimoto, Suguru Sasaki, and Teruo Higashino. A timed failure equivalence preserving abstraction for parametric time-interval automata. International Journal of Foundations of Computer Science (IJFCS), 17(4):833– ??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nakano:2003:P

[NW03] Koji Nakano and Jie Wu.

Preface. International Journal of Foundations of Computer Science (IJFCS), 14 (1):1–??, February 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Nakano:2004:P

[NW04] Koji Nakano and Jie Wu.

Preface. International Journal of Foundations of Computer Science (IJFCS), 15 (3):459–??, June 2004. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ni:2022:HEB

[NWHL22] Baixiu Ni, Ying Wang, Jingfu Huang, and Guocheng Li. Hybrid enhanced binary honey badger algorithm with quadratic programming for cardinality constrained portfolio optimization. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):787–803, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420151.

Ngassam:2005:FDI

[NWK05] Ernest Ketcha Ngassam, Bruce W. Watson, and Derrick G. Kourie. A framework for the dynamic implementation of finite automata for performance enhancement. International Journal of Foundations of Computer Science (IJFCS), 16(6):1193–??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ngassam:2006:DAF

[NWK06] Ernest Ketcha Ngassam, Bruce W. Watson, and Derrick G. Kourie. Dynamic allocation of finite automata states for fast string recognition. International Journal of Foundations of Computer Science (IJFCS), 17(6):1307–1323, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Namjoshi:2010:P

[NY10] Kedar Namjoshi and Tomohiro Yoneda. Preface. International Journal of Foundations of Computer Science (IJFCS), 21(2):113– 114, April 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ni:2022:RPC

[NZH22] Zeqiang Ni, Lei Zhang, and Junqing He. Recomputation of public capital based on PIM and the effect on China regional economic growth. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):735–753, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420126.

Niu:2024:WDF

[NZT^{+}24] Baohua Niu, Shuming Zhou, Tao Tian, , and Qifan Zhang. The wide diameter and fault diameter of exchanged crossed cube. International Journal of Foundations of Computer Science (IJFCS), 35(04):435– 451, June 2024. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123500107.

Niu:2024:INI

[NZZ24] Baohua Niu, Shuming Zhou, , and Hong Zhang. Influential node identification of network based on agglomeration operation. International Journal of Foundations of Computer Science (IJFCS), 35(03):271– 295, April 2024. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123500028.

Obtulowicz:2001:MCO

[Obt01] Adam Obtulowicz. Membrane computing and oneway functions. International Journal of Foundations of Computer Science (IJFCS), 12(4):551–558, 2001. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Obtulowicz:2006:GPM

[Obt06] Adam Obtułowicz. Gandy’s principles for mechanisms and membrane computing. International Journal of Foundations of Computer Science (IJFCS), 17(1):167– ??, February 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ogata:2000:CCL

[Oga00] I. Ogata. Constructive classical logic as CPS-calculus. International Journal of Foundations of Computer Science (IJFCS), 11(1):89– 112, 2000. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ogihara:1994:SL

[Ogi94] M. Ogihara. On serializable languages. International Journal of Foundations of Computer Science (IJFCS), 5(3/4):??, 1994. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Okawa:1998:PGN

[Oka98] Satoshi Okawa. The permutational graph: a new network topology. International Journal of Foundations of Computer Science (IJFCS), 9(1):3–??, March 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Okadome:1999:SFL

[Oka99] T. Okadome. Simple flat languages: a learnable class in the limit from positive data. International Journal of Foundations of Computer Science (IJFCS), 10 (4):483–502, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Okadome:2000:SSC

[Oka00] T. Okadome. Some sufficient conditions of learnability in the limit from positive data. International Journal of Foundations of Computer Science (IJFCS), 11 (3):515–524, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Okhotin:2003:EAB

[Okh03] Alexander Okhotin. Efficient automaton-based recognition for linear conjunctive languages. International Journal of Foundations of Computer Science (IJFCS), 14(6):1103– ??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Okhotin:2005:CAH

[Okh05] Alexander Okhotin. A characterization of the arithmetical hierarchy by language equations. International Journal of Foundations of Computer Science (IJFCS), 16(5):985– ??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Okhotin:2006:GLP

[Okh06] Alexander Okhotin. Generalized LR parsing algorithm for Boolean grammars. International Journal of Foundations of Computer Science (IJFCS), 17(3):629– ??, June 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Okhotin:2007:NDC

[Okh07] Alexander Okhotin. Notes on dual concatenation. International Journal of Foundations of Computer Science (IJFCS), 18(6):1361– 1370, December 2007. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Oles:1992:WCM

[Ole92] F. J. Oles. When is a category of many-sorted algebras Cartesian closed? International Journal of Foundations of Computer Science (IJFCS), 3(2):225–??, June 1992. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Oliveira:2013:WAC

[Oli13] José N. Oliveira. Weighted automata as coalgebras in categories of matrices. International Journal of Foundations of Computer Science (IJFCS), 24(6):709–??, September 2013. CODEN IFCSEN. ISSN 0129-0541.

Orlandic:1996:SOT

[OM96] R. Orlandic and H. M. Mahmoud. Storage overhead of O-trees, B-trees and prefix B-trees: a comparative analysis. International Journal of Foundations of Computer Science (IJFCS), 7 (3):209–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

ONeil:2015:CCS

[O’N15] Thomas E. O’Neil. Complement, complexity, and symmetric representation. International Journal of Foundations of Computer Science (IJFCS), 26(5):557–??, August 2015. CODEN IFC-SEN. ISSN 0129-0541.

Ochem:2008:AAS

[ORS08] Pascal Ochem, Narad Rampersad, and Jeffrey Shallit. Avoiding approximate squares. International Journal of Foundations of Computer Science (IJFCS), 19 (3):633–648, June 2008. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Olariu:1993:NCU

[OS93] S. Olariu and I. A. Stewart. A new characterization of unbreakable graphs. International Journal of Foundations of Computer Science (IJFCS), 4(3):193–196, September 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Oida:2001:CDO

[OS01] K. Oida and K. Shinjo.

Characteristics of deterministic optimal routing for two heterogeneous parallel servers. International Journal of Foundations of Computer Science (IJFCS), 12(6):775–790, December 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Okhotin:2019:SCQ

[OS19] Alexander Okhotin and Kai Salomaa. State complexity of the quotient operation on inputdriven pushdown automata. International Journal of Foundations of Computer Science (IJFCS), 30(6– 7):1217–1235, September– November 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400367.

Olariu:1992:OPE

[OSZ92] S. Olariu, J. L. Schwing, and J. Zhang. Optimal parallel encoding and decoding algorithms for trees. International Journal of Foundations of Computer Science (IJFCS), 3(1):1–10, March 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Otto:2013:CPC

[Ott13] Friedrich Otto. On centralized parallel communicating grammar systems with context-sensitive components. International Journal of Foundations of Computer Science (IJFCS), 24 (4):453–??, June 2013. CO-DEN IFCSEN. ISSN 0129-0541.

Otto:2015:APC

[Ott15]  Friedrich Otto.  Asyn-

chronous parallel communicating systems of pushdown automata. International Journal of Foundations of Computer Science (IJFCS), 26(5):643–??, August 2015. CODEN IFC-SEN. ISSN 0129-0541.

Ottmann:1992:UBT

[OW92] T. Ottmann and D. Wood. Updating binary trees with constant linkage cost. International Journal of Foundations of Computer Science (IJFCS), 3(4):479–502, December 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Okubo:2011:MCL

[OY11] Fumiya Okubo and Takashi Yokomori. Morphic characterizations of language families in terms of insertion systems and star languages. International Journal of Foundations of Computer Science (IJFCS), 22(1):247– 260, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Pavel:1998:ISR

[PA98] Sandy Pavel and Selim G.

Akl. Integer sorting and routing in arrays with reconfigurable optical buses. International Journal of Foundations of Computer Science (IJFCS), 9(1):99–??, March 1998. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Palis:2001:PSP

[Pal01a] M. A. Palis. Part 1 (selected papers from ISPAN ’99). International Journal of Foundations of Computer Science (IJFCS), 12 (3):245–??, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Palis:2001:SIP

[Pal01b] Michael A. Palis. Special issue on parallel and distributed computing. International Journal of Foundations of Computer Science (IJFCS), 12(3):245– 247, 2001. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Palis:2003:COR

[Pal03] Michael A. Palis. On the competitiveness of online real-time scheduling with rate of progress guarantees. International Journal of Foundations of Computer Science (IJFCS), 14(3):359– ??, June 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Palano:2008:RCC

[Pal08] Beatrice Palano. A regularity condition for context-free grammars. International Journal of Foundations of Computer Science (IJFCS), 19(4):845–857, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Panti:1991:SNT

[Pan91] G. Panti. Solution of a number theoretic problem involving knowledge. International Journal of Foundations of Computer Science (IJFCS), 2(4):419–424, December 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Paraponiari:2023:FPC

[Par23a] Paulina Paraponiari. Fuzzy propositional configuration logics. International Journal of Foundations of Computer Science (IJFCS), 34(06):603–631, September 2023. ISSN 0129-

0541. URL https://www.
worldscientific.com/doi/
10.1142/S012905412345003X.

Park:2023:PDP

[Par23b] Jung-Heum Park. Paired 3-disjoint path covers in bipartite torus-like graphs with edge faults. International Journal of Foundations of Computer Science (IJFCS), 34(04):429– 441, June 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500241.

Peterlongo:2008:IGF

[PAS08] Pierre Peterlongo, Julien Allali, and Marie-France Sagot. Indexing gappedfactors using a tree. International Journal of Foundations of Computer Science (IJFCS), 19(1):71–87, February 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Patrignani:2006:EPS

[Pat06] Maurizio Patrignani. On extending a partial straightline drawing. International Journal of Foundations of Computer Science (IJFCS), 17(5):1061– ??, October 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Paun:2000:CMS

[Pău00] G. Păun. Computing with membranes (P systems): a variant. International Journal of Foundations of Computer Science (IJFCS), 11 (1):167–182, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Paul:2024:EUS

[Pau24] Erik Paul. Equivalence, unambiguity, and sequentiality of finitely ambiguous max-plus tree automata. International Journal of Foundations of Computer Science (IJFCS), 35(1–2):23–49, January– February 2024. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123480027.

Paun:2020:USP

[PB20] Andrei Păun and Florin-Daniel Bîlbîe. Universality of SNQ P systems using one type of spikes and restrictive rule application. International Journal of Foundations of Computer Science (IJFCS), 31(1):117– 132, January 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120400080.

Pescini:2006:DPS

[PBMZ06] Dario Pescini, Daniela Besozzi, Giancarlo Mauri, and Claudio Zandron. Dynamical probabilistic P systems. International Journal of Foundations of Computer Science (IJFCS), 17(1):183– ??, February 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Pan:2011:CRN

[PDPPJ11] Linqiang Pan, Daniel Díaz-Pernil, and Mario J. Pérez-Jiménez. Computation of Ramsey numbers by P systems with active membranes. International Journal of Foundations of Computer Science (IJFCS), 22 (1):29–38, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Penczek:1993:TLT

[Pen93] W. Penczek. Temporal logics for trace systems: On automated verification. International Journal of Foundations of Computer Science (IJFCS), 4(1):31–68, March 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Petersen:2011:STB

[Pet11]  Holger Petersen.  Simu-

lations by time-bounded counter machines. International Journal of Foundations of Computer Science (IJFCS), 22(2):395– 409, February 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Priore:2001:DSM

[PFG^{+}01] Paolo Priore, D. D. L. Fuente, A. Gomez, et al. Dynamic scheduling of manufacturing systems with machine learning. International Journal of Foundations of Computer Science (IJFCS), 12(6):751–762, December 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Perez-Hurtado:2011:PAA

[PHPJRN^{+}11] Ignacio Pérez-Hurtado, Mario J. Pérez-Jiménez, Agustín Riscos-Núñez, Miguel A. Gutiérrez-Naranjo, and Miquel Rius-Font. On a partial affirmative answer for a Păun’s Conjecture. International Journal of Foundations of Computer Science (IJFCS), 22(1):55–64, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Peng:1995:NTP

[PI95] W. Peng and S. P. Iyer. A new type of pushdown automata on infinite trees. International Journal of Foundations of Computer Science (IJFCS), 6(2):169–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Pighizzini:2009:DPA

[Pig09] Giovanni Pighizzini. Deterministic pushdown automata and unary languages. International Journal of Foundations of Computer Science (IJFCS), 20 (4):629–645, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Pighizzini:2015:IAL

[Pig15] Giovanni Pighizzini. Investigations on automata and languages over a unary alphabet. International Journal of Foundations of Computer Science (IJFCS), 26 (7):827–??, November 2015. CODEN IFCSEN. ISSN 0129-0541.

Pin:2012:EDL

[Pin12] Jean-Éric Pin. Equational descriptions of languages. International Journal of Foundations of Computer Science (IJFCS), 23(6): 1227–??, September 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Prieur:2006:STS

[PL06] Élise Prieur and Thierry Lecroq. From suffix trees to suffix vectors. International Journal of Foundations of Computer Science (IJFCS), 17(6):1385–1402, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Pape-Lange:2023:TUB

[PL23] Julian Pape-Lange. Tight upper bounds on distinct maximal (sub-)Repetitions in highly compressible strings. International Journal of Foundations of Computer Science (IJFCS), 34(2–3): 321–345, February–April 2023. ISSN 0129-0541. URL

                                        https://www.worldscientific.
                                        com/doi/10.1142/S0129054122440075.

Plaza:1996:PSR

[Pla96] J. A. Plaza. On the propositional SLDNF-resolution. International Journal of Foundations of Computer Science (IJFCS), 7(4):359– ??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Porreca:2011:SAM

[PLMZ11] Antonio E. Porreca, Alberto Leporati, Giancarlo Mauri, and Claudio Zandron. P systems with active membranes working in polynomial space. International Journal of Foundations of Computer Science (IJFCS), 22(1):65–73, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Prrusa:2013:RTA

[PM13] Daniel Průša and František Mráz. Restarting tiling automata. International Journal of Foundations of Computer Science (IJFCS), 24 (6):863–??, September 2013. CODEN IFCSEN. ISSN 0129-0541.

Paun:2010:ICS

[PNN^{+}10] Mihaela Păun, Nichamon Naksinehaboon, Raja Nassar, Chokchai Leangsuksun, Stephen L. Scott, and Narate Taerat. Incremental checkpoint schemes for Weibull failure distribution. International Journal of Foundations of Computer Science (IJFCS), 21(3):329– 344, June 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Polak:2005:MNU

[Pol05] Libor Polák. Minimalizations of NFA using the universal automaton. International Journal of Foundations of Computer Science (IJFCS), 16(5):999– ??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Platek:2022:OWR

[POM22] Martin Plátek, Friedrich Otto, and František Mráz. One-way restarting automata and their sensitivities. International Journal of Foundations of Computer Science (IJFCS), 33(3–4):371–387, April– June 2022. ISSN 0129-

0541. URL https://www.

                                        worldscientific.com/doi/
                                        10.1142/S0129054122410106.

Poon:2004:ORM

[Poo04] Chung Keung Poon. Optimal range max datacube for fixed dimensions. International Journal of Foundations of Computer Science (IJFCS), 15(5):773– ??, October 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Paquette:2006:FBB

[PP06] Michel Paquette and Andrzej Pelc. Fast broadcasting with Byzantine faults. International Journal of Foundations of Computer Science (IJFCS), 17(6): 1423–1439, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Poinsot:2011:NBA

[PP11] Laurent Poinsot and Alexander Pott. Non-Boolean almost perfect nonlinear functions on non-Abelian groups. International Journal of Foundations of Computer Science (IJFCS), 22 (6):1351–1367, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Pighizzini:2014:LAR

[PP14] Giovanni Pighizzini and Andrea Pisoni. Limited automata and regular languages. International Journal of Foundations of Computer Science (IJFCS), 25 (7):897–??, November 2014. CODEN IFCSEN. ISSN 0129-0541.

Paun:2006:P

[PPJ06] Gheorghe Păun and Mario J.

Pérez-Jiménez. Preface. International Journal of Foundations of Computer Science (IJFCS), 17(1):1–??, February 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Paun:2006:STS

[PPJR06] Gheorghe Păun, Mario J.

Pérez-Jiménez, and Grzegorz Rozenberg. Spike trains in spiking neural P systems. International Journal of Foundations of Computer Science (IJFCS), 17(4):975– ??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Paun:2007:CMS

[PPJR07] Gheorghe Păun, Mario J.

Pérez-Jiménez, and Grzegorz Rozenberg. Computing morphisms by spiking neural P systems. International Journal of Foundations of Computer Science (IJFCS), 18(6):1371–1382, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Paun:2007:SNS

[PPJS07] Gheorghe Păun, Mario J. Pérez-Jiménez, and Arto Salomaa. Spiking neural P systems: an early survey. International Journal of Foundations of Computer Science (IJFCS), 18(3):435– 455, June 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Paun:2008:RCL

[PPJY08] Gheorghe Păun, Mario J. Pérez-Jiménez, and Takashi Yokomori. Representations and characterizations of languages in Chomsky hierarchy by means of insertiondeletion systems. International Journal of Foundations of Computer Science (IJFCS), 19(4):859– 871, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Paun:2002:CCN

[PPR02] Andrei Păun, Gheorghe Păun, and Grzegorz Rozenberg. Computing by communication in networks of membranes. International Journal of Foundations of Computer Science (IJFCS), 13(6):779–??, December 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Poyias:2018:MBP

[PPR18] Andreas Poyias, Simon J.

Puglisi, and Rajeev Raman. m-Bonsai: A practical compact dynamic trie. International Journal of Foundations of Computer Science (IJFCS), 29(8):1257– 1278, December 2018. ISSN 0129-0541.

Paun:2011:SPM

[PPRPS11] Andrei Păun, Mihaela Păun, Alfonso Rodríguez-Patón, and Manuela Sidoroff. P systems with proteins on membranes: a survey. International Journal of Foundations of Computer Science (IJFCS), 22(1):39– 53, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Peled:2006:ECT

[PQ06] Doron Peled and Hongyang Qu. Enforcing concurrent temporal behaviors. International Journal of Foundations of Computer Science (IJFCS), 17(4):743– ??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Perkovic:2000:IAF

[PR00] L. Perkovic and B. Reed. An improved algorithm for finding tree decompositions of small width. International Journal of Foundations of Computer Science (IJFCS), 11(3):365–372, 2000. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Pribavkina:2011:SCC

[PR11] Elena Pribavkina and Emanuele Rodaro. State complexity of code operators. International Journal of Foundations of Computer Science (IJFCS), 22(7):1669–1681, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Piatkowski:2012:ABM

[PR12] Marcin Piatkowski and Wojciech Rytter. Asymptotic behaviour of the maximal number of squares in standard Sturmian words. International Journal of Foundations of Computer Science (IJFCS), 23(2):303– 321, February 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Pittou:2023:MUA

[PR23] Maria Pittou and George Rahonis. Modelling uncertainty in architectures of parametric componentbased systems. International Journal of Foundations of Computer Science (IJFCS), 34(06):559– 601, September 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123450028.

Preparata:1990:PPL

[Pre90] Franco P. Preparata. Planar point location revisited. International Journal of Foundations of Computer Science (IJFCS), 1(1):71– ??, March 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Preface:2001:SIF

[Pre01] Preface. Special issue on functional and logic programming — part 1. International Journal of Foundations of Computer Science (IJFCS), 12(1):1–??, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Pribavkina:2006:SPL

[Pri06] Elena V. Pribavkina. On some properties of the language of 2-collapsing words. International Journal of Foundations of Computer Science (IJFCS), 17(3):665– ??, June 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Pathak:2013:EEA

[PRN13] Sudipta Pathak, Sanguthevar Rajasekaran, and Marius Nicolae. Ems1: an elegant algorithm for edit distance based Motif search. International Journal of Foundations of Computer Science (IJFCS), 24(4):473– ??, June 2013. CODEN IFCSEN. ISSN 0129-0541.

Prodinger:1996:DPL

[Pro96] H. Prodinger. Depth and path length of heap ordered trees. International Journal of Foundations of Computer Science (IJFCS), 7 (3):293–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Paik:1998:VSD

[PRS98] Doowon Paik, Sudhakar Reddy, and Sartaj Sahni. Vertex splitting in DAGs and applications to partial scan designs and lossy circuits. International Journal of Foundations of Computer Science (IJFCS), 9(4):377– ??, December 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Prusa:2017:CMS

[Prů17] Daniel Průša. Complexity of matching sets of two-dimensional patterns by two-dimensional on-line tessellation automaton. International Journal of Foundations of Computer Science (IJFCS), 28(5):623–??, August 2017. CODEN IFC-SEN. ISSN 0129-0541.

Paun:2001:HL

[PRY01] Gheorghe Păun, G. Rozenberg, and T. Yokomori. Hairpin languages. International Journal of Foundations of Computer Science (IJFCS), 12(6):837– 848, December 2001. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Pighizzini:2002:ULO

[PS02] G. Pighizzini and J. Shallit. Unary language operations, state complexity and Jacobsthal’s function. International Journal of Foundations of Computer Science (IJFCS), 13(1):145–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Pelantova:2012:ARW

[PS12a] Edita Pelantová and Štépán Starosta. Almost rich words as morphic images of rich words. International Journal of Foundations of Computer Science (IJFCS), 23 (5):1067–??, August 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Petrova:2012:CPB

[PS12b] Elena A. Petrova and Arseny M. Shur. Constructing premaximal binary cubefree words of any level. International Journal of Foundations of Computer Science (IJFCS), 23(8):1595– ??, December 2012. CODEN IFCSEN. ISSN 0129-0541.

Potapov:2018:P

[PS18] Igor Potapov and Pavel Semukhin. Preface. International Journal of Foundations of Computer Science (IJFCS), 29(2):139–??, February 2018. CODEN IFCSEN. ISSN 0129-0541.

Priya:2022:IQS

[PS22] R. Priya and N. Sivakumar. Improving the quality of service (QoS) and resource allocation in vehicular platoon using metaheuristic optimization algorithm. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):625–647, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420072.

Palioudakis:2017:WCB

[PSA17] Alexandros Palioudakis, Kai Salomaa, and Selim G. Akl. Worst case branching and other measures of nondeterminism. International Journal of Foundations of Computer Science (IJFCS), 28 (3):195–??, April 2017. CO-DEN IFCSEN. ISSN 0129-

Prusinkiewicz:2012:SGM

[PSS12] Przemysław Prusinkiewicz, Mitra Shirmohammadi, and Faramarz Samavati. Lsystems in geometric modeling. International Journal of Foundations of Computer Science (IJFCS), 23(1):133– 146, January 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Piperno:1990:RSE

[PT90] Adolfo Piperno and Enrico Tronci. Regular systems of equations in lambda-calculus. International Journal of Foundations of Computer Science (IJFCS), 1(3):325–??, September 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Peled:2007:P

[PT07] Doron A. Peled and Yih-Kuen Tsay. Preface. International Journal of Foundations of Computer Science (IJFCS), 18(1):1–3, February 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Pelc:2014:EGE

[PT14] Andrzej Pelc and Anas Tiane. Efficient Grid exploration with a stationary token. International Journal of Foundations of Computer Science (IJFCS), 25 (3):247–??, April 2014. CO-DEN IFCSEN. ISSN 0129-0541.

Poovanandran:2018:MES

[PT18] Ghajendran Poovanandran and Wen Chean Teh. On Mequivalence and strong Mequivalence for Parikh matrices. International Journal of Foundations of Computer Science (IJFCS), 29(1):123– ??, January 2018. CODEN IFCSEN. ISSN 0129-0541.

Poovanandran:2019:SST

[PT19] Ghajendran Poovanandran and Wen Chean Teh. Strong (2 · t) and strong (3 · t) transformations for strong M-equivalence. International Journal of Foundations of Computer Science (IJFCS), 30(5):719– 733, August 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119500187.

Petrini:1998:PAW

[PV98] Fabrizio Petrini and Marco Vanneschi. Performance analysis of wormhole routed K-ary N-trees. International Journal of Foundations of Computer Science (IJFCS), 9(2):157–??, June 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Potanin:2013:P

[PV13] Alex Potanin and Taso Viglas. Preface. International Journal of Foundations of Computer Science (IJFCS), 24(1):1–??, January 2013. CODEN IFCSEN. ISSN 0129-0541.

Poon:2004:MTC

[PY04] Chung Keung Poon and Wenci Yu. On minimizing total completion time in batch machine scheduling. International Journal of Foundations of Computer Science (IJFCS), 15(4):593– ??, August 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Pym:1992:UAL

[Pym92] D. Pym. A unification algorithm for the lambdaPicalculus. International Journal of Foundations of Computer Science (IJFCS), 3(3):333–378, September 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Peng:2010:AAS

[PYTH10] Yung-Hsing Peng, Chang-Biau Yang, Kuo-Tsung Tseng, and Kuo-Si Huang. An algorithm and applications to sequence alignment with weighted constraints. International Journal of Foundations of Computer Science (IJFCS), 21 (1):51–59, February 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Poon:2007:DBB

[PZX07] Chung Keung Poon, Feifeng Zheng, and Yinfeng Xu. Ondemand bounded broadcast scheduling with tight deadlines. International Journal of Foundations of Computer Science (IJFCS), 18(2):251– 262, April 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Qiu:2003:INT

[QD03] Ke Qiu and Sajal K. Das. Interconnection networks and their eigenvalues. International Journal of Foundations of Computer Science (IJFCS), 14(3):371– ??, June 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Qi:2015:LED

[QFL^{+}15] Xingqin Qi, Edgar Fuller, Rong Luo, Guodong Guo, and Cunquan Zhang. Laplacian energy of digraphs and a minimum Laplacian energy algorithm. International Journal of Foundations of Computer Science (IJFCS), 26(3):367–??, April 2015. CODEN IFC-SEN. ISSN 0129-0541.

Qi:2006:SSP

[QLWL06] Xingqin Qi, Guojun Li, Jichang Wu, and Bingqiang Liu. Sorting signed permutations by fixed-length reversals. International Journal of Foundations of Computer Science (IJFCS), 17 (4):933–??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Quaglia:2007:SDB

[Qua07] Francesco Quaglia. Software diversity-based active replication as an approach for enhancing the performance of advanced simulation systems. International Journal of Foundations of Computer Science (IJFCS), 18(3):495– 515, June 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rajabi-Alni:2015:CPS

[RAB15] Fatemeh Rajabi-Alni and Alireza Bagheri. Constrained point set embedding of a balanced binary tree. International Journal of Foundations of Computer Science (IJFCS), 26(2):195– ??, February 2015. CODEN IFCSEN. ISSN 0129-0541.

Rampersad:2005:WAP

[Ram05] Narad Rampersad. Words avoiding \frac{7}{3}-powers and the Thue–Morse morphism. International Journal of Foundations of Computer Science (IJFCS), 16(4):755– ??, August 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rao:2008:GCR

[Rao08] M. V. Panduranga Rao. Generalized counters and reversal complexity. International Journal of Foundations of Computer Science (IJFCS), 19(6):1403–1412, December 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ravikumar:2008:BND

[Rav08]  Bala Ravikumar.  The

Benford–Newcomb distribution and unambiguous context-free languages. International Journal of Foundations of Computer Science (IJFCS), 19(3):717– 727, June 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rispal:2005:CRS

[RC05] Chloé Rispal and Olivier Carton. Complementation of rational sets on countable scattered linear orderings. International Journal of Foundations of Computer Science (IJFCS), 16(4):767– ??, August 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rosenberg:2011:HCI

[RC11] Arnold L. Rosenberg and Ron C. Chiang. Heterogeneity in computing: Insights from a worksharing scheduling problem. International Journal of Foundations of Computer Science (IJFCS), 22(6):1471–1493, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Romero-Campero:2009:MAC

[RCTC^{+}09] Francisco J. Romero-Campero, Jamie Twycross, Miguel Cámara, Malcolm Bennett, Marian Gheorghe, and Natalio Krasnogor. Modular assembly of cell systems biology models using P systems. International Journal of Foundations of Computer Science (IJFCS), 20(3):427– 442, June 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Reinhardt:2007:THH

[Rei07] Klaus Reinhardt. A treeheight hierarchy of contextfree languages. International Journal of Foundations of Computer Science (IJFCS), 18(6):1383–1394, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rahul:2011:DSR

[RGR11] Saladi Rahul, Prosenjit Gupta, and K. S. Rajan. Data structures for range-aggregation over categories. International Journal of Foundations of Computer Science (IJFCS), 22 (7):1707–1728, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Raghavendra:2022:MSP

[RHN^{+}22] S. Raghavendra, A. Harshavardhan, S. Neelakandan, R. Partheepan, Ranjan Walia, and V. Chandra Shekhar Rao. Multilayer stacked probabilistic belief network-based brain tumor segmentation and classification. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):559–582, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420047.

Recalde:2010:CPN

[RHS10] Laura Recalde, Serge Haddad, and Manuel Silva. Continuous Petri nets: Expressive power and decidability issues. International Journal of Foundations of Computer Science (IJFCS), 21(2):235–256, April 2010.

CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Richomme:2019:CIL

[Ric19] Gwenaël Richomme. Characterization of infinite LSP words and endomorphisms preserving the LSP property. International Journal of Foundations of Computer Science (IJFCS), 30(1):171– 196, January 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400082.

Rivals:2004:SAA

[Riv04] Eric Rivals. A survey on algorithmic aspects of tandem repeats evolution. International Journal of Foundations of Computer Science (IJFCS), 15(2):225– ??, April 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rajasekaran:2009:SBT

[RK09] Sanguthevar Rajasekaran and Vamsi Kundeti. Spectrum based techniques for graph isomorphism. International Journal of Foundations of Computer Science (IJFCS), 20(3):479– 499, June 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Roychoudhury:2002:BTS

[RKRR02] A. Roychoudhury, K. N. Kumar, C. R. Ramakrishnan, and I. V. Ramakrishnan. Beyond Tamaki– Sato style unfold/fold transformations for normal logic programs. International Journal of Foundations of Computer Science (IJFCS), 13(3):387–??, 2002. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ripphausen-Lipa:1996:LTA

[RLWW96] H. Ripphausen-Lipa, D. Wagner, and K. Weihe. Lineartime algorithms for disjoint two-face paths problems in planar graphs. International Journal of Foundations of Computer Science (IJFCS), 7(2):95–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rajasekaran:1998:PRS

[RM98] Sanguthevar Rajasekaran and Theodore McKendall. Permutation routing and sorting on the reconfigurable mesh. International Journal of Foundations of Computer Science (IJFCS), 9(2): 199–??, June 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rezaeibagha:2019:PSB

[RMZW19] Fatemeh Rezaeibagha, Yi Mu, Shiwei Zhang, and Xiaofen Wang. Provably secure (broadcast) homomorphic signcryption. International Journal of Foundations of Computer Science (IJFCS), 30(4):511– 529, June 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400100.

Raju:2022:SEM

[RN22] Konduru Upendra Raju and Amutha Prabha Nagarajan. A steganography embedding method based on CDF-DWT technique for reversible data hiding application using Elgamal algorithm. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):489–512, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420011.

Rogojin:2009:SEG

[Rog09] Vladimir Rogojin. Successful elementary gene assembly strategies. International Journal of Foundations of Computer Science (IJFCS), 20(3):455–477, June 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Roth:2008:SRK

[ROK08] Camille Roth, Sergei Obiedkov, and Derrick G. Kourie. On succinct representation of knowledge community taxonomies with formal concept analysis. International Journal of Foundations of Computer Science (IJFCS), 19(2):383–404, April 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rosolini:1990:AMS

[Ros90] G. Rosolini. About modest sets. International Journal of Foundations of Computer Science (IJFCS), 1 (3):341–??, September 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rosenberg:2000:GDP

[Ros00] A. L. Rosenberg. Guidelines for data-parallel cyclestealing in networks of workstations II: On maximizing guaranteed output. International Journal of Foundations of Computer Science (IJFCS), 11(1):183– 204, 2000. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rosenberg:2003:EPF

[Ros03] Arnold L. Rosenberg. Efficient pairing functions — and why you should care.

International Journal of Foundations of Computer Science (IJFCS), 14(1):3– ??, February 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Roversi:2000:LAL

[Rov00] L. Roversi. Light affine logic as a programming language: a first contribution. International Journal of Foundations of Computer Science (IJFCS), 11 (1):113–152, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Roussel:1999:HDM

[RR99] Florian Roussel and Irena Rusu. Holes and dominoes in Meyniel graphs. International Journal of Foundations of Computer Science (IJFCS), 10(2):127–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rauber:2004:PBL

[RR04] Thomas Rauber and Gudula Rünger. Program-based locality measures for scientific computing. International Journal of Foundations of Computer Science (IJFCS), 15(3):535–??, June 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rauber:2006:DRD

[RR06] Thomas Rauber and Gudula Rünger. A data redistribution library for multi-processor task programming. International Journal of Foundations of Computer Science (IJFCS), 17(2):251–??, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Roayaei:2018:PCD

[RR18] Mehdy Roayaei and MohammadReza Razzazi. Parameterized complexity of directed Steiner network with respect to shared vertices and arcs. International Journal of Foundations of Computer Science (IJFCS), 29(7):1215– 1230, November 2018. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054118500302.

Roussel:1999:GLN

[RRT99] F. Roussel, I. Rusu, and H. Thuillier. On graphs with limited number of P_{4}-partners. International Journal of Foundations of Computer Science (IJFCS), 10(1):103–??, 1999. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Raja:1995:QBC

[RS95] N. Raja and R. K. Shyamasundar. The Quine-Bernays combinatory calculus. International Journal of Foundations of Computer Science (IJFCS), 6(4):417–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rajasekaran:2000:SIR

[RS00] S. Rajasekaran and S. K.

Sahni. Special issue on randomized computing. International Journal of Foundations of Computer Science (IJFCS), 11(2):205– 206, 2000. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Roberts:2001:RNC

[RS01] Alan Roberts and Antonios Symvonis. On the routing number of complete d-ary trees. International Journal of Foundations of Computer Science (IJFCS), 12 (4):411–434, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Richomme:2004:CRM

[RS04] Gwénaël Richomme and Patrice Séébold. Conjectures and results on morphisms generating k-power-free words. International Journal of Foundations of Computer Science (IJFCS), 15(2):307–??, April 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ravikumar:2007:ELD

[RS07] Bala Ravikumar and Nicolae Santean. On the existence of lookahead delegators for NFA. International Journal of Foundations of Computer Science (IJFCS), 18(5):949– 973, October 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Reghizzi:2012:RSL

[RS12] Stefano Crespi Reghizzi and Pierluigi San Pietro. From regular to strictly locally testable languages. International Journal of Foundations of Computer Science (IJFCS), 23(8):1711–??, December 2012. CODEN IFC-SEN. ISSN 0129-0541.

Reghizzi:2013:SLT

[RS13] Stefano Crespi Reghizzi and Pierluigi San Pietro. Strict local testability with consensus equals regularity, and other properties. International Journal of Foundations of Computer Science (IJFCS), 24(6):747–??, September 2013. CODEN IFCSEN. ISSN 0129-0541.

Rowland:2015:ASR

[RS15] Eric Rowland and Jeffrey Shallit. Automatic sets of rational numbers. International Journal of Foundations of Computer Science (IJFCS), 26(3):343–??, April 2015. CODEN IFC-SEN. ISSN 0129-0541.

Russell:2017:GSL

[RS17] Benjamin Russell and Susan Stepney. The geometry of speed limiting resources in physical models of computation. International Journal of Foundations of Computer Science (IJFCS), 28 (4):321–??, June 2017. CO-DEN IFCSEN. ISSN 0129-0541.

Renjith:2022:HCK

[RS22] P. Renjith and N. Sadagopan.

Hamiltonian cycle in K_{1,r}- free split graphs — a dichotomy. International Journal of Foundations of Computer Science (IJFCS), 33(01):1–32, January 2022. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500337.

Reynier:2016:VPT

[RT16] Pierre-Alain Reynier and Jean-Marc Talbot. Visibly pushdown transducers with well-nested outputs. International Journal of Foundations of Computer Science (IJFCS), 27(2):235–??, February 2016. CODEN IFCSEN. ISSN 0129-0541.

Rudy:2015:DRA

[Rud15] Jarosław Rudy. Dynamic random-access stored-program machine for runtime code modification. International Journal of Foundations of Computer Science (IJFCS), 26(4):441–??, June 2015. CODEN IFCSEN. ISSN 0129-0541.

Ruohonen:1996:ECP

[Ruo96] K. Ruohonen. An effective Cauchy-Peano existence theorem for unique solutions. International Journal of Foundations of Computer Science (IJFCS), 7 (2):151–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rovan:2022:FAS

[RV22] Branislav Rovan and András Varga. Finite approximations and similarity of languages. International Journal of Foundations of Computer Science (IJFCS), 33(08):967– 1003, December 2022. CO-DEN IFCSEN. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122500113.

Roy:2006:RMW

[RVT06] Krishnendu Roy, Ramachandran Vaidyanathan, and Jerry L. Trahan. Routing multiple width communications on the circuit switched tree. International Journal of Foundations of Computer Science (IJFCS), 17(2):271– ??, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rigo:2011:LCR

[RW11] Michel Rigo and Laurent Waxweiler. Logical characterization of recognizable sets of polynomials over a finite field. International Journal of Foundations of Computer Science (IJFCS), 22(7):1549–1563, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Rhodes:2001:TCC

[RWZ01] David L. Rhodes, Wayne Wolf, and Albert Zomaya. Two CoNP-complete schedule analysis problems. International Journal of Foundations of Computer Science (IJFCS), 12(5):565– 580, October 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ryabko:2015:CAF

[Rya15] Boris Ryabko. Complexity of approximating functions on real-life computers. International Journal of Foundations of Computer Science (IJFCS), 26(1):153– ??, January 2015. CODEN IFCSEN. ISSN 0129-0541.

Ryabko:2021:PRG

[Rya21] Boris Ryabko. A pseudorandom generator whose output is a normal sequence. International Journal of Foundations of Computer Science (IJFCS), 32(08):981–989, December 2021. ISSN 0129-

0541. URL https://www.
worldscientific.com/doi/
10.1142/S0129054121500325.

Ranjan:2012:VIP

[RZ12] Desh Ranjan and Mohammad Zubair. Vertex isoperimetric parameter of a computation graph. International Journal of Foundations of Computer Science (IJFCS), 23(4):941–??, June 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Saad:2022:SSC

[SA22] Aldosary Saad and Abdullah Alharbi. Securing smart city services in cyber-physical systems using the computation annealed selection process. International Journal of Foundations of Computer Science (IJFCS), 33(6– 7):531–557, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420035.

Sahni:2001:MAO

[Sah01] Sartaj Sahni. Models and algorithms for optical and optoelectronic parallel computers. International Journal of Foundations of Computer Science (IJFCS), 12 (3):249–264, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sakurai:2001:CMC

[Sak01] Takafumi Sakurai. Categorical model construction for proving syntactic properties. International Journal of Foundations of Computer Science (IJFCS), 12 (2):213–244, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Salomaa:2007:CSO

[Sal07] Arto Salomaa. Comparing subword occurrences in binary D0L sequences. International Journal of Foundations of Computer Science (IJFCS), 18(6):1395–1406, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Salomaa:2011:PSA

[Sal11] Arto Salomaa. Power sums associated with certain recursive procedures on words. International Journal of Foundations of Computer Science (IJFCS), 22(1):261– 272, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Salomaa:2013:FCB

[Sal13] Arto Salomaa. Functional constructions between reaction systems and propositional logic. International Journal of Foundations of Computer Science (IJFCS), 24(1):147–??, January 2013. CODEN IFCSEN. ISSN 0129-0541.

Salikhov:2018:ICD

[Sal18] Kamil Salikhov. Improved compression of DNA sequencing data with cascading Bloom filters. International Journal of Foundations of Computer Science (IJFCS), 29(8):1249– 1255, December 2018. ISSN 0129-0541.

Santhosh:2013:SSD

[San13] Prabhu M. Santhosh. Self stabilization in distributed knot detection. International Journal of Foundations of Computer Science (IJFCS), 24(6):913–??, September 2013. CODEN IFCSEN. ISSN 0129-0541.

Saoudi:1992:PAI

[Sao92] A. Saoudi. Pushdown automata on infinite trees and nondeterministic contextfree programs. International Journal of Foundations of Computer Science (IJFCS), 3(1):21–40, March 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Schopf:2001:USI

[SB01] Jennifer M. Schopf and Francine Berman. Using stochastic information to predict application behavior on contended resources. International Journal of Foundations of Computer Science (IJFCS), 12(3):341– 363 (or 341–364??), 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Seredynski:2012:DRB

[SB12] Marcin Seredynski and Pascal Bouvry. Direct reciprocity-based cooperation in mobile ad hoc networks. International Journal of Foundations of Computer Science (IJFCS), 23 (2):501–521, February 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sharma:2017:BST

[SB17] Gokarna Sharma and Costas Busch. The bursty Steiner tree problem. International Journal of Foundations of Computer Science (IJFCS), 28(7):869–??, November 2017. CODEN IFCSEN. ISSN 0129-0541.

Sburlan:2006:FRS

[Sbu06] Dragoş Sburlan. Further results on P systems with promoters/inhibitors. International Journal of Foundations of Computer Science (IJFCS), 17(1):205–??, February 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Schmidhuber:2002:HGK

[Sch02] J. Schmidhuber. Hierarchies of generalized Kolmogorov complexities and nonenumerable universal measures computable in the limit. International Journal of Foundations of Computer Science (IJFCS), 13(4):587–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Schnoor:2010:CMC

[Sch10] Henning Schnoor. The complexity of model checking for Boolean formulas. International Journal of Foundations of Computer Science (IJFCS), 21(3):289– 309, June 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Schmid:2013:ICR

[Sch13] Markus L. Schmid. Inside the class of regex languages. International Journal of Foundations of Computer Science (IJFCS), 24(7):1117–??, November 2013. CODEN IFCSEN. ISSN 0129-0541.

Salomaa:2015:AA

[SCIS15] Arto Salomaa, Francis Chin, Oscar Ibarra, and Sartaj Sahni. Alberto Apostolico. International Journal of Foundations of Computer Science (IJFCS), 26(5):iii, August 2015. CODEN IFC-SEN. ISSN 0129-0541.

Shtrakov:2016:CCF

[SD16] Slavcho Shtrakov and Ivo Damyanov. On the computational complexity of finite operations. International Journal of Foundations of Computer Science (IJFCS), 27(1):15–??, January 2016. CODEN IFC-SEN. ISSN 0129-0541.

Sarac:1999:DTS

[SEE99] K. Saraç, Ö. Egecioğlu, and A. El Abbadi. DFT techniques for size estimation of database join operations. International Journal of Foundations of Computer Science (IJFCS), 10 (1):81–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Seki:2020:SID

[Sek20] Shinnosuke Seki. Special issue developments in language theory 2018 — preface. International Journal of Foundations of Computer Science (IJFCS), 31(06): 663–665, September 2020. CODEN IFCSEN. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120020025.

Sellink:1998:CLE

[Sel98] M. P. A. Sellink. On the conservativity of Leibniz equality. International Journal of Foundations of Computer Science (IJFCS), 9(4):431– ??, December 1998. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Selivanov:2008:FHR

[Sel08] Victor Selivanov. Fine hierarchy of regular aperiodic ω-languages. International Journal of Foundations of Computer Science (IJFCS), 19(3):649–675, June 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sempere:2020:CLG

[Sem20]  José M. Sempere.  On

compensation loops in genomic duplications. International Journal of Foundations of Computer Science (IJFCS), 31(1):133– 142, January 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120400092.

Serbanuta:2009:PMA

[Şer09] Virgil Nicolae Şerbănuţă.

On Parikh matrices, ambiguity, and prints. International Journal of Foundations of Computer Science (IJFCS), 20(1):151– 165, February 2009. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Seth:2008:ACS

[Set08] Anil Seth. An alternative construction in symbolic reachability analysis of second order pushdown systems. International Journal of Foundations of Computer Science (IJFCS), 19(4):983– 998, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Schewe:2007:SAD

[SF07] Sven Schewe and Bernd Finkbeiner. Semi-automatic distributed synthesis. International Journal of Foundations of Computer Science (IJFCS), 18(1):113– 138, February 2007. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sun:2017:CAL

[SFL17] Lei Sun, Fangwei Fu, and Jian Liu. On the conjecture about the linear structures of rotation symmetric Boolean functions. International Journal of Foundations of Computer Science (IJFCS), 28(7):819–??, November 2017. CODEN IFCSEN. ISSN 0129-0541.

Skulrattanakulchai:2004:CAS

[SG04] San Skulrattanakulchai and Harold N. Gabow. Coloring algorithms on subcubic graphs. International Journal of Foundations of Computer Science (IJFCS), 15 (1):21–??, February 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Shu:2002:NNA

[SGZ02] Jiwu Shu, Yonggeng Gu, and Weimin Zheng. A novel numerical approach of computing American option. International Journal of Foundations of Computer Science (IJFCS), 13(5):685– ??, October 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sun:2017:SCB

[SH17] Zhiqiang Sun and Lei Hu. Several classes of Boolean functions with four-valued Walsh spectra. International Journal of Foundations of Computer Science (IJFCS), 28(4):357–??, June 2017. CODEN IFCSEN. ISSN 0129-0541.

Shi:2022:FDO

[SH22] Feng Shi and Xiaomei Hu. Fuzzy dynamic obstacle avoidance algorithm for basketball robot based on multi-sensor data fusion technology. International Journal of Foundations of Computer Science (IJFCS), 33(6–7):649– 666, September–November 2022. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054122420084.

Shallit:2004:SAL

[Sha04] Jeffrey Shallit. Simultaneous avoidance of large squares and fractional powers in infinite binary words. International Journal of Foundations of Computer Science (IJFCS), 15(2):317– ??, April 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Subramanian:2009:PM

[SHN09] K. G. Subramanian, Ang Miin Huey, and Atulya K. Nagar. On Parikh matrices. International Journal of Foundations of Computer Science (IJFCS), 20(2):211– 219, April 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Shur:2007:RAP

[Shu07] Arseny M. Shur. Rational approximations of polynomial factorial languages. International Journal of Foundations of Computer Science (IJFCS), 18(3):655– 665, June 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Shur:2011:EMP

[Shu11] Arseny M. Shur. On the existence of minimal βpowers. International Journal of Foundations of Computer Science (IJFCS), 22 (7):1683–1696, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Shur:2014:LFA

[Shu14] Arseny M. Shur. Languages with a finite antidictionary: Some growth questions. International Journal of Foundations of Computer Science (IJFCS), 25(8):937–??, December 2014. CODEN IFCSEN. ISSN 0129-0541.

Shur:2016:P

[Shu16] Arseny Shur. Preface. International Journal of Foundations of Computer Science (IJFCS), 27(2):101–??, February 2016. CODEN IFCSEN. ISSN 0129-0541.

Sibeyn:1997:RTT

[Sib97] Jop F. Sibeyn. Routing on triangles, tori and honeycombs. International Journal of Foundations of Computer Science (IJFCS), 8 (3):269–??, September 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sirakoulis:2015:CPC

[Sir15] Georgios Ch. Sirakoulis.

The computational paradigm of cellular automata in crowd evacuation. International Journal of Foundations of Computer Science (IJFCS), 26(7):851–??, November 2015. CODEN IFCSEN. ISSN 0129-0541.

Shi:2004:FAD

[SJ04] Qingmin Shi and Joseph Jájá. Fast algorithms for 3-D dominance reporting and counting. International Journal of Foundations of Computer Science (IJFCS), 15(4):673–??, August 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Saidane:2001:MPE

[SK01] Leila Azouz Saidane and F. Kamoun. Modelling and performance evaluation of the circulating multisequencer, the multi-tokens and the consensus algorithms in a real time distributed transactional system. International Journal of Foundations of Computer Science (IJFCS), 12(6):719– 750, December 2001. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Schmollinger:2003:DPA

[SK03] Martin Schmollinger and Michael Kaufmann. Designing parallel algorithms for hierarchical SMP clusters. International Journal of Foundations of Computer Science (IJFCS), 14(1):59– ??, February 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sahni:2004:EDL

[SK04] Sartaj Sahni and Kun Suk Kim. Efficient dynamic lookup for bursty access patterns. International Journal of Foundations of Computer Science (IJFCS), 15(4):567– ??, August 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Saikia:2020:DAA

[SK20] Parikshit Saikia and Sushanta Karmakar. Distributed approximation algorithms for Steiner tree in the CON-GESTED CLIQUE. International Journal of Foundations of Computer Science (IJFCS), 31(07):941– 968, November 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120500367.

Sahni:2003:DSO

[SKL03] Sartaj Sahni, Kun Suk Kim, and Haibin Lu. Data structures for one-dimensional packet classification using most-specific-rule matching. International Journal of Foundations of Computer Science (IJFCS), 14(3):337– ??, June 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Strauss:2008:CSB

[SKW08] Tinus Strauss, Derrick G.

Kourie, and Bruce W. Watson. A concurrent specification of Brzozowski’s DFA construction algorithm. International Journal of Foundations of Computer Science (IJFCS), 19(1):125– 135, February 2008. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sulzmann:2017:DBD

[SL17] Martin Sulzmann and Kenny Zhuo Ming Lu. Derivativebased diagnosis of regular expression ambiguity. International Journal of Foundations of Computer Science (IJFCS), 28(5):543–??, August 2017. CODEN IFC-SEN. ISSN 0129-0541.

Shen:2021:NHA

[SL21] Chenli Shen and Wensong Lin. NP-hardness and approximation algorithms for iterative pricing on social networks with externalities. International Journal of Foundations of Computer Science (IJFCS), 32(08):957– 979, December 2021. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054121500313.

Slobodova:1995:POW

[Slo95] A. Slobodova. On the power of one-way globally deterministic synchronized alternating Turing machines and multihead automata. International Journal of Foundations of Computer Science (IJFCS), 6(4):431–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Supol:2005:ACP

[ŠM05] Jan Šupol and Bořivoj Melichar. Arithmetic coding in parallel. International Journal of Foundations of Computer Science (IJFCS), 16(6):1207–??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Simunek:2007:BFA

[ŠM07] Martin Šimůnek and Bořivoj Melichar. Borders and finite automata. International Journal of Foundations of Computer Science (IJFCS), 18(4):859–871, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Subramanian:2013:TDD

[SMAN13] K. G. Subramanian, Kalpana Mahalingam, Rosni Abdullah, and Atulya K. Nagar. Two-dimensional digitized picture arrays and Parikh matrices. International Journal of Foundations of Computer Science (IJFCS), 24(3):393–??, April 2013. CODEN IFC-SEN. ISSN 0129-0541.

Smith:1995:HPT

[Smi95] S. F. Smith. Hybrid partialtotal type theory. International Journal of Foundations of Computer Science (IJFCS), 6(3):235–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Saoudi:1990:COT

[SMS90] A. Saoudi, D. E. Muller, and P. E. Schupp. On the complexity of omegatree sets and Nerode theorem. International Journal of Foundations of Computer Science (IJFCS), 1(1): 11–??, March 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Saoudi:1992:FSP

[SMS92] A. Saoudi, D. E. Muller, and P. E. Schupp. Finite state processes, Z-temporal logic and the monadic theory of the integers. International Journal of Foundations of Computer Science (IJFCS), 3(3):233–244, September 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Smyczynski:2012:CMI

[Smy12] Sebastian Smyczyński. Constantmemory iterative generation of special strings representing binary trees. International Journal of Foundations of Computer Science (IJFCS), 23(2):375– 387, February 2012. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Skrypnyuk:2013:RFS

[SN13] Nataliya Skrypnyuk and Flemming Nielson. Reachability for finite-state process algebras using Horn clauses. International Journal of Foundations of Computer Science (IJFCS), 24 (2):283–??, February 2013. CODEN IFCSEN. ISSN 0129-0541.

Salami:2024:SEM

[SNB24] Hosein Salami and Mostafa Nouri-Baygi. A simple and efficient method for accelerating construction of the gap-greedy spanner. International Journal of Foundations of Computer Science (IJFCS), 35(04):453– 481, June 2024. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054123500119.

Safavi-Naini:2011:USC

[SNJ11] Reihaneh Safavi-Naini and Shaoquan Jiang. Unconditionally secure conference key distribution: Security notions, bounds and constructions. International Journal of Foundations of Computer Science (IJFCS), 22(6):1369–1393, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Safavi-Naini:2006:RLS

[SNWW06] Reihaneh Safavi-Naini, Huaxiong Wang, and Duncan S. Wong. Resilient LKH: Secure multicast key distribution schemes. International Journal of Foundations of Computer Science (IJFCS), 17(5):1205– ??, October 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sosik:2009:P

[Sos09] Petr Sosík. Preface. International Journal of Foundations of Computer Science (IJFCS), 20(3):379– 380, June 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Singh:2004:HMD

[SP04] Mitali Singh and Viktor K. Prasanna. A hierarchical model for distributed collaborative computation in wireless sensor networks. International Journal of Foundations of Computer Science (IJFCS), 15(3):485– ??, June 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sprojcar:2009:PSM

[Špr09] Josef Šprojcar. Proposal of a semiformal model of anonymous communication. International Journal of Foundations of Computer Science (IJFCS), 20(3):523– 548, June 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Shyamasundar:2000:LRS

[SR00a] R. K. Shyamasundar and S. Ramesh. Languages for reactive specifications: Synchrony vs asynchrony. International Journal of Foundations of Computer Science (IJFCS), 11(2):283– 314, 2000. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Shyamasundar:2000:PRP

[SR00b] R. K. Shyamasundar and S. Ramesh. Part 2 (regular papers). International Journal of Foundations of Computer Science (IJFCS), 11 (2):283–??, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sharma:2021:AAO

[SR21] Amit Sharma and P. Venkata Subba Reddy. Algorithmic aspects of outer-independent total Roman domination in graphs. International Journal of Foundations of Computer Science (IJFCS), 32 (03):331–339, April 2021. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054121500180.

Sadhu:2020:EAC

[SRN^{+}20] Sanjib Sadhu, Sasanka Roy, Soumen Nandi, Subhas C. Nandy, and Suchismita Roy. Efficient algorithm for computing the triangle maximizing the length of its smallest side inside a convex polygon. International Journal of Foundations of Computer Science (IJFCS), 31(04):421–443, June 2020. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500173.

Sosik:2011:PFR

[SRPC11] Petr Sosík, Alfonso Rodríguez-Patón, and Luděk Cienciala. On the power of families of recognizer spiking neural P systems. International Journal of Foundations of Computer Science (IJFCS), 22(1):75–88, January 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Saha:2015:NRF

[SRR15] Subrata Saha, Sanguthevar Rajasekaran, and Rampi Ramprasad. Novel randomized feature selection algorithms. International Journal of Foundations of Computer Science (IJFCS), 26 (3):321–??, April 2015. CO-DEN IFCSEN. ISSN 0129-0541.

Sajith:1999:PVC

[SS99] G. Sajith and S. Saxena. Parallel vertex colouring of interval graphs. International Journal of Foundations of Computer Science (IJFCS), 10(1):19–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sitharam:2001:DLB

[SS01] M. Sitharam and Timothy Straney. Derandomized learning of Boolean functions over finite Abelian groups. International Journal of Foundations of Computer Science (IJFCS), 12 (4):491–516, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Salomaa:2007:SCA

[SS07a] Kai Salomaa and Paul Schofield. State complexity of additive weighted finite automata. International Journal of Foundations of Computer Science (IJFCS), 18(6):1407–1416, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Soper:2007:IAA

[SS07b] Alan J. Soper and Vitaly A.

Strusevich. An improved approximation algorithm for the two-machine flow shop scheduling problem with an interstage transporter. International Journal of Foundations of Computer Science (IJFCS), 18(3):565– 591, June 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Schaeffer:2012:CEC

[SS12a] Luke Schaeffer and Jeffrey Shallit. The critical exponent is computable for automatic sequences. International Journal of Foundations of Computer Science (IJFCS), 23(8):1611–??, December 2012. CODEN IFC-SEN. ISSN 0129-0541.

Switalski:2012:EMS

[SS12b] Piotr Switalski and Franciszek Seredynski. An effective multiprocessor scheduling with use of geo metaheuristic. International Journal of Foundations of Computer Science (IJFCS), 23(2):465–481, February 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Salomaa:2024:VEM

[SS24] A. Salomaa and M. Steinby. Volume edited by Magnus Steinby. International Journal of Foundations of Computer Science (IJFCS), 35(1–2):7– 10, January–February 2024. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054123760018.

Shi:2020:NCC

[SSF20] Zexia Shi, Lei Sun, and Fang-Wei Fu. New constructions of codebooks nearly meeting the Welch bound. International Journal of Foundations of Computer Science (IJFCS), 31(07): 875–889, November 2020. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412050032X.

Sankoff:1996:SPS

[SSK96] D. Sankoff, G. Sundaram, and J. Kececioglu. Steiner points in the space of genome rearrangements. International Journal of Foundations of Computer Science (IJFCS), 7(1):1–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Shakhlevich:2009:SMS

[SSS09] Natalia V. Shakhlevich, Akiyoshi Shioura, and Vitaly A. Strusevich. Single machine scheduling with controllable processing times by submodular optimization. International Journal of Foundations of Computer Science (IJFCS), 20(2):247–269, April 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Salomaa:2013:GKD

[SSS13] Arto Salomaa, Kai T. Salomaa, and Andrew L. Szilard. Goodby to the Kindhearted Dragon Prof. Sheng Yu, 1950–2012. International Journal of Foundations of Computer Science (IJFCS), 24(7):945–??, November 2013. CODEN IFCSEN. ISSN 0129-0541.

Symvonis:1993:SPS

[ST93] A. Symvonis and S. Tragoudas.

Searching a pseudo 3-sided solid orthoconvex grid. International Journal of Foundations of Computer Science (IJFCS), 4(4):325–354, December 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sprague:1999:QTA

[ST99] A. P. Sprague and T. Takaoka. O(1) query time algorithm for all pairs shortest distances on interval graphs. International Journal of Foundations of Computer Science (IJFCS), 10(4):465– 472, 1999. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Shi:2001:LBH

[ST01] Weisong Shi and Zhimin Tang. Load balancing in home-based software DSMS. International Journal of Foundations of Computer Science (IJFCS), 12(3):307– 324, 2001. CODEN IFC-SEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Saifullah:2011:SSC

[ST11] Abusayeed Saifullah and Yung H. Tsin. Selfstabilizing computation of 3-edge-connected components. International Journal of Foundations of Computer Science (IJFCS), 22 (5):1161–1185, August 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Shikishima-Tsuji:2016:RIH

[ST16] Kayoko Shikishima-Tsuji. Regularity of iterative hairpin completions of crossing (2, 2)-words. International Journal of Foundations of Computer Science (IJFCS), 27(3):375–??, April 2016. CODEN IFCSEN. ISSN 0129-0541.

Staiger:2005:IIF

[Sta05] Ludwig Staiger. Infinite iterated function systems in Cantor space and the Hausdorff measure of ω-power languages. International Journal of Foundations of Computer Science (IJFCS), 16(4):787–??, August 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Staiger:2007:PFL

[Sta07] Ludwig Staiger. Prefix-free Łukasiewicz languages. International Journal of Foundations of Computer Science (IJFCS), 18(6):1417– 1423, December 2007. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Stewart:1993:TAA

[Ste93] I. A. Stewart. On two approximation algorithms for the clique problem. International Journal of Foundations of Computer Science (IJFCS), 4(2):117– 134, June 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Steinberg:2011:ATC

[Ste11] Benjamin Steinberg. The averaging trick and the Černý Conjecture. International Journal of Foundations of Computer Science (IJFCS), 22(7):1697–1706, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Steinby:2024:MMS

[Ste24] Paula Steinby. In memoriam: Magnus Steinby (1941–2021). International Journal of Foundations of Computer Science (IJFCS), 35(1–2):3– 5, January–February 2024. ISSN 0129-0541. URL

                                        https://www.worldscientific.
                                        com/doi/10.1142/S0129054123200010.

Stolboushkin:1992:CPP

[Sto92] A. P. Stolboushkin. On the computing power of programs with sets. International Journal of Foundations of Computer Science (IJFCS), 3(2):161– 180, June 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Subrahmanian:1990:RTBa

[Sub90a] V. S. Subrahmanian. A ring-theoretic basis for logic programming. International Journal of Foundations of Computer Science (IJFCS), 1(1):23–??, March 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Subrahmanian:1990:RTBb

[Sub90b] V. S. Subrahmanian. A ring-theoretic basis for logic programming. International Journal of Foundations of Computer Science (IJFCS), 1(4):465–??, December 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Subramani:2005:CRW

[Sub05] K. Subramani. Cascading random walks. International Journal of Foundations of Computer Science (IJFCS), 16(3):599–??, June 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Suchenek:1990:ALH

[Suc90] Marek A. Suchenek. Applications of Lyndon homomorphism theorems to the theory of minimal models. International Journal of Foundations of Computer Science (IJFCS), 1(1):49– ??, March 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sun:2000:DRR

[Sun00] Hung-Min Sun. On the dealer’s randomness required in perfect secret sharing schemes with access structures of constant rank. International Journal of Foundations of Computer Science (IJFCS), 11 (2):263–282, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sunckel:2005:DCM

[Sun05] Bettina Sunckel. On the descriptional complexity of metalinear CD grammar systems. International Journal of Foundations of Computer Science (IJFCS), 16 (5):1011–??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sun:2011:PSM

[Sun11] Wei Sun. Population size modeling for Ga in timecritical task scheduling. International Journal of Foundations of Computer Science (IJFCS), 22(3):603– 620, April 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Sutner:2003:RPA

[Sut03] Klaus Sutner. Reduced power automata and sofic systems. International Journal of Foundations of Computer Science (IJFCS), 14(6):1117–??, December 2003. CODEN IFCSEN.

ISSN 0129-0541 (print), 1793-6373 (electronic).

Sutner:2014:IIT

[Sut14] K. Sutner. Iteration of invertible transductions. International Journal of Foundations of Computer Science (IJFCS), 25(7):857–??, November 2014. CODEN IFCSEN. ISSN 0129-0541.

Suzuki:2013:EET

[SUZ13] Akira Suzuki, Kei Uchizawa, and Xiao Zhou. Energyefficient threshold circuits computing mod functions. International Journal of Foundations of Computer Science (IJFCS), 24(1):15– ??, January 2013. CODEN IFCSEN. ISSN 0129-0541.

Stuber:2009:DWM

[SVF09] Torsten Stüber, Heiko Vogler, and Zoltán Fülöp. Decomposition of weighted multioperator tree automata. International Journal of Foundations of Computer Science (IJFCS), 20(2):221– 245, April 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Santoso:2001:SHR

[SVSN01] J. Santoso, G. D. Van Albada, P. M. A. Sloot, and B. A. A. Nazief. Simulation of hierarchical resource management for meta-computing systems.

International Journal of Foundations of Computer Science (IJFCS), 12(5):629– 643 (or 629–644??), October 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Smyth:2009:AHP

[SW09] W. F. Smyth and Shu Wang. An adaptive hybrid patternmatching algorithm on indeterminate strings. International Journal of Foundations of Computer Science (IJFCS), 20(6):985–1004, December 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Seki:2017:CFH

[SW17] Shinnosuke Seki and Andrew Winslow. The complexity of fixed-height patterned tile self-assembly. International Journal of Foundations of Computer Science (IJFCS), 28(5):465–??, August 2017. CODEN IFC-SEN. ISSN 0129-0541.

Smith:1997:CPL

[SWZ97] Carl H. Smith, Rolf Wiehagen, and Thomas Zeugmann. Classifying predicates and languages. International Journal of Foundations of Computer Science (IJFCS), 8(1):15–41, March 1997. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Salomaa:2005:P

[SY05] Kai Salomaa and Sheng Yu. Preface. International Journal of Foundations of Computer Science (IJFCS), 16(3):399–??, June 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Salomaa:2007:SCC

[SY07] Kai Salomaa and Sheng Yu. On the state complexity of combined operations and their estimation. International Journal of Foundations of Computer Science (IJFCS), 18(4):683– 698, August 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Salomaa:2010:SOP

[SY10] Arto Salomaa and Sheng Yu. Subword occurrences, Parikh matrices and Lyndon images. International Journal of Foundations of Computer Science (IJFCS), 21 (1):91–111, February 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Say:2012:QCA

[SY12] A. C. Cem Say and Abuzer Yakaryilmaz. Quantum counter automata. International Journal of Foundations of Computer Science (IJFCS), 23(5):1099– ??, August 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Salehi:2019:NRV

[SYS19] Özlem Salehi, Abuzer Yakaryılmaz, and A. C. Cem Say. New results on vector and Homing vector automata. International Journal of Foundations of Computer Science (IJFCS), 30(8):1335– 1361, December 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119500291.

Sun:2022:RPH

[SZ22] Ziyi Sun and Jing Zhang. Research on prediction of housing prices based on GA-PSO-BP neural network model: Evidence from Chongqing, China. International Journal of Foundations of Computer Science (IJFCS), 33(6– 7):805–818, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420163.

Sun:2020:CSR

[SZFX20] Yinxia Sun, Futai Zhang, Anmin Fu, and Zhe Xia. CCA-secure and revocable certificateless encryption with ciphertext evolution. International Journal of Foundations of Computer Science (IJFCS), 31(2):175– 191, February 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500021.

Sun:2017:CPP

[SZQ^{+}17] Jiameng Sun, Binrui Zhu, Jing Qin, Jiankun Hu, and Qianhong Wu. Confidentialitypreserving publicly verifiable computation. International Journal of Foundations of Computer Science (IJFCS), 28(6):799–??, September 2017. CODEN IFCSEN. ISSN 0129-0541.

Shi:2018:SDF

[SZQS18] Minjia Shi, Hongwei Zhu, Liqin Qian, and Patrick Solé. On self-dual four circulant codes. International Journal of Foundations of Computer Science (IJFCS), 29(7):1143– 1150, November 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118500259.

Szwast:1995:NAP

[Szw95] W. Szwast. A note on the asymptotic probabilities of existential second-order minimal Goedel sentences with equality. International Journal of Foundations of Computer Science (IJFCS), 6(4):339–??, 1995. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Teh:2017:MRS

[TA17] Wen Chean Teh and Adrian Atanasiu. Minimal reaction systems revisited and reaction system rank. International Journal of Foundations of Computer Science (IJFCS), 28(3):247–??, April 2017. CODEN IFC-SEN. ISSN 0129-0541.

Tamm:2008:TMB

[Tam08] Hellis Tamm. On transition minimality of bideterministic automata. International Journal of Foundations of Computer Science (IJFCS), 19(3):677–690, June 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Tirnauca:2020:CSB

[TBGP20] Cristina Tîrnăucă, José L. Balcázar, and Domingo Gómez-Pérez. Closed-set-based discovery of representative association rules. International Journal of Foundations of Computer Science (IJFCS), 31(1):143– 156, January 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120400109.

Ti:2010:SIS

[TCLS10] Yen-Wu Ti, Ching-Lueh Chang, Yuh-Dauh Lyuu, and Alexander Shen. Sets of k-independent strings. International Journal of Foundations of Computer Science (IJFCS), 21(3):321– 327, June 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Tang:2014:CRB

[TCT14] Deng Tang, Claude Carlet, and Xiaohu Tang. A class of 1-resilient Boolean functions with optimal algebraic immunity and good behavior against fast algebraic attacks. International Journal of Foundations of Computer Science (IJFCS), 25(6):763– ??, September 2014. CO-DEN IFCSEN. ISSN 0129-0541.

Teh:2015:CWP

[Teh15] Wen Chean Teh. On core words and the Parikh matrix mapping. International Journal of Foundations of Computer Science (IJFCS), 26(1):123–??, January 2015. CODEN IFCSEN. ISSN 0129-0541.

Teh:2016:PMS

[Teh16a] Wen Chean Teh. Parikh matrices and strong Mequivalence. International Journal of Foundations of Computer Science (IJFCS), 27(5):545–??, August 2016. CODEN IFCSEN. ISSN 0129-0541.

Teh:2016:SME

[Teh16b] Wen Chean Teh. Separability of M-equivalent words by morphisms. International Journal of Foundations of Computer Science (IJFCS), 27(1):39–??, January 2016. CODEN IFC-SEN. ISSN 0129-0541.

Teh:2018:CFP

[Teh18] Wen Chean Teh. Compositions of functions and permutations specified by minimal reaction systems. International Journal of Foundations of Computer Science (IJFCS), 29(7):1165– 1179, November 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118500272.

Teichmann:2017:RAW

[Tei17] Markus Teichmann. Regular approximation of weighted linear context-free tree languages. International Journal of Foundations of Computer Science (IJFCS), 28 (5):523–??, August 2017. CODEN IFCSEN. ISSN 0129-0541.

Teixeira:2018:SEM

[TFF18] Andreia Sofia Teixeira, Francisco Fernandes, and Alexandre P. Francisco. SpliceTAPyR — an efficient method for transcriptome alignment. International Journal of Foundations of Computer Science (IJFCS), 29(8):1297–1310, December 2018. ISSN 0129-0541.

Tseng:2019:AMR

[TFS19] Yi-Fan Tseng, Chun-I Fan, and Cheng-Wei Sung. On the anonymity of multireceiver identity-based encryption based on Fujisaki– Okamoto transformation. International Journal of Foundations of Computer Science (IJFCS), 30(4):493– 509, June 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400094.

Touyama:2001:PEP

[TH01] Takayoshi Touyama and Susumu Horiguchi. Performance evaluation of practical parallel computer model LogPQ. International Journal of Foundations of Computer Science (IJFCS), 12 (3):325–340, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Tran:2022:SCW

[TH22] Tien Tran and Dung T.

Huynh. Symmetric connectivity in wireless sensor networks with π/3 directional antennas. International Journal of Foundations of Computer Science (IJFCS), 33(02):119– 140, February 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500022.

Thanh:1991:RMD

[Tha91] Tung Nguyen Thanh. A relational model of demonic nondeterministic programs. International Journal of Foundations of Computer Science (IJFCS), 2(2):101– 132, June 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Thomas:2006:MSS

[Tho06] Fernique Thomas. Multidimensional Sturmian sequences and generalized substitutions. International Journal of Foundations of Computer Science (IJFCS), 17(3):575–??, June 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Tu:2013:COV

[TJZ13] Ziran Tu, Yupeng Jiang, and Xiangyong Zeng. Constructing odd variable Boolean functions with optimal algebraic immunity. International Journal of Foundations of Computer Science (IJFCS), 24(3):409–??, April 2013. CODEN IFC-SEN. ISSN 0129-0541.

Tokat:2019:DI

[TK19] Hüseyin Tokat and Alpay Kirlangiç. On the domination integrity. International Journal of Foundations of Computer Science (IJFCS), 30(6–7):811–826, September–November 2019. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119500230.

Tse:1999:CSA

[TL99] S. S. H. Tse and F. C. M.

Lau. On the complexity of some adaptive polling algorithms in general networks. International Journal of Foundations of Computer Science (IJFCS), 10 (2):211–??, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Torkestani:2013:LAB

[Tor13] Javad Akbari Torkestani. A learning automata-based algorithm to the stochastic min-degree constrained minimum spanning tree problem. International Journal of Foundations of Computer Science (IJFCS), 24(3):329– ??, April 2013. CODEN IFCSEN. ISSN 0129-0541.

Torkestani:2015:ASC

[Tor15] Javad Akbari Torkestani.

Algorithms for Steiner connected dominating set problem based on learning automata theory. International Journal of Foundations of Computer Science (IJFCS), 26(6):769–??, September 2015. CODEN IFCSEN. ISSN 0129-0541.

Tosic:2006:CCF

[Toš06] Predrag T. Tošić. On the complexity of counting fixed points and Gardens of Eden in sequential dynamical systems on planar bipartite graphs. International Journal of Foundations of Computer Science (IJFCS), 17 (5):1179–??, October 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Tran:2002:UPC

[Tra02] Nicholas Tran. On universally polynomial contextfree languages. International Journal of Foundations of Computer Science (IJFCS), 13(6):829–??, December 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Truthe:2008:RCF

[Tru08] Bianca Truthe. Remarks on context-free parallel communicating grammar systems generating crossed agreements. International Journal of Foundations of Computer Science (IJFCS), 19(4):873–886, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Tse:2016:BAT

[Tse16] Savio S. H. Tse. Belated analyses of three creditbased adaptive polling algorithms. International Journal of Foundations of Computer Science (IJFCS), 27 (5):579–??, August 2016. CODEN IFCSEN. ISSN 0129-0541.

Trejo-Sanchez:2017:SSA

[TSFZRP17] Joel Antonio Trejo-Sánchez, José Alberto Fernández-Zepeda, and Julio César Ramírez-Pacheco. A selfstabilizing algorithm for a maximal 2-packing in a cactus graph under any scheduler. International Journal of Foundations of Computer Science (IJFCS), 28(8): 1021–??, December 2017. CODEN IFCSEN. ISSN 0129-0541.

Tsin:2006:EDA

[Tsi06] Yung H. Tsin. An efficient distributed algorithm for 3-edge-connectivity. International Journal of Foundations of Computer Science (IJFCS), 17(3):677– ??, June 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Tiwari:2013:CPR

[TSS13] S. P. Tiwari, Shambhu Sharan, and Anupam K. Singh. On coverings of products of rough transformation semigroups. International Journal of Foundations of Computer Science (IJFCS), 24 (3):375–??, April 2013. CO-DEN IFCSEN. ISSN 0129-0541.

Takahashi:2001:LRF

[TST01a] Masako Takahashi, M. Sato, and Y. Toyama. Lambdarepresentable functions over term algebras. International Journal of Foundations of Computer Science (IJFCS), 12(1):3–30 (or 3–29??), 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Tsukada:2001:MLT

[TST01b] Yasuyuki Tsukada, M. Sato, and Y. Toyama. Martin-Löf’s type theory as an open-ended framework. International Journal of Foundations of Computer Science (IJFCS), 12(1):31–67 (or 31–68??), 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic). See errata [Tsu01].

Tsukada:2001:EPM

[Tsu01] Yasuyuki Tsukada. Errata: The paper: Martin-Löf’s Type Theory as an Open-Ended Framework. International Journal of Foundations of Computer Science (IJFCS), 12(5):695– ??, October 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic). See [TST01b].

Tian:2016:SFL

[TSZ16] Jing Tian, Yong Shao, and Xianzhong Zhao. Out subword-free languages and its subclasses. International Journal of Foundations of Computer Science (IJFCS), 27(3):305–??, April 2016. CODEN IFCSEN. ISSN 0129-0541.

Trahan:1994:API

[TV94] J. L. Trahan and S. Vedantham. Analysis of PRAM instruction sets from a log cost perspective. International Journal of Foundations of Computer Science (IJFCS), 5(3/4):??, 1994. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Takaoka:2007:FLA

[TV07] Tadao Takaoka and Stephen Violich. Fusing loopless algorithms for combinatorial generation. International Journal of Foundations of Computer Science (IJFCS), 18(2):263–293, April 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Tan:2014:REQ

[TV14] Tony Tan and Domagoj Vrgoč. Regular expressions for querying data graphs. International Journal of Foundations of Computer Science (IJFCS), 25(8):971– ??, December 2014. CODEN IFCSEN. ISSN 0129-0541.

Tsay:2009:ACR

[TW09] Yih-Kuen Tsay and Bow-Yaw Wang. Automated compositional reasoning of intuitionistically closed regular properties. International Journal of Foundations of Computer Science (IJFCS), 20(4):747– 762, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Tartary:2011:EIT

[TWZ11] Christophe Tartary, Huaxiong Wang, and Yun Zhang. An efficient and information theoretically secure rational secret sharing scheme based on symmetric bivariate polynomials. International Journal of Foundations of Computer Science (IJFCS), 22(6):1395–1416, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Thiagarajan:2002:SI

[TY02] P. S. Thiagarajan and R. Yap. Special issue. International Journal of Foundations of Computer Science (IJFCS), 13(3):313–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Tang:2003:IIV

[TY03] Peiyi Tang and Pen-Chung Yew. Interprocedural induction variable analysis. International Journal of Foundations of Computer Science (IJFCS), 14(3):405– ??, June 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Turker:2015:CSP

[TY15] Uraz Cengiz Türker and Hüsnü Yenigün. Complexities of some problems related to synchronizing, nonsynchronizing and monotonic automata. International Journal of Foundations of Computer Science (IJFCS), 26(1):99–??, January 2015. CODEN IFC-SEN. ISSN 0129-0541.

Touili:2023:RAS

[TY23] Tayssir Touili and Xin Ye.

Reachability analysis of self modifying code. International Journal of Foundations of Computer Science (IJFCS), 34(05):507– 536, August 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500290.

Tian:2017:ORA

[TYM^{+}17] Yangguang Tian, Guomin Yang, Yi Mu, Shiwei Zhang, Kaitai Liang, and Yong Yu. One-round attributebased key exchange in the multi-party setting. International Journal of Foundations of Computer Science (IJFCS), 28(6):725–??, September 2017. CODEN IFCSEN. ISSN 0129-0541.

Tucker:1991:PSR

[TZ91] J. V. Tucker and J. I. Zucker. Projections of semicomputable relations on abstract data types. International Journal of Foundations of Computer Science (IJFCS), 2(3):267–??, September 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Taheri:2011:PSD

[TZ11] Javid Taheri and Albert Y. Zomaya. On the performance of static and dynamic location management strategies in mobile computing. International Journal of Foundations of Computer Science (IJFCS), 22(3):519– 546, April 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Uehara:1999:MLF

[Ueh99] R. Uehara. A measure for the Lexicographically First Maximal Independent Set Problem and its limits. International Journal of Foundations of Computer Science (IJFCS), 10 (4):473–482, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ueno:2013:BRB

[Uen13]  Kenya Ueno.  Breaking

the rectangle bound barrier against formula size lower bounds. International Journal of Foundations of Computer Science (IJFCS), 24(8):1339–??, December 2013. CODEN IFC-SEN. ISSN 0129-0541.

Urgaonkar:2007:APC

[URS07] Bhuvan Urgaonkar, Arnold L.

Rosenberg, and Prashant Shenoy. Application placement on a cluster of servers. International Journal of Foundations of Computer Science (IJFCS), 18(5): 1023–1041, October 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ungor:2002:PTS

[ÜS02] A. Üngör and A. Sheffer.

Pitching tents in space-time: Mesh generation for discontinuous Galerkin method. International Journal of Foundations of Computer Science (IJFCS), 13(2):201– ??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Uehara:2007:CLP

[UU07] Ryuhei Uehara and Yushi Uno. On computing longest paths in small graph classes. International Journal of Foundations of Computer Science (IJFCS), 18(5):911– 930, October 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

VanZijl:2005:MNS

[Van05] Lynette Van Zijl. Magic numbers for symmetric difference NFAs. International Journal of Foundations of Computer Science (IJFCS), 16(5):1027– ??, October 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

vanderHoek:1992:MSI

[vdHM92] W. van der Hoek and J.- J. Ch. Meyer. Making some issues of implicit knowledge explicit. International Journal of Foundations of Computer Science (IJFCS), 3(2):193–224, June 1992. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

vanderMeyden:2000:PBL

[vdM00] R. van der Meyden. Predicate boundedness of linear monadic datalog is in PSPACE. International Journal of Foundations of Computer Science (IJFCS), 11(4):591–614, 2000. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Vergnaud:2009:NEP

[Ver09] Damien Vergnaud. New extensions of pairing-based signatures into universal (multi) designated verifier signatures. International Journal of Foundations of Computer Science (IJFCS), 20(1):109–133, February 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Verbitsky:2001:RQB

[VG01] Oleg Verbitsky and Shafi Goldwasser. Remarks on a query-based variant of the parallel repetition theorem. International Journal of Foundations of Computer Science (IJFCS), 12(4):517– 531 (or 517–532??), 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Viksna:1996:IIL

[Vik96] J. Viksna. Inductive inference of limiting programs with bounded number of mind changes. International Journal of Foundations of Computer Science (IJFCS), 7(3):187–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Vinodchandran:2005:NCM

[Vin05] N. V. Vinodchandran. Nondeterministic circuit minimization problem and derandomizing Arthur–Merlin games. International Journal of Foundations of Computer Science (IJFCS), 16(6):1297–??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Vermeulen-Jourdan:2005:LDS

[VJDT05] L. Vermeulen-Jourdan, C. Dhaenens, and E-G. Talbi. Linkage disequilibrium study with a parallel adaptive GA. International Journal of Foundations of Computer Science (IJFCS), 16(2):241– ??, April 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

vanLeeuwen:2015:SCR

[vLW15] Jan van Leeuwen and Jiří Wiedermann. Separating the classes of recursively enumerable languages based on machine size. International Journal of Foundations of Computer Science (IJFCS), 26(6):677–??, September 2015. CODEN IFCSEN. ISSN 0129-0541.

Vorel:2016:SSC

[Vor16] Vojtěch Vorel. Subset synchronization and careful synchronization of binary finite automata. International Journal of Foundations of Computer Science (IJFCS), 27(5):557–??, August 2016. CODEN IFC-SEN. ISSN 0129-0541.

Vorel:2018:BPJ

[Vor18] Vojtěch Vorel. On basic properties of jumping finite automata. International Journal of Foundations of Computer Science (IJFCS), 29(1):1–??, January 2018. CODEN IFC-SEN. ISSN 0129-0541.

Voisin:1999:SCP

[VP99] F. Voisin and G. R. Perrin.

Sparse computation with PEI. International Journal of Foundations of Computer Science (IJFCS), 10 (4):425–442, 1999. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Vincent:1993:RJF

[VS93] M. W. Vincent and B. Srinivasan. Redundancy and the justification for fourth normal form in relational databases. International Journal of Foundations of Computer Science (IJFCS), 4(4):355–366, December 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Vollmer:1993:CFM

[VW93] H. Vollmer and K. W. Wagner. The complexity of finding middle elements. International Journal of Foundations of Computer Science (IJFCS), 4(4):293–308, December 1993. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wan:2003:SHM

[WAF03] Peng-Jun Wan, Khaled M. Alzoubi, and Ophir Frieder. A simple heuristic for minimum connected dominating set in graphs. International Journal of Foundations of Computer Science (IJFCS), 14(2):323–??, April 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wernicke:2006:CCA

[WAG^{+}06] Sebastian Wernicke, Jochen Alber, Jens Gramm, Jiong Guo, and Rolf Niedermeier. The computational complexity of avoiding forbidden submatrices by row deletions. International Journal of Foundations of Computer Science (IJFCS), 17 (6):1467–1484, December 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wang:2004:ICN

[Wan04]  Farn Wang.  Inductive

composition of numbers with maximum, minimum, and addition: a new theory for program executiontime analysis. International Journal of Foundations of Computer Science (IJFCS), 15(6):865–??, December 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wang:2006:P

[Wan06] Farn Wang. Preface. International Journal of Foundations of Computer Science (IJFCS), 17(4):731– ??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wang:2014:DRG

[Wan14] Pai-Chou Wang. Dynamic reducts generation using cascading hashes. International Journal of Foundations of Computer Science (IJFCS), 25(2):219–??, February 2014. CODEN IFCSEN. ISSN 0129-0541.

Wang:2021:EDH

[Wan21] Shiying Wang. The rextra diagnosability of hyper Petersen graphs. International Journal of Foundations of Computer Science (IJFCS), 32(04):405– 416, June 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500234.

Wang:2004:HCD

[WC04] Dajin Wang and Jiannong Cao. On hierarchical configuration of distributed systems on mesh and hypercube. International Journal of Foundations of Computer Science (IJFCS), 15(3):517– ??, June 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wei:2013:GCK

[WC13] Yunchao Wei and Fuguang Chen. Generalized connectivity of (n, k)-star graphs. International Journal of Foundations of Computer Science (IJFCS), 24(8): 1235–??, December 2013. CODEN IFCSEN. ISSN 0129-0541.

Wang:2014:ZKB

[WCD^{+}14] Eric Wang, Cewei Cui, Zhe Dang, Thomas R. Fischer, and Linmin Yang. Zeroknowledge blackbox testing: Where are the faults? International Journal of Foundations of Computer Science (IJFCS), 25(2):195–??, February 2014. CODEN IFCSEN. ISSN 0129-0541.

Wan-Di:1990:PAP

[WD90] Wei Wan-Di. On a personnel assignment problem. International Journal of Foundations of Computer Science (IJFCS), 1(1):61–??, March 1990. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wu:2003:BAH

[WD03] Jie Wu and Fei Dai. Broadcasting in ad hoc networks based on self-pruning. International Journal of Foundations of Computer Science (IJFCS), 14(2):201– ??, April 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wang:2020:LCB

[WD20] Lianhua Wang and Xiaoni Du. Linear complexity of binary threshold sequences derived from generalized polynomial quotient with primepower modulus. International Journal of Foundations of Computer Science (IJFCS), 31(05):569– 581, August 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500264.

Wang:2021:IFD

[WDFN21] Rong Wang, Xiaoni Du, Cuiling Fan, and Zhihua Niu. Infinite families of 2-designs from a class of linear codes related to Dembowski–Ostrom functions. International Journal of Foundations of Computer Science (IJFCS), 32(03): 253–267, April 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500143.

Wang:2017:DCC

[WF17] Xiang Wang and Fang-Wei Fu. Deterministic construction of compressed sensing matrices from codes. International Journal of Foundations of Computer Science (IJFCS), 28(2):99–109, February 2017. CODEN IFCSEN. ISSN 0129-0541.

Wen:2021:CMC

[WF21] Jiejing Wen and Fang-Wei Fu. On the construction of multiply constantweight codes. International Journal of Foundations of Computer Science (IJFCS), 32(07):861– 870, November 2021. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S012905412150026X.

Wang:2015:SCM

[WFG15] Shiying Wang, Kai Feng, and Yubao Guo. Sufficient conditions for maximally kisoperimetric edge connectivity of graphs. International Journal of Foundations of Computer Science (IJFCS), 26(5):583–??, August 2015. CODEN IFC-SEN. ISSN 0129-0541.

Wadhwa:2017:AAR

[WG17] Vaishali M. Wadhwa and Deepak Garg. Approximation algorithm for resource allocation problems with time dependent penalties. International Journal of Foundations of Computer Science (IJFCS), 28(7):931– ??, November 2017. CO-DEN IFCSEN. ISSN 0129-0541.

Wang:2018:ERS

[WGD18] Hong Wang, Jie Guan, and Lin Ding. On equivalence relations of state diagram of cascade connection of an LFSR into an NFSR. International Journal of Foundations of Computer Science (IJFCS), 29(7):1107– 1117, November 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118500235.

Wang:2016:SSS

[WGF16] Xianfang Wang, Jian Gao, and Fang-Wei Fu. Secret sharing schemes from linear codes over F_{p} + νF_{p}. International Journal of Foundations of Computer Science (IJFCS), 27(5):595–??, August 2016. CODEN IFC-SEN. ISSN 0129-0541.

Wu:2003:SOS

[WH03] Chin-Hsiung Wu and Shi-Jinn Horng. Scalable and optimal speed-up parallel algorithms for template matching on arrays with reconfigurable optical buses. International Journal of Foundations of Computer Science (IJFCS), 14(1):79– ??, February 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wei:2017:CES

[WHLH17] Jianghong Wei, Xinyi Huang, Wenfen Liu, and Xuexian Hu. Cost-effective and scalable data sharing in cloud storage using hierarchical attribute-based encryption with forward security. International Journal of Foundations of Computer Science (IJFCS), 28(7):843–??, November 2017. CODEN IFCSEN. ISSN 0129-0541.

Widmer:2012:PCL

[Wid12] Steven Widmer. Permutation complexity and the letter doubling map. International Journal of Foundations of Computer Science (IJFCS), 23(8):1653–??, December 2012. CODEN IFC-SEN. ISSN 0129-0541.

Wilmes:1991:FPS

[Wil91] Thomas Wilmes. Functional production systems viewed as grammars. International Journal of Foundations of Computer Science (IJFCS), 2(1):23–??, March 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Watson:2008:EAC

[WKS^{+}08] Bruce W. Watson, Derrick G. Kourie, Tinus Strauss, Ernest Ketcha, and Loek Cleophas. Efficient automata constructions and approximate automata. International Journal of Foundations of Computer Science (IJFCS), 19 (1):185–193, February 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wei:2012:IRK

[WLC12] Yuechuan Wei, Chao Li, and Dan Cao. Improved related-key rectangle attack on the full HAS-160 encryption mode. International Journal of Foundations of Computer Science (IJFCS), 23(3):733–??, April 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wang:2003:DSB

[WLF03] Yu Wang, Xiang-Yang Li, and Ophir Frieder. Distributed spanners with bounded degree for wireless ad hoc networks. International Journal of Foundations of Computer Science (IJFCS), 14(2):183– ??, April 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wang:2021:SBS

[WLZT21] Yihong Wang, Cheng-Kuan Lin, Shuming Zhou, and Tao Tian. Subgraph-based strong Menger connectivity of hypercube and exchanged hypercube. International Journal of Foundations of Computer Science (IJFCS), 32(03):305– 330, April 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500179.

Wilson:2005:CPP

[WM05] Lucas A. Wilson and Michelle D. Moore. Crosspollinating parallel genetic algorithms for multiobjective search and optimization. International Journal of Foundations of Computer Science (IJFCS), 16(2):261–??, April 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ware:2013:CVG

[WM13] Simon Ware and Robi Malik. Compositional verification of the generalized nonblocking property using abstraction and canonical automata. International Journal of Foundations of Computer Science (IJFCS), 24(8):1183–??, December 2013. CODEN IFC-SEN. ISSN 0129-0541.

Wang:2019:BSC

[WNF19] Gang Wang, Min-Yao Niu, and Fang-Wei Fu. Bounds on subspace codes based on orthogonal space over finite fields of characteristic 2. International Journal of Foundations of Computer Science (IJFCS), 30(5):735– 757, August 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119500199.

Wang:2020:CRL

[WNF20] Gang Wang, Min-Yao Niu, and Fang-Wei Fu. Constructions of (r, t)-LRC based on totally isotropic subspaces in symplectic space over finite fields. International Journal of Foundations of Computer Science (IJFCS), 31(03):327– 339, April 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500112.

Wu:2003:COM

[WO03] Jie Wu and Stephen Olariu.

On cost-optimal merge of two intransitive sorted sequences. International Journal of Foundations of Computer Science (IJFCS), 14 (1):99–??, February 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wong:1996:AFM

[Won96] P.-K. Wong. An algorithm for finding a maximum cycle of bipartite graphs with large degrees. International Journal of Foundations of Computer Science (IJFCS), 7(4):301–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wong:2001:AFL

[Won01] Pak-Ken Wong. An algorithm for finding longest cycles in certain bipartite graphs. International Journal of Foundations of Computer Science (IJFCS), 12 (4):445–454, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wiedermann:2008:WMC

[WP08] Jiří Wiedermann and Dana Pardubská. Wireless mobile computing and its links to descriptive complexity. International Journal of Foundations of Computer Science (IJFCS), 19(4):887– 913, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wen:2021:SLA

[WPX^{+}21] Tangliu Wen, Jie Peng, Jinyun Xue, Zhen You, and Lan Song. Strict linearizability and abstract atomicity. International Journal of Foundations of Computer Science (IJFCS), 32 (01):1–35, January 2021. CODEN IFCSEN. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054121500015.

Wang:2016:NCO

[WPZ16] Changyuan Wang, Daiyuan Peng, and Limengnan Zhou.

New constructions of optimal frequency-hopping sequence sets with low-hit-zone. International Journal of Foundations of Computer Science (IJFCS), 27(1):53– ??, January 2016. CODEN IFCSEN. ISSN 0129-0541.

Wu:1997:MCN

[WQ97] Jie Wu and Haifeng Qian.

Multitriangle: a constant node degree interconnection network. International Journal of Foundations of Computer Science (IJFCS), 8(2):187–??, June 1997. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wei:2016:VNS

[WQY16] Zongtian Wei, Nannan Qi, and Xiaokui Yue. Vertexneighbor-scattering number of bipartite graphs. International Journal of Foundations of Computer Science (IJFCS), 27(4):501–??, June 2016. CODEN IFCSEN. ISSN 0129-0541.

Wu:2016:ESS

[WR16] Peng Wu and Minsoo Ryu. EDZL scheduling and schedulability analysis for performance asymmetric multiprocessors. International Journal of Foundations of Computer Science (IJFCS), 27(1):1–??, January 2016. CODEN IFC-SEN. ISSN 0129-0541.

Wei:2003:EAS

[WRNK03] David S. L. Wei, Sanguthevar Rajasekaran, Kshirasagar Naik, and Sy-Yen Kuo. Efficient algorithms for selection and sorting of large distributed files on de Bruijn and hypercube structures. International Journal of Foundations of Computer Science (IJFCS), 14(6):1129– ??, December 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wang:2020:CBS

[WWT20] Zhongxiao Wang, Xiangyu Wang, and Tian Tian. Constructing de Bruijn sequences based on a new necessary condition. International Journal of Foundations of Computer Science (IJFCS), 31(03):301– 312, April 2020. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054120500094.

Wang:2016:RRU

[WXF16] Yongjia Wang, Xi Xiong, and Haining Fan. GF(2^{n}) redundant representation using matrix embedding for irreducible trinomials. International Journal of Foundations of Computer Science (IJFCS), 27(4):463–??, June 2016. CODEN IFC-SEN. ISSN 0129-0541.

Wu:2005:EEN

[WY05] Jie Wu and Shuhui Yang.

Energy-efficient node scheduling models in sensor networks with adjustable ranges. International Journal of Foundations of Computer Science (IJFCS), 16(1):3– ??, February 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Wu:2022:GCN

[WY22] Jian Wu and Chaoyu Yang. Graph convolutional network-guided mine gas concentration predictor. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):771–785, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905412242014X.

Wei:2015:CPK

[WZ15] Puwen Wei and Yuliang Zheng. On the construction of public key encryption with sender recovery. International Journal of Foundations of Computer Science (IJFCS), 26(1):1–??, January 2015. CODEN IFC-SEN. ISSN 0129-0541.

Wu:2019:PFI

[WZCH19] Libing Wu, Yubo Zhang, Kim-Kwang Raymond Choo, and Debiao He. Pairingfree identity-based encryption with authorized equality test in online social networks. International Journal of Foundations of Computer Science (IJFCS), 30(4):647– 664, June 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400185.

Xu:2002:MHM

[XBE02] G. Xu, C. L. Bajaj, and S. Evans. C^{1} modeling with hybrid multiple-sided Apatches. International Journal of Foundations of Computer Science (IJFCS), 13 (2):261–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Xu:2015:CNP

[XC15] Guangkui Xu and Xiwang Cao. Constructing new piecewise differentially 4uniform permutations from known APN functions. International Journal of Foundations of Computer Science (IJFCS), 26(5):599–??, August 2015. CODEN IFC-SEN. ISSN 0129-0541.

Xu:2016:FOR

[XCC16] Hui-Jie Xu, Wan-Dong Cai, and Gui-Rong Chen. Forums-oriented research on the spreading and inhibition of rumors. International Journal of Foundations of Computer Science (IJFCS), 27(7):845–??, November 2016. CODEN IFCSEN. ISSN 0129-0541.

Xu:2020:SBF

[XCMT20] Shanding Xu, Xiwang Cao, Jiafu Mi, and Chunming Tang. Simplified bounds on FHSs set and its strictly optimal construction. International Journal of Foundations of Computer Science (IJFCS), 31(04):499– 513, June 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500215.

Xu:2016:OFH

[XCX16] Shanding Xu, Xiwang Cao, and Guangkui Xu. Optimal frequency-hopping sequence sets based on cyclotomy. International Journal of Foundations of Computer Science (IJFCS), 27 (4):443–??, June 2016. CO-DEN IFCSEN. ISSN 0129-0541.

Xu:2017:SCQ

[XCX17] Guangkui Xu, Xiwang Cao, and Shanding Xu. Several classes of quadratic ternary bent, near-bent and 2-plateaued functions. International Journal of Foundations of Computer Science (IJFCS), 28(1):1–18, January 2017. CODEN IFC-SEN. ISSN 0129-0541.

Xuan:2003:CSF

[XFJ03] B. Bui Xuan, A. Ferreira, and A. Jarry. Computing shortest, fastest, and foremost journeys in dynamic networks. International Journal of Foundations of Computer Science (IJFCS), 14(2):267–??, April 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Xu:2002:LSB

[XHLF02] W. Xu, R. Hammersley, K. Lu, and D. Fussell. Lossless subdivision-based multiresolution representation of arbitrary triangle meshes using kite trees. International Journal of Foundations of Computer Science (IJFCS), 13(2):243–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Xu:2004:MBU

[XLC^{+}04] Yinlong Xu, Li Lin, Guoliang Chen, Yingyu Wan, and Weijun Guo. Multicasting and broadcasting in undirected WDM networks and QoS extensions of multicasting. International Journal of Foundations of Computer Science (IJFCS), 15(1):187–??, February 2004. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Xu:2019:NPF

[XLZ19] Xiaofang Xu, Chunlei Li, and Xiangyong Zeng. Nonsingular polynomials from feedback shift registers. International Journal of Foundations of Computer Science (IJFCS), 30(3):469– 487, 2019. ISSN 0129-0541.

Xu:2006:SSA

[XS06] Zhenyu Xu and Pradip K.

Srimani. Self-stabilizing anonymous leader election in a tree. International Journal of Foundations of Computer Science (IJFCS), 17(2):323–??, April 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Xu:2011:DAA

[XS11] Shihong Xu and Hong Shen.

A distributed approximation algorithm for faulttolerant metric facility location. International Journal of Foundations of Computer Science (IJFCS), 22 (5):1019–1034, August 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Xu:2022:ADA

[XWL^{+}22] Li-Jun Xu, Shou-Yu Wei, Xiao-Qing Lu, Ze-Hua He, and Jia-Ming Zhu. Algorithm design for asset trading under multiple factors. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):867–886, September– November 2022. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122420199.

Xie:2022:SSP

[XWY^{+}22] Kai Xie, Zijian Wei, Kang Yin, Songsong Li, Xinyan Yao, and Xiaoyu Zhou. Structural synthesis of PLC program for realtime specification patterns. International Journal of Foundations of Computer Science (IJFCS), 33(6– 7):903–929, September– November 2022. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122420217.

Xu:2019:PPS

[XZL^{+}19] Zifeng Xu, Fucai Zhou, Yuxi Li, Jian Xu, and Qiang Wang. Privacy-preserving subgraph matching protocol for two parties. International Journal of Foundations of Computer Science (IJFCS), 30(4):571– 588, June 2019. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054119400136.

Xiao:2016:ACT

[XZS16] Zibi Xiao, Xiangyong Zeng, and Zhimin Sun. 2-adic complexity of two classes of generalized cyclotomic binary sequences. International Journal of Foundations of Computer Science (IJFCS), 27(7):879–??, November 2016. CODEN IFCSEN. ISSN 0129-0541.

Xu:2021:FTE

[XZW^{+}21] Xirong Xu, Huifeng Zhang, Ziming Wang, Qiang Zhang, and Peng Zhang. (n − 2)- fault-tolerant edge-pancyclicity of crossed cubes CQn. International Journal of Foundations of Computer Science (IJFCS), 32(03):289– 304, April 2021. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500167.

Xu:2019:FTM

[XZY19] Liqiong Xu, Shuming Zhou, and Weihua Yang. Faulttolerant maximal localconnectivity on Cayley graphs generated by transpositions. International Journal of Foundations of Computer Science (IJFCS), 30(8):1301–1315, December 2019. ISSN 0129-

0541. URL https://www.
worldscientific.com/doi/
10.1142/S0129054119500278.

Xu:2019:FTP

[XZZY19] Xirong Xu, Huifeng Zhang, Sijia Zhang, and Yuansheng Yang. Fault-tolerant panconnectivity of augmented cubes AQ_{n}. International Journal of Foundations of Computer Science (IJFCS), 30(8):1247– 1278, December 2019. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054119500254.

Yahalom:2012:TFP

[Yah12] Orly Yahalom. Testing for forbidden posets in ordered rooted forests. International Journal of Foundations of Computer Science (IJFCS), 23(6):1405–??, September 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yamakami:2003:AQF

[Yam03] Tomoyuki Yamakami. Analysis of quantum functions. International Journal of Foundations of Computer Science (IJFCS), 14(5):815– ??, October 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yang:2021:SKC

[Yan21] Yuxing Yang. Super C_{k} and sub-C_{k} connectivity of kary n-cube networks. International Journal of Foundations of Computer Science (IJFCS), 32(02):151– 162, February 2021. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054121500088.

Ye:2006:CTS

[YB06] Yuli Ye and Janusz Brzozowski. Covering of transient simulation of feedbackfree circuits by binary analysis. International Journal of Foundations of Computer Science (IJFCS), 17(4):949– ??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yigit:2019:NLR

[YB19] Esin Yiğit and Zeynep Nihan Berberler. A note on the link residual closeness of graphs under join operation. International Journal of Foundations of Computer Science (IJFCS), 30(3):417– 424, 2019. ISSN 0129-0541.

Yildirim:2022:IRT

[YB22] Halil İbrahim Yildirim and Zeynep Nihan Berberler. Isolated rupture in thorny networks. International Journal of Foundations of Computer Science (IJFCS), 33(05):429– 438, August 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500083.

Yildirim:2023:IRC

[YB23] Halil İbrahim Yildirim and Zeynep Nihan Berberler. Isolated rupture in composite networks. International Journal of Foundations of Computer Science (IJFCS), 34(04):363– 371, June 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500204.

Yu:2011:RSV

[YBI11] Fang Yu, Tevfik Bultan, and Oscar H. Ibarra. Relational string verification using multi-track automata. International Journal of Foundations of Computer Science (IJFCS), 22(8): 1909–1924, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yamauchi:2011:RCE

[YBM11] Yukiko Yamauchi, Doina Bein, and Toshimitsu Masuzawa. Reliable communication on emulated channels resilient to transient faults. International Journal of Foundations of Computer Science (IJFCS), 22 (5):1099–1122, August 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yuan:2011:LMF

[YCL11] Allen Yuan, Eddie Cheng, and László Lipták. Linearly many faults in (n, k)- star graphs. International Journal of Foundations of Computer Science (IJFCS), 22(7):1729–1745, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yang:2010:CMI

[YCTW10] Jinn-Shyong Yang, Jou-Ming Chang, Shyue-Ming Tang, and Yue-Li Wang. Constructing multiple independent spanning trees on recursive circulant graphs G(2^{m}, 2). International Journal of Foundations of Computer Science (IJFCS), 21(1):73–90, February 2010. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yang:2008:SAS

[YDI08] Linmin Yang, Zhe Dang, and Oscar H. Ibarra. On stateless automata and P systems. International Journal of Foundations of Computer Science (IJFCS), 19(5):1259–1276, October 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yen:2008:DCA

[Yen08] Hsu-Chun Yen. Decidability and complexity analysis of forbidden state problems for discrete event systems. International Journal of Foundations of Computer Science (IJFCS), 19(4):999– 1013, August 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yen:2009:PDS

[Yen09] Hsu-Chun Yen. Path decomposition and semilinearity of Petri nets. International Journal of Foundations of Computer Science (IJFCS), 20(4):581– 596, August 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Ye:2011:WCP

[YH11] Deshi Ye and Qinming He. Worst-case performance evaluation on multiprocessor task scheduling with resource augmentation. International Journal of Foundations of Computer Science (IJFCS), 22(4):971– 982, June 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yamamoto:2014:TIV

[YHK14] Yoshiyuki Yamamoto, Kouichi Hirata, and Tetsuji Kuboyama. Tractable and intractable variations of unordered tree edit distance. International Journal of Foundations of Computer Science (IJFCS), 25(3):307–??, April 2014. CODEN IFCSEN. ISSN 0129-0541.

Yen:2013:P

[YI13] Hsu-Chun Yen and Oscar H. Ibarra. Preface. International Journal of Foundations of Computer Science (IJFCS), 24(7):941–??, November 2013. CODEN IFCSEN. ISSN 0129-0541.

Yli-Jyra:2005:ADG

[YJ05] Anssi Yli-Jyrä. Approximating dependency grammars through intersection of star-free regular languages. International Journal of Foundations of Computer Science (IJFCS), 16(3):565– ??, June 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yang:2023:RGU

[YKCW23] Zhiyao Yang, Pinhui Ke, Zhixiong Chen, and Chenhuang Wu. Results on the Gowers U2 norm of generalized Boolean functions. International Journal of Foundations of Computer Science (IJFCS), 34(04):373– 393, June 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500216.

Yu:2022:SAP

[YLX22] Liting Yu, Dongyan Liu, and Nairu Xu. Special aquatic products supply chain coordination considering bilateral green input in the context of high-quality development. International Journal of Foundations of Computer Science (IJFCS), 33(6– 7):819–844, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420175.

Yang:2014:PAG

[YLZ14] Haijun Yang, Minqiang Li, and Qinghua Zheng. Performance analysis of Grid architecture via queueing theory. International Journal of Foundations of Computer Science (IJFCS), 25(6):697– ??, September 2014. CO-DEN IFCSEN. ISSN 0129-0541.

Yao:2019:ACC

[YM19] Zhongyuan Yao and Yi Mu. ACE with compact ciphertext size and decentralized sanitizers. International Journal of Foundations of Computer Science (IJFCS), 30(4):531– 549, June 2019. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054119400112.

Yu:2017:ACA

[YMC^{+}17] Gang Yu, Xiaoxiao Ma, Zhenfu Cao, Guang Zeng, and Wenbao Han. Accountable CP-ABE with public verifiability: How to effectively protect the outsourced data in cloud. International Journal of Foundations of Computer Science (IJFCS), 28(6):705–??, September 2017. CODEN IFCSEN. ISSN 0129-0541.

Yahia:2008:P

[YN08] Sadok Ben Yahia and Engelbert Mephu Nguifo. Preface. International Journal of Foundations of Computer Science (IJFCS), 19(2):251– 254, April 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yakaryilmaz:2013:TBS

[YS13] Abuzer Yakaryilmaz and A. C. Cem Say. Tight bounds for the space complexity of nonregular language recognition by realtime machines. International Journal of Foundations of Computer Science (IJFCS), 24(8):1243–??, December 2013. CODEN IFC-SEN. ISSN 0129-0541. Yakaryilmaz:2016:DST

[YSD16] Abuzer Yakaryilmaz, A. C. Cem Say, and H. Gökalp Demirci. Debates with small transparent quantum verifiers. International Journal of Foundations of Computer Science (IJFCS), 27(2):283– ??, February 2016. CODEN IFCSEN. ISSN 0129-0541.

Yang:2000:GMC

[YSM^{+}00a] Z.-H. Yang, C.-Z. Sun, Y. Miao, A. Sattar, and Y. Y. Yang. Guaranteed mutually consistent checkpointing in distributed computations. International Journal of Foundations of Computer Science (IJFCS), 11(1):153–166, 2000. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yang:2000:PRP

[YSM^{+}00b] Z.-H. Yang, C.-Z. Sun, Y. Miao, A. Sattar, and Y. Y. Yang. Part 2 (regular papers). International Journal of Foundations of Computer Science (IJFCS), 11 (1):153–??, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yang:2002:CSB

[YTLC02] Xiaoguang Yang, Shuo Tao, Rongjun Liu, and Maocheng Cai. Complexity of scenariobased portfolio optimization problem with var objective. International Journal of Foundations of Computer Science (IJFCS), 13(5):671– ??, October 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yamazaki:2001:CPP

[YTN01] Koich Yamazaki, Sei’ichi Tani, and Tetsuro Nishino. A characterization of k-th powers P_{n,k} of paths in terms of k-trees. International Journal of Foundations of Computer Science (IJFCS), 12(4):435–443 (or 435–444??), 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yang:2011:ACD

[YTP11] Yang Yang, Xiaohu Tang, and Udaya Parampalli. Authentication codes from difference balanced functions. International Journal of Foundations of Computer Science (IJFCS), 22(6): 1417–1429, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yu:2002:SI

[Yu02] S. Yu. Special issue. International Journal of Foundations of Computer Science (IJFCS), 13(1):1–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yu:2011:P

[Yu11] Sheng Yu. Preface. International Journal of Foundations of Computer Science (IJFCS), 22(7):1495–1498, November 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yue:2013:CIE

[Yue13] Houguang Yue. From computing to interaction: on the expressiveness of asynchronous pi-calculus. International Journal of Foundations of Computer Science (IJFCS), 24(3):349–??, April 2013. CODEN IFC-SEN. ISSN 0129-0541.

Yu:2006:SBM

[YW06] Fang Yu and Bow-Yaw Wang. SAT-based model checking for region automata. International Journal of Foundations of Computer Science (IJFCS), 17 (4):775–??, August 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yang:2020:KEL

[YW20] Minghui Yang and Jiejing Wen. On the k-error linear complexity of subsequences of d-ary Sidel’nikov sequences over prime field F_{d}. International Journal of Foundations of Computer Science (IJFCS), 31 (03):293–300, April 2020. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500082.

Yang:2022:IHD

[YW22] Bin Yang and Jie Wang. An improved helmet detection algorithm based on YOLO V4. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):887–902, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420205.

Yen:1994:SCR

[YWY94] H.-C. Yen, B.-Y. Wang, and M.-S. Yang. Some complexity results for rings of Petri nets. International Journal of Foundations of Computer Science (IJFCS), 5(3/4):??, 1994. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yu:2019:PSI

[YYW19] Yong Yu, Guomin Yang, and Huaxiong Wang. Preface: Special issue cryptography and provable security. International Journal of Foundations of Computer Science (IJFCS), 30 (4):489–492, June 2019. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119020027.

Yu:2007:SEO

[YZ07] Sheng Yu and Qing Zhao.

SC-expressions in objectoriented languages. International Journal of Foundations of Computer Science (IJFCS), 18(6):1441–1452, December 2007. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Yao:2021:TCC

[YZP21] Ting Yao, Shixin Zhu, and Binbin Pang. Triple cyclic codes over Fq + uFq. International Journal of Foundations of Computer Science (IJFCS), 32(02):115– 135, February 2021. CO-DEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500064.

Yan:2018:DAB

[YZY^{+}18] Shuo Yan, Yunyong Zhang, Binfeng Yan, Lin Yan, and Jinfeng Kou. Data associations between two hierarchy trees. International Journal of Foundations of Computer Science (IJFCS), 29(7):1181– 1201, November 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118500284.

Yu:2022:FTS

[YZZ22] Zhengqin Yu, Shuming Zhou, and Hong Zhang. Fault-tolerant strong Menger (edge) connectivity of DCC linear congruential graphs. International Journal of Foundations of Computer Science (IJFCS), 33(08): 1019–1032, December 2022. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054122500137.

Zajicek:2009:NSP

[Zaj09] Ondřej Zajíček. A note on scheduling parallel unit jobs on hypercubes. International Journal of Foundations of Computer Science (IJFCS), 20(2):341– 349, April 2009. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zanko:1991:PCM

[Zan91]  Viktoria Zanko.  #P-

completeness via many-one reductions. International Journal of Foundations of Computer Science (IJFCS), 2(1):77–??, March 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zantema:2000:FSE

[ZB00] H. Zantema and H. L. Bodlaender. Finding small equivalent decision trees is hard. International Journal of Foundations of Computer Science (IJFCS), 11 (2):343–354, 2000. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zantema:2002:SOD

[ZB02] H. Zantema and H. L. Bodlaender. Sizes of ordered decision trees. International Journal of Foundations of Computer Science (IJFCS), 13(3):445–??, 2002. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zhang:2005:AAC

[ZBS05] Lili Zhang and F. Blanchet-Sadri. Algorithms for approximate K-covering of strings. International Journal of Foundations of Computer Science (IJFCS), 16(6):1231–??, December 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zomaya:2005:ECP

[ZC05] Albert Y. Zomaya and Gerard Chan. Efficient clustering for parallel tasks execution in distributed systems. International Journal of Foundations of Computer Science (IJFCS), 16(2):281– ??, April 2005. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zhang:2013:RMS

[ZC13] En Zhang and Yongquan Cai. Rational multi-secret sharing scheme in standard point-to-point communication networks. International Journal of Foundations of Computer Science (IJFCS), 24(6):879–??, September 2013. CODEN IFCSEN. ISSN 0129-0541.

Zhou:2012:CFD

[ZCX12] Shuming Zhou, Lanxiang Chen, and Jun-Ming Xu. Conditional fault diagnosability of dual-cubes. International Journal of Foundations of Computer Science (IJFCS), 23(8):1729–??, December 2012. CODEN IFC-SEN. ISSN 0129-0541.

Zetzsche:2011:TUG

[Zet11] Georg Zetzsche. Toward understanding the generative capacity of erasing rules in matrix grammars. International Journal of Foundations of Computer Science (IJFCS), 22(2):411– 426, February 2011. CO-DEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zhou:2017:CEE

[ZFL^{+}17] Dongfang Zhou, Jianxi Fan, Cheng-Kuan Lin, Jingya Zhou, and Xi Wang. Cycles embedding in exchanged crossed cube. International Journal of Foundations of Computer Science (IJFCS), 28(1):61–76, January 2017. CODEN IFCSEN. ISSN 0129-0541.

Zhou:2013:NUB

[ZG13] Guangyan Zhou and Zongsheng Gao. A new upper bound for random (2 + p)- SAT by flipping two variables. International Journal of Foundations of Computer Science (IJFCS), 24(6):899– ??, September 2013. CO-DEN IFCSEN. ISSN 0129-0541.

Zhou:2018:CBG

[ZGCZ18] Caixue Zhou, Guangyong Gao, Zongmin Cui, and Zhiqiang Zhao. Certificatebased generalized ring signcryption scheme. International Journal of Foundations of Computer Science (IJFCS), 29(6):1063– 1088, September 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054118500211.

Zhuang:2022:CDG

[ZGL^{+}22] Hongbin Zhuang, Wenzhong Guo, Xiaoyan Li, Ximeng Liu, and Cheng-Kuan Lin. The component diagnosability of general networks. International Journal of Foundations of Computer Science (IJFCS), 33(01):67– 89, January 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500374.

Zhang:2006:AWO

[ZH06] Huaming Zhang and Xin He. An application of wellorderly trees in graph drawing. International Journal of Foundations of Computer Science (IJFCS), 17 (5):1129–??, October 2006. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zha:2013:CNA

[ZH13] Zhengbang Zha and Lei Hu. Constructing new APN functions from known PN functions. International Journal of Foundations of Computer Science (IJFCS), 24(8):1209–??, December 2013. CODEN IFC-SEN. ISSN 0129-0541.

Zhao:2019:GCB

[ZH19] Shu-Li Zhao and Rong-Xia Hao. The generalized connectivity of bubblesort star graphs. International Journal of Foundations of Computer Science (IJFCS), 30(5):793– 809, August 2019. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119500229.

Zhou:2022:SCO

[ZH22] Hongyu Zhou and Xinmin Hou. Strongly connected orientation with minimum lexicographic order of indegrees. International Journal of Foundations of Computer Science (IJFCS), 33(02):149– 153, February 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500046.

Zhang:2017:FCP

[Zha17] Meng Zhang. Fast convolutions of packed strings and pattern matching with wildcards. International Journal of Foundations of Computer Science (IJFCS), 28 (3):289–??, April 2017. CO-DEN IFCSEN. ISSN 0129-0541.

Zhong:2002:RCO

[Zho02]  N. Zhong.  Representa-

tion and construction of ontologies for Web intelligence. International Journal of Foundations of Computer Science (IJFCS), 13 (4):555–??, 2002. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zhang:2011:WAF

[ZHZ11] Meng Zhang, Liang Hu, and Yi Zhang. Weighted automata for full-text indexing. International Journal of Foundations of Computer Science (IJFCS), 22(4):921– 943, June 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zhou:2019:LBS

[ZK19] Guangyan Zhou and Rui Kang. On the lower bounds of (1, 0)-super solutions for random k-SAT. International Journal of Foundations of Computer Science (IJFCS), 30(2):247– 254, February 2019. ISSN 0129-0541.

Zhang:2022:ICO

[ZL22] Wei Zhang and Zhiguang Li. Identifying the configurations to operating efficiency in China’s life insurance industry using fuzzy-set qualitative comparative analysis. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):845–865, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420187.

Zhou:2021:ERV

[ZLG21] Bo Zhou, Zhenan Li, and Haiyan Guo. Extremal results on vertex and link residual closeness. International Journal of Foundations of Computer Science (IJFCS), 32(08):921– 941, December 2021. CO-DEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054121500295.

Zhang:2011:NBF

[ZLL11] Yin Zhang, Meicheng Liu, and Dongdai Lin. On the nonexistence of bent functions. International Journal of Foundations of Computer Science (IJFCS), 22 (6):1431–1438, September 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zhang:2020:ECD

[ZLL20] Shunzhe Zhang, Dong Li, and Huiqing Liu. On g-extra conditional diagnosability of twisted hypercubes under MM^{⋆} model. International Journal of Foundations of Computer Science (IJFCS), 31(04):445– 459, June 2020. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054120500185.

Zhang:2023:EEC

[ZLL23] Mingzu Zhang, Hongxi Liu, and Pingping Li. Embedded edge-connectivity reliability evaluation of augmented hypercube interconnection networks. International Journal of Foundations of Computer Science (IJFCS), 34(01):1–10, January 2023. CODEN IFCSEN. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122500150.

Zhu:2017:PSN

[ZLW^{+}17] Youwen Zhu, Xingxin Li, Jian Wang, Yining Liu, and Zhiguo Qu. Practical secure naïve Bayesian classification over encrypted big data in cloud. International Journal of Foundations of Computer Science (IJFCS), 28(6):683– ??, September 2017. CO-DEN IFCSEN. ISSN 0129-0541.

Zdarek:2011:TBI

[ŽM11] Jan Žďárek and Bořivoj Melichar. Tree-based 2D indexing. International Journal of Foundations of Computer Science (IJFCS), 22 (8):1893–1907, December 2011. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zomaya:2001:SIP

[Zom01a] A. Y. Zomaya. Special issue: Part 1. International Journal of Foundations of Computer Science (IJFCS), 12 (5):559–??, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zomaya:2001:S

[Zom01b] Albert Y. Zomaya. Scheduling. International Journal of Foundations of Computer Science (IJFCS), 12 (5):559–564, 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zomaya:2001:STA

[Zom01c] Albert Y. Zomaya. Scheduling: Theory and applications: Guest Editor’s preface. International Journal of Foundations of Computer Science (IJFCS), 12(5):559– 564, October 2001. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zomaya:2003:MCO

[Zom03] Albert Y. Zomaya. Mobile computing: Opportunities for parallel algorithms research. International Journal of Foundations of Computer Science (IJFCS), 14 (1):19–??, February 2003. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zhou:2017:IBB

[ZPXX17] Fucai Zhou, Su Peng, Jian Xu, and Zifeng Xu. Identity-based batch provable data possession with detailed analyses. International Journal of Foundations of Computer Science (IJFCS), 28(6):743–??, September 2017. CODEN IFCSEN. ISSN 0129-0541.

Zheng:2012:SLR

[ZQL12] Shenggen Zheng, Daowen Qiu, and Lvzhou Li. Some languages recognized by two-way finite automata with quantum and classical states. International Journal of Foundations of Computer Science (IJFCS), 23 (5):1117–??, August 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zheng:2022:OOL

[ZSG^{+}22] Yangfeng Zheng, Zheng Shao, Zhanghao Gao, Mingming Deng, and Xuesong Zhai. Optimizing the online learners’ verbal intention classification efficiency based on the multi-head attention mechanism algorithm. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):717–733, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420114.

Zhou:2014:NWC

[ZSW14] Junping Zhou, Weihua Su, and Jianan Wang. New worst-case upper bound for counting exact satisfiability. International Journal of Foundations of Computer Science (IJFCS), 25(6):667– ??, September 2014. CO-DEN IFCSEN. ISSN 0129-0541.

Zhang:2022:SACa

[ZWC^{+}22] Qi Zhang, Guang Wen, Zhixin Chen, Qin Zhou, Guoqi Xiang, Guangchun Yang, and Xuegang Zhang. Sensitivity analysis contact reliability of VH-CATT cylindrical gear and its reliability with material strength degradation. International Journal of Foundations of Computer Science (IJFCS), 33 (6–7):691–716, September– November 2022. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122420102.

Zheng:2014:CMV

[ZWCL14] Jia Zheng, Baofeng Wu, Yufu Chen, and Zhuojun Liu. Constructing 2mvariable Boolean functions with optimal algebraic immunity based on polar decomposition of F_{2^{2m}}^{∗}. International Journal of Foundations of Computer Science (IJFCS), 25(5):537–??, August 2014. CODEN IFC-SEN. ISSN 0129-0541.

Zhang:1996:EDB

[ZWS96] K.-Z. Zhang, J. T. L. Wang, and D. Shasha. On the editing distance between undirected acyclic graphs. International Journal of Foundations of Computer Science (IJFCS), 7(1):43–??, 1996. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zhou:2014:OSN

[ZWW^{+}14] Yu Zhou, Lin Wang, Weiqiong Wang, Xinfeng Dong, and Xiaoni Du. One sufficient and necessary condition on balanced Boolean functions with σ_{f} = 2^{2n} + 2^{n+3}(n ≥ 3). International Journal of Foundations of Computer Science (IJFCS), 25(3):343–??, April 2014. CODEN IFCSEN. ISSN 0129-0541.

Zhou:2012:PTE

[ZYLW12] Junping Zhou, Minghao Yin, Xiangtao Li, and Jinyan Wang. Phase transitions of expspace-complete problems: a further step. International Journal of Foundations of Computer Science (IJFCS), 23(1):173– 184, January 2012. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zhang:2014:CGC

[ZYYH14] Jing Zhang, Xiaofan Yang, Cui Yu, and Li He. The congestion of generalized cube communication pattern in linear array network. International Journal of Foundations of Computer Science (IJFCS), 25(3):263–??, April 2014. CODEN IFC-SEN. ISSN 0129-0541.

Zhang:2019:LRH

[ZYZ^{+}19] Yinghui Zhang, Menglei Yang, Dong Zheng, Tiantian Zhang, Rui Guo, and Fang Ren. Leakage-resilient hierarchical identity-based encryption with recipient anonymity. International Journal of Foundations of Computer Science (IJFCS), 30(5):665–681, August 2019. CODEN IFCSEN. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054119400197.

Zhao:2018:CEC

[ZYZX18] Shuli Zhao, Weihua Yang, Shurong Zhang, and Liqiong Xu. Component edge connectivity of hypercubes. International Journal of Foundations of Computer Science (IJFCS), 29(6):995– 1001, September 2018. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S012905411850017X.

Zhang:2018:SER

[ZZ18] Meng Zhang and Yi Zhang.

Space-efficient representations for Glushkov automata. International Journal of Foundations of Computer Science (IJFCS), 29 (7):1089–1105, November 2018. CODEN IFCSEN. ISSN 0129-0541. URL

https://www.worldscientific.
com/doi/10.1142/S0129054118500223.

Zhu:2015:SBT

[ZZC15] Xishun Zhu, Xiangyong Zeng, and Yuan Chen. Some binomial and trinomial differentially 4-uniform permutation polynomials. International Journal of Foundations of Computer Science (IJFCS), 26(4):487–??, June 2015. CODEN IFCSEN. ISSN 0129-0541.

Zhang:2022:SACb

[ZZC22] Jiajing Zhang, Tingting Zhang, and Jinlan Chen. Sentiment analysis of Chinese reviews based on BiTCN-attention model. International Journal of Foundations of Computer Science (IJFCS), 33(6– 7):755–770, September– November 2022. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122420138.

Zhang:2023:FTS

[ZZN23] Hong Zhang, Shuming Zhou, , and Baohua Niu. Fault-tolerance of star graph based on subgraph fault pattern. International Journal of Foundations of Computer Science (IJFCS), 34(05): 469–485, August 2023. ISSN 0129-0541. URL https:

                                        //www.worldscientific.
                                        com/doi/10.1142/S0129054122500277.

Zhi-Zhong:1991:CCO

[ZZT91] Chen Zhi-Zhong and Seinosuke Toda. On the complexity of computing optimal solutions. International Journal of Foundations of Computer Science (IJFCS), 2(3): 207–220, September 1991. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Zhang:2016:OTS

[ZZZ16] Wenming Zhang, E. Zhang, and Feifeng Zheng. Online two stage k-search problem and its competitive analysis. International Journal of Foundations of Computer Science (IJFCS), 27(6):653– ??, September 2016. CO-DEN IFCSEN. ISSN 0129-0541.

Zhang:2023:CCA

[ZZZ23] Hong Zhang, Shuming Zhou, and Qifan Zhang. Component connectivity of alternating group networks and godan graphs. International Journal of Foundations of Computer Science (IJFCS), 34(04):395– 410, June 2023. ISSN 0129-0541. URL https:

//www.worldscientific.
com/doi/10.1142/S0129054122500228.